### v1

# A Survey on Self-play Methods in Reinforcement Learning

Abstract
Self-play, characterized by agents’ interactions with copies or past versions of itself, has recently gained prominence in reinforcement learning. This paper first clarifies the preliminaries of self-play, including the multi-agent reinforcement learning framework and basic game theory concepts. Then it provides a unified framework and classifies existing self-play algorithms within this framework. Moreover, the paper bridges the gap between the algorithms and their practical implications by illustrating the role of self-play in different scenarios. Finally, the survey highlights open challenges and future research directions in self-play. This paper is an essential guide map for understanding the multifaceted landscape of self-play in RL.
Index Terms:
Self-play, reinforcement learning, game theory, multi-agent
I Introduction
Reinforcement learning (RL) represents a significant paradigm [1] within machine learning, concerned with the optimization of decision-making processes through interaction with an environment. It’s fundamentally modeled using a Markov decision process (MDP), a mathematical framework that describes an environment in terms of states, actions, transitions, and rewards. Within an MDP, agents operate by observing states, executing actions according to defined policies, receiving subsequent rewards, and transitioning to subsequent states. The primary goal of RL algorithms is to derive the optimal policy that yields the maximum expected accumulated reward over time. Deep RL extends traditional RL by employing deep neural networks as function approximators [2]. This fusion of deep learning with RL has been instrumental in handling high-dimensional state spaces, contributing to breakthroughs in various complex tasks.
Moreover, the transition from single-agent to multi-agent reinforcement learning (MARL) introduces complex dynamics [3, 4, 5]. In MARL, the interdependence of agents’ actions introduces significant challenges, as the environment appears non-stationary to each agent. The main issues in MARL are coordination, communication, and equilibrium selection, particularly in competitive scenarios. These challenges often lead to difficulties in achieving convergence, maintaining stability, and efficiently exploring the solution space.
With the help of game theory, a mathematical framework that models the interactions between multiple decision-makers, self-play emerges as an elegant solution to some inherent challenges in MARL. By addressing issues such as non-stationarity and coordination, self-play offers an approach where an agent interacts with copies or past versions of itself [6, 7]. This method promises a more stable and manageable learning process. The capabilities of self-play extend to a wide range of scenarios, including its high-profile applications in Go [8, 9, 10, 11], chess [10, 11], poker [12, 13], and video games [14, 15]. In these scenarios, it has developed strategies that surpass human expertise. Although the application of self-play is extensive and promising, it is accompanied by limitations, such as the potential convergence to suboptimal strategies and significant computational requirements[8, 10].
Although some research takes a broad perspective through empirical game-theoretic analysis (EGTA) [16], it is important to note that there are relatively few comprehensive surveys focusing exclusively on self-play. Among these, some studies address the theoretical safety of self-play [17], while others develop an algorithmic framework for self-play that unfortunately does not accommodate the Policy-Space Response Oracle (PSRO) series of algorithms [18]. Furthermore, another study concentrates exclusively on PSRO [19]. Although these varied studies are valuable, they do not offer a perspective that fully captures the breadth and depth of self-play. Therefore, this survey aims to bridge this gap.
The survey is organized as follows. Sec. II introduces the background of self-play, including the RL framework and basic game theory concepts. Sec. III proposes a unified framework and then categorizes existing self-play algorithms into four categories based on this framework, clarifying the self-play landscape. In Sec. IV, a comprehensive analysis is performed to illustrate how self-play is applied in various scenarios. Sec. V describes open problems in self-play and explores future research directions. Finally, Sec. VI concludes the survey on self-play.
II Preliminaries
In this section, we first introduce the framework of RL. Next, we present the basic game theory concepts and the typical evaluation metrics used in self-play.
II-A RL Framework
[…]
In the RL context, agents interact with the environment based on the subsequent protocol: At each discrete time step $t$, every agent $i$ receives an observation $o_{i,t}$ from the environment and selects an action based on a stochastic policy $\pi_{\theta_{i}}:\mathcal{O}_{i}\times\mathcal{A}_{i}\rightarrow[0,1]$, where $\theta_{i}$ is the parameters. After receiving the joint actions $\mathbf{a}_{t}=(a_{1,t},\cdots,a_{n,t})$, the environment undergoes a transition from the current state $s_{t}$ to a subsequent state $s_{t+1}$ according to the transition function $\mathcal{P}$ and sends a reward $r_{i,t+1}$ to every agent $i$. The ultimate goal of agent $i$ is to maximize the expected discount accumulated rewards: $\mathbb{E}_{\pi_{\theta_{i}}}[\sum_{t=0}^{\infty}\gamma^{t}r_{i,t}]$.
II-B Game Theory Concepts
II-B1 (Im)Perfect Information and (In)Complete Information
In a game characterized by perfect information, only one player moves at a time. Each player has a comprehensive understanding of the current game state, the full history of moves that have been made, and all potential future developments. If these conditions are not met, the game is considered to have imperfect information [25, 26]. In a game of incomplete information, there exists at least one player who is unaware of the payoff of another player; otherwise, it is a game of complete information [27].
For instance, Go is a game of both perfect and complete information. Players have full awareness of the entire game structure, including all possible moves, and they can see every move made by their opponent as they take turns to act (perfect information). Furthermore, if the outcomes are considered binary, such as win or loss, the payoff for the players is known to both sides (complete information).
II-B2 Normal-Form and Extensive-Form
The normal-form and extensive-form are two different ways of representing games in game theory. If a game $\mathcal{G}$ is represented in the normal-form, it can be expressed by $\mathcal{G}=(\mathcal{N},\mathbf{\Pi},\mathbf{u})$. $\mathcal{N}=\{1,2,\cdots,n\}$ denotes the players. $\mathbf{\Pi}=\Pi_{1}\times\cdots\times\Pi_{n}$ is pure strategy space of all players. A vector $\bm{\pi}=(\pi_{1},\cdots,\pi_{n})\in\mathbf{\Pi}$ is called a strategy profile. A pure strategy defines a specific and deterministic action for a player in a game, while a mixed strategy designates a probability distribution over the set of pure strategies, allowing randomized actions. A mixed strategy for the player $i$ is a probability distribution $\sigma_{i}\in\Delta(\Pi_{i})$, where $\Delta$ is a probability simplex. $\mathbf{u}=(u_{1},\cdots,u_{n})$, where $u_{i}:\mathbf{\Pi}\rightarrow\mathbb{R}$, is a utility function that assigns a real-valued payoff to each player $i$. If $\forall\bm{\pi}\in\mathbf{\Pi},\sum_{i}u_{i}(\bm{\pi})=0$, the game is a zero-sum game, otherwise it is a general-sum game. If $\Pi_{1}=\cdots=\Pi_{n}$ and the payoffs are invariant under any permutation of the players’ strategies, the game is a symmetric game. If finite players (especially two players) are involved and each player has a finite set of strategies, a game in the normal-form can be directly depicted in a matrix.
Specifically, in two-player zero-sum symmetric normal-form games, the pure strategy space for both player 1 and player 2 is identical, denoted by $\Pi$, such that $\Pi=\Pi_{1}=\Pi_{2}$. As the utility function $u_{1}(\pi_{i},\pi_{j})=-u_{2}(\pi_{i},\pi_{j})$, for simplicity, we can use only one utility function $u$ such that $\forall\pi_{i},\pi_{j}\in\Pi$, if $\pi_{i}$ beats $\pi_{j}$, then $u(\pi_{i},\pi_{j})=-u(\pi_{j},\pi_{i})>0$. The evaluation matrix captures the outcomes of the game by detailing the results of different strategies when they are played against each other: $A_{\Pi}=\{u(\pi_{i},\pi_{j}):\pi_{i},\pi_{j}\in\Pi\times\Pi\}$.
If a game is represented in the extensive-form, it is expressed sequentially, illustrating the sequence of moves, choices made by the players, and the information available to each player during decision-making. Typically, a game in the extensive-form is represented by a game tree. This tree demonstrates the sequential and potentially conditional nature of decisions. Moreover, if player $i$ has perfect recall, it means that player $i$ remembers which action they have taken in the past. A game $\mathcal{G}$ represented in the extensive-form can be expressed by $\mathcal{G}=(\mathcal{N}\cup\{c\},H,Z,P,\mathcal{I},A,\mathbf{u})$. $\mathcal{N}=\{1,2,\cdots,n\}$ denotes a set of players. $c$ is chance and can be regarded as a special agent. $H$ represents a set of possible histories and $Z\subseteq H$ is a set of terminal histories. Order of moves is represented by a function $P(h)\in\mathcal{N}\cup\{c\}$ to indicate which player is to move, where $h\in H$. $\mathcal{I}$ denotes information set partitions and $\mathcal{I}_{i}$ denotes the information set partitions for player $i$. This implies that in an imperfect information game, if player $i$ reaches a history $h\in I_{i}$, where $I_{i}\in\mathcal{I}_{i}$ is a specific information set, player $i$ cannot distinguish which particular history $h\in I_{i}$ it is encountering. Action space is represented by $A(h)$ for a non-terminal history $h\in H$. For all non-terminal histories $h$ within an information set $I_{i}$, the available actions are the same; otherwise, they are distinguishable. Therefore we use $A(I_{i})$ to represent the available actions for the information set $I_{i}$. Utility functions is denoted by $\mathbf{u}=(u_{1},\cdots,u_{n})$, where $u_{i}:Z\rightarrow\mathbb{R}$. Together, these components define the structure and dynamics of an extensive-form game. Moreover, A strategy profile in an extensive-form game can be expressed by $\bm{\pi}=(\pi_{1},\cdots,\pi_{n})$, where $\pi_{i}$ maps each $I_{i}\in\mathcal{I}_{i}$ to a probability distribution over $A(I_{i})$. A subgame of an extensive-form game is a portion of the game that starts from a single initial node, includes all successors of any node within the subgame, and contains all nodes in the same information set as any node in the subgame.
The Prisoner’s Dilemma serves as a classic example to illustrate various concepts in game theory. In a modified version of the dilemma, the outcomes are as follows:
  • If one player confesses (C) and the other lies (L), the confessor will serve 1 year in jail, while the liar will serve 8 years.
  • If both players choose to confess, they will each serve 7 years behind bars.
  • If both players choose to lie, they will each serve only 2 years behind bars.
The classic scenario is known as the simultaneous Prisoner’s Dilemma, where two players must decide simultaneously whether to confess or lie, without knowledge of the other’s choice. Games played in this manner are referred to as static games. The normal-form representation is particularly suitable for static games, as it captures the simultaneous nature of decision-making (as depicted in Fig. 1a). Another variant is the sequential Prisoner’s Dilemma, where the second player makes their decision with knowledge of the first player’s action. Games of this nature are called dynamic games. The extensive-form representation is well-suited for dynamic games, as it can clearly illustrate the sequence of moves and the information available to each player at each decision point (as shown in Fig. 1d).
Furthermore, a normal-form representation can be transformed into an extensive-form representation. For instance, the transformation from Fig. 1a to Fig. 1b involves representing the simultaneous decisions in a tree structure. In this extensive representation, the dotted lines indicate that certain statuses belong to the same information set. Conversely, an extensive-form representation can also be transformed into a normal-form representation. For example, the transformation from Fig. 1d to Fig. 1c involves condensing the sequential decisions into a matrix format. In this case, the pure strategy space for player 2 is $2\times 2=4$, reflecting the need for player 2 to have a strategy for each of player 1’s possible actions (confess or lie). For instance, the notation for one of player 2’s pure strategies $C\rightarrow C,L\rightarrow C$ in Fig. 1c indicates that player 2 chooses to confess when he knows player 1 has chosen to confess and player 2 also chooses to confess when he knows player 1 has chosen to lie. Other symbols in Fig. 1c follow the same logic.
(a) Matrix Representation of simultaneous Prisoner’s Dilemma in normal-form.
(b) Game Tree Representation of simultaneous Prisoner’s Dilemma in extensive-form.
(c) Matrix Representation of sequential Prisoner’s Dilemma in normal-form.
(d) Game Tree Representation of sequential Prisoner’s Dilemma in extensive-form.
Figure 1: The example of Prisoner’s Dilemma.
Compared to normal-form games, extensive-form games introduce sequential decision-making, adding complexity to the game structure. Extensive-form games also have a close relationship with MGs. In MGs with simultaneous moves, agents’ actions are unknown to each other, creating various histories that are condensed into a single information set. The game’s utility is the sum of rewards discounted over time [28]. Beyond normal-form and extensive-form games, as well as MGs, the analysis of complex Markov or extensive-form games often employs a higher-level abstraction: the meta-game. The meta-game facilitates the exploration of policy learning within these games, focusing not on isolated actions but on broader strategies that arise from the game’s dynamics. In this advanced normal-form context, the policy population consists of the strategies currently employed by players. Meta-strategies are mixed strategies that assign probabilities over the policy population in the meta-game.
II-B3 Transitive Game and Non-transitive Game
For the sake of simplicity, we restrict our focus to two-player zero-sum symmetric games. In a transitive game, the strategies or outcomes adhere to a transitive relationship. Formally, $\forall\pi_{i},\pi_{j},\pi_{k}\in\Pi$, if $u(\pi_{i},\pi_{j})>0$ and $u(\pi_{j},\pi_{k})>0$, then it must follow that $u(\pi_{i},\pi_{k})>0$. This transitive property simplifies the strategic landscape, allowing for an ordinal ranking of strategies. Conversely, in a non-transitive game, $\exists\leavevmode\nobreak\ \pi_{i},\pi_{j},\pi_{k}\in\Pi$ such that $u(\pi_{i},\pi_{j})>0$ and $u(\pi_{j},\pi_{k})>0$, but $u(\pi_{i},\pi_{k})\leq 0$. This introduces a cyclic relationship among strategies, thereby complicating the game. The complexity often results in a mixed-strategy equilibrium, where players randomize their choices among multiple strategies to maximize their expected payoff. A quintessential example of a non-transitive game is Rock-Paper-Scissors, in which no single strategy uniformly dominates all others. In real-world settings, games exhibit complexities that extend beyond theoretical models. [29] argues that real-world games have two salient features: first, practice usually leads to performance improvements; and second, there are a plethora of qualitatively distinct strategies, each with unique advantages and disadvantages. In such games, the strategies form a geometric topology resembling a spinning top, where the vertical axis represents the performance of the strategy, and the radial axis represents the length of the longest cycle.
II-B4 Stage Game and Repeated Game
A stage game (or one-shot game) is a game that is played only once, namely a one-shot interaction between players. A famous example of a stage game is the Prisoner’s Dilemma. A repeated game is derived from a stage game that is played multiple times. Formally, a repeated game based on a stage game $\mathcal{G}$ is defined by playing $\mathcal{G}$ for $T$ periods, where $T$ can be finite or infinite. The strategies in a repeated game are history-contingent, meaning that they can depend on the entire sequence of past plays. It’s important to note that a stage game or a repeated game can be either represented in the normal-form or extensive-form.
II-B5 Nash Equilibrium
For simplicity, $\pi_{i}$ denotes the strategy of player $i$, and $\pi_{-i}$ denotes the strategies of all players other than player $i$. Given $\pi_{-i}$, player $i$’s best response (BR) is the strategy that maximizes player $i$’s payoff:
$BR_{i}(\pi_{-i})=\arg\max_{\pi_{i}}u_{i}(\pi_{i},\pi_{-i}).$ (1)
A strategy $\pi_{i}^{*}$ is an $\bm{\epsilon}$-BR to strategies $\pi_{-i}$ if:
$u_{i}(\pi_{i}^{*},\pi_{-i})\geq u_{i}(BR_{i}(\pi_{-i}),\pi_{-i})-\epsilon,$ (2)
where $\epsilon$ is a pre-specified threshold.
A strategy profile $(\pi_{1}^{*},\pi_{2}^{*},...,\pi_{n}^{*})$ is a Nash equilibrium (NE) if, for every player $i$:
$u_{i}(\pi_{i}^{*},\pi_{-i}^{*})\geq u_{i}(\pi_{i},\pi_{-i}^{*}),\forall\pi_{i},$ (3)
meaning that no player can benefit by changing their strategy unilaterally, given the strategies of all other players. In other words, an NE is a situation where the strategy chosen by each player is a BR to the strategies chosen by all other players.
A strategy profile $(\pi_{1}^{*},\pi_{2}^{*},...,\pi_{n}^{*})$ is an $\bm{\epsilon}$-NE if, for every player $i$:
$u_{i}(\pi_{i}^{*},\pi_{-i}^{*})\geq u_{i}(\pi_{i},\pi_{-i}^{*})-\epsilon,\forall\pi_{i},$ (4)
meaning that no player can increase their payoff by more than $\epsilon$ by unilaterally changing their strategy.
However, computing NE is generally intractable in complex games, leading some researchers to utilize $\alpha$-Rank [30] and Correlated Equilibrium (CE) [31] as alternatives. Additionally, some studies resort to Replicator Dynamics [32] as a method to analyze and understand the evolution of strategies within these games.
II-B6 Team Games
The framework of a two-player zero-sum game can be naturally extended to encompass team-based zero-sum games. Von Stengel and Koller analyzed zero-sum normal-form games involving a single team competing against an adversary [33]. In this type of team game, consider a team denoted by $\mathcal{T}=\{1,2,\cdots,n-1\}$. The player $n$ is the adversary $(\mathcal{D})$. In this kind of zero-sum normal-form team games, for any player $i,j\in\mathcal{T}$, the utility functions satisfy $u_{i}(\bm{\pi})=u_{j}(\bm{\pi})=u_{\mathcal{T}}(\bm{\pi})$ and $u_{\mathcal{D}}(\bm{\pi})=-(n-1)u_{\mathcal{T}}(\bm{\pi})$. A zero-sum single-team single-adversary normal-form game can also be extended to the domain of extensive games [34]. For any player $i,j\in\mathcal{T}$ and all terminal nodes $z\in Z$, the utility functions satisfy $u_{i}(z)=u_{j}(z)=u_{\mathcal{T}}(z)$ and $u_{\mathcal{D}}(z)=-(n-1)u_{\mathcal{T}}(z)$. Let $I_{\mathcal{T}}$ denote the information set defined as $\bigcup_{i\in\mathcal{T}}I_{i}$, and let $A_{\mathcal{T}}$ represent the set of actions accessible in the information sets within $I_{\mathcal{T}}$.
In scenarios where teammates are unable to coordinate their strategies, the team-maxmin equilibrium (TME) emerges as the most suitable solution concept [33]. We denote the collection of sequences of player $i$ by $Q_{i}$, representing the sequence-form actions undertaken by player $i$. The sequence-form strategy is encapsulated by a function $p_{i}:Q_{i}\rightarrow\mathbb{R}$, which maps each sequence $q\in Q_{i}$ to its associated probability of execution. Formally, the TME is articulated as:
$\arg\max_{p_{1},\ldots,p_{n-1}}\min_{p_{n}}u_{\mathcal{T}}\prod_{i=1}^{n}p_{i}.$ (5)
Similar to the arguments in normal-form games [33], it can also be inferred that in extensive-form games, a TME exists uniquely, barring any degeneracy and this TME aligns with the team’s utility maximization of the NE [34].
II-C Evaluation Metrics in Self-play
In this section, we introduce various self-play evaluation metrics, including NASHCONV (Section II-C1), Elo (Section II-C2), Glicko (Section II-C3), WHR (Section II-C4), and TrueSkill (Section II-C5). While NASHCONV measures the distance from Nash equilibrium, the other four metrics evaluate relative skill levels and are compared in Table I. It’s important to note that although numerous other evaluation metrics exist, the metrics highlighted here are among the most widely used in the field.
Table I: Comparison of Relative Skill Evaluation Metrics.
Elo Glicko WHR TrueSkill
Uncertainty Modeling $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Ratings At Any Time $\times$ $\times$ $\checkmark$ $\times$
Multiplayer In One Team $\times$ $\times$ $\times$ $\checkmark$
Bayesian Foundation $\times$ $\times$ $\checkmark$ $\checkmark$
II-C1 NASHCONV
Nash convergence (NASHCONV) serves as a metric to measure the deviation of a particular strategy from an NE. A lower NASHCONV value suggests that the strategy is closer to an NE, implying that no player would benefit from unilateral deviation from the strategy. Formally, it is defined as:
$\text{NASHCONV}(\bm{\pi})=\sum_{i}\max_{\pi_{i}\in\Pi_{i}}u_{i}(\pi_{i},\pi_{-i})-u_{i}(\bm{\pi}),$ (6)
where $\bm{\pi}$ denotes the combined strategy profile of all participating agents. In particular, in the context of two players, this deviation is commonly referred to as exploitability.
II-C2 Elo
The Elo system [35] operates under the assumption that the performance of each player in each game is a normally distributed random variable, with the mean being the player’s current rating. In a match between player A and player B, $R_{A}$ and $R_{B}$ are the current ratings of player A and player B. The probability density functions for the performance of player A and player B are given by:
$f_{A}(x)=\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(x-R_{A})^{2}}{2\sigma^{2}}},$ (7)
$f_{B}(x)=\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(x-R_{B})^{2}}{2\sigma^{2}}}.$ (8)
$E_{A}$ and $E_{B}$ denote the expected score (or the probability of winning) for player A and player B:
[…]
For convenience, the logistic function is used to approximate the probability (using the logistic curve with base 10 and choosing 400 to scale and normalize the difference in ratings):
$E_{A}$ $=\frac{1}{1+10^{(R_{B}-R_{A})/400}},$ (11)
$E_{B}$ $=\frac{1}{1+10^{(R_{A}-R_{B})/400}}.$ (12)
Note that $E_{A}+E_{B}=1$. $S_{A}$ is the actual outcome of the match for player A (1 for a win, 0.5 for a draw, 0 for a loss). $S_{B}=1-S_{A}$. After a match, the ratings are updated based on the difference between the actual outcome and the expected outcome. The adjustment is given by:
$\Delta R_{A}=K(S_{A}-E_{A}),$ (13)
$\Delta R_{B}=K(S_{B}-E_{B}).$ (14)
$K$ is a scaling factor, often determined by the specific domain of the application, and controls the maximum possible rating change for a single match. If two players have nearly identical ratings, the expected outcome will be close to 0.5. A win for either player will result in a moderate increase in their rating. In contrast, if there is a significant rating difference, the expected outcome will be heavily skewed. If the higher-rated player wins, their rating will increase by a value much smaller than $K$, reflecting the anticipated nature of their victory. However, if the lower-rated player secures an unexpected win, their rating will surge by a value approaching $K$, signifying the upset. The updated ratings become:
$R_{A}^{\prime}=R_{A}+\Delta R_{A},$ (15)
$R_{B}^{\prime}=R_{B}+\Delta R_{B}.$ (16)
However, there are challenges and limitations in Elo. First, the Elo system assumes that all matches are equally important. This might not be the case in all domains. Second, $K$ is often kept constant. However, a dynamic $K$ factor might be more appropriate, especially for players new to the rating pool or in scenarios where the importance of the match varies. Third, the standard Elo system does not incorporate a decay mechanism to account for the potential degradation or improvement of skills during periods of inactivity. Fourth, the basic Elo system is designed for one-on-one competitions. Adapting it to team or multi-player scenarios, such as team sports or online multiplayer games, can be challenging. Lastly, an important limitation of the Elo rating system is its unsuitability for games that exhibit high levels of non-transitivity [36].
II-C3 Glicko
The Glicko system refines the Elo system by introducing a measure of uncertainty or reliability in a player’s rating, termed rating deviation [37]. The primary motivation is to account for the variability in a player’s performance and the potential changes in skill over time. The Glicko-2 system, an extension of the original Glicko system, further refines these concepts and introduces the rating volatility $\sigma$, indicating the degree of expected fluctuation in the player’s rating [38].
In the Glicko-2 system, $r$ denotes the current rating of the player. $RD$ is the player’s rating deviation, and $\sigma$ is the player’s volatility. Convert the rating and rating deviation to the Glicko-2 scale:
$\mu$ $=\frac{r-1500}{173.7178},$ (17)
$\phi$ $=\frac{RD}{173.7178}.$ (18)
Then, calculate $v$ representing the estimated variance of the player’s rating based on game outcomes, $E(\mu,\mu_{j},\phi_{j})$ representing the probability of a player with rating $\mu$ defeating an opponent player $j$ and $\Delta$ representing the estimated improvement based only on game outcomes $s_{j}$:
$v$ $=\left[\sum_{j=1}^{m}g(\phi_{j})^{2}\{1-E(\mu,\mu_{j},\phi_{j})\}\right]^{-1},$ (19)
$E(\mu,\mu_{j},\phi_{j})$ $=\frac{1}{1+\exp(-g(\phi_{j})(\mu-\mu_{j}))},$ (20)
$\Delta$ $=v\sum_{j=1}^{m}g(\phi_{j})\{s_{j}-E(\mu,\mu_{j},\phi_{j})\},$ (21)
where:
$g(\phi)$ $=\frac{1}{\sqrt{1+3\phi^{2}/\pi^{2}}}.$ (22)
The update for $\sigma$ is more involved and requires an iterative procedure to solve for its new value $\sigma^{\prime}$. Then, calculate new $\mu^{\prime}$ and $\phi^{\prime}$:
$\mu^{\prime}$ $=\mu+\phi^{\prime 2}\sum_{j=1}^{m}g(\phi_{j})\{s_{j}-E(\mu,\mu_{j},\phi_{j})\},$ (23)
$\phi^{\prime}$ $=\frac{1}{\sqrt{\frac{1}{{\phi^{*}}^{2}}+\frac{1}{v}}},$ (24)
where:
$\phi^{*}=\sqrt{\phi^{2}+\sigma^{\prime 2}}.$ (25)
After calculations, the rating and rating deviation are converted back from the Glicko-2 scale:
$r^{\prime}$ $=173.7178\mu^{\prime}+1500,$ (26)
$RD^{\prime}$ $=173.7178\phi^{\prime}.$ (27)
II-C4 WHR
The Whole-History Rating (WHR) system [39] is a Bayesian rating system designed to estimate players’ skills from their entire game history. It is particularly adopted in handling the temporal dynamics of player skills. $R_{i}(t)$ denotes Elo rating of the player $i$ at time $t$. Similar to Equ. (11) and Equ. (12), $E_{A}(t)$ and $E_{B}(t)$ denote the expected score (or the probability of winning) for player A and player B at time $t$:
$E_{A}(t)$ $=\frac{1}{1+10^{(R_{B}(t)-R_{A}(t))/400}},$ (28)
$E_{B}(t)$ $=\frac{1}{1+10^{(R_{A}(t)-R_{B}(t))/400}}.$ (29)
Moreover, WHR assumes that ratings of each player $i$ is a Wiener process:
$r_{i}(t_{2})-r_{i}(t_{1})\sim\mathcal{N}(0,|t_{2}-t_{1}|\omega^{2}),$ (30)
where $\omega$ is a parameter representing the variability of ratings in time. Thus, if $r^{\prime}=R_{A}(t_{1}),r^{\prime\prime}=R_{A}(t_{2})$:
$p(r^{\prime\prime}|r^{\prime})=\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{1}{2}(\frac{r^{\prime\prime}-r^{\prime}}{\sigma})^{2}},$ (31)
where $\sigma=\omega\sqrt{|t_{2}-t_{1}|}$. In addition, a Wiener process is memoryless (or Markovian).
$\mathbf{R}(t)$ denotes player A’s ratings, $\mathbf{G}$ denotes the observation of game results. Thanks to the Bayes formula, the WHR system is to solve:
$\arg\max_{\mathbf{R}}p(\mathbf{R}|\mathbf{G})=\arg\max_{\mathbf{R}}\frac{p(\mathbf{G}|\mathbf{R})p(\mathbf{R})}{p(\mathbf{G})}.$ (32)
$P(\mathbf{G}|\mathbf{R})$ can be derived from the product of Equ. (28) and $p(\mathbf{R})$ can be derived from the product of Equ. (31). Utilizing the iterative process of Newton’s method, we can ascertain the solution to the given problem.
II-C5 TrueSkill
TrueSkill [40] is based on a probabilistic graphical model that employs Bayesian inference and adapts to multiple players in multiple teams. TrueSkill 2 [41] extends the original TrueSkill model with several enhancements by taking into account the experience of a player, the affiliation with a team, and some game-specific factors such as the number of kills. For simplicity, TrueSkill is introduced using the following scenario: team 1 has player 1, while team 2 has player 2 and player 3. Ultimately, team 1 defeated team 2. This is a simpler case than the one presented in the original paper [40]. It can be described in an undirected probabilistic graphic model (Fig. 2). The skill of each player, denoted by $s_{i}$, is represented by a Gaussian distribution $N(s_{i};\mu_{i},var_{i})$. The performance of each player, denoted by $p_{i}$, is also represented by a Gaussian distribution $N(p_{i};s_{i},\beta^{2})$. The performance of each team is denoted by $t_{i}$. The difference in team performance is denoted by $d$. If $d>\epsilon$, where $\epsilon$ is a small positive threshold, then team 1 beats team 2. The TrueSkill algorithm’s updating mechanism within the context of undirected probabilistic graphs uses the sum-product message-passing algorithm. This ensures that the skill estimates are refined iteratively, leading to accurate and reliable ratings for each player.
Figure 2: An undirected probabilistic graphic model example of TrueSkill.
III Algorithms
Based on existing self-play work [18, 42, 43, 44], we propose a self-play framework (Algo. 1) that boasts enhanced expressivity and superior generalization capabilities. The framework is adept at handling multi-homogeneous-player general-sum games. It’s important to note that although homogeneous players represent a specific subset of heterogeneous players, the latter can be reformulated as the former by expanding the dimensions of the input vector, which essentially entails embedding agent identity information. Moreover, as players are homogeneous, we assume each player shares the same policy population.
To clarify our framework, we describe it in a more easily understandable way. All players share a common policy population with a fixed maximum size. In each iteration, a new policy is initialized for training, and opponent policies are sampled from the existing policy population. During this iteration, the opponent policies typically remain fixed, while only the policy being trained is updated. After the training process, the new policy replaces one of the policies in the policy population. An evaluation metric is then used to assess the performance of the updated policy population. Based on this performance, the strategy for sampling opponents is adjusted for the next iteration. This process is repeated. For a more detailed and precise description, please refer to Sec. III-A.
In addition, we divide self-play algorithms into four primary groups: traditional self-play algorithms (Sec. III-B), the PSRO series (Sec. III-C), the ongoing-training-based series (Sec. III-D), and the regret-minimization-based series (Sec. III-E). We analyze how these four categories align with our framework in their respective sections and introduce representative algorithms in each category. Moreover, in Sec. III-F, we discuss the differences among these four categories and their connections to our proposed framework. We also explain why our framework is more general than existing frameworks. Furthermore, we summarize the major elements of key algorithms within our framework for each category in Table II.
III-A Framework Definition
Algorithm 1 the Framework of Self-play
0:  $\Pi:=\{\pi_{i}(\cdot|h(i))\}_{i=1}^{N}$$\triangleright$ Policy population.
0:  $\Sigma:=\{\sigma_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{1}}$$\triangleright$ Interaction matrix.
0:  $\mathcal{F}:\mathbb{R}^{N\times C_{2}}\rightarrow\mathbb{R}^{N\times C_{1}}$$\triangleright$ MSS.
1:  for $e\in[[E]]$ do
2:     for $\sigma_{i}\in\Sigma$ do
3:        Initialize $\pi_{i}^{h}$
4:        $\pi_{i}^{h}\leftarrow\text{ORACLE}(\pi_{i}^{h},\sigma_{i},\Pi)$$\triangleright$ Compute the oracle.
5:        $\mathcal{P}\leftarrow\text{EVAL}(\Pi)$triangleright Get policies’ performance.
6:        $\Sigma\leftarrow\mathcal{F}(\mathcal{P})$$\triangleright$ Update the interaction matrix.
7:     end for
8:  end for
9:  return  $\Pi,\Sigma$
In Algo. 1, we define a unified framework of self-play based on [43, 42, 44, 18]. First, the input of the framework is defined as follows:
  • $\Pi$: Each policy $\pi_{i}$ in the policy population $\Pi$ is conditioned on a policy condition function $h(i)$, which is determined by specific algorithms. We will further introduce this policy condition function in the following sections. For simplicity, we denote each policy $\pi_{i}(\cdot|h(i))$ by $\pi_{i}^{h}$. It’s important to note that $i$ refers to the $i$th policy in the policy population, not to a player. Additionally, the hyper-parameter $N$ denotes the policy population size of $\Pi$, not the number of game players. Furthermore, $\Pi$ can be initialized in two different ways. First, $\Pi$ can be considered as initialized with $N$ placeholder policies, meaning that there is no actual initialization of the policies in practice. Instead, each policy will be actually initialized in the training iterations (Line 3 in Algo. 1). We refer to this approach as the placeholder initialization. Second, $\Pi$ can be initialized with $N$ real policies, which may include random initialization or pretrained models. We refer to this approach as the actual initialization. A policy $\pi_{i}^{h}$ is considered an ineffective policy if it serves as a placeholder policy; otherwise, it is deemed an effective policy. Additionally, Table II provides a clearer illustration of the initialization of $\Pi$ and the expression of $h(i)$ across different categories of the main self-play algorithms.
  • $\Sigma:=\{\sigma_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{1}}$: The interaction matrix of the policy population. $\sigma_{i}\in\mathbb{R}^{C_{1}}$ represents the opponent sampling strategy of policy $i$. Namely, $\sigma_{i}$ illustrates how to sample opponent(s)’ policies against policy $i$. For instance, let $\sigma_{i}$ represent the probability of policies for each opponent. In this context, $C_{1}=N^{n-1}$, where n denotes the number of players. Alternatively, $\sigma_{i}$ can also be considered as sampling parameters within a sampling network. Especially, in a two-player game, if $C_{1}=N$ and $\sigma_{ij}$ represents the probability that policy $i$ is optimized against policy $j$, $\Sigma$ can be depicted in directed interaction graphs (Fig. 3). It’s important to note that, unlike the original PSRO framework [42], $\sigma_{i}$ here is not the meta-strategy of policy $i$. Instead, in the two-player setting, if $\sigma_{i}$ in our framework is the opponent’s meta-strategy against policy i, our framework can reduce to the original PSRO framework.
  • $\mathcal{F}:\mathbb{R}^{N\times C_{2}}\rightarrow\mathbb{R}^{N\times C_{1}}$: A meta-strategy solver (MSS) $\mathcal{F}$ takes the performance matrix $\mathcal{P}:=\{p_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{2}}$ as its input and produces a new interaction matrix $\Sigma\in\mathbb{R}^{N\times C_{1}}$ as its output. $p_{i}$ is the performance of policy i. For instance, $p_{i}$ can be depicted as the relative skill like Elo ratings ($C_{2}=1$) or can be depicted as the payoff tensor ($C_{2}=N^{n-1}$, where n is the number of players). Specially, in two-player symmetric zero-sum games, the expected payoffs can serve as the evaluation metric. In such cases, $\Sigma$ is a square matrix ($C_{2}=N$). In addition, Table II concludes the MSS of representative self-play algorithms across these four categories.
Figure 3: When $C_{1}=N$ and $\sigma_{ij}$ represents the probability that policy $i$ is optimized against policy $j$, we can consider various examples of conventional self-play algorithms: In the Top section, we define $\Sigma\in\mathbb{R}^{3\times 3}$ as $\{\sigma_{i}\}_{i=1}^{3}$. In the Bottom section, we present directed interaction graphs where the outgoing edges from each node are equally weighted, and their weights collectively sum to one. The relationship between the Top and Bottom sections is established through directed edges: an edge directed from node $i$ to node $j$ with a weight of $\sigma_{ij}$ signifies that policy $i$ is optimized against policy $j$ with a probability of $\sigma_{ij}$. Note that this figure is reproduced from [43] and this concept is initially proposed by [44].
Next, the core process within the framework is as follows:
  • For epoch $e\in[[E]]$ (Line 1 in Algo. 1): $E$ denotes the total number of epochs for the entire policy population. For example, if the algorithm only introduces new policies into the population without updating existing ones, then $E=1$. This implies that only the ineffective policy is likely to be chosen for training in each iteration and turned into an effective one, while effective policies remain unchanged. Conversely, if effective policies are updated multiple times throughout the algorithm, then $E>1$. Indeed, $E$ accurately rechoflects the number of updates performed. Additionally, Table II summarizes the values of $E$ for different categories of self-play algorithms.
  • Initialize $\pi_{i}^{h}$ (Line 3 in Algo. 1): The initialization of $\pi_{i}^{h}$ can vary depending on the algorithm being used. For instance, it may be initialized randomly, by leveraging pre-trained models [8, 15] or through a recently updated policy. We provide detailed descriptions of the initialization process for each algorithm series. Also, these are summarized in Table II.
  • $\text{ORACLE}(\pi_{i},\sigma_{i},\Pi)$ (Line 4 in Algo. 1): The ORACLE is an abstract computational entity that returns a new policy adhering to specific criteria. Here, we divide ORACLE into three types. (1) One type is the BR oracle, which is designed to identify the optimal counter-strategies against an opponent’s strategy, including finding NE [45]. However, it often requires considerable computational effort. This computational demand can be a limitation, particularly in complex environments or with large action spaces. (2) To alleviate the computational demands, the approximate best response (ABR) oracle is introduced, which can be calculated using techniques such as RL (Algo. 2), evolution-theory-based methods (Algo. 3) or regret minimization methods (Algo. 4). (3) Additionally, other specially crafted ORACLEs are either tailored to new MSSes [46, 47] or introduced to enhance diversity [48, 49, 50].
    Some details of Algo. 23 and 4 need to be mentioned. $\bm{\pi_{opp}}$ denotes opponents’ policies. Moreover, in Algo. 2, off-policy RL algorithms [2, 51, 52, 53], typically require a replay buffer to gather samples; however, for simplicity, they are not explicitly included. Furthermore, Line 1 in Algo. 2 suggests that RL may not always yield an ABR that meets specific criteria. For instance, AlphaGo Zero [9] stipulates that the new policy must achieve a win rate exceeding 55 percent against its predecessor. Therefore, validation of the new policy is sometimes required. In addition, it is important to note that Algo. 3 is specifically devised for R-NaD [54], while Algo. 4 is expressly designed for a series based on regret minimization. They will be introduced in detail in Sec. III-C5 and Sec. III-E respectively.
  • $\text{EVAL}(\Pi)$ (Line 5 in Algo. 1): Evaluating the policy population $\Pi$. There are multiple evaluation metrics available to gauge the performance of each policy. The performance matrix is represented as $\mathcal{P}:=\{p_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{2}}$. Moreover, only effective policies will be evaluated to save on computation. Notably, if the MSS produces a constant matrix irrespective of the input, the evaluation step can be skipped to reduce computations.
Algorithm 2 Compute the Oracle in RL
0:  $\pi_{i}^{h}$$\triangleright$ Policy $i$ is being trained.
0:  $\sigma_{i}$$\triangleright$ Opponent sampling strategy of policy $i$.
0:  $\Pi$$\triangleright$ Policy population.
1:  while $\pi_{i}^{h}$ is not valid do
2:     for trajectory $\tau\in[[\tau_{max}]]$ do
3:        sample $\bm{\pi_{opp}^{h}}\sim P(\sigma_{i})$$\triangleright$ Policies of opponents.
4:        $\bm{\pi}=(\pi_{i}^{h},\bm{\pi_{opp}^{h}})$
5:        $s_{0},\bm{o_{0}}\sim\rho$triangleright Initial state.
6:        for $t\in[[t_{max}]]$ do
7:           $\bm{a_{t}}\sim\bm{\pi}(\bm{o_{t}})$
8:           $s_{t+1},\bm{o_{t+1}}\sim P(s_{t},\bm{a_{t}})$
9:           $\bm{r_{t}}\leftarrow\bm{R}(s_{t},\bm{a_{t}})$
10:        end for
11:        $\pi_{i}^{h}\leftarrow\text{update}(\pi_{i}^{h})$$\triangleright$ Using RL algorithms
12:     end for
13:  end while
14:  return  $\pi_{i}^{h}$
Algorithm 3 Compute the Oracle in Evolution Theory
0:  $\pi_{i}^{h}$$\triangleright$ Policy $i$ is being trained.
0:  $\sigma_{i}$$\triangleright$ Opponent sampling strategy of policy $i$.
0:  $\Pi$$\triangleright$ Policy population.
1:  sample $\bm{\pi_{opp}^{h}}\sim P(\sigma_{i})$$\triangleright$ Policies of opponents.
2:  $\bm{\pi_{\text{reg}}}=(\pi_{i}^{h},\bm{\pi_{opp}^{h}})$$\triangleright$ Regularization policies.
3:  Transformed the reward according to $\bm{\pi_{\text{reg}}}$.
4:  $\pi_{i}^{h}\leftarrow$ Use replicator dynamics to play the reward-transformed game until convergence.
5:  return  $\pi_{i}^{h}$
Algorithm 4 Compute the Oracle in Regret Matching
0:  $\pi_{i}^{h}$$\triangleright$ Policy $i$ is being trained.
0:  $\sigma_{i}$$\triangleright$ Opponent sampling strategy of policy $i$.
0:  $\Pi$$\triangleright$ Policy population.
1:  for each player j do
2:     sample $\bm{\pi_{opp}^{h}}\sim P(\sigma_{i})$$\triangleright$ Policies of opponents.
3:     $\bm{\pi}=(\pi_{i}^{h}(j),\bm{\pi_{opp}^{h}}(-j))$
4:     Use regret matching to play the game and obtain new regret minimization information added to $h(i)$.
5:  end for
6:  return  $\pi_{i}^{h}$
Moving forward, we categorize self-play algorithms into four primary groups. We provide a detailed analysis of how each algorithm in these categories integrates and aligns with our proposed framework.
III-B Traditional Self-play Algorithms
Traditional self-play algorithms involve agents improving their strategies by repeatedly playing against themselves, allowing them to explore various strategies and enhance their decision-making abilities without external input. These algorithms can start with agents training against their most recent version, helping to identify and exploit weaknesses. Additionally, other approaches involve training against a set of strategies from different iterations, enabling agents to develop robust and adaptive strategies. In this section, we will explain how traditional self-play algorithms fit into our framework and introduce representative traditional self-play methods, ranging from simpler forms to more complex ones.
III-B1 Integration into Our Framework
We can incorporate the traditional self-play algorithms into our proposed framework (Algo. 1), using the following settings. First, the policy population $\Pi$ utilizes placeholder initialization, meaning that initially, the policies are placeholders rather than actually initialized strategies. This initialization approach is used because the policy population in traditional self-play algorithms is intended to grow with each iteration. Second, we set $E=1$ because in traditional self-play algorithms, only an ineffective policy is likely to be chosen for training in each iteration, thereby turning it into an effective policy. Here, the policy population size $N$ serves as the upper limit for the number of effective policies in the population. In other words, we use $N$ iterations to optimize the policy. Third, the strategy being trained $\pi_{i}^{h}$ can be initialized in a general manner. For instance, the strategy can be initiated randomly, which signifies starting the learning process from scratch. More often, $\pi_{i}^{h}$ is initialized by $\pi_{i-1}(\cdot|h(i-1))$, allowing incremental learning and adaptation based on the most current trained policy to accelerate convergence. Fourth, since the policies in traditional self-play algorithms are not conditioned, we simply set $h(i)=\emptyset$.
Next, we outline the traditional self-play schemes. For the sake of simplicity, we operate under the following assumption:
Assumption 1.
In a two-player symmetric game, $C_{1}=N$, with $\sigma_{ij}$ denoting the probability that policy $i$ is optimized in response to policy $j$ which leads to $\sum_{j=1}^{N}\sigma_{ij}=1,\leavevmode\nobreak\ \forall i$.
Based on Assumption 1, we can further deduce the following important corollary:
Corollary 1.
In traditional self-play algorithms, the interaction matrix $\Sigma$ is a lower triangular matrix.
Proof.
The policy population gradually increases over time. Consequently, when policy $i$ is selected for training, only policy $j$ $(j\leq i)$ has already been trained and holds meaningful outcomes. Other policies are ineffective policies. As a result, we exclusively select policy $j$, where $j\leq i$, to serve as the opponent for policy $i$. ∎
III-B2 Vanilla Self-play
In vanilla self-play [6], agents are trained by competing against their latest versions, ensuring consistent and aligned learning. The MSS of vanilla self-play:
$\mathcal{F}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}1,\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ j=i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right..$ (33)
No matter what $\mathcal{P}$ is, the MSS produces the same interaction matrix, similar to the iterative refinement process of the policy. Although this MSS is simple, it is utilized in many further works, so we refer to it as vanilla MSS. Although vanilla self-play is effective in transitive games, it can lead the agent to cyclic learning patterns in non-transitive games like Rock-Paper-Scissors.
III-B3 Fictitious Self-play
Fictitious Play (FP) [55] is a learning algorithm in game theory where each player best responds to the empirical frequency of the strategies used by the opponent. If the opponent’s strategy is static, FP can find the NE. Based on FP intuition, Fictitious Self-play (FSP) [13] is introduced to make agents play against past versions of themselves to learn optimal strategies to improve the robustness of vanilla self-play. Neural Fictitious Self-play (NFSP) [56] is a modern variant that combines FSP with deep learning techniques. It uses neural networks to approximate the BRs. In the original versions of NFSP and FSP, two distinct types of agent memory are used: one to record the agent’s own behavior and the other to capture the opponent’s behavior. However, in more recent approaches [42], random sampling is frequently employed to approximate the opponents’ average strategy, eliminating the need to maintain two separate types of agent memory. Thus, the MSS of FSP:
$\mathcal{F}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{i-1},\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ j\leq i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right..$ (34)
In FSP, the MSS continues to generate a constant interaction matrix. Compared to vanilla self-play, this approach enhances the robustness of the policy by sampling older versions of its own policies to be used as opponent strategies.
III-B4 $\delta$-uniform Self-play
$\delta$-uniform self-play is introduced by [7]. The hyper-parameter $\delta$ ranges between 00 and $1$ used to select the most recent $\delta$ (percentage) of policies for uniform sampling to generate opponent policies. When policy $i$ is in the training phase, following Corollary 1, only the preceding $i-1$ policies hold significance. As opponents for policy $i$, we select policies from the discrete uniform distribution $[\delta(i-1),i-1]$. When $\delta=0$, the system retains the complete historical memory, whereas $\delta=1$ implies that only the most recent policy is utilized. Thus, we can get the MSS of $\delta$-uniform self-play:
$\mathcal{F}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{f(i)},\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \delta(i-1)\leq j\leq i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right.,$ (35)
$f(i)=\max\{\lceil(1-\delta)(i-1)\rceil,1\},$ (36)
[…]
III-B5 Prioritized Fictitious Self-play
Prioritized Fictitious Self-play (PFSP)  [15] leverages a priority function to allocate a higher probability of selection to agents with higher priorities. Here, $\mathcal{P}$ represents the winning rates, specifically defined as $\mathcal{P}_{ij}=\text{Prob}(\pi_{i}\leavevmode\nobreak\ \text{beats}\leavevmode\nobreak\ \pi_{j})$. The MSS of PFSP is given by Algo. 5.
Algorithm 5 the PFSP Meta-strategy Solver
1:  function $\mathcal{F}(\mathcal{P})$
2:   $\sigma_{i+1,j}\leftarrow\frac{f(\mathcal{P}_{i,j})}{\sum_{j\leq i}f(\mathcal{P}_{i,j})},\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ j\leq i$
3:   Append zeros to $\sigma_{i+1}$ until its length is N.
4:  return  $\Sigma$
The function $f:[0,1]\rightarrow[0,\infty)$ is a priority function. For example, $f(x)=(1-x)^{p}$ with $p>0$ indicates that stronger policies against the currently being trained policy have a higher chance of being chosen as opponents. Alternatively, $f=x(1-x)$ implies that players of similar levels are more likely to be chosen as opponents. Additionally, in broader terms, $\mathcal{P}$ can be assessed by other metrics as well. A similar MSS in  [14] can be utilized to allocate higher probabilities to strategies that perform better or are harder to defeat.
III-B6 Independent RL
Independent RL is a fully decentralized method in MARL. This simplifies the learning process and is useful in competitive or adversarial settings, while it may lead to suboptimal outcomes in situations requiring cooperation. Independent RL can be seen as a special case of self-play algorithms. In each iteration, the policy to be trained, $\pi_{i}^{h}$, is initialized using the previous policy $\pi_{i-1}(\cdot|h(i-1))$. The MSS of independent RL simplifies to an identity matrix:
$\mathcal{F}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}1,\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ i=j\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right..$ (37)
The difference between independent RL and vanilla self-play lies in how the opponents’ policies are handled during training. In vanilla self-play, the opponents’ policies are often fixed, making the training process stationary. In contrast, in independent RL, the opponents’ policies evolve along with the policy being trained, resulting in a non-stationary training process. Moreover, if off-policy RL methods are used in independent RL, the samples collected can still be useful for training, even if they are generated by a different policy. This allows the agent to leverage past experiences more effectively and learn from a wider range of scenarios.
III-C PSRO Series of Algorithms
Similar to traditional self-play algorithms, the PSRO series of algorithms starts with a single policy and gradually expands the policy space by incorporating new oracles. These oracles are policies that approximate optimal responses to the current meta-strategies of other agents. Additionally, PSRO employs EGTA to update meta-strategy distributions, thereby incorporating a level of exploration in policy selection to mitigate the risk of overfitting.
III-C1 Integration into Our Framework
The PSRO series of algorithms can also be integrated into our proposed framework (Algo. 1). First, similar to traditional self-play algorithms, we also utilize placeholder initialization to initialize $\Pi$. Second, we also set $E=1$ and $N$ can be considered as the upper limit for the policy population size in the original PSRO algorithms. Third, in the context of the PSRO series of algorithms, the strategy of our player $\pi_{i}^{h}$ can also be initialized in a general manner. Fourth, we simply set $h(i)=\emptyset$ since the PSRO series of algorithms do not use any conditioning function for their policies. Fifth, it’s crucial to highlight that our framework diverges from the traditional PSRO model [42] in how $\sigma_{i}$ is defined. In contrast to being the meta-strategy for policy, in our framework, $\sigma_{i}$ is the opponent sampling strategy. It means that $\sigma_{i}$ here represents the opponent’s meta-strategy against policy $i$ for the PSRO series of algorithms. Sixth, compared with traditional self-play methods, the MSSes of the PSRO series are often more complex. For example, some MSSes incorporate concepts from different types of game equilibria [45, 46, 47].
For simplicity, we also follow Assumption 1. Similar to traditional self-play algorithms, we can derive the Corollary 2 using a similar proof as Corollary 1.
Corollary 2.
In the PSRO series of algorithms, the interaction matrix $\Sigma$ is a lower triangular matrix.
III-C2 Double Oracle
Double oracle (DO) [45] is traditionally only applied to two-player normal-form games. In this context, we can utilize the payoff matrix as the evaluation matrix. The interaction matrix can be initialized with all zeros, reflecting the initial absence of interactions between strategies. The MSS of DO can then be outlined as described in Algo. 6. The opponent sampling strategy $\sigma_{i}$ corresponds to the opponent’s NE strategy of the restricted game. Therefore, the oracle in DO is a BR rather than an ABR, computing the BR against the current NE opponent strategy of the restricted game. In the context of two-player normal-form games, DO theoretically can achieve the NE of the full game.
Algorithm 6 the NE-based Meta-strategy Solver
1:  function $\mathcal{F}(\mathcal{P})$
2:   $\sigma_{i+1}\leftarrow$ SOLVE-NASH$(\mathcal{P}_{1:i,1:i})$triangleright Opponent’s NE meta-strategy.
3:   Append zeros to $\sigma_{i+1}$ until its length is N.
4:  return  $\Sigma$
III-C3 PSRO
PSRO [42] extends DO to more complex games beyond just two-player normal-form games. This approach first introduces the concept of the MSS, which is a broader concept than simply computing NE. The MSS framework allows for a more flexible representation of strategic solutions in various game settings. Many variants of PSRO focus on designing new MSSes to better capture different aspects of strategic play in these more complex games. In the original PSRO framework, the oracle is computed using RL techniques, similar to those described in Algorithm 2. This allows the algorithm to effectively handle large and intricate strategy spaces, making it applicable to a wide range of game scenarios.
III-C4 PSRO Variants
The traditional PSRO algorithm has been the subject of numerous extensions in recent research. Some studies focus on making PSRO more computationally tractable in different scenarios or achieving fast convergence. Through the establishment of a hierarchical structure of RL workers, Pipeline PSRO [57] achieves the parallelization of PSRO and simultaneously provides guarantees of convergence. Efficient PSRO [58] introduces the unrestricted-restricted (URR) game to narrow down the selection of opponent policies, thereby obviating the need for meta-game simulations in traditional PSRO. And, similar to Pipeline PSRO, Efficient PSRO is equipped to solve URR in parallel. In addition, unlike traditional PSRO, which confines the integration of population strategies to the game’s initial state, XDO [59] allows this integration across all information states. It ensures a linear convergence to an approximate NE based on the number of information states, thereby enhancing its tractability in extensive-form games. ODO [60] integrates the no-regret analysis of online learning with the Double Oracle approach to improve both the convergence rate to an NE and the average payoff. Anytime PSRO [61] and Self-play PSRO [62] are designed to incorporate less exploitable policies into the policy population, thereby facilitating faster convergence. Moreover, as the number of agents grows, determining BRs becomes exponentially challenging. Mean-field PSRO [63] has been introduced to address this complexity in mean-field games. In addition, due to the computational intractability of solving NE in multi-player general-sum games [64, 65] and the selection problem in NE [66], Müller et al. put forward $\alpha$-PSRO [46]. Instead of NE, $\alpha$-rank [30], which is unique and polynomial-time solvable in multi-player general-sum games, is introduced in the MSS and a preference-based best response (PBR) oracle is incorporated in the approach. Similar to $\alpha$-PSRO, [47] proposes JPSRO to tackle multi-player general-sum games. It utilizes concepts of CE and coarse correlated equilibrium (CCE) as alternatives to NE to make it computationally tractable in multi-player general-sum games. It also puts forward the CE and CCE-based MSS. Expanding on JPSRO, NeuPL-JPSRO [67] takes advantage of transfer learning. In addition to a shared parameter $\theta$, each strategy in NeuPL-JPSRO is characterized by a strategy embedding vector $v_{i}$. This approach avoids the need for training from scratch in each iteration, thereby enhancing efficiency.
Other research focuses on policy diversification, given that deriving a singular policy in highly transitive games often lacks significance. [48] introduces an open-ended framework in two-player zero-sum games. This framework enhances the diversity of the strategy population and introduces a gamescape, which geometrically represents the latent objectives in games for open-ended learning. The study proposed two algorithms: Nash response $\text{PSRO}_{\text{N}}$ and rectified Nash response $\text{PSRO}_{\text{rN}}$. Both algorithms utilize an asymmetric payoff matrix as their performance evaluation metric. Similarly to DO, they employ the Nash-based MSS (Algo. 6). Compared to $\text{PSRO}_{\text{N}}$, $\text{PSRO}_{\text{rN}}$ incorporates an additional step within the ABR oracle to focus on the opponents they beat or tie and ignore the opponents they lose to. [49] uses determinantal point process [68] to assess diversity and introduces diverse PSRO by incorporating a diversity term into the PSRO oracle. This modification can also be easily implemented in FP and $\alpha$-PSRO as well. Similarly,  [50] formulates behavioral diversity and response diversity and incorporates them into the PSRO oracle. Policy Space Diversity PSRO [69] defines a diversity metric named population exploitability that helps to achieve a full-game NE.
III-C5 R-NaD
Although R-NaD [54] is initially described as leveraging evolutionary game theory with a regularization component. Here, we categorized it into the PSRO series with a special oracle computation technique (Algo. 3), which is executed in two stages: the first stage involves transforming the reward based on the regulation policy to make it policy-dependent, and the second stage applies replicator dynamics [32] until convergence to a fixed point. Crucially, in each iteration, the oracle added to the policy population $\Pi$ is derived from the reward-transformed game, not the original problem’s oracle. Nonetheless, this approach ensures that the policy will converge to the NE of the original game, provided that the game is monotone. The MSS of R-NaD is the vanilla MSS, as described in Equ. (33), the same as the MSS of vanilla self-play. This equation illustrates that the fixed point reached in each iteration, the oracle, is utilized as the regularization policy for the next iteration.
III-D Ongoing-training-based Series of Algorithms
In the PSRO series of algorithms, two key challenges arise. First, when operating with a limited budget, it is often necessary to truncate the ABR operators during each iteration. This can introduce sub-optimally trained responses into the population. Second, the redundant process of relearning basic skills in every iteration is not only inefficient, but it also becomes untenable when confronted with increasingly formidable opponents [43]. To address these challenges, the ongoing-training-based series of algorithms promote the ongoing training of all policies repeatedly. Namely, all effective policies are likely to be selected for training.
III-D1 Integration into Our Framework
We can incorporate these ongoing-training-based series of algorithms into our proposed framework (Algo. 1) using the following settings: First, we use actual initialization to initialize $\Pi$ because in the ongoing-training-based series, all policies in the policy population are trained together, rather than the policy population growing with each iteration. Second, we set $E=E_{max}>1$, which represents the maximum number of epochs to optimize each policy within the policy population. In other words, each unique policy undergoes iterative training for a total of $E_{max}$ times. Third, since each policy undergoes training for $E_{max}$ times, we utilize $\pi_{i}(\cdot|h(i))$ to initialize $\pi_{i}^{h}$. This means that policy updates are self-referential.
For the sake of simplicity, we also adopt Assumption 1. Different from Collory 1 and Collory 2, due to the continuous training process of all policies, we derive Collory 3.
Corollary 3.
In the ongoing-training-based series of algorithms, the interaction matrix $\Sigma$ is generally not a lower triangular matrix.
Proof.
When policy $i$ is selected for training, policy $k$ $(k\geq i)$ has already been actually initialized and is therefore considered an effective policy. Furthermore, in epoch $e$ $(e>1)$, policy $k$ has already been updated and holds significant meaning. As a result, policy $k$, where $k\geq i$, is likely to be chosen as the opponent for policy $i$. Consequently, the interaction matrix $\Sigma$ is generally not a lower triangular matrix. ∎
III-D2 FTW
Quake III Arena Capture the Flag is a renowned 3D multiplayer first-person video game where two teams vie to capture as many flags as possible. The For The Win (FTW) agent [70] is designed to perform at human-level proficiency in this game. A pivotal aspect of FTW is its employment of the ongoing-training-based self-play in RL. Specifically, it trains a population of $N$ different policies in parallel which compete and collaborate with each other. When policy $i$ is undergoing training, FTW samples both its teammates and adversaries from the population. Specially, in scenarios where each team comprises a single member, it can be seamlessly integrated into our framework using the subsequent MSS:
$\mathcal{F}(\mathcal{P})_{ij}=\frac{1}{N}.$ (38)
This essentially means that the interaction graph is densely connected. Moreover, all policies draw upon a unified policy network parameterized by $\phi$. Hence, $\pi_{i}(\cdot|h(i))$ can be aptly depicted as $\pi_{\phi}(\cdot|h(i))$. Furthermore, since these policies are not conditioned on any external parameters, it is straightforward to represent the conditioning function $h(i)=\emptyset$.
III-D3 NeuPL
NeuPL [43] introduces another key innovation: it employs a unified conditional network, where each policy is adjusted against a specific meta-game mixture strategy. This is instrumental in enabling transfer learning across policies. Owing to NeuPL’s reliance on a unified conditional network parameterized by $\theta$, $\pi_{i}(\cdot|h(i))$ can be succinctly represented as $\pi_{\theta}(\cdot|h(i))$. Moreover, given that the policies in NeuPL are contingent on the opponent sampling strategy $\sigma_{i}$, it is apt to define $h(i)=\sigma_{i}$.
III-D4 Simplex-NeuPL
Simplex-NeuPL [71], which builds on NeuPL, is designed to achieve any-mixture optimality, which signifies that the formulated policy should exhibit flexibility across a diverse spectrum of adversaries, including those that might not present equivalent competitive prowess. To model the population learning process from a geometric perspective, Simplex-NeuPL introduces the concept of the population simplex. Analogously to its predecessor, Simplex-NeuPL integrates a conditional network to characterize the policy, represented as $\pi_{\theta}(\cdot|h(i))$ conditioned on the opponent sampling strategy $h(i)=\sigma_{i}$. Intriguingly, there is a slight possibility that $\sigma_{i}$ does not originate from the MSS. Instead, it is drawn as a sample from the population simplex. This sampling mechanism results in greater robustness.
III-E Regret-minimization-based Series of Algorithms
Another line of self-play algorithms is based on regret minimization. The key distinction between regret-minimization-based algorithms and the other categories is that they prioritize accumulated payoff over time, rather than focusing solely on a single episode. This approach leads to more aggressive and adaptable strategies, which are essential to avoid being exploited by opponents as time progresses. Additionally, these algorithms require players to deduce and adapt to opponents’ strategies over several rounds. This scenario is commonly observed in repeated games rather than stage games. For example, in games like Texas Hold’em or Werewolf, players must use deception, concealment, and bluffing to aim for overall victories rather than just winning a single game. It’s important to note that while traditional regret-minimization-based self-play doesn’t typically use RL, many subsequent research efforts have combined regret minimization with RL to achieve strong performance. In this section, we will also thoroughly discuss traditional regret-minimization-based methods to provide a foundation for understanding how integrating regret minimization with RL can lead to enhanced performance.
III-E1 Integration into Our Framework
We can also integrate the regret-minimization-based series of algorithms into our proposed framework (Algo. 1) with the following settings: First, similar to traditional self-play algorithms and the PSRO series, we use placeholder initialization to initialize the policy population $\Pi$. Second, we set $E=1$, and $N$ is regarded as the maximum iteration to optimize the policy. Third, we initialize $\pi_{i}^{h}$ using $\pi_{i-1}(\cdot|h(i-1))$ to utilize the most current trained policy. More specifically, $h(i)=h(i-1)$ and $\pi_{i}^{h}=\pi_{i-1}^{h}$. Fourth, in each iteration $i$, $h(i)$ represents the specific elements that regret-minimization-based self-play algorithms need to store. It’s important to note that the regret-minimization-based series relies heavily on the information contained in $h(i)$. For instance, in vanilla Counterfactual Regret Minimization (CFR) [72], it is necessary to store counterfactual regrets for every action in every information set for each player within $h(i)$ and once $h(i)$ is determined, the corresponding policy is also defined through regret matching. We will discuss vanilla CFR in detail in Sec. III-E2. Fifth, the ABR operator is described in Algo. 4, with the key point being to incorporate new regret minimization-based information into $h(i)$. Notably, while the original CFR updates regrets for all players simultaneously, this oracle (Algo. 4) updates the regrets of each player sequentially. This means that the regrets of player 2 are updated after considering the already-updated regrets of player 1. Namely, $h(i)$ is changing during the iteration $i$. This adjustment has not only been shown to empirically accelerate convergence but also possesses a theoretical error bound [73]. Additionally, each $\pi_{i}^{h}$ represents the strategies for all players, and in iteration $i$, player $j$ uses the strategy $\pi_{i}^{h}(j)$.
Moreover, the MSS of the regret-minimization-based series is the vanilla MSS, as described in Equ. (33). We can derive Collory 4.
Corollary 4.
In the regret-minimization-based series of algorithms, the interaction matrix $\Sigma$ is a lower triangular matrix. More specifically, it is a unit lower shift matrix, with ones only on the subdiagonal and zeroes elsewhere.
Proof.
Regret minimization-based algorithms always use the latest strategy for training. In other words, in iteration $i$, $\pi_{i-1}^{h}$ is always chosen as the opponent policy. Consequently, the interaction matrix $\Sigma$ is a unit lower shift matrix. ∎
III-E2 Vanilla CFR
Regret measures the difference between the actual payoff and the best possible payoff that could have been achieved with a different strategy. The regret matching algorithm [74] in game theory optimizes decisions in iterative games by learning from past outcomes. More specifically, it involves selecting strategies based on accumulated positive overall regrets. Strategies with higher overall regret are generally more likely to be chosen, as players seek to correct past underperformance. After each round, players update the overall regret values for each strategy. When each player aims to reduce their average overall regret, their average strategies will converge towards an approximation of the NE over time. However, this traditional regret minimization algorithm applies primarily to normal-form games because computing overall regret in extensive-form games is challenging. Although extensive-form games can theoretically be converted into normal-form games, this conversion results in an exponential increase in states, rendering it impractical for complex extensive-form games.
[72] proposes counterfactual regret (CFR) minimization for extensive-form games with incomplete information. Here, it is referred to as vanilla CFR to distinguish it from subsequent advancements in the field. Vanilla CFR involves maintaining the strategy and the counterfactual regrets for each information set. Theoretically, the immediate counterfactual regret is specific to each information set, and the aggregation of these immediate counterfactual regrets can act as an upper bound for the overall regret. Thus, the problem can be simplified from minimizing the overall regret to minimizing the immediate counterfactual regrets in each information set. This simplification significantly reduces the computational complexity.
Next, there is a more detailed description of vanilla CFR. In this survey, $\pi_{i}$ is used to denote the strategy at iteration $i$, whereas the original paper uses $\sigma$. This change is implemented to maintain consistency throughout the discussion in this survey. Furthermore, we use $j$ to represent player $j$, and $\pi_{i}(j)$ denotes the policy used by player $j$ at iteration $i$. The counterfactual value $v_{j}(\pi,I)$, which represents the expected value upon reaching the information set $I$ when all players, except for player $j$, adhere to the strategy $\pi$:
$v_{j}(\pi,I)=\sum_{h\in I,h^{\prime}\in Z}\text{Prob}_{-j}^{\pi}(h)\text{Prob}^{\pi}(h,h^{\prime})u_{j}(h^{\prime}).$ (39)
It is an unnormalized form of the counterfactual utility in the original paper. Based on this counterfactual value, the immediate counterfactual regret is defined as:
$R_{j,\text{imm}}^{T}(I)=\frac{1}{T}\max_{a\in A(I)}\sum_{i=1}^{T}\left(v_{j}(\pi_{i}|_{I\to a},I)-v_{j}(\pi_{i},I)\right),$ (40)
where $\pi_{i}|_{I\to a}$ denotes player $j$ choose action $a$ with probability $1$ at iteration $i$. Moreover, the positive immediate counterfactual regret is:
$R_{j,\mathrm{imm}}^{T,+}(I)=\mathrm{max}(R_{j,\mathrm{imm}}^{T}(I),0).$ (41)
The player $j$’s average overall regret $R_{j}^{T}$ is defined as:
$R_{j}^{T}=\dfrac{1}{T}\max_{\pi(j)^{*}\in\Pi(j)}\sum_{i=1}^{T}\left(u_{j}(\pi(j)^{*},\pi_{i}(-j))-u_{j}(\pi_{i})\right).$ (42)
And theoretically, average overall regret $R_{j}^{T}$ is bounded by the sum of positive immediate counterfactual regrets:
$R_{j}^{T}\leq\sum_{I\in\mathcal{I}_{j}}R_{j,\text{imm}}^{T,+}(I).$ (43)
Therefore, based on the theory discussed, the overall problem of regret minimization can be decomposed into numerous smaller regret minimization problems. This approach makes the problem manageable for extensive-form games of not excessively large sizes. In practical applications, the focus is primarily on immediate counterfactual regrets. For simplicity, we often drop "immediate" in discussions, thereby referring directly to counterfactual regrets. Consequently, the counterfactual regret associated with every action $a$ within every information set $I$ is recorded as follows:
$R_{j}^{T}(I,a)=\frac{1}{T}\sum_{i=1}^{T}\left(v_{j}(\pi_{i}|_{I\to a},I)-v_{j}(\pi_{i},I)\right),$ (44)
$R_{j}^{T,+}(I,a)=\max\{R_{j}^{T}(I,a),0\}.$ (45)
The regret matching algorithm [74] is used to decide the strategy in each information set:
$\left.\pi_{j}^{T+1}(I)(a)=\left\{\begin{array}[]{ll}\frac{R_{j}^{T,+}(I,a)}{\sum_{a\in A(I)}R_{j}^{T,+}(I,a)}&\text{ if denominator}>0\\ \frac{1}{|A(I)|}&\text{ otherwise}\end{array}\right.\right..$ (46)
It’s worth noting that in some studies, the normalization factor $\frac{1}{T}$ is omitted in Equ. (40), (42) and (44).
Vanilla CFR has several shortcomings. Firstly, it requires traversing the entire game tree in each iteration, which becomes computationally intractable for larger game trees. Although some efforts have focused on game abstraction to reduce the size of the game tree, greater abstraction can lead to decreased performance. Second, it requires storing counterfactual regrets $R^{i}(I,a)$ for every action $a$ in every information $I$ at each iteration $i$. These values are stored in $h(i)$ within our proposed framework (Algo. 1). This requirement leads to significant storage challenges.
III-E3 Time-saving Variants of CFR
Therefore, numerous studies focus on enhancing the time efficiency of CFR primarily through two main approaches. The first approach involves modifying the regret calculation to increase its speed. For example, CFR+ [75] implements regret-matching+ by storing non-negative regret-like values $R_{j}^{+,i}(I,a)$, rather than $R_{j}^{i}(I,a)$. Note that $R_{j}^{+,i}(I,a)$ is computed slightly differently from $R_{j}^{i,+}(I,a)$ in Equ. (45). Additionally, CFR+ updates the regrets of each player sequentially. Furthermore, CFR+ adopts a weighted average strategy instead of a uniform average strategy, assigning a lower weight to the initial strategy for accelerated convergence. Beyond employing a weighted average strategy, [76] introduces the concept of weighted regrets and developed Discounted CFR (DCFR). Linear CFR (LCFR) is a specific case of DCFR, where LCFR assigns a linear weight of $i$ to regrets at iteration $i$.
The other approach involves adopting sampling methods. Although this approach requires more iterations, the duration of each iteration is reduced, ultimately decreasing the overall time to convergence. Monte Carlo CFR (MCCFR) [77] outlines a framework to incorporate sampling into CFR. This method divides terminal histories into blocks, with each iteration sampling from these blocks instead of traversing the entire game tree. This allows for the calculation of sampled counterfactual values for each player, leading to immediate counterfactual regrets that, in expectation, match those of the Vanilla CFR. Vanilla CFR represents a specific case where all histories are divided into just one block. MCCFR typically manifests in three forms: outcome-sampling MCCFR, where each block corresponds to a single history; external-sampling MCCFR, which samples opponent and chance nodes; and chance-sampling MCCFR, focusing on chance nodes alone. Moreover, [78] expands on chance-sampling into naive chance sampling, opponent/public chance sampling, self/public chance sampling, and public chance sampling. These methods differ in how they handle the sampling of public and private chance nodes. Some studies also focus on learning how to reduce the variance of MCCFR to speed up convergence [79, 80].
In addition to these two primary approaches, other studies have identified that warm starting [81] and pruning techniques [82, 83, 84] can accelerate CFR convergence as well.
III-E4 Space-saving Variants of CFR
Storing the strategies and regrets of each information set at each iteration requires substantial storage capacity. In games with perfect information, decomposition is employed to reduce the scale of problem-solving; that is, we solve subgames instead of the entire game. However, this approach faces challenges in imperfect information games. The difficulty arises in defining subgames, as they may intersect with the boundaries of information sets, complicating their delineation. CFR-D [85] is a pioneering method for decomposing imperfect information games with a theoretical guarantee. The game is divided into a main component, called the trunk, and several subgames. It posits that in an imperfect information game, a subgame can be conceptualized as a forest of trees that does not divide any information sets. In each iteration, CFR is applied within the trunk for both players, accompanied by the use of a solver to determine the counterfactual BR in the subgame. The process includes updating the trunk with the counterfactual values from the subgame’s root and updating the average counterfactual values at this root. Following these updates, the solution to the subgame is discarded. CFR-D minimizes storage needs by only saving values $R^{i}(I_{*},a)$ for information sets located in the trunk and at each subgame’s root, which is denoted by $I_{*}$ at each iteration $i$, trading off storage efficiency against the time required to resolve subgames. Similar thoughts are echoed in the Continue-Resolving technique used by DeepStack [12] and the Safe and Nested Subgame Solving technique used by Libratus [86]. We will discuss these approaches in Sec. IV-B1, where we explore their application to the specific context of Texas Hold’em.
III-E5 Estimation Variants of CFR
Although the above variants of CFR advance the field, they still fail to solve large imperfect-information extensive-form games directly. This limitation largely stems from the reliance on tabular representations for strategies. Typically, the process involves abstracting the original large game into a simpler form, applying the above CFR variants to this abstracted game, and then translating the developed strategies back to the original game. However, this abstraction process is highly game-specific and is heavily based on domain knowledge. Moreover, abstracting the game to a smaller scale often leads to suboptimal results compared to a more comprehensive abstraction, making the choice of abstraction scale critical for performance. Given these challenges, there is a shift towards estimation techniques. [87] introduces Regression CFR (RCFR), which employs a shared regressor $\varphi(I,a)$ to estimate counterfactual regrets. Nevertheless, the use of regression trees as the regressor limits RCFR’s applicability to smaller games, and the necessity for manually crafted features remains a drawback.
After advantage-based regret minimization (ARM) [88] merges CFR with deep RL in single-agent scenarios, a growing body of research has focused on applying CFR in conjunction with neural networks to multi-agent scenarios. Double Neural CFR [89] utilizes two neural networks: one for estimating counterfactual regrets and another for approximating the average strategy. In a similar vein, Deep CFR [90] leverages an advantage network $V(I,a|\theta_{p})$ to estimate counterfactual regrets with each player having a distinct hyperparameter $\theta_{p}$ and employs $\pi(I,a|\theta_{\pi})$ for strategy estimation after the training process of the advantage network. Since these two networks are trained in sequence rather than concurrently, the strategy for each intermediate iteration remains conditioned on the output of the advantage network: $h(i)=V(I,a|\theta_{p})$. Despite similarities, Deep CFR distinguishes itself from Double Neural CFR through its data collection and its proven effectiveness in larger-scale poker games. Moreover, Single Deep CFR (SD-CFR) [91] demonstrates that training an average strategy network is unnecessary, with only an advantage network required. Building on the foundation of SD-CFR, DREAM [92] utilizes a learned baseline to maintain low variance in a model-free setting when only one action is sampled at each decision point. Moreover, advantage regret-matching actor-critic (ARMAC) [93] combines the thought of retrospective policy improvement with CFR.
III-F Reassessment of the Framework
After introducing these four categories of self-play algorithms, we will further compare them in this section and illustrate the differences between our framework and other frameworks to demonstrate why our proposed framework is more general. We also summarize the key points in Table II.
Traditional self-play algorithms and the PSRO series share many similarities. Initially, they require only one randomly initiated policy, and as training progresses, the policy population expands. Therefore, in our framework, we use placeholder initialization to initialize the policy population and set $E=1$ for these two categories. The interaction matrix is typically a lower triangular matrix (Corollary 1 and Corollary 2). The primary difference between the PSRO series and traditional self-play algorithms is that the PSRO series employs more complex MSSes, designed to handle more intricate tasks. For example, $\alpha$-PSRO [46] specifically utilizes an $\alpha$-rank-based MSS to address multi-player general-sum games.
Unlike the two previously mentioned categories, the ongoing-training-based series employs a different paradigm by training the entire policy population together. This approach aims to strengthen all policies simultaneously at each epoch, rather than expanding the policy population and relying on newer policies to be stronger. As a result, actual initialization is used for the policy population, and $\pi_{i}(\cdot|h(i))$ is utilized to initialize $\pi_{i}^{h}$ to ensure that policy updates are self-referential. Consequently, the interaction matrix is generally not a lower triangular matrix (Corollary 3).
Lastly, the regret-minimization-based series focuses on the overall performance over time rather than on a single episode. For example, it is well-suited for games like Texas Hold’em, which require strategies involving deception, concealment, and bluffing. The main goal of the training process in this series is to update the regrets associated with different strategies. In our framework, we use $h(i)$ to store this information. Since the policies are determined by $h(i)$, only the most recent policy is relevant. Therefore, the interaction matrix is a unit lower shift matrix (Corollary 4). We also do not need actually to initialize the whole policy population and only need to use $\pi_{i-1}(\cdot|h(i-1))$ to initialize $\pi_{i}^{h}$ in the training process.
Our framework is built upon PSRO [42] and NeuPL [43]. Here, we outline the differences between our framework and these two existing frameworks. The primary distinction between our framework and PSRO is the use of an interaction matrix $\Sigma:=\{\sigma_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{1}}$ to represent the opponent sampling strategy, allowing for the integration of more complex competitive dynamics. Moreover, in our framework, $\sigma_{i}$ denotes the opponent sampling strategy, which specifies how to sample opponents’ policies against policy $i$, rather than being the meta-strategy of policy $i$. Additionally, our framework incorporates a policy condition function $h(i)$, making it more general than NeuPL, where policies are conditioned on $\sigma_{i}$. This enhancement gives our framework greater expressiveness. Furthermore, we describe how to compute the oracle (Line 4 in Alg. 1) in three different ways (Alg. 2, Alg. 3 and Alg. 4) to provide a clearer understanding. Also, to the best of our knowledge, our framework is the first self-play framework to integrate the regret-minimization-based series, which is a significant self-play paradigm.
Table II: Overview of Main Self-play Algorithms within Our Framework.
Algorithms MSS h(i) Categories $E$ Initialization of $\Pi$ Initialization of $\pi_{i}^{h}$
Vanilla Self-play [6] Equ. (33) $\emptyset$ Traditional Self-play $E=1$ Placeholder General
Fictitious Self-play [13] Equ. (34)
$\delta$-Uniform Self-play [7] Equ. (35)
PFSP [15] Algo. 5
Independent RL Equ. (37)
DO [45] NE-based (Algo. 6) $\emptyset$ PSRO Series $E=1$ Placeholder General
PSRO [42] General
$\alpha$-PSRO [46] $\alpha$-rank-based
JPSRO [47] (C)CE-based
R-NaD [54] Equ. (33)
FTW [70] Equ. (38) $\emptyset$ Ongoing-training-based $E>1$ Actual $\pi_{i}^{h}\leftarrow\pi_{i}(\cdot|h(i))$
NeuPL [43] General $\sigma_{i}$
Simplex-NeuPL [71] General $\sigma_{i}$
Vanilla CFR [72] Equ. (33) $R^{i}(I,a)$ Regret-minimization-based $E=1$ Placeholder $\pi_{i}^{h}\leftarrow\pi_{i-1}(\cdot|h(i-1))$
CFR+ [75] $R^{+,i}(I,a)$
CFR-D [85] $R^{i}(I_{*},a)$
RCFR[87] $\varphi(I,a)$
Deep CFR[87] $V(I,a|\theta_{p})$
IV Empirical Analysis
In this section, we introduce iconic applications of self-play by categorizing the scenarios into three distinct groups: board games, which typically involve perfect information; card games and Mahjong, which usually involve imperfect information; and video games, which feature real-time actions rather than turn-based play. We then illustrate how self-play is applied in each of these complex scenarios and provide a comparative analysis of these applications in Table III.
IV-A Board Games
The landscape of board games, the majority of which are perfect information games, was previously revolutionized by the introduction of two key techniques: position evaluation and Monte Carlo tree search (MCTS) [94, 95]. These methodologies, with minor modifications, demonstrated superhuman effectiveness in solving board games such as chess [96], checkers [97], othello [98], backgammon [99], and Scrabble [100]. In contrast, the application of these techniques to the game of Go with an estimated $2.1\times 10^{170}$ legal board configurations, only enabled performance at the amateur level [101, 102, 103, 104, 105]. In light of this, our discussion will specifically focus on the game of Go to illustrate the application of self-play. In addition to Go, we will broaden our exploration to include Stratego, a board game characterized by imperfect information, contrasting the majority of board games that are based on perfect information.
IV-A1 Go
The paradigm of Go is revolutionized with the launch of DeepMind’s AlphaGo series [8, 9, 10, 11], which leveraged the power of self-play to significantly elevate performance, setting a new benchmark in the field of Go.
In AlphaGo [8], the training regime can be split into three stages. In the first stage, supervised learning with expert data trains a fast policy network $p_{\pi}(a|s)$ for rollouts in MCTS expansion and a precise policy network $p_{\sigma}(a|s)$. The second stage employs RL to refine the policy network $p_{\rho}(a|s)$ based on $p_{\sigma}(a|s)$ and subsequently trains a value network $v_{\theta}(s)$. Self-play is instrumental here. More specifically, $p_{\rho}(a|s)$ is refined by competing against a randomly chosen historical version $p_{\rho^{-}}(a|s)$, akin to the MSS shown in Equ. (34). Afterwards, $v_{\theta}(s)$ is trained using the samples from the games in which the trained policy network $p_{\rho}(a|s)$ competes against itself. In the final stage, MCTS integrates the policy and value networks to inform action selection.
Unlike AlphaGo, AlphaGo Zero [9] does not require any expert data except game rules and instead learns through self-play. It utilizes only one network $f_{\theta}(s)$ to concurrently predict the action probabilities $p=\text{Pr}(a|s)$ and the state value $v$. Each move is generated by MCTS with guidance from $f_{\theta}(s)$ to aid MCTS expansion, as opposed to the rollouts used in AlphaGo. Self-play is employed to generate data and refine $f_{\theta}(s)$ with the current best policy competing against itself, a process analogous to the MSS referenced in Equ. (37). Furthermore, for a new policy to be incorporated into the policy pool, it must surpass a 55 percent win rate against its predecessor, aligning with the stipulations set in Algo. 2 at Line 1.
AlphaZero [10] extends AlphaGo Zero to include games beyond Go, such as Chess and Shogi, with some modifications. Notably, a draw is introduced as an additional expected outcome, and data augmentation is omitted due to the asymmetry of Chess and Shogi. Concerning the self-play procedure, the only difference between AlphaZero and AlphaGo Zero is that AlphaZero utilizes the newly updated network without the validation process present in AlphaGo Zero.
Building upon AlphaZero, MuZero [11] takes the concept of learning from scratch to the next level, even operating without predefined game rules. MuZero incorporates ideas from model-based RL to model the dynamics of games. More concretely, in addition to the prediction network $f$ (similar to the networks in AlphaGo Zero and AlphaZero), MuZero introduces a dynamics network $g$ to model the MDP and a representation network $h$ to map observations to hidden states. These three networks are trained jointly. Similarly to AlphaGo Zero and AlphaZero, MuZero employs MCTS guided by the three aforementioned networks to make decisions. The self-play process in MuZero operates similarly to that in AlphaZero. In practice, in addition to excelling in board games like Go, MuZero also achieves state-of-the-art performance in Atari games. Furthermore, several studies have extended the capabilities of MuZero. For example, Stochastic MuZero [106] learns a stochastic model instead of a deterministic model to enhance its performance in more complex scenarios. Sampled MuZero [107] makes learning feasible in games with intricate action spaces.
IV-A2 Stratego
Unlike most board games which are perfect information games, Stratego, a two-player imperfect information board game, distinguishes itself by incorporating elements of memory, deduction, and bluffing. This complexity is further amplified by the game’s long episode length and a large number of potential game states, estimated $10^{535}$ [108]. The game is divided into two phases: the deployment phase, where players secretly arrange their units, setting the stage for strategic depth, and the game-play phase, where the objective is to deduce the opponent’s setup and capture their flag. The depth and computational complexity of the game remained a challenge until breakthroughs such as DeepNash [108] showed promising advances in AI’s ability to tackle it.
DeepNash scales up R-NaD [54] to neural R-NaD. It employs a neural network with four heads: one for value prediction, one for the deployment phase, one for piece selection, and one for piece displacement. In the dynamics stage, Neural Replicator Dynamics [109] is utilized to obtain the approximate fixed-point policy effectively. Additionally, DeepNash incorporates train-time fine-tuning and test-time post-processing methods to eliminate unreliable actions, thereby enhancing the robustness and accuracy of its output. DeepNash holds the third-place ranking among all Gravon Stratego players and secures victories in almost every match against existing Stratego bots.
IV-B Card Games and Mahjong
IV-B1 Texas Hold’em
Texas Hold’em, a popular poker game with 2-10 players, is known for its strategic depth and bluffing elements. In its two-player variant, it’s referred to as heads-up Texas Hold’em. The game becomes more complex with additional players. The gameplay begins with each player receiving two private cards (hole cards), followed by a round of betting. Subsequently, three community cards (the flop) are revealed, leading to another betting round. This is followed by the dealing of a fourth (the turn) and a fifth community card (the river), each accompanied by further betting rounds. The objective is to construct the best five-card poker hand from any combination of hole cards and community cards. The betting proceeds until a showdown occurs, where the remaining players disclose their cards. The individual with the superior hand claims the pot. Moreover, Texas Hold’em is distinguished by its three betting formats: no-limit, fixed-limit, and pot-limit. Each format offers a different risk profile and complexity in terms of game state sizes. Among these, the no-limit format is particularly noted for its complexity, allowing players to bet any amount up to their entire stack of chips. Its popularity is partly due to its frequent appearance in tournaments and media, making it a staple in modern poker culture. While simplified versions such as Kuhn Poker and Leduc Poker serve valuable roles in theoretical analysis and educational contexts within game theory, our focus will focus on the development of algorithms designed to compete in real-world Texas Hold’em.
In poker, especially Texas Hold’em, abstraction is crucial to simplifying the inherent complexity of the game [110]. This methodology reduces the vast state space of the game to a manageable size, thus allowing algorithms to calculate strategies more efficiently. There are two main types of abstraction in Texas Hold’em: action abstraction, which simplifies the range of possible moves, and information abstraction, which groups similar hand strengths or pot sizes, streamlining the decision-making process.
Although heads-up limit Texas Hold’em (HULHE), the simplest form of Texas Hold’em with approximately $3.16\times 10^{17}$ game states, was not solved until the introduction of Cepheus [111]. It utilizes fixed-point arithmetic with compression and CFR+ [75] to address the issues of storage and computation, respectively, thus solving HULHE without involving any neural networks. This resulted in superhuman performance and surpassed previous agents. Moreover, NSFP [56] introduces the use of end-to-end RL through self-play training, achieving competitive performance in HULHE. Similarly, Poker-CNN [112], which utilizes a 3D-vector representation of HULHE combined with convolutional networks, learned through self-play and also achieved competitive performance. Deep CFR [90], combines deep neural networks with CFR to surpass the performance of NFSP. However, it’s noteworthy that despite advancements, network-based methods haven’t surpassed the regret-based method used by Cepheus, despite its substantial computational demands.
After the solution of HULHE, research focus shifted to heads-up no-limit Texas Hold’em (HUNL), representing a more complex challenge due to its significantly larger space of game states, approximately $10^{164}$ [113]. Thus, it is impossible to traverse the game tree as Cepheus does in HULHE. The primary technique employed by DeepStack [12] is continual re-solving. This method involves dynamically computing the strategy in real-time when it is the program’s turn to act, focusing on a subtree of limited depth and breadth. To estimate the outcomes of the furthest reaches of this subtree, DeepStack utilizes neural networks. These networks take into account both the player’s range and the opponent’s counterfactual values, which are monitored throughout the entire game. The neural networks employed include the turn network, the flop network, and the auxiliary network, all of which are pre-trained using 10 million randomly generated poker games. The target counterfactual values for these generated poker situations are derived through thousands of iterations using CFR+ [75]. DeepStack has demonstrated competitive performance against professional poker players by integrating these approaches. Libratus [86] employs self-play to develop its blueprint strategy based on an abstraction of the game, leveraging an enhanced version of MCCFR [77]. Initially, the blueprint strategy is deployed for decision-making in the first two rounds, benefiting from the detailed abstraction. In the later rounds, the blueprint strategy aids in nested safe subgame solving, akin to DeepStack’s continual re-solving aiming to get a more precise strategy for subgames in real-time. Additionally, Libratus features a self-improvement component designed to enhance the blueprint strategy by addressing any missing branches. Through these strategies, Libratus has proven capable of defeating professional HUNL players. ReBel [114] expands the concept of a game state in poker into a public belief state (PBS), thereby transforming the imperfect-information game into a perfect-information game with continuous state space. ReBel adopts the methodology of AlphaZero [10], which combines RL with a search algorithm to train both the value network and the policy network by self-play. To mitigate heavy computation and storage demands, AlphaHoldem [115] employs an end-to-end self-play RL method augmented with several additional techniques, to outperform its reimplementation of DeepStack. During training, the system employs the $K$-best Self-play, where it selects the top $K$ agents based on their ELO ratings [35] and the newly initialized agent. It then generates experience replays from competitions among these $K+1$ agents.
When there are more than two players in Texas Hold’em, the game becomes significantly more complex. Pluribus [116] based on the framework of Libratus [86], addresses this complexity in six-player no-limit Texas Hold’em. It employs both action abstraction and information abstraction to simplify the intricate dynamics of the original game. Similarly to Libratus, Pluribus utilizes MCCFR [77] in self-play to develop its blueprint strategy. Building on this blueprint strategy, it conducts a depth-limited search before executing actions. More specifically, diverging from Libratus, Pluribus maintains a streamlined policy pool consisting of only four strategies. This approach assumes that its opponents might adjust their strategies during gameplay among these four strategies, which allows Pluribus to manage complexity more efficiently. Despite its strategy not adapting to the observed tendencies of its opponents, and lacking theoretical proofs of achieving an NE in six-player scenarios, Pluribus has demonstrated its capability by outperforming top professional players in six-player no-limit Texas Hold’em.
IV-B2 DouDiZhu
DouDizhu (a.k.a. Fight the Landlord) is a strategic Chinese card game for three players. In this game, one player takes on the role of the landlord and competes against the other two players, who are the peasants. These two roles have distinct objectives: the landlord’s goal is to outplay the peasants, whereas the peasants work together to defeat the landlord. The essence of the game lies in its blend of strategy, cooperation among peasants, and competition between the two factions. The game is played in two main stages: the bidding stage and the card play stage. During the bidding stage, players vie to become the landlord through a process that involves assessing the strength of their cards. The outcome of this stage determines the game’s dynamics, with the landlord playing against the peasants. The card play stage is where the main action unfolds. Players take turns playing cards in various combinations, intending to be the first to empty their hands. The game’s complexity arises from its requirement for strategic thinking, both in playing cards effectively and in the peasants’ need to cooperate seamlessly. DouDiZhu is characterized by imperfect information, as players can only see their own cards and the cards played, adding a layer of unpredictability and strategic depth to the game. Moreover, with an estimated $10^{76}\sim 10^{106}$ possible game states [117] and an action space comprising $27472$ possible moves [118], DouDiZhu presents a vast landscape of strategic possibilities.
DeltaDou [119], built upon the AlphaZero framework [10], is the first algorithm to achieve expert-level performance in DouDizhu although it only considers the card play stage. It enhances the original framework by incorporating asymmetric MCTS to improve exploration in games characterized by imperfect information. Additionally, DeltaDou employs policy-value networks specially designed with input and output encodings to accelerate the training process. A notable feature of DeltaDou is its ability to infer the cards of other players using Bayesian methods with the help of the policy-value networks. DouZero [120] reduces training costs by opting for a sampling method over the tree search approach. It utilizes Monte Carlo methods combined with deep neural networks. More specifically, it works similarly to independent RL self-play to generate data, which is then stored in buffers. Subsequently, data is sampled from the buffers to update the neural networks. Remarkably, DouZero surpasses DeltaDou in performance, demonstrating the effectiveness of its methodology in mastering. Additionally, DouZero employs supervised learning to train a bidding network used during the bidding stage, utilizing expert data for this purpose. Based on DouZero, DouZero+ [121] enhances the original algorithm by incorporating opponent modeling, which assists the agent in making more informed decisions. Additionally, DouZero+ introduces coach-guided learning, aiming to ensure a balanced intensity of initial hand cards among the three players to further accelerate the training process. PerfectDou [118] leverages perfect information distillation within a Perfect-Training-Imperfect-Execution (PTIE) framework. In this approach, the value network is trained using both perfect and imperfect features, while the policy network is trained solely on imperfect features. The networks are refined through independent RL self-play, utilizing Proximal Policy Optimization (PPO) [122] with Generalized Advantage Estimation (GAE) [123]. Notably, PerfectDou does not incorporate the bidding stage into its model, thereby eliminating the need for expert data in this phase. PerfectDou achieves better performance than DouZero.
IV-B3 Mahjong
Mahjong, renowned for its complex interplay of skill, strategy, and chance, has evolved into various global variants, including the popular Japanese version known as Riichi Mahjong. This game is typically played by four participants who must navigate both the visible aspects of the game, such as discarded tiles, and the hidden elements, like their own hand and the unseen public tiles. The strategic depth and complexity of Mahjong pose significant challenges for artificial intelligence. Despite ongoing research, many AI implementations have yet to reach expert human levels [124, 125, 126]. AI systems must make decisions based on incomplete information and adapt strategies dynamically in response to the actions of multiple opponents, which complicates the decision-making process. Additionally, the game’s intricate winning rules and the enormous number of possible game states, estimated at around $10^{169}$ [117], add layers of complexity. These factors make Mahjong a particularly challenging domain for AI development, requiring sophisticated approaches to handle its unique demands effectively.
Suphx [127] is recognized as one of the first algorithms to master Mahjong, achieving a performance level comparable to expert human players, specifically 10 dan on Tenhou [128], the most popular online Mahjong platform. Initially, Suphx employs supervised learning, utilizing expert data to train its model. It then advances its capabilities through self-play RL, employing several innovative techniques to enhance its effectiveness. These techniques include global reward prediction, which provides a continuous reward throughout the game, and oracle guiding, which offers perfect information at the start of self-play to guide strategy development. Additionally, parametric Monte Carlo policy adaptation is used to refine the policy further. Similarly, NAGA [129], developed by Dwango Media Village, and LuckyJ [130], designed by Tencent, have also achieved the rank of 10 dan on Tenhou. Furthermore, LuckyJ has even defeated human professional players,
IV-C Video Games
In contrast to traditional board games and card games, video games often feature real-time actions, long time horizons and a higher level of complexity due to the broader range of possible actions and observations. Here, we illustrate some representative video games to demonstrate how self-play has advanced AI within these games.
IV-C1 StarCraft II
StarCraft is a real-time strategy (RTS) game developed and published by Blizzard Entertainment. It showcases three distinct species: the Terrans, Zerg, and Protoss, each with unique units and strategic options that enhance the gameplay experience. Renowned for its balanced gameplay, strategic depth, and challenging multiplayer features, the game tasks players with gathering resources, constructing bases, and building armies. Victory requires meticulous planning and tactical execution, with defeat occurring when a player loses all their buildings.
AlphaStar [15], a significant advancement in AI, dominates the 1v1 mode competitions in StarCraft II and has defeated professional players. Its framework bears similarities to that of AlphaGo [8], utilizing supervised learning initially to train the policy with data from human experts. Subsequently, it uses end-to-end RL and a hierarchical self-play method to further train the networks. More specifically, it divides all the agents into three types: main agents, league exploiters, and main exploiters. Additionally, it maintains a policy pool of past players that records all these types of agents. Main agents engage in both FSP and PFSP, competing against main agents themselves and other agents in the policy pool. They are periodically added to the pool and never reset. League exploiters use PFSP to play against all policy pool agents, added to the pool if they show a high win rate and potentially reset to expose global blind spots. Main exploiters only compete with main agents to improve their robustness and are added to the pool after achieving a high win rate or certain training steps, and are reset upon each addition. Most importantly, among those three agent types, the main agent is the core agent and embodies the final AlphaStar strategy. However, the training computation for AlphaStar is extensive. Further studies [131, 132, 133] have enhanced the league self-play training procedure by introducing innovative techniques to reduce computations.
IV-C2 MOBA Games
Multiplayer Online Battle Arena games (MOBA) are a popular genre of video games that blend RTS with role-playing elements. In typical MOBA games, two teams of players control their own unique characters, known as heroes and compete to destroy the opposing team’s main structure, often called the base. Each hero has distinct abilities and plays a specific role within the team, such as Warrior, Tank or Support. Managing multiple lanes and battling under the fog of war, which obscures parts of the map are key aspects of gameplay. Popular examples of MOBA games include League of Legends, Dota 2 and Honor of Kings. These games are known for their complex strategic depth, decision-making under conditions of imperfect information and emphasis on teamwork and player skill.
OpenAI Five [14] defeated the world champion team in a simplified version of Dota 2 that featured a limited pool of heroes and certain banned items. It employs distributed RL using PPO [122] along with GAE [123] to scale up training. Each player on a team shares the same policy network and receives shared observation inputs, with the only difference being a unique identity input for each player. The training process utilizes self-play, with an 80% chance of engaging in vanilla self-play and a 20% chance of using a self-play technique similar to PFSP used in AlphaStar [15]. Here, the probabilities of selecting each policy from the policy pool vary according to its quality score, which is continually updated based on competition results throughout the training. Higher quality scores increase the likelihood of a policy being selected. A new agent is added to the pool every 10 iterations. Similarly to AlphaStar, OpenAI Five requires extensive training resources. Thus, they developed surgery tools to resume training with minimal performance loss to manage continual changes in the environment and codes during development. Additionally, during training, OpenAI Five employs TrueSkill [40] to evaluate agent performance.
Another notable MOBA game, especially popular in China, is Honor of Kings. The 1v1 mode is conquered by [134], which boasts a significant winning rate against top professional players. This success can be attributed to well-designed systems tailored for large-scale RL training and sophisticated algorithm designs that handle complex control challenges effectively. A key element of its RL training methodology is the use of $\delta$-uniform Self-play. The 5v5 mode was later mastered by [135] as well. Unlike OpenAI Five, this work expands the hero pool to 40 heroes, which exponentially increases the possible combinations of lineups and poses a learning collapse problem in RL training. It proposes curriculum self-play learning (CSPL) to mitigate this issue. Specifically, the training process is divided into three stages, drawing on insights from curriculum learning. The first stage involves training fixed lineups through self-play and utilizing human data to keep the two teams balanced to aid policy improvements. The second stage employs multi-teacher policy distillation to produce a distilled model. The final stage uses this distilled model as the initial model for self-play with randomly picked lineups. This approach defeats professional player teams. Additionally, the self-play generated data is used to learn effective lineup drafting by utilizing MCTS and neural networks [136].
IV-C3 Google Research Football
Google Research Football (GRF) [137] is an open-source football simulator that emphasizes high-level actions. It initially offers two scenarios: the football benchmark and the football academy with 11 specific tasks. Here, we focus exclusively on the football benchmark because it presents a more complex scenario that better demonstrates the effects of self-play, as opposed to testing on specially designed tasks. GRF is particularly challenging due to the need for cooperation among teammates and competition against opposing teams. Additionally, it features long trajectories with 3000 steps per round, stochastic transitions, and sparse rewards, all of which contribute to its complexity.
WeKick [138] claimed victory in the GRF competition on Kaggle [139], which modified the game dynamics by allowing competitors to control just one player, either the ball carrier on offense or the nearest defender on defense. WeKick employed self-play strategies similar to those used in league training [15]. It initializes its opponent policy pool using strategies developed through reward shaping via RL, and through Generative Adversarial Imitation Learning [140], which facilitated learning from the tactics of other teams.
Further research delves into the full football game rather than controlling only one player in the team. Team-PSRO [141] extends PSRO [42] to team games, demonstrating approximate TMECor [34] and outperforming self-play in the 4v4 version of the full GRF. In the context of the 11v11 version of the full GRF, where the goalkeeper is rule-based controlled, TiKick [142] employs imitation learning to glean insights from data gathered by WeKick’s self-play, resulting in a multi-agent AI through distributed offline RL. Moreover, Fictitious Cross-Play (FXP) [143] proposes two populations: the main population and the counter population. Policies in the counter population improve solely by cross-playing with policies in the main population as opponents, while policies in the main population engage in playing with policies from both populations. FXP achieves a win rate of over 94% against TiKick in the 11v11 version of the full GRF. TiZero [144], a follow-up to TiKick, combines curriculum learning with FSP and PFSP [15] to avoid expert data reliance, achieving a higher TrueSkill rating [40] than TiKick.
Category Game Number of Perfect Number of Algorithms Self-play Method Search Expert Data
Players Information Game States
Board Game Chess 2 $\checkmark$ $10^{45}$ AlphaZero [10] Independent RL $\checkmark$ $\times$
MuZero [11] Independent RL $\checkmark$ $\times$
Go 2 $\checkmark$ $10^{360}$ AlphaGo [8] FSP $\checkmark$ $\checkmark$
AlphaGo Zero [9] Independent RL $\checkmark$ $\times$
AlphaZero [10] Independent RL $\checkmark$ $\times$
MuZero [11] Independent RL $\checkmark$ $\times$
Stratego 2 $\times$ $10^{535}$ DeepNash [108] R-NaD $\times$ $\times$
Card Game HULHE 2 $\times$ $10^{17}$ Cepheus [111] CFR+ $\checkmark$ $\times$
NFSP [56] NFSP $\times$ $\times$
HUNL 2 $\times$ $10^{164}$ DeepStack [12] CFR+ $\checkmark$ $\times$
Libratus [86] MCCFR $\checkmark$ $\times$
ReBel [114] CFR-D / FSP $\checkmark$ $\times$
Six-player No-limit Texas Hold’em 6 $\times$ $>10^{164}$ Pluribus [116] MCCFR $\checkmark$ $\times$
DouDiZhu 3 $\times$ $10^{76}\sim 10^{106}$ DeltaDou [119] Independent RL $\checkmark$ $\times$
DouZero [120] Independent RL $\times$ $\checkmark$
PerfectDou [118] Independent RL $\times$ $\times$
Mahjong 4 $\times$ $10^{169}$ Suphx [127] Independent RL $\checkmark$ $\checkmark$
Video Game StarCraft II 2 $\times$ / AlphaStar [15] FSP & PFSP $\times$ $\checkmark$
Dota 2 10 $\times$ OpenAI Five [14] Naive SP & PFSP $\times$ $\times$
Honor of Kings 10 $\times$ MOBA AI [135] $\delta$-uniform SP $\checkmark$ $\checkmark$
Google Research Football 22 $\checkmark$ TiZero [144] FSP & PFSP $\times$ $\times$
Table III: Overview of Representative Studies in Empirical Analysis.
V Open Problems and Future Work
Self-play approaches have demonstrated superior performance due to their unique iterative learning processes and ability to adapt to complex environments. However, there are still several areas that require further research and development.
V-A Theoretical Foundation
Although NE has been shown to exist in games with finite players and finite actions [145], computing NE in larger games remains challenging and consequently, many studies aim to achieve approximate NE [146]. However, in some cases, even computing an approximate NE is difficult [65]. Some research has resorted to higher levels of equilibrium, such as CE [47] and $\alpha$-rank [46].
Although many algorithms are developed with theoretical foundations in game theory, there is often a gap when it comes to applying these algorithms to complex real-world scenarios. For example, although AlphaGo [8], AlphaStar [147] and OpenAI Five [14] have achieved empirical success, they lack formal game theory proofs behind their effectiveness. Future work should focus on bridging this gap by combining empirical success with theoretical validation. This could involve developing new algorithms that are both theoretically sound and practically effective in complex scenarios or proving the theoretical underpinnings of existing successful algorithms in complex environments.
V-B Non-stationarity of the Environment
The strategies of the opponent players evolve as training progresses, and the opponents are a vital factor of the environment in the self-play framework. This evolution can cause the same strategy to lead to different results over time, creating a non-stationary environment. This problem is also shared by the MARL area. Future research should aim to develop algorithms that are more robust and can adapt to changing conditions. For example, incorporating opponent modeling into self-play [121] can help agents anticipate changes in opponent strategies and adjust their own strategies proactively, making them more robust to environmental changes.
V-C Scalability and Training Efficiency
The scalability of self-play methods faces significant challenges as the number of teams and players within those teams increases. As the number of participants grows, the complexity of interactions explodes. For example, in OpenAI Five [14], the hero pool size is limited to only 17 heroes. MOBA AI [135] extends this to a 40-hero pool with the help of curriculum learning, but it still cannot cover the entire hero pool available in the actual game. One potential solution is to leverage the inherent connections among players to optimize the learning process. For instance, using graph-based models to represent and exploit the relationships between players can help manage and reduce the complexity of large-scale multi-agent environments.
These scalability issues are fundamentally rooted in the limited training efficiency of self-play methods, which involve both computational and storage aspects. Computational efficiency is a limiting factor due to the iterative nature of self-play, where agents repeatedly play against themselves or past versions, requiring extensive computational resources. Moreover, although forming more complex populations and competitive mechanisms [147] can enhance the intensity and quality of training, it further exacerbates the demand for computational resources. Techniques such as parallel computing, distributed learning, and more efficient neural network architectures could be explored to address these challenges. Additionally, self-play is storage-intensive, as it requires maintaining a policy pool. Even when using a shared network architecture, storing the parameters of large models can be problematic. This issue is particularly pronounced in regret-minimization-based self-play algorithms, which need to store the regrets for each information set and potential action. Managing both the computational load and storage requirements is essential for improving the overall training efficiency and scalability of self-play methods.
V-D With Large Language Models
With their remarkable capability and emergent generalizability, large language models (LLMs) have been regarded as a potential foundation for achieving human-level intelligence [148], and self-play methods have been proposed to fine-tune LLMs, enhance LLMs’ reasoning performance, and build LLM-based agents with strong decision-making abilities. Post-training fine-tuning [149, 150] is a key step in aligning LLM with more desired behaviors, but requires a huge amount of human-annotated data. To reduce reliance on human-annotated data, Self-play fIne-tuNing (SPIN) [151] introduces a self-play mechanism to generate training data using the LLM itself and fine-tuning the LLM by distinguishing self-generated responses from human-annotated data. Some other work [152, 153] also utilizes model-generated data to fine-tune LLMs with minimal human annotations. The idea of self-improvement has also been applied to improve the reasoning ability of LLMs. SPAG [154] observes that self-play in a two-player adversarial language game called Adversarial Taboo can boost the LLM’s performance on a wide range of reasoning benchmarks. Besides improving the capability of LLMs, self-play methods have also contributed to building LLM-based agents with strong strategic abilities. A representative work is Cicero [155] which achieves human-level play in the game of Diplomacy by combining language models with an RL policy trained by self-play. Cicero uses the self-play policy to produce an intended action and prompts the language model to generate languages conditioned on the policy’s intent. Another work [156] also combines LLM with self-play policy but takes a different approach by first prompting the LLM to propose multiple action candidates and then using the self-play policy to produce the action distribution over these candidates. Despite recent progress, applying self-play with LLMs is still underexplored and requires further research efforts.
V-E Reality Applications
Self-play is a powerful technique with widespread applications across various domains. It is particularly effective in addressing some problems abstracted from real-world situations through its iterative learning approach. In the field of economics, for instance, self-play is used to enhance supervised learning models in multi-issue bargaining tasks [157]. Furthermore, self-play proves beneficial in solving combinatorial optimization problems (COPs) such as the traveling salesman problem (TSP) and the capacitated vehicle routing problem (CVRP) [158]. Within the domain of traffic, self-play facilitates the development of human-like autonomous driving behaviors [159] and enables vehicles to learn negotiation strategies to merge on or off roads [160], although currently within 2D simulators.
However, a notable challenge with self-play is its impracticality for direct application in real-world scenarios. Its iterative approach requires extensive trial and error, which is computationally demanding and often involves actions that are impractical or unsafe outside of a controlled simulation. Therefore, self-play is often carried out in simulators. The effective deployment of self-play in real-world applications boils down to overcoming the Sim2Real gap. For example, for tasks where the Sim2Real gap is less pronounced, self-play can be well employed to enhance video streaming capabilities [161], and to address the image retargeting problem [162]. Moreover, EvoPlay [163] leverages self-play to design protein sequences, utilizing the advanced AlphaFold2 simulator [164] to narrow the Sim2Real gap. Similarly, in heterogeneous multi-robot systems, self-play is utilized for adversarial catching tasks, with significant efforts dedicated to Sim2Real transitions for real-world success [165].
VI Conclusion
The burgeoning field of self-play in RL has been systematically explored in this survey. The core idea of self-play, in which agents interact with copies or past versions of themselves, has proven to be a powerful approach for developing robust strategies across various domains. Before delving into the specifics of self-play, this paper first elucidates the MARL framework and introduces the background of self-play. Moreover, the paper presents a unified framework and categorizes existing self-play algorithms into four main categories: traditional self-play algorithms, the PSRO series, the ongoing-training-based series, and the regret-minimization-based series. Details are meticulously provided on how these four groups are seamlessly integrated into our proposed framework, ensuring a comprehensive understanding of their functionalities. The transition from theory to application is underscored by a rigorous analysis of self-play’s role within complex scenarios, such as board games, card games, and video games. Despite the successes of self-play in many areas, challenges remain, such as convergence to suboptimal strategies and substantial computational demands. Future work may focus on solving these problems, integrating self-play with LLMs, and achieving real-world applications. In conclusion, self-play stands as a cornerstone of modern RL research, offering profound insights and tools for developing advanced AI systems. This survey serves as an essential guide for researchers and practitioners, paving the way for further advancements in this dynamic and evolving field.
Acknowledgment
This research was supported by National Natural Science Foundation of China (No.62325405, U19B2019, M-0248), Tsinghua University Initiative Scientific Research Program, Tsinghua-Meituan Joint Institute for Digital Life, Beijing National Research Center for Information Science, Technology (BNRist) and Beijing Innovation Center for Future Chips, Postdoctoral Fellowship Program of CPSF under Grant Number GZC20240830, China Postdoctoral Science Special Foundation 2024T170496.

### v2

@@
-A Survey on Self-play Methods in Reinforcement Learning
+
 Abstract
-Self-play, characterized by agents’ interactions with copies or past versions of itself, has recently gained prominence in reinforcement learning.
+Self-play, characterized by agents’ interactions with copies or past versions of themselves, has recently gained prominence in reinforcement learning (RL).
 This paper first clarifies the preliminaries of self-play, including the multi-agent reinforcement learning framework and basic game theory concepts.
-Then it provides a unified framework and classifies existing self-play algorithms within this framework.
+Then, it provides a unified framework and classifies existing self-play algorithms within this framework.
 Moreover, the paper bridges the gap between the algorithms and their practical implications by illustrating the role of self-play in different scenarios.
@@
 This paper is an essential guide map for understanding the multifaceted landscape of self-play in RL.
-Index Terms:
-Self-play, reinforcement learning, game theory, multi-agent
-I Introduction
-Reinforcement learning (RL) represents a significant paradigm [1] within machine learning, concerned with the optimization of decision-making processes through interaction with an environment.
-It’s fundamentally modeled using a Markov decision process (MDP), a mathematical framework that describes an environment in terms of states, actions, transitions, and rewards.
-Within an MDP, agents operate by observing states, executing actions according to defined policies, receiving subsequent rewards, and transitioning to subsequent states.
+keywords:
+[…]
+1 Introduction
+Reinforcement learning (RL) represents a significant paradigm [120] within machine learning (ML), focused on optimizing decision processes through interaction with an environment.
+In RL, the environment is modeled as a Markov decision process (MDP), where agents observe states, take actions, receive rewards, and cause transitions to new states.
 The primary goal of RL algorithms is to derive the optimal policy that yields the maximum expected accumulated reward over time.
-Deep RL extends traditional RL by employing deep neural networks as function approximators [2].
-This fusion of deep learning with RL has been instrumental in handling high-dimensional state spaces, contributing to breakthroughs in various complex tasks.
-Moreover, the transition from single-agent to multi-agent reinforcement learning (MARL) introduces complex dynamics [3, 4, 5].
-In MARL, the interdependence of agents’ actions introduces significant challenges, as the environment appears non-stationary to each agent.
+Deep RL extends traditional RL by employing deep neural networks as function approximators to handle high-dimensional state spaces, contributing to breakthroughs in various complex tasks [94].
+Moreover, transitioning from single-agent to multi-agent reinforcement learning (MARL) introduces complex dynamics [107, 85, 145].
+In MARL, the interdependence of agents’ actions presents significant challenges, as the environment appears non-stationary to each agent.
 The main issues in MARL are coordination, communication, and equilibrium selection, particularly in competitive scenarios.
@@
 With the help of game theory, a mathematical framework that models the interactions between multiple decision-makers, self-play emerges as an elegant solution to some inherent challenges in MARL.
-By addressing issues such as non-stationarity and coordination, self-play offers an approach where an agent interacts with copies or past versions of itself [6, 7].
+Self-play provides an approach where an agent interacts with copies or past versions of itself [108, 5].
 This method promises a more stable and manageable learning process.
@@
 Although the application of self-play is extensive and promising, it is accompanied by limitations, such as the potential convergence to suboptimal strategies and significant computational requirements[115, 116].
-Although some research takes a broad perspective through empirical game-theoretic analysis (EGTA) [16], it is important to note that there are relatively few comprehensive surveys focusing exclusively on self-play.
-Among these, some studies address the theoretical safety of self-play [17], while others develop an algorithmic framework for self-play that unfortunately does not accommodate the Policy-Space Response Oracle (PSRO) series of algorithms [18].
-Furthermore, another study concentrates exclusively on PSRO [19].
-Although these varied studies are valuable, they do not offer a perspective that fully captures the breadth and depth of self-play.
-Therefore, this survey aims to bridge this gap.
+Although some research adopts a broad perspective through empirical game-theoretic analysis (EGTA)[134], relatively few comprehensive surveys focus solely on self-play algorithms.
+Among them, a study develops an algorithmic framework for self-play, but it does not incorporate the Policy-Space Response Oracle (PSRO) series of algorithms [54], while a separate study focuses exclusively on PSRO [8], without considering other self-play algorithms.
+Another study examines the theoretical safety of self-play [34].
+While these studies are valuable, they do not provide a comprehensive perspective that fully captures the breadth and depth of self-play.
+Therefore, this survey aims to fill this gap.
 The survey is organized as follows.
 Sec.
-II introduces the background of self-play, including the RL framework and basic game theory concepts.
+2 introduces the background of self-play, including the RL framework and game theory concepts.
 Sec.
-III proposes a unified framework and then categorizes existing self-play algorithms into four categories based on this framework, clarifying the self-play landscape.
+3 proposes a unified framework and then categorizes existing self-play algorithms into four categories based on this framework.
 In Sec.
-IV, a comprehensive analysis is performed to illustrate how self-play is applied in various scenarios.
+4, a comprehensive analysis illustrates how self-play is applied in various scenarios.
 Sec.
-V describes open problems in self-play and explores future research directions.
+5 describes open problems in self-play and explores future research directions.
 Finally, Sec.
-VI concludes the survey on self-play.
-II Preliminaries
-In this section, we first introduce the framework of RL. Next, we present the basic game theory concepts and the typical evaluation metrics used in self-play.
-II-A RL Framework
-[…]
-In the RL context, agents interact with the environment based on the subsequent protocol: At each discrete time step $t$, every agent $i$ receives an observation $o_{i,t}$ from the environment and selects an action based on a stochastic policy $\pi_{\theta_{i}}:\mathcal{O}_{i}\times\mathcal{A}_{i}\rightarrow[0,1]$, where $\theta_{i}$ is the parameters.
-After receiving the joint actions $\mathbf{a}_{t}=(a_{1,t},\cdots,a_{n,t})$, the environment undergoes a transition from the current state $s_{t}$ to a subsequent state $s_{t+1}$ according to the transition function $\mathcal{P}$ and sends a reward $r_{i,t+1}$ to every agent $i$.
-The ultimate goal of agent $i$ is to maximize the expected discount accumulated rewards: $\mathbb{E}_{\pi_{\theta_{i}}}[\sum_{t=0}^{\infty}\gamma^{t}r_{i,t}]$.
-II-B Game Theory Concepts
-II-B1 (Im)Perfect Information and (In)Complete Information
-In a game characterized by perfect information, only one player moves at a time.
-Each player has a comprehensive understanding of the current game state, the full history of moves that have been made, and all potential future developments.
+6 concludes the survey on self-play.
+A more detailed overview of our survey is depicted in Fig.
+1, which further illustrates the relationships among the subsections.
+Figure 1: Overview of our survey.
+2 Preliminaries
+This section first introduces the framework of RL, where an agent learns to make decisions by interacting with an environment to maximize cumulative rewards.
+We then present basic game theory concepts, which analyze strategic interactions between rational decision-makers and provide tools to study outcomes based on players’ strategies.
+2.1 RL Framework
+In RL, the environment is modeled as an MDP, where the Markovian assumption states that the environment’s evolution is fully determined by its current state, eliminating the need to consider past states.
+An agent interacts with the environment by taking actions, which result in different states and associated rewards.
+MDPs can be extended to multi-agent settings, known as Markov games (MGs) [79] or stochastic games [113].
+We focus on the most general form: partially observable Markov games (POMGs), where multiple agents interact with the environment, and each agent receives only individual observations instead of the full state, meaning they have limited access to the environment’s state.
+A POMG $\mathcal{G}$ is defined by $\mathcal{G}=(\mathcal{N},\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{P},\mathcal{R},\gamma,\rho)$.
+$\mathcal{N}=\{1,\cdots,n\}$ denotes n agents.
+$\mathcal{S}$ is the state space.
+$\mathcal{A}=\prod_{i=1}^{n}\mathcal{A}_{i}$ is the product of the action space of each agent.
+Similarly, $\mathcal{O}=\prod_{i=1}^{n}\mathcal{O}_{i}$ is the product of the observation space of each agent.
+$\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ denotes the transition probability from one state to another given the actions of each agent.
+=$\mathcal{R}=\{\mathcal{R}_{1},\cdots,\mathcal{R}_{n}\}$, where $\mathcal{R}_{i}:\mathcal{S}\times\mathcal{A}_{i}\rightarrow\mathbb{R}$ denotes the reward function of agent $i$.
+$\gamma\in[0,1]$ is the discount factor.
+$\rho:\mathcal{S}\rightarrow[0,1]$ describes initial state distribution.
+If it is a fully cooperative setting, agents can share the same reward function [145, 106, 39, 68].
+When $n=1,\mathcal{O}_{i}=\mathcal{S}$, the POMG returns to the MDP.
+More concretely, in RL, agents interact with the environment through the following procedure: At each discrete time step $t$, each agent $i$ receives an observation $o_{i,t}$ from the environment and selects an action based on a stochastic policy $\pi_{\theta_{i}}:\mathcal{O}_{i}\times\mathcal{A}_{i}\rightarrow[0,1]$, where $\theta_{i}$ represents the parameters.
+After receiving the joint actions $\mathbf{a}_{t}=(a_{1,t},\cdots,a_{n,t})$, the environment transitions from the current state $s_{t}$ to a subsequent state $s_{t+1}$ according to the transition function $\mathcal{P}$ and sends a reward $r_{i,t+1}$ to every agent $i$.
+The ultimate goal of agent $i$ is to maximize the expected discounted cumulative rewards: $\mathbb{E}_{\pi_{\theta_{i}}}[\sum_{t=0}^{\infty}\gamma^{t}r_{i,t}]$.
+2.2 Game Theory Concepts
+2.2.1 Perfect Information and Imperfect Information
+In a game with perfect information, only one player moves at a time.
+Each player comprehensively understands the current game state, the complete history of moves made, and the potential future outcomes.
 If these conditions are not met, the game is considered to have imperfect information [84, 99].
-In a game of incomplete information, there exists at least one player who is unaware of the payoff of another player; otherwise, it is a game of complete information [27].
-For instance, Go is a game of both perfect and complete information.
-Players have full awareness of the entire game structure, including all possible moves, and they can see every move made by their opponent as they take turns to act (perfect information).
-Furthermore, if the outcomes are considered binary, such as win or loss, the payoff for the players is known to both sides (complete information).
-II-B2 Normal-Form and Extensive-Form
-The normal-form and extensive-form are two different ways of representing games in game theory.
+An example of a game with perfect information is Go.
+In Go, both players have full access to the game state at all times, including the positions of all stones and the identity of the next player.
+In contrast, Texas Hold’em poker is a game with imperfect information.
+Players cannot see their opponents’ private cards, creating uncertainty and requiring decisions based on probabilistic reasoning, bluffing, and interpreting others’ actions.
+2.2.2 Normal-Form and Extensive-Form
+The normal form and extensive form are two distinct representations of games in game theory.
 If a game $\mathcal{G}$ is represented in the normal-form, it can be expressed by $\mathcal{G}=(\mathcal{N},\mathbf{\Pi},\mathbf{u})$.
 $\mathcal{N}=\{1,2,\cdots,n\}$ denotes the players.
-$\mathbf{\Pi}=\Pi_{1}\times\cdots\times\Pi_{n}$ is pure strategy space of all players.
-A vector $\bm{\pi}=(\pi_{1},\cdots,\pi_{n})\in\mathbf{\Pi}$ is called a strategy profile.
-A pure strategy defines a specific and deterministic action for a player in a game, while a mixed strategy designates a probability distribution over the set of pure strategies, allowing randomized actions.
-A mixed strategy for the player $i$ is a probability distribution $\sigma_{i}\in\Delta(\Pi_{i})$, where $\Delta$ is a probability simplex.
+The set $\mathbf{\Pi}=\Pi_{1}\times\cdots\times\Pi_{n}$ represents the pure strategy space for all players.
+A pure strategy specifies a particular and deterministic action for a player in the game.
+A vector $\bm{\pi}=(\pi_{1},\dots,\pi_{n})\in\mathbf{\Pi}$ is called a strategy profile.
+A mixed strategy assigns a probability distribution over the set of pure strategies.
+For player $i$, a mixed strategy is represented by $\sigma_{i}\in\Delta(\Pi_{i})$, where $\Delta$ denotes the probability simplex.
 $\mathbf{u}=(u_{1},\cdots,u_{n})$, where $u_{i}:\mathbf{\Pi}\rightarrow\mathbb{R}$, is a utility function that assigns a real-valued payoff to each player $i$.
@@
 If $\Pi_{1}=\cdots=\Pi_{n}$ and the payoffs are invariant under any permutation of the players’ strategies, the game is a symmetric game.
-If finite players (especially two players) are involved and each player has a finite set of strategies, a game in the normal-form can be directly depicted in a matrix.
-Specifically, in two-player zero-sum symmetric normal-form games, the pure strategy space for both player 1 and player 2 is identical, denoted by $\Pi$, such that $\Pi=\Pi_{1}=\Pi_{2}$.
-As the utility function $u_{1}(\pi_{i},\pi_{j})=-u_{2}(\pi_{i},\pi_{j})$, for simplicity, we can use only one utility function $u$ such that $\forall\pi_{i},\pi_{j}\in\Pi$, if $\pi_{i}$ beats $\pi_{j}$, then $u(\pi_{i},\pi_{j})=-u(\pi_{j},\pi_{i})>0$.
-The evaluation matrix captures the outcomes of the game by detailing the results of different strategies when they are played against each other: $A_{\Pi}=\{u(\pi_{i},\pi_{j}):\pi_{i},\pi_{j}\in\Pi\times\Pi\}$.
+If a finite set of players (especially two players) are involved and each player has a finite set of strategies, a normal-form game can be directly depicted in a matrix.
+In two-player zero-sum symmetric normal-form games, both players share the same pure strategy space, denoted by $\Pi$, such that $\Pi=\Pi_{1}=\Pi_{2}$.
+Since the utility function satisfies $u_{1}(\pi_{i},\pi_{j})=-u_{2}(\pi_{i},\pi_{j})$, we can simplify the utility to a single function $u$, where for $\pi_{i},\pi_{j}\in\Pi$, if $\pi_{i}$ beats $\pi_{j}$, then $u(\pi_{i},\pi_{j})=-u(\pi_{j},\pi_{i})>0$.
+The evaluation matrix captures the game outcomes by detailing the results of different strategies when they are played against each other: $A_{\Pi}=\{u(\pi_{i},\pi_{j}):\pi_{i},\pi_{j}\in\Pi\times\Pi\}$.
 If a game is represented in the extensive-form, it is expressed sequentially, illustrating the sequence of moves, choices made by the players, and the information available to each player during decision-making.
@@
 $\mathcal{N}=\{1,2,\cdots,n\}$ denotes a set of players.
-$c$ is chance and can be regarded as a special agent.
+$c$ is chance node and can be regarded as a special player.
 $H$ represents a set of possible histories and $Z\subseteq H$ is a set of terminal histories.
@@
 For all non-terminal histories $h$ within an information set $I_{i}$, the available actions are the same; otherwise, they are distinguishable.
-Therefore we use $A(I_{i})$ to represent the available actions for the information set $I_{i}$.
+Therefore, we use $A(I_{i})$ to represent the available actions for the information set $I_{i}$.
 Utility functions is denoted by $\mathbf{u}=(u_{1},\cdots,u_{n})$, where $u_{i}:Z\rightarrow\mathbb{R}$.
@@
 A subgame of an extensive-form game is a portion of the game that starts from a single initial node, includes all successors of any node within the subgame, and contains all nodes in the same information set as any node in the subgame.
-The Prisoner’s Dilemma serves as a classic example to illustrate various concepts in game theory. In a modified version of the dilemma, the outcomes are as follows:
+The Prisoner’s Dilemma is a classic example in game theory. The game outcomes are as follows:
   • If one player confesses (C) and the other lies (L), the confessor will serve 1 year in jail, while the liar will serve 8 years.
@@
   • If both players choose to lie, they will each serve only 2 years behind bars.
-The classic scenario is known as the simultaneous Prisoner’s Dilemma, where two players must decide simultaneously whether to confess or lie, without knowledge of the other’s choice.
+The classic scenario is known as the simultaneous Prisoner’s Dilemma, where two players must simultaneously decide whether to confess or lie without knowing the other’s choice.
 Games played in this manner are referred to as static games.
-The normal-form representation is particularly suitable for static games, as it captures the simultaneous nature of decision-making (as depicted in Fig.
-1a).
-Another variant is the sequential Prisoner’s Dilemma, where the second player makes their decision with knowledge of the first player’s action.
+The normal-form representation is suitable for them, as it captures the simultaneous nature of decision-making (as depicted in Fig.
+2a).
+Another variant is the sequential Prisoner’s Dilemma, where the second player decides with knowledge of the first player’s action.
 Games of this nature are called dynamic games.
-The extensive-form representation is well-suited for dynamic games, as it can clearly illustrate the sequence of moves and the information available to each player at each decision point (as shown in Fig.
-1d).
-Furthermore, a normal-form representation can be transformed into an extensive-form representation.
-For instance, the transformation from Fig.
-1a to Fig.
-1b involves representing the simultaneous decisions in a tree structure.
-In this extensive representation, the dotted lines indicate that certain statuses belong to the same information set.
-Conversely, an extensive-form representation can also be transformed into a normal-form representation.
-For example, the transformation from Fig.
-1d to Fig.
-1c involves condensing the sequential decisions into a matrix format.
-In this case, the pure strategy space for player 2 is $2\times 2=4$, reflecting the need for player 2 to have a strategy for each of player 1’s possible actions (confess or lie).
-For instance, the notation for one of player 2’s pure strategies $C\rightarrow C,L\rightarrow C$ in Fig.
-1c indicates that player 2 chooses to confess when he knows player 1 has chosen to confess and player 2 also chooses to confess when he knows player 1 has chosen to lie.
-Other symbols in Fig.
-1c follow the same logic.
+The extensive-form representation is well-suited for them, as it can clearly illustrate the sequence of moves and the information available to each player at each decision point (as shown in Fig.
+2b).
+Figure 2: The example of Prisoner’s Dilemma.
 (a) Matrix representation of simultaneous Prisoner’s Dilemma in normal-form.
-(b) Game Tree Representation of simultaneous Prisoner’s Dilemma in extensive-form.
-(c) Matrix Representation of sequential Prisoner’s Dilemma in normal-form.
-(d) Game Tree Representation of sequential Prisoner’s Dilemma in extensive-form.
-Figure 1: The example of Prisoner’s Dilemma.
-Compared to normal-form games, extensive-form games introduce sequential decision-making, adding complexity to the game structure.
-Extensive-form games also have a close relationship with MGs.
-In MGs with simultaneous moves, agents’ actions are unknown to each other, creating various histories that are condensed into a single information set.
-The game’s utility is the sum of rewards discounted over time [28].
-Beyond normal-form and extensive-form games, as well as MGs, the analysis of complex Markov or extensive-form games often employs a higher-level abstraction: the meta-game.
-The meta-game facilitates the exploration of policy learning within these games, focusing not on isolated actions but on broader strategies that arise from the game’s dynamics.
-In this advanced normal-form context, the policy population consists of the strategies currently employed by players.
-Meta-strategies are mixed strategies that assign probabilities over the policy population in the meta-game.
-II-B3 Transitive Game and Non-transitive Game
+(b) Game tree representation of sequential Prisoner’s Dilemma in extensive-form.
+Beyond normal-form and extensive-form games, the analysis of complex games often involves a higher-level abstraction: the meta-game.
+The meta-game can be viewed as an advanced form of a normal-form game, but it focuses on the selection of policies (also called strategies) rather than isolated actions.
+Each player chooses from a set of policies, rather than a set of actions.
+This set of policies is referred to as the policy population.
+Meta-strategies are mixed strategies that assign probabilities to the policies within the policy population in the meta-game.
+2.2.3 Transitive Game and Non-transitive Game
 For the sake of simplicity, we restrict our focus to two-player zero-sum symmetric games.
-In a transitive game, the strategies or outcomes adhere to a transitive relationship.
+In a transitive game, the strategies or outcomes follow a transitive relationship.
 Formally, $\forall\pi_{i},\pi_{j},\pi_{k}\in\Pi$, if $u(\pi_{i},\pi_{j})>0$ and $u(\pi_{j},\pi_{k})>0$, then it must follow that $u(\pi_{i},\pi_{k})>0$.
@@
 The complexity often results in a mixed-strategy equilibrium, where players randomize their choices among multiple strategies to maximize their expected payoff.
-A quintessential example of a non-transitive game is Rock-Paper-Scissors, in which no single strategy uniformly dominates all others.
-In real-world settings, games exhibit complexities that extend beyond theoretical models.
-[29] argues that real-world games have two salient features: first, practice usually leads to performance improvements; and second, there are a plethora of qualitatively distinct strategies, each with unique advantages and disadvantages.
-In such games, the strategies form a geometric topology resembling a spinning top, where the vertical axis represents the performance of the strategy, and the radial axis represents the length of the longest cycle.
-II-B4 Stage Game and Repeated Game
+A classical example of a non-transitive game is Rock-Paper-Scissors, in which no single strategy uniformly dominates all others.
+2.2.4 Stage Game and Repeated Game
 A stage game (or one-shot game) is a game that is played only once, namely a one-shot interaction between players.
 A famous example of a stage game is the Prisoner’s Dilemma.
-A repeated game is derived from a stage game that is played multiple times.
+A repeated game is derived from a stage game played multiple times.
 Formally, a repeated game based on a stage game $\mathcal{G}$ is defined by playing $\mathcal{G}$ for $T$ periods, where $T$ can be finite or infinite.
-The strategies in a repeated game are history-contingent, meaning that they can depend on the entire sequence of past plays.
-It’s important to note that a stage game or a repeated game can be either represented in the normal-form or extensive-form.
-II-B5 Nash Equilibrium
+The strategies in a repeated game are history-contingent, meaning they can depend on the entire sequence of past plays.
+An example of a repeated game is Texas Hold’em, where players engage in multiple rounds, and each player’s strategy may evolve based on previous rounds, betting patterns, and the history of interactions among players.
+2.2.5 Nash Equilibrium
 For simplicity, $\pi_{i}$ denotes the strategy of player $i$, and $\pi_{-i}$ denotes the strategies of all players other than player $i$.
@@
 where $\epsilon$ is a pre-specified threshold.
-A strategy profile $(\pi_{1}^{*},\pi_{2}^{*},...,\pi_{n}^{*})$ is a Nash equilibrium (NE) if, for every player $i$:
+A strategy profile $(\pi_{1}^{*},\pi_{2}^{*},...,\pi_{n}^{*})$ is a Nash equilibrium (NE) if, for each player $i$:
 $u_{i}(\pi_{i}^{*},\pi_{-i}^{*})\geq u_{i}(\pi_{i},\pi_{-i}^{*}),\forall\pi_{i},$ (3)
-meaning that no player can benefit by changing their strategy unilaterally, given the strategies of all other players.
-In other words, an NE is a situation where the strategy chosen by each player is a BR to the strategies chosen by all other players.
-A strategy profile $(\pi_{1}^{*},\pi_{2}^{*},...,\pi_{n}^{*})$ is an $\bm{\epsilon}$-NE if, for every player $i$:
+meaning that no player can benefit by changing their strategy unilaterally, given the strategies of others.
+In other words, an NE is a situation where each player’s strategy is a BR to the others’ strategies.
+A strategy profile $(\pi_{1}^{*},\pi_{2}^{*},...,\pi_{n}^{*})$ is an $\bm{\epsilon}$-NE if, for each player $i$:
 $u_{i}(\pi_{i}^{*},\pi_{-i}^{*})\geq u_{i}(\pi_{i},\pi_{-i}^{*})-\epsilon,\forall\pi_{i},$ (4)
 meaning that no player can increase their payoff by more than $\epsilon$ by unilaterally changing their strategy.
-However, computing NE is generally intractable in complex games, leading some researchers to utilize $\alpha$-Rank [30] and Correlated Equilibrium (CE) [31] as alternatives.
-Additionally, some studies resort to Replicator Dynamics [32] as a method to analyze and understand the evolution of strategies within these games.
-II-B6 Team Games
-The framework of a two-player zero-sum game can be naturally extended to encompass team-based zero-sum games.
-Von Stengel and Koller analyzed zero-sum normal-form games involving a single team competing against an adversary [33].
-In this type of team game, consider a team denoted by $\mathcal{T}=\{1,2,\cdots,n-1\}$.
-The player $n$ is the adversary $(\mathcal{D})$.
-In this kind of zero-sum normal-form team games, for any player $i,j\in\mathcal{T}$, the utility functions satisfy $u_{i}(\bm{\pi})=u_{j}(\bm{\pi})=u_{\mathcal{T}}(\bm{\pi})$ and $u_{\mathcal{D}}(\bm{\pi})=-(n-1)u_{\mathcal{T}}(\bm{\pi})$.
-A zero-sum single-team single-adversary normal-form game can also be extended to the domain of extensive games [34].
-For any player $i,j\in\mathcal{T}$ and all terminal nodes $z\in Z$, the utility functions satisfy $u_{i}(z)=u_{j}(z)=u_{\mathcal{T}}(z)$ and $u_{\mathcal{D}}(z)=-(n-1)u_{\mathcal{T}}(z)$.
-Let $I_{\mathcal{T}}$ denote the information set defined as $\bigcup_{i\in\mathcal{T}}I_{i}$, and let $A_{\mathcal{T}}$ represent the set of actions accessible in the information sets within $I_{\mathcal{T}}$.
-In scenarios where teammates are unable to coordinate their strategies, the team-maxmin equilibrium (TME) emerges as the most suitable solution concept [33].
-We denote the collection of sequences of player $i$ by $Q_{i}$, representing the sequence-form actions undertaken by player $i$.
-The sequence-form strategy is encapsulated by a function $p_{i}:Q_{i}\rightarrow\mathbb{R}$, which maps each sequence $q\in Q_{i}$ to its associated probability of execution.
-Formally, the TME is articulated as:
-$\arg\max_{p_{1},\ldots,p_{n-1}}\min_{p_{n}}u_{\mathcal{T}}\prod_{i=1}^{n}p_{i}.$ (5)
-Similar to the arguments in normal-form games [33], it can also be inferred that in extensive-form games, a TME exists uniquely, barring any degeneracy and this TME aligns with the team’s utility maximization of the NE [34].
-II-C Evaluation Metrics in Self-play
-In this section, we introduce various self-play evaluation metrics, including NASHCONV (Section II-C1), Elo (Section II-C2), Glicko (Section II-C3), WHR (Section II-C4), and TrueSkill (Section II-C5).
-While NASHCONV measures the distance from Nash equilibrium, the other four metrics evaluate relative skill levels and are compared in Table I.
-It’s important to note that although numerous other evaluation metrics exist, the metrics highlighted here are among the most widely used in the field.
-Table I: Comparison of Relative Skill Evaluation Metrics.
-Elo Glicko WHR TrueSkill
-Uncertainty Modeling $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-Ratings At Any Time $\times$ $\times$ $\checkmark$ $\times$
-Multiplayer In One Team $\times$ $\times$ $\times$ $\checkmark$
-Bayesian Foundation $\times$ $\times$ $\checkmark$ $\checkmark$
-II-C1 NASHCONV
-Nash convergence (NASHCONV) serves as a metric to measure the deviation of a particular strategy from an NE.
-A lower NASHCONV value suggests that the strategy is closer to an NE, implying that no player would benefit from unilateral deviation from the strategy.
-Formally, it is defined as:
-$\text{NASHCONV}(\bm{\pi})=\sum_{i}\max_{\pi_{i}\in\Pi_{i}}u_{i}(\pi_{i},\pi_{-i})-u_{i}(\bm{\pi}),$ (6)
-where $\bm{\pi}$ denotes the combined strategy profile of all participating agents. In particular, in the context of two players, this deviation is commonly referred to as exploitability.
-II-C2 Elo
-The Elo system [35] operates under the assumption that the performance of each player in each game is a normally distributed random variable, with the mean being the player’s current rating.
-In a match between player A and player B, $R_{A}$ and $R_{B}$ are the current ratings of player A and player B.
-The probability density functions for the performance of player A and player B are given by:
-$f_{A}(x)=\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(x-R_{A})^{2}}{2\sigma^{2}}},$ (7)
-$f_{B}(x)=\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(x-R_{B})^{2}}{2\sigma^{2}}}.$ (8)
-$E_{A}$ and $E_{B}$ denote the expected score (or the probability of winning) for player A and player B:
-[…]
-[…]
-For convenience, the logistic function is used to approximate the probability (using the logistic curve with base 10 and choosing 400 to scale and normalize the difference in ratings):
-$E_{A}$ $=\frac{1}{1+10^{(R_{B}-R_{A})/400}},$ (11)
-$E_{B}$ $=\frac{1}{1+10^{(R_{A}-R_{B})/400}}.$ (12)
-Note that $E_{A}+E_{B}=1$.
-$S_{A}$ is the actual outcome of the match for player A (1 for a win, 0.5 for a draw, 0 for a loss).
-$S_{B}=1-S_{A}$.
-After a match, the ratings are updated based on the difference between the actual outcome and the expected outcome.
-The adjustment is given by:
-$\Delta R_{A}=K(S_{A}-E_{A}),$ (13)
-$\Delta R_{B}=K(S_{B}-E_{B}).$ (14)
-$K$ is a scaling factor, often determined by the specific domain of the application, and controls the maximum possible rating change for a single match.
-If two players have nearly identical ratings, the expected outcome will be close to 0.5.
-A win for either player will result in a moderate increase in their rating.
-In contrast, if there is a significant rating difference, the expected outcome will be heavily skewed.
-If the higher-rated player wins, their rating will increase by a value much smaller than $K$, reflecting the anticipated nature of their victory.
-However, if the lower-rated player secures an unexpected win, their rating will surge by a value approaching $K$, signifying the upset.
-The updated ratings become:
-$R_{A}^{\prime}=R_{A}+\Delta R_{A},$ (15)
-$R_{B}^{\prime}=R_{B}+\Delta R_{B}.$ (16)
-However, there are challenges and limitations in Elo.
-First, the Elo system assumes that all matches are equally important.
-This might not be the case in all domains.
-Second, $K$ is often kept constant.
-However, a dynamic $K$ factor might be more appropriate, especially for players new to the rating pool or in scenarios where the importance of the match varies.
-Third, the standard Elo system does not incorporate a decay mechanism to account for the potential degradation or improvement of skills during periods of inactivity.
-Fourth, the basic Elo system is designed for one-on-one competitions.
-Adapting it to team or multi-player scenarios, such as team sports or online multiplayer games, can be challenging.
-Lastly, an important limitation of the Elo rating system is its unsuitability for games that exhibit high levels of non-transitivity [36].
-II-C3 Glicko
-The Glicko system refines the Elo system by introducing a measure of uncertainty or reliability in a player’s rating, termed rating deviation [37].
-The primary motivation is to account for the variability in a player’s performance and the potential changes in skill over time.
-The Glicko-2 system, an extension of the original Glicko system, further refines these concepts and introduces the rating volatility $\sigma$, indicating the degree of expected fluctuation in the player’s rating [38].
-In the Glicko-2 system, $r$ denotes the current rating of the player.
-$RD$ is the player’s rating deviation, and $\sigma$ is the player’s volatility.
-Convert the rating and rating deviation to the Glicko-2 scale:
-$\mu$ $=\frac{r-1500}{173.7178},$ (17)
-$\phi$ $=\frac{RD}{173.7178}.$ (18)
-Then, calculate $v$ representing the estimated variance of the player’s rating based on game outcomes, $E(\mu,\mu_{j},\phi_{j})$ representing the probability of a player with rating $\mu$ defeating an opponent player $j$ and $\Delta$ representing the estimated improvement based only on game outcomes $s_{j}$:
-$v$ $=\left[\sum_{j=1}^{m}g(\phi_{j})^{2}\{1-E(\mu,\mu_{j},\phi_{j})\}\right]^{-1},$ (19)
-$E(\mu,\mu_{j},\phi_{j})$ $=\frac{1}{1+\exp(-g(\phi_{j})(\mu-\mu_{j}))},$ (20)
-$\Delta$ $=v\sum_{j=1}^{m}g(\phi_{j})\{s_{j}-E(\mu,\mu_{j},\phi_{j})\},$ (21)
-where:
-$g(\phi)$ $=\frac{1}{\sqrt{1+3\phi^{2}/\pi^{2}}}.$ (22)
-The update for $\sigma$ is more involved and requires an iterative procedure to solve for its new value $\sigma^{\prime}$. Then, calculate new $\mu^{\prime}$ and $\phi^{\prime}$:
-$\mu^{\prime}$ $=\mu+\phi^{\prime 2}\sum_{j=1}^{m}g(\phi_{j})\{s_{j}-E(\mu,\mu_{j},\phi_{j})\},$ (23)
-$\phi^{\prime}$ $=\frac{1}{\sqrt{\frac{1}{{\phi^{*}}^{2}}+\frac{1}{v}}},$ (24)
-where:
-$\phi^{*}=\sqrt{\phi^{2}+\sigma^{\prime 2}}.$ (25)
-After calculations, the rating and rating deviation are converted back from the Glicko-2 scale:
-$r^{\prime}$ $=173.7178\mu^{\prime}+1500,$ (26)
-$RD^{\prime}$ $=173.7178\phi^{\prime}.$ (27)
-II-C4 WHR
-The Whole-History Rating (WHR) system [39] is a Bayesian rating system designed to estimate players’ skills from their entire game history.
-It is particularly adopted in handling the temporal dynamics of player skills.
-$R_{i}(t)$ denotes Elo rating of the player $i$ at time $t$.
-Similar to Equ.
-(11) and Equ.
-(12), $E_{A}(t)$ and $E_{B}(t)$ denote the expected score (or the probability of winning) for player A and player B at time $t$:
-$E_{A}(t)$ $=\frac{1}{1+10^{(R_{B}(t)-R_{A}(t))/400}},$ (28)
-$E_{B}(t)$ $=\frac{1}{1+10^{(R_{A}(t)-R_{B}(t))/400}}.$ (29)
-Moreover, WHR assumes that ratings of each player $i$ is a Wiener process:
-$r_{i}(t_{2})-r_{i}(t_{1})\sim\mathcal{N}(0,|t_{2}-t_{1}|\omega^{2}),$ (30)
-where $\omega$ is a parameter representing the variability of ratings in time. Thus, if $r^{\prime}=R_{A}(t_{1}),r^{\prime\prime}=R_{A}(t_{2})$:
-$p(r^{\prime\prime}|r^{\prime})=\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{1}{2}(\frac{r^{\prime\prime}-r^{\prime}}{\sigma})^{2}},$ (31)
-where $\sigma=\omega\sqrt{|t_{2}-t_{1}|}$. In addition, a Wiener process is memoryless (or Markovian).
-$\mathbf{R}(t)$ denotes player A’s ratings, $\mathbf{G}$ denotes the observation of game results. Thanks to the Bayes formula, the WHR system is to solve:
-$\arg\max_{\mathbf{R}}p(\mathbf{R}|\mathbf{G})=\arg\max_{\mathbf{R}}\frac{p(\mathbf{G}|\mathbf{R})p(\mathbf{R})}{p(\mathbf{G})}.$ (32)
-$P(\mathbf{G}|\mathbf{R})$ can be derived from the product of Equ.
-(28) and $p(\mathbf{R})$ can be derived from the product of Equ.
-(31).
-Utilizing the iterative process of Newton’s method, we can ascertain the solution to the given problem.
-II-C5 TrueSkill
-TrueSkill [40] is based on a probabilistic graphical model that employs Bayesian inference and adapts to multiple players in multiple teams.
-TrueSkill 2 [41] extends the original TrueSkill model with several enhancements by taking into account the experience of a player, the affiliation with a team, and some game-specific factors such as the number of kills.
-For simplicity, TrueSkill is introduced using the following scenario: team 1 has player 1, while team 2 has player 2 and player 3.
-Ultimately, team 1 defeated team 2.
-This is a simpler case than the one presented in the original paper [40].
-It can be described in an undirected probabilistic graphic model (Fig.
-2).
-The skill of each player, denoted by $s_{i}$, is represented by a Gaussian distribution $N(s_{i};\mu_{i},var_{i})$.
-The performance of each player, denoted by $p_{i}$, is also represented by a Gaussian distribution $N(p_{i};s_{i},\beta^{2})$.
-The performance of each team is denoted by $t_{i}$.
-The difference in team performance is denoted by $d$.
-If $d>\epsilon$, where $\epsilon$ is a small positive threshold, then team 1 beats team 2.
-The TrueSkill algorithm’s updating mechanism within the context of undirected probabilistic graphs uses the sum-product message-passing algorithm.
-This ensures that the skill estimates are refined iteratively, leading to accurate and reliable ratings for each player.
-Figure 2: An undirected probabilistic graphic model example of TrueSkill.
-III Algorithms
+However, computing NE is generally intractable in complex games, so some researchers utilize $\alpha$-Rank [101] and Correlated Equilibrium (CE) [2] as alternatives.
+Some studies also resort to Replicator Dynamics [124] to analyze strategy evolution.
+3 Algorithms
 Based on existing self-play work [54, 73, 82, 42], we propose a self-play framework (Algo.
 1) that boasts enhanced expressivity and superior generalization capabilities.
-The framework is adept at handling multi-homogeneous-player general-sum games.
-It’s important to note that although homogeneous players represent a specific subset of heterogeneous players, the latter can be reformulated as the former by expanding the dimensions of the input vector, which essentially entails embedding agent identity information.
-Moreover, as players are homogeneous, we assume each player shares the same policy population.
-To clarify our framework, we describe it in a more easily understandable way.
-All players share a common policy population with a fixed maximum size.
-In each iteration, a new policy is initialized for training, and opponent policies are sampled from the existing policy population.
-During this iteration, the opponent policies typically remain fixed, while only the policy being trained is updated.
-After the training process, the new policy replaces one of the policies in the policy population.
-An evaluation metric is then used to assess the performance of the updated policy population.
-Based on this performance, the strategy for sampling opponents is adjusted for the next iteration.
-This process is repeated.
-For a more detailed and precise description, please refer to Sec.
-III-A.
-In addition, we divide self-play algorithms into four primary groups: traditional self-play algorithms (Sec.
-III-B), the PSRO series (Sec.
-III-C), the ongoing-training-based series (Sec.
-III-D), and the regret-minimization-based series (Sec.
-III-E).
-We analyze how these four categories align with our framework in their respective sections and introduce representative algorithms in each category.
-Moreover, in Sec.
-III-F, we discuss the differences among these four categories and their connections to our proposed framework.
-We also explain why our framework is more general than existing frameworks.
-Furthermore, we summarize the major elements of key algorithms within our framework for each category in Table II.
-III-A Framework Definition
-Algorithm 1 the Framework of Self-play
-0:  $\Pi:=\{\pi_{i}(\cdot|h(i))\}_{i=1}^{N}$$\triangleright$ Policy population.
-0:  $\Sigma:=\{\sigma_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{1}}$$\triangleright$ Interaction matrix.
-0:  $\mathcal{F}:\mathbb{R}^{N\times C_{2}}\rightarrow\mathbb{R}^{N\times C_{1}}$$\triangleright$ MSS.
-1:  for $e\in[[E]]$ do
-2:     for $\sigma_{i}\in\Sigma$ do
-3:        Initialize $\pi_{i}^{h}$
-4:        $\pi_{i}^{h}\leftarrow\text{ORACLE}(\pi_{i}^{h},\sigma_{i},\Pi)$$\triangleright$ Compute the oracle.
-5:        $\mathcal{P}\leftarrow\text{EVAL}(\Pi)$triangleright Get policies’ performance.
-6:        $\Sigma\leftarrow\mathcal{F}(\mathcal{P})$$\triangleright$ Update the interaction matrix.
-7:     end for
+For simplicity, our framework is illustrated for symmetric games.
+All players share a policy population with a fixed maximum size.
+In each iteration, a newly initialized policy is trained, while opponent policies are sampled from the population.
+After training, the new policy is added to the population.
+The updated policy population is then evaluated, and then the opponent sampling strategy is recalculated for the next iteration.
+This section is organized as follows: In Sec.3.1, we provide a formalized description of our framework.
+We then categorize self-play algorithms into four primary groups: traditional self-play algorithms (Sec.
+3.2), the PSRO series (Sec.
+3.3), the ongoing-training-based series (Sec.
+3.4), and the regret-minimization-based series (Sec.
+3.5).
+We analyze how these four categories align with our framework and introduce corresponding algorithms for each category.
+To make it more straightforward, we highlight the representative classic algorithms from these and present them in Table 1 for comparison within our framework.
+In Sec.
+3.6, we compare the four categories, explain our classification rationale, and highlight the distinctions between our framework and existing ones to demonstrate its greater generality.
+3.1 Framework Definition
+Figure 3: Illustration of our framework.
+The top row shows two different initialization methods for the policy population $\Pi$ and the interaction matrix $\Sigma$: lazy initialization and immediate initialization (corresponding to Line 1 in Algo.
+1).
+The bottom row illustrates two training paradigms: standard training and ongoing training (corresponding to Lines 2~9 in Algo.
+1).
+We categorize self-play algorithms into four types: traditional self-play, the PSRO series, the ongoing-training-based series, and the regret-minimization-based series.
+Among these, traditional self-play, the PSRO series, and the regret-minimization-based series utilize lazy initialization and standard training, whereas the ongoing-training-based series employs immediate initialization and ongoing training.
+For a more detailed description and analysis, please refer to Sec.3 and Table1.
+Algorithm 1 the Framework of Self-play.
+1:  Initialize $\Pi,\Sigma$
+2:  for $e\in[[E]]$ do
+3:     for $\sigma_{i}\in\Sigma$ do
+4:        Initialize $\pi_{i}^{h}$
+5:        $\pi_{i}^{h}\leftarrow\textit{ORACLE}(\pi_{i}^{h},\sigma_{i},\Pi)$$\triangleright$ Compute the oracle.
+6:        $\mathcal{P}\leftarrow\textit{EVAL}(\Pi)$triangleright Get policies’ performance.
+7:        $\Sigma\leftarrow\textit{MSS}(\mathcal{P})$$\triangleright$ Update the interaction matrix.
 8:     end for
-9:  return  $\Pi,\Sigma$
-In Algo. 1, we define a unified framework of self-play based on [43, 42, 44, 18]. First, the input of the framework is defined as follows:
-  • $\Pi$: Each policy $\pi_{i}$ in the policy population $\Pi$ is conditioned on a policy condition function $h(i)$, which is determined by specific algorithms.
-We will further introduce this policy condition function in the following sections.
-For simplicity, we denote each policy $\pi_{i}(\cdot|h(i))$ by $\pi_{i}^{h}$.
-It’s important to note that $i$ refers to the $i$th policy in the policy population, not to a player.
-Additionally, the hyper-parameter $N$ denotes the policy population size of $\Pi$, not the number of game players.
-Furthermore, $\Pi$ can be initialized in two different ways.
-First, $\Pi$ can be considered as initialized with $N$ placeholder policies, meaning that there is no actual initialization of the policies in practice.
-Instead, each policy will be actually initialized in the training iterations (Line 3 in Algo.
-1).
-We refer to this approach as the placeholder initialization.
-Second, $\Pi$ can be initialized with $N$ real policies, which may include random initialization or pretrained models.
-We refer to this approach as the actual initialization.
-A policy $\pi_{i}^{h}$ is considered an ineffective policy if it serves as a placeholder policy; otherwise, it is deemed an effective policy.
-Additionally, Table II provides a clearer illustration of the initialization of $\Pi$ and the expression of $h(i)$ across different categories of the main self-play algorithms.
-  • $\Sigma:=\{\sigma_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{1}}$: The interaction matrix of the policy population.
-$\sigma_{i}\in\mathbb{R}^{C_{1}}$ represents the opponent sampling strategy of policy $i$.
-Namely, $\sigma_{i}$ illustrates how to sample opponent(s)’ policies against policy $i$.
-For instance, let $\sigma_{i}$ represent the probability of policies for each opponent.
-In this context, $C_{1}=N^{n-1}$, where n denotes the number of players.
-Alternatively, $\sigma_{i}$ can also be considered as sampling parameters within a sampling network.
-Especially, in a two-player game, if $C_{1}=N$ and $\sigma_{ij}$ represents the probability that policy $i$ is optimized against policy $j$, $\Sigma$ can be depicted in directed interaction graphs (Fig.
-3).
-It’s important to note that, unlike the original PSRO framework [42], $\sigma_{i}$ here is not the meta-strategy of policy $i$.
-Instead, in the two-player setting, if $\sigma_{i}$ in our framework is the opponent’s meta-strategy against policy i, our framework can reduce to the original PSRO framework.
-  • $\mathcal{F}:\mathbb{R}^{N\times C_{2}}\rightarrow\mathbb{R}^{N\times C_{1}}$: A meta-strategy solver (MSS) $\mathcal{F}$ takes the performance matrix $\mathcal{P}:=\{p_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{2}}$ as its input and produces a new interaction matrix $\Sigma\in\mathbb{R}^{N\times C_{1}}$ as its output.
-$p_{i}$ is the performance of policy i.
-For instance, $p_{i}$ can be depicted as the relative skill like Elo ratings ($C_{2}=1$) or can be depicted as the payoff tensor ($C_{2}=N^{n-1}$, where n is the number of players).
-Specially, in two-player symmetric zero-sum games, the expected payoffs can serve as the evaluation metric.
-In such cases, $\Sigma$ is a square matrix ($C_{2}=N$).
-In addition, Table II concludes the MSS of representative self-play algorithms across these four categories.
-Figure 3: When $C_{1}=N$ and $\sigma_{ij}$ represents the probability that policy $i$ is optimized against policy $j$, we can consider various examples of conventional self-play algorithms: In the Top section, we define $\Sigma\in\mathbb{R}^{3\times 3}$ as $\{\sigma_{i}\}_{i=1}^{3}$.
+9:  end for
+10:  return  $\Pi,\Sigma$
+In Algo.
+1, we define a unified self-play framework based on [82, 73, 42, 54].
+We combine visual illustrations (Fig.
+3) with detailed textual descriptions to explain the key processes of our framework.
+Next, we provide an in-depth explanation of these processes:
+Figure 4: Interaction matrix $\Sigma$ examples.
+When $C_{1}=N$ and the opponent sampling strategy $\sigma_{ij}$ represents the probability that policy $i$ is optimized against policy $j$, the interaction matrix $\Sigma$ can be depicted in directed interaction graphs.
+Here, we consider three representative self-play algorithms.
+In the Top section, we define the interaction matrix $\Sigma\in\mathbb{R}^{3\times 3}$ as $\{\sigma_{i}\}_{i=1}^{3}$.
 In the Bottom section, we present directed interaction graphs where the outgoing edges from each node are equally weighted, and their weights collectively sum to one.
@@
 Note that this figure is reproduced from [82] and this concept is initially proposed by [42].
-Next, the core process within the framework is as follows:
-  • For epoch $e\in[[E]]$ (Line 1 in Algo.
+  • $\Pi:=\{\pi_{i}(\cdot|h(i))\}_{i=1}^{N}$: Each policy $\pi_{i}$ in the policy population $\Pi$ is conditioned on a policy condition function $h(i)$, which provides supplementary information for certain algorithms.
+We denote the policy $\pi_{i}(\cdot|h(i))$ as $\pi_{i}^{h}$.
+Note that $i$ refers to the $i$th policy in the population, not the $i$th player, and $N$ denotes the policy population size, not the number of players.
+    $\Pi$ can be initialized (Line 1 in Algo.
+1) in two ways: lazy initialization and immediate initialization.
+In lazy initialization, $\Pi$ starts with $N$ placeholder policies, with the actual policies being initialized during the training iterations (Line 4 in Algo.
+1).
+In immediate initialization, $\Pi$ is initialized upfront with $N$ real policies, which may be randomly generated or pre-trained models.
+A policy $\pi_{i}^{h}$ is considered inactive if it is a placeholder; otherwise, it is considered active.
+  • $\Sigma:=\{\sigma_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{1}}$: The interaction matrix $\Sigma$ consists of rows $\sigma_{i}$, where each $\sigma_{i}\in\mathbb{R}^{C_{1}}$ represents the opponent sampling strategy of policy $i$, indicating how to sample opponents’ policies.
+The dimension $C_{1}$ varies according to the sampling method.
+For instance, $\sigma_{i}$ can represent the probability distribution over opponent policies in the policy population $\Pi$, where $C_{1}=N^{n-1}$ and $n$ denotes the number of players.
+Alternatively, $\sigma_{i}$ can be viewed as parameters of a sampling network.
+Specially, in a two-player game, if $C_{1}=N$ and $\sigma_{ij}$ represents the probability that policy $i$ is optimized against policy $j$, $\Sigma$ can be depicted in directed interaction graphs (Fig.
+4).
+    $\Sigma$ will be initialized (Line 1 in Algo.
+1) based on the initialization of $\Pi$.
+If $\Pi$ uses lazy initialization, $\Sigma$ will be initialized as an empty matrix.
+If $\Pi$ uses immediate initialization, $\Sigma$ will be calculated based on $\Pi$.
+We will discuss this in more detail later.
+  • For epoch $e\in[[E]]$ (Line 2 in Algo.
 1): $E$ denotes the total number of epochs for the entire policy population.
 For example, if the algorithm only introduces new policies into the population without updating existing ones, then $E=1$.
-This implies that only the ineffective policy is likely to be chosen for training in each iteration and turned into an effective one, while effective policies remain unchanged.
-Conversely, if effective policies are updated multiple times throughout the algorithm, then $E>1$.
-Indeed, $E$ accurately rechoflects the number of updates performed.
-Additionally, Table II summarizes the values of $E$ for different categories of self-play algorithms.
-  • Initialize $\pi_{i}^{h}$ (Line 3 in Algo.
+This implies that only the inactive policy is likely to be chosen for training in each iteration and turned into an active one, while activa policies remain unchanged.
+Conversely, if active policies are updated multiple times throughout the algorithm, then $E>1$.
+Indeed, $E$ accurately reflects the number of updates performed.
+  • Initialize $\pi_{i}^{h}$ (Line 4 in Algo.
 1): The initialization of $\pi_{i}^{h}$ can vary depending on the algorithm being used.
-For instance, it may be initialized randomly, by leveraging pre-trained models [8, 15] or through a recently updated policy.
-We provide detailed descriptions of the initialization process for each algorithm series.
-Also, these are summarized in Table II.
-  • $\text{ORACLE}(\pi_{i},\sigma_{i},\Pi)$ (Line 4 in Algo.
+It can be initialized randomly or by leveraging pre-trained models [115, 129] or through a recently updated policy.
+  • $\text{ORACLE}(\pi_{i},\sigma_{i},\Pi)$ (Line 5 in Algo.
 1): The ORACLE is an abstract computational entity that returns a new policy adhering to specific criteria.
@@
 However, it often requires considerable computational effort.
-This computational demand can be a limitation, particularly in complex environments or with large action spaces.
-(2) To alleviate the computational demands, the approximate best response (ABR) oracle is introduced, which can be calculated using techniques such as RL (Algo.
+(2) To alleviate the computational demands, the approximate best response (ABR) oracle is introduced, using techniques such as RL (Algo.
 2), evolution-theory-based methods (Algo.
@@
 4).
-(3) Additionally, other specially crafted ORACLEs are either tailored to new MSSes [46, 47] or introduced to enhance diversity [48, 49, 50].
+(3) Other specially crafted ORACLEs are either tailored to specific meta-strategy solvers (MSSes) [96, 86] or introduced to enhance diversity [4, 103, 83].
     Some details of Algo.
 23 and 4 need to be mentioned.
-$\bm{\pi_{opp}}$ denotes opponents’ policies.
-Moreover, in Algo.
-2, off-policy RL algorithms [2, 51, 52, 53], typically require a replay buffer to gather samples; however, for simplicity, they are not explicitly included.
-Furthermore, Line 1 in Algo.
-2 suggests that RL may not always yield an ABR that meets specific criteria.
-For instance, AlphaGo Zero [9] stipulates that the new policy must achieve a win rate exceeding 55 percent against its predecessor.
-Therefore, validation of the new policy is sometimes required.
-In addition, it is important to note that Algo.
-3 is specifically devised for R-NaD [54], while Algo.
-4 is expressly designed for a series based on regret minimization.
+Algo.
+2 is a RL method that can be widely used across different categories.
+Moreover, off-policy RL algorithms typically require a replay buffer to gather samples; however, they are not explicitly included for simplicity.
+Algo.
+3 is specifically utilized by Regularized Nash Dynamics (R-NaD) [105] and Algo.
+4 is specifically utilized by the regret-minimization-based series.
 They will be introduced in detail in Sec.
-III-C5 and Sec.
-III-E respectively.
-  • $\text{EVAL}(\Pi)$ (Line 5 in Algo.
+3.3.5 and Sec.
+3.5 respectively.
+  • $\text{EVAL}(\Pi)$ (Line 6 in Algo.
 1): Evaluating the policy population $\Pi$.
-There are multiple evaluation metrics available to gauge the performance of each policy.
+There are multiple evaluation metrics available to assess the performance of each policy.
 The performance matrix is represented as $\mathcal{P}:=\{p_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{2}}$.
-Moreover, only effective policies will be evaluated to save on computation.
-Notably, if the MSS produces a constant matrix irrespective of the input, the evaluation step can be skipped to reduce computations.
-Algorithm 2 Compute the Oracle in RL
+$p_{i}$ is the performance of policy i, and the dimension $C_{2}$ depends on the evaluation metric used.
+For instance, $p_{i}$ can be depicted as the relative skill like Elo ratings ($C_{2}=1$) or can be depicted as the payoff tensor ($C_{2}=N^{n-1}$, where n is the number of players).
+Specially, in two-player symmetric zero-sum games, the expected payoffs can serve as the evaluation metric.
+In such cases, $\mathcal{P}$ is a square matrix ($C_{2}=N$).
+  • $\textit{MSS}:\mathbb{R}^{N\times C_{2}}\rightarrow\mathbb{R}^{N\times C_{1}}$ (Line 7 in Algo.
+1): A meta-strategy solver (MSS) MSS takes the performance matrix $\mathcal{P}$ as its input and produces a new interaction matrix $\Sigma$ as its output.
+If the MSS produces a constant matrix irrespective of the input, the evaluation step (Line 6 in Algo.
+1) can be skipped to reduce computations.
+    Moreover, if $\Pi$ uses immediate initialization, $\Sigma$ will be initialized (Line 1 in Algo. 1) by first evaluating the performance and then applying the MSS.
+Algorithm 2 Compute the Oracle in RL.
 0:  $\pi_{i}^{h}$$\triangleright$ Policy $i$ is being trained.
@@
 14:  return  $\pi_{i}^{h}$
-Algorithm 3 Compute the Oracle in Evolution Theory
+Algorithm 3 Compute the Oracle in Evolution Theory.
 0:  $\pi_{i}^{h}$$\triangleright$ Policy $i$ is being trained.
@@
 5:  return  $\pi_{i}^{h}$
-Algorithm 4 Compute the Oracle in Regret Matching
+Algorithm 4 Compute the Oracle in Regret Matching.
 0:  $\pi_{i}^{h}$$\triangleright$ Policy $i$ is being trained.
@@
 6:  return  $\pi_{i}^{h}$
-Moving forward, we categorize self-play algorithms into four primary groups. We provide a detailed analysis of how each algorithm in these categories integrates and aligns with our proposed framework.
-III-B Traditional Self-play Algorithms
+3.2 Traditional Self-play Algorithms
 Traditional self-play algorithms involve agents improving their strategies by repeatedly playing against themselves, allowing them to explore various strategies and enhance their decision-making abilities without external input.
-These algorithms can start with agents training against their most recent version, helping to identify and exploit weaknesses.
-Additionally, other approaches involve training against a set of strategies from different iterations, enabling agents to develop robust and adaptive strategies.
-In this section, we will explain how traditional self-play algorithms fit into our framework and introduce representative traditional self-play methods, ranging from simpler forms to more complex ones.
-III-B1 Integration into Our Framework
-We can incorporate the traditional self-play algorithms into our proposed framework (Algo.
-1), using the following settings.
-First, the policy population $\Pi$ utilizes placeholder initialization, meaning that initially, the policies are placeholders rather than actually initialized strategies.
-This initialization approach is used because the policy population in traditional self-play algorithms is intended to grow with each iteration.
-Second, we set $E=1$ because in traditional self-play algorithms, only an ineffective policy is likely to be chosen for training in each iteration, thereby turning it into an effective policy.
-Here, the policy population size $N$ serves as the upper limit for the number of effective policies in the population.
+The simplest form involves agents training against their most recent version to identify weaknesses.
+Other approaches involve training against a set of strategies from different iterations, enabling agents to develop robust and adaptive strategies.
+This section will explain how traditional self-play algorithms fit into our framework and introduce representative traditional self-play methods, ranging from simpler forms to more complex ones.
+3.2.1 Integration into Our Framework
+Traditional self-play algorithms can be incorporated into our proposed framework (Algo.
+1) with the following settings.
+First, the policy population $\Pi$ utilizes lazy initialization because the policy population in traditional self-play algorithms is intended to grow with each iteration.
+Second, we set $E=1$ because only the inactive policy can be trained in each iteration, then turning into an active policy.
+Here, the policy population size $N$ serves as the upper limit for the number of active policies in the population.
 In other words, we use $N$ iterations to optimize the policy.
 Third, the strategy being trained $\pi_{i}^{h}$ can be initialized in a general manner.
-For instance, the strategy can be initiated randomly, which signifies starting the learning process from scratch.
+For instance, the strategy can be initialized randomly, learning from scratch.
 More often, $\pi_{i}^{h}$ is initialized by $\pi_{i-1}(\cdot|h(i-1))$, allowing incremental learning and adaptation based on the most current trained policy to accelerate convergence.
-Fourth, since the policies in traditional self-play algorithms are not conditioned, we simply set $h(i)=\emptyset$.
-Next, we outline the traditional self-play schemes. For the sake of simplicity, we operate under the following assumption:
+Fourth, the policies in traditional self-play algorithms don’t need supplementary information, we set the policy condition function $h(i)=\emptyset$.
+Fifth, the MSSes of traditional self-play algorithms are straightforward, often yielding a constant interaction matrix $\Sigma$ that eliminates performance evaluation.
+Only the MSS for prioritized fictitious self-play (described in Sec.
+3.2.5) requires the performance matrix; nonetheless, it does not need to solve for complex game outcomes such as NE.
+Next, we outline the traditional self-play schemes. For simplicity, we operate under the following assumption:
 Assumption 1.
@@
 Consequently, when policy $i$ is selected for training, only policy $j$ $(j\leq i)$ has already been trained and holds meaningful outcomes.
-Other policies are ineffective policies.
+Other policies are inactive policies.
 As a result, we exclusively select policy $j$, where $j\leq i$, to serve as the opponent for policy $i$.
 ∎
-III-B2 Vanilla Self-play
-In vanilla self-play [6], agents are trained by competing against their latest versions, ensuring consistent and aligned learning. The MSS of vanilla self-play:
-$\mathcal{F}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}1,\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ j=i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right..$ (33)
-No matter what $\mathcal{P}$ is, the MSS produces the same interaction matrix, similar to the iterative refinement process of the policy.
-Although this MSS is simple, it is utilized in many further works, so we refer to it as vanilla MSS.
-Although vanilla self-play is effective in transitive games, it can lead the agent to cyclic learning patterns in non-transitive games like Rock-Paper-Scissors.
-III-B3 Fictitious Self-play
+3.2.2 Vanilla Self-play
+In vanilla self-play [108], agents are trained by competing against their latest versions. Thus, the MSS of vanilla self-play:
+$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}1,\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ j=i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right..$ (5)
+Regardless of $\mathcal{P}$, the MSS produces the same interaction matrix.
+Although this MSS is straightforward, it is utilized in many other algorithms, so we refer to it as vanilla MSS.
+Vanilla self-play is effective in transitive games, but it can lead the agent to cyclic learning patterns in non-transitive games.
+It may also lead to an unstationary problem, causing the system to get stuck in a local optimum.
+A more modern version of vanilla self-play is to combine it with off-policy RL.
+In each iteration, the policy to be trained, $\pi_{i}^{h}$, is initialized using the previous policy $\pi_{i-1}(\cdot|h(i-1))$.
+Typically, each player’s policy is fixed, and samples are collected into a buffer by vanilla self-play.
+These samples remain useful for training, even if they were generated by an earlier policy.
+This allows the agent to leverage past experiences and maintain more stationary training.
+Once the policy is updated, it will be added to the policy population.
+3.2.3 Fictitious Self-play
 Fictitious Play (FP) [11] is a learning algorithm in game theory where each player best responds to the empirical frequency of the strategies used by the opponent.
@@
 In the original versions of NFSP and FSP, two distinct types of agent memory are used: one to record the agent’s own behavior and the other to capture the opponent’s behavior.
-However, in more recent approaches [42], random sampling is frequently employed to approximate the opponents’ average strategy, eliminating the need to maintain two separate types of agent memory.
+In more recent approaches [73], random sampling is frequently employed to approximate the opponents’ average strategy, eliminating the need to maintain two separate types of agent memory.
 Thus, the MSS of FSP:
-$\mathcal{F}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{i-1},\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ j\leq i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right..$ (34)
+$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{i-1},\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ j\leq i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right..$ (6)
 In FSP, the MSS continues to generate a constant interaction matrix.
 Compared to vanilla self-play, this approach enhances the robustness of the policy by sampling older versions of its own policies to be used as opponent strategies.
-III-B4 $\delta$-uniform Self-play
-$\delta$-uniform self-play is introduced by [7].
-The hyper-parameter $\delta$ ranges between 00 and $1$ used to select the most recent $\delta$ (percentage) of policies for uniform sampling to generate opponent policies.
-When policy $i$ is in the training phase, following Corollary 1, only the preceding $i-1$ policies hold significance.
-As opponents for policy $i$, we select policies from the discrete uniform distribution $[\delta(i-1),i-1]$.
-When $\delta=0$, the system retains the complete historical memory, whereas $\delta=1$ implies that only the most recent policy is utilized.
-Thus, we can get the MSS of $\delta$-uniform self-play:
-$\mathcal{F}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{f(i)},\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \delta(i-1)\leq j\leq i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right.,$ (35)
-$f(i)=\max\{\lceil(1-\delta)(i-1)\rceil,1\},$ (36)
+3.2.4 $\delta$-uniform Self-play
 […]
-III-B5 Prioritized Fictitious Self-play
+$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{f(i)},\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \lceil\delta(i-1)\rceil\leq j\leq i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right.,$ (7)
+$f(i)=i-\lceil\delta(i-1)\rceil.$ (8)
+In $\delta$-uniform self-play, The MSS generates a constant interaction matrix. Specially, if $\delta=0$, it corresponds to FSP, and if $\delta=1$, it corresponds to vanilla self-play.
+3.2.5 Prioritized Fictitious Self-play
 Prioritized Fictitious Self-Play (PFSP)  [129] leverages a priority function to allocate a higher probability of selection to agents with higher priorities.
@@
 5.
-Algorithm 5 the PFSP Meta-strategy Solver
+The function $f:[0,1]\rightarrow[0,\infty)$ is a priority function.
+For example, $f(x)=(1-x)^{p}$ with $p>0$ indicates that stronger policies against the currently being trained policy have a higher chance of being chosen as opponents.
+Alternatively, $f=x(1-x)$ implies that players of similar levels are more likely to be chosen as opponents.
+Additionally, in broader terms, $\mathcal{P}$ can also be assessed by other metrics.
+A similar MSS in  [7] can be utilized to allocate higher probabilities to strategies that perform better or are more challenging to defeat.
+Algorithm 5 the PFSP Meta-strategy Solver.
 1:  function $\mathcal{F}(\mathcal{P})$
@@
 4:  return  $\Sigma$
-The function $f:[0,1]\rightarrow[0,\infty)$ is a priority function.
-For example, $f(x)=(1-x)^{p}$ with $p>0$ indicates that stronger policies against the currently being trained policy have a higher chance of being chosen as opponents.
-Alternatively, $f=x(1-x)$ implies that players of similar levels are more likely to be chosen as opponents.
-Additionally, in broader terms, $\mathcal{P}$ can be assessed by other metrics as well.
-A similar MSS in  [14] can be utilized to allocate higher probabilities to strategies that perform better or are harder to defeat.
-III-B6 Independent RL
-Independent RL is a fully decentralized method in MARL.
-This simplifies the learning process and is useful in competitive or adversarial settings, while it may lead to suboptimal outcomes in situations requiring cooperation.
-Independent RL can be seen as a special case of self-play algorithms.
-In each iteration, the policy to be trained, $\pi_{i}^{h}$, is initialized using the previous policy $\pi_{i-1}(\cdot|h(i-1))$.
-The MSS of independent RL simplifies to an identity matrix:
-$\mathcal{F}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}1,\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ i=j\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right..$ (37)
-The difference between independent RL and vanilla self-play lies in how the opponents’ policies are handled during training.
-In vanilla self-play, the opponents’ policies are often fixed, making the training process stationary.
-In contrast, in independent RL, the opponents’ policies evolve along with the policy being trained, resulting in a non-stationary training process.
-Moreover, if off-policy RL methods are used in independent RL, the samples collected can still be useful for training, even if they are generated by a different policy.
-This allows the agent to leverage past experiences more effectively and learn from a wider range of scenarios.
-III-C PSRO Series of Algorithms
+3.3 PSRO Series
 Similar to traditional self-play algorithms, the PSRO series of algorithms starts with a single policy and gradually expands the policy space by incorporating new oracles.
 These oracles are policies that approximate optimal responses to the current meta-strategies of other agents.
-Additionally, PSRO employs EGTA to update meta-strategy distributions, thereby incorporating a level of exploration in policy selection to mitigate the risk of overfitting.
-III-C1 Integration into Our Framework
+3.3.1 Integration into Our Framework
 The PSRO series of algorithms can also be integrated into our proposed framework (Algo.
 1).
-First, similar to traditional self-play algorithms, we also utilize placeholder initialization to initialize $\Pi$.
+First, similar to traditional self-play algorithms, we also utilize lazy initialization to initialize $\Pi$.
 Second, we also set $E=1$ and $N$ can be considered as the upper limit for the policy population size in the original PSRO algorithms.
 Third, in the context of the PSRO series of algorithms, the strategy of our player $\pi_{i}^{h}$ can also be initialized in a general manner.
-Fourth, we simply set $h(i)=\emptyset$ since the PSRO series of algorithms do not use any conditioning function for their policies.
+Fourth, we set $h(i)=\emptyset$ since the PSRO series of algorithms do not use any policy condition function for their policies.
 Fifth, it’s crucial to highlight that our framework diverges from the traditional PSRO model [73] in how $\sigma_{i}$ is defined.
@@
 For example, some MSSes incorporate concepts from different types of game equilibria [92, 96, 86].
-For simplicity, we also follow Assumption 1. Similar to traditional self-play algorithms, we can derive the Corollary 2 using a similar proof as Corollary 1.
+Lastly, we also simply follow Assumption 1.
+We can derive the Corollary 2 using a similar proof as Corollary 1.
 Corollary 2.
 In the PSRO series of algorithms, the interaction matrix $\Sigma$ is a lower triangular matrix.
-III-C2 Double Oracle
-Double oracle (DO) [45] is traditionally only applied to two-player normal-form games.
+3.3.2 Double Oracle
+Double Oracle (DO) [92] can be only applied to two-player normal-form games.
 In this context, we can utilize the payoff matrix as the evaluation matrix.
@@
 In the context of two-player normal-form games, DO theoretically can achieve the NE of the full game.
-Algorithm 6 the NE-based Meta-strategy Solver
+Algorithm 6 the NE-based Meta-strategy Solver.
 1:  function $\mathcal{F}(\mathcal{P})$
@@
 4:  return  $\Sigma$
-III-C3 PSRO
+We take rock-paper-scissors as an example, a symmetric two-player game.
+The policy condition function $h(i)=\emptyset$.
+The policy population size $N=4$.
+The policy population $\Pi$ uses lazy initialization and the interaction matrix $\Sigma$ is initialized with all zeros.
+Assume that in the first iteration, the policy $\pi_{1}$ is set to "rock".
+Obviously, the NE-based MSS returns $\sigma_{2}=(1,0,0,0)$.
+In the second iteration, the BR oracle adds "paper" to $\Pi$ as $\pi_{2}$.
+Consequently, the NE-based MSS yields $\sigma_{3}=(0,1,0,0)$, since "paper" always beats "rock." In the third iteration, noting that $\sigma_{3}$ selects only "paper" and never "rock," the BR oracle adds "scissors" as $\pi_{3}$.
+The NE-based MSS then solves the game and returns $\sigma_{4}=(1/3,1/3,1/3,0)$.
+In the final iteration, the BR oracle adds a mixed strategy that assigns a uniform probability over "rock," "paper," and "scissors" as $\pi_{4}$.
+The final interaction matrix is:
+$\Sigma=\begin{bmatrix}0&0&0&0\\ 1&0&0&0\\ 0&1&0&0\\ \frac{1}{3}&\frac{1}{3}&\frac{1}{3}&0\end{bmatrix}$ (9)
+This process demonstrates how the DO algorithm iteratively expands the strategy set within our framework.
+Also, using DO as an example, we show that, compared with traditional self-play algorithms, the PSRO series employs more sophisticated MSSes to better capture the essence of the game, albeit at a higher computational cost.
+3.3.3 PSRO
 PSRO [73] extends DO to more complex games beyond just two-player normal-form games.
@@
 In the original PSRO framework, the oracle is computed using RL techniques, similar to those described in Algorithm 2.
-This allows the algorithm to effectively handle large and intricate strategy spaces, making it applicable to a wide range of game scenarios.
-III-C4 PSRO Variants
+This allows the algorithm to effectively handle large and intricate strategy spaces, making it applicable to various game scenarios.
+3.3.4 PSRO Variants
 The traditional PSRO algorithm has been the subject of numerous extensions in recent research.
 Some studies focus on making PSRO more computationally tractable in different scenarios or achieving fast convergence.
-Through the establishment of a hierarchical structure of RL workers, Pipeline PSRO [57] achieves the parallelization of PSRO and simultaneously provides guarantees of convergence.
-Efficient PSRO [58] introduces the unrestricted-restricted (URR) game to narrow down the selection of opponent policies, thereby obviating the need for meta-game simulations in traditional PSRO.
-And, similar to Pipeline PSRO, Efficient PSRO is equipped to solve URR in parallel.
-In addition, unlike traditional PSRO, which confines the integration of population strategies to the game’s initial state, XDO [59] allows this integration across all information states.
-It ensures a linear convergence to an approximate NE based on the number of information states, thereby enhancing its tractability in extensive-form games.
-ODO [60] integrates the no-regret analysis of online learning with the Double Oracle approach to improve both the convergence rate to an NE and the average payoff.
-Anytime PSRO [61] and Self-play PSRO [62] are designed to incorporate less exploitable policies into the policy population, thereby facilitating faster convergence.
+By establishing a hierarchical structure of RL workers, Pipeline PSRO [87] achieves the parallelization of PSRO and simultaneously provides guarantees of convergence.
+Efficient PSRO [151] introduces the unrestricted-restricted (URR) game to narrow the selection of opponent policies, thereby preventing the need for meta-game simulations in traditional PSRO.
+Similar to Pipeline PSRO, Efficient PSRO is equipped to solve URR in parallel.
+In addition, unlike traditional PSRO, which confines the integration of population strategies to the game’s initial state, Extensive-Form Double Oracle (XDO) [89] allows this integration across all information states.
+It ensures a linear convergence to an approximate NE based on the number of information states, enhancing its tractability in extensive-form games.
+Online Double Oracle (ODO) [35] integrates the no-regret analysis of online learning with DO to improve both the convergence rate to an NE and the average payoff.
+Anytime PSRO [90] and Self-play PSRO [88] are designed to incorporate less exploitable policies into the policy population, facilitating faster convergence.
 Moreover, as the number of agents grows, determining BRs becomes exponentially challenging.
@@
 Instead of NE, $\alpha$-rank [101], which is unique and polynomial-time solvable in multi-player general-sum games, is introduced in the MSS and a preference-based best response (PBR) oracle is incorporated in the approach.
-Similar to $\alpha$-PSRO, [47] proposes JPSRO to tackle multi-player general-sum games.
-It utilizes concepts of CE and coarse correlated equilibrium (CCE) as alternatives to NE to make it computationally tractable in multi-player general-sum games.
-It also puts forward the CE and CCE-based MSS.
-Expanding on JPSRO, NeuPL-JPSRO [67] takes advantage of transfer learning.
-In addition to a shared parameter $\theta$, each strategy in NeuPL-JPSRO is characterized by a strategy embedding vector $v_{i}$.
-This approach avoids the need for training from scratch in each iteration, thereby enhancing efficiency.
-Other research focuses on policy diversification, given that deriving a singular policy in highly transitive games often lacks significance.
+Similar to $\alpha$-PSRO, [86] proposes Joint Policy-Space Response Oracle (JPSRO) to tackle multi-player general-sum games.
+It utilizes CE and coarse correlated equilibrium (CCE) as alternatives to NE to make it computationally tractable in multi-player general-sum games and designs the CE and CCE-based MSS.
+Other research focuses on policy diversification, as deriving a single policy in transitive games is less meaningful.
 [4] introduces an open-ended framework in two-player zero-sum games.
@@
 [103] uses determinantal point process [69] to assess diversity and introduces diverse PSRO by incorporating a diversity term into the PSRO oracle.
-This modification can also be easily implemented in FP and $\alpha$-PSRO as well.
-Similarly,  [50] formulates behavioral diversity and response diversity and incorporates them into the PSRO oracle.
+This modification can also be easily implemented in FP and $\alpha$-PSRO.
+Similarly,  [83] introduces the concepts of behavioral diversity and response diversity and incorporates them into the PSRO oracle.
 Policy Space Diversity PSRO [142] defines a diversity metric named population exploitability that helps to achieve a full-game NE.
-III-C5 R-NaD
+3.3.5 R-NaD
 Although R-NaD [105] is initially described as leveraging evolutionary game theory with a regularization component.
-Here, we categorized it into the PSRO series with a special oracle computation technique (Algo.
-3), which is executed in two stages: the first stage involves transforming the reward based on the regulation policy to make it policy-dependent, and the second stage applies replicator dynamics [32] until convergence to a fixed point.
-Crucially, in each iteration, the oracle added to the policy population $\Pi$ is derived from the reward-transformed game, not the original problem’s oracle.
-Nonetheless, this approach ensures that the policy will converge to the NE of the original game, provided that the game is monotone.
+Here, we categorized it into the PSRO series with a unique oracle computation technique (Algo.
+3), executed in two stages: the first stage involves transforming the reward based on the regularization policy to make it policy-dependent, and the second stage applies replicator dynamics [124] until convergence to a fixed point.
+The oracle added to the policy population $\Pi$ in each iteration is derived from the reward-transformed game, not the original problem’s oracle.
+Nonetheless, this approach ensures that the policy converges to the NE of the original game when the game is monotone.
 The MSS of R-NaD is the vanilla MSS, as described in Equ.
-(33), the same as the MSS of vanilla self-play.
+(5).
 This equation illustrates that the fixed point reached in each iteration, the oracle, is utilized as the regularization policy for the next iteration.
-III-D Ongoing-training-based Series of Algorithms
+3.4 Ongoing-training-based Series
 In the PSRO series of algorithms, two key challenges arise.
-First, when operating with a limited budget, it is often necessary to truncate the ABR operators during each iteration.
-This can introduce sub-optimally trained responses into the population.
-Second, the redundant process of relearning basic skills in every iteration is not only inefficient, but it also becomes untenable when confronted with increasingly formidable opponents [43].
-To address these challenges, the ongoing-training-based series of algorithms promote the ongoing training of all policies repeatedly.
-Namely, all effective policies are likely to be selected for training.
-III-D1 Integration into Our Framework
+First, truncating the ABR operators during each iteration is often necessary when operating with a limited budget, introducing sub-optimally trained responses into the population.
+Second, relearning basic skills in every iteration is redundant and becomes untenable when confronted with increasingly formidable opponents [82].
+To address these challenges, the ongoing-training-based series of algorithms emphasizes the repeated ongoing training of all policies.
+In other words, instead of only inactive policies being selected for training, all active policies are also likely to be chosen for training.
+3.4.1 Integration into Our Framework
 We can incorporate these ongoing-training-based series of algorithms into our proposed framework (Algo.
-1) using the following settings: First, we use actual initialization to initialize $\Pi$ because in the ongoing-training-based series, all policies in the policy population are trained together, rather than the policy population growing with each iteration.
+1) using the following settings: First, we use immediate initialization to initialize $\Pi$ because, in the ongoing-training-based series, all policies in the policy population are trained together, rather than the policy population growing with each iteration.
 Second, we set $E=E_{max}>1$, which represents the maximum number of epochs to optimize each policy within the policy population.
-In other words, each unique policy undergoes iterative training for a total of $E_{max}$ times.
+In other words, each unique policy undergoes iterative training for $E_{max}$ times.
 Third, since each policy undergoes training for $E_{max}$ times, we utilize $\pi_{i}(\cdot|h(i))$ to initialize $\pi_{i}^{h}$.
 This means that policy updates are self-referential.
-For the sake of simplicity, we also adopt Assumption 1. Different from Collory 1 and Collory 2, due to the continuous training process of all policies, we derive Collory 3.
+Lastly, under Assumption 1, different from Collory 1 and 2, due to the continuous training process of all policies, we derive Collory 3.
 Corollary 3.
@@
 Proof.
-When policy $i$ is selected for training, policy $k$ $(k\geq i)$ has already been actually initialized and is therefore considered an effective policy.
+When policy $i$ is selected for training, policy $k$ $(k\geq i)$ has already been actually initialized and is therefore considered an active policy.
 Furthermore, in epoch $e$ $(e>1)$, policy $k$ has already been updated and holds significant meaning.
@@
 ∎
-III-D2 FTW
-Quake III Arena Capture the Flag is a renowned 3D multiplayer first-person video game where two teams vie to capture as many flags as possible.
+3.4.2 FTW
+Quake III Arena Capture the Flag is a renowned 3D multi-player first-person video game where two teams vie to capture as many flags as possible.
 The For The Win (FTW) agent [60] is designed to perform at human-level proficiency in this game.
 A pivotal aspect of FTW is its employment of the ongoing-training-based self-play in RL.
-Specifically, it trains a population of $N$ different policies in parallel which compete and collaborate with each other.
+Specifically, it trains a population of $N$ different policies in parallel, which compete and collaborate with each other.
 When policy $i$ is undergoing training, FTW samples both its teammates and adversaries from the population.
 Specially, in scenarios where each team comprises a single member, it can be seamlessly integrated into our framework using the subsequent MSS:
-$\mathcal{F}(\mathcal{P})_{ij}=\frac{1}{N}.$ (38)
+$\textit{MSS}(\mathcal{P})_{ij}=\frac{1}{N}.$ (10)
 This essentially means that the interaction graph is densely connected.
@@
 Hence, $\pi_{i}(\cdot|h(i))$ can be aptly depicted as $\pi_{\phi}(\cdot|h(i))$.
-Furthermore, since these policies are not conditioned on any external parameters, it is straightforward to represent the conditioning function $h(i)=\emptyset$.
-III-D3 NeuPL
-NeuPL [43] introduces another key innovation: it employs a unified conditional network, where each policy is adjusted against a specific meta-game mixture strategy.
+Furthermore, since these policies are not conditioned on external parameters, it is straightforward to represent the conditioning function $h(i)=\emptyset$.
+3.4.3 NeuPL
+Neural Population Learning (NeuPL) [82] introduces another critical innovation: it employs a unified conditional network, where each policy is adjusted against a specific meta-game mixture strategy.
 This is instrumental in enabling transfer learning across policies.
@@
 Moreover, given that the policies in NeuPL are contingent on the opponent sampling strategy $\sigma_{i}$, it is apt to define $h(i)=\sigma_{i}$.
-III-D4 Simplex-NeuPL
+3.4.4 Simplex-NeuPL
 Simplex-NeuPL [80], which builds on NeuPL, is designed to achieve any-mixture optimality, which signifies that the formulated policy should exhibit flexibility across a diverse spectrum of adversaries, including those that might not present equivalent competitive prowess.
@@
 This sampling mechanism results in greater robustness.
-III-E Regret-minimization-based Series of Algorithms
+3.5 Regret-minimization-based Series
 Another line of self-play algorithms is based on regret minimization.
-The key distinction between regret-minimization-based algorithms and the other categories is that they prioritize accumulated payoff over time, rather than focusing solely on a single episode.
-This approach leads to more aggressive and adaptable strategies, which are essential to avoid being exploited by opponents as time progresses.
-Additionally, these algorithms require players to deduce and adapt to opponents’ strategies over several rounds.
-This scenario is commonly observed in repeated games rather than stage games.
+Unlike other categories, this series prioritizes accumulated payoff over time rather than focusing solely on a single episode, which is common in repeated games.
 For example, in games like Texas Hold’em or Werewolf, players must use deception, concealment, and bluffing to aim for overall victories rather than just winning a single game.
-It’s important to note that while traditional regret-minimization-based self-play doesn’t typically use RL, many subsequent research efforts have combined regret minimization with RL to achieve strong performance.
-In this section, we will also thoroughly discuss traditional regret-minimization-based methods to provide a foundation for understanding how integrating regret minimization with RL can lead to enhanced performance.
-III-E1 Integration into Our Framework
+Traditional regret-minimization-based self-play typically doesn’t use RL, but later studies have combined them for improved performance.
+This section will also discuss traditional regret-minimization methods to lay the groundwork for understanding their integration with RL.
+3.5.1 Integration into Our Framework
 We can also integrate the regret-minimization-based series of algorithms into our proposed framework (Algo.
-1) with the following settings: First, similar to traditional self-play algorithms and the PSRO series, we use placeholder initialization to initialize the policy population $\Pi$.
+1) with the following settings: First, similar to traditional self-play algorithms and the PSRO series, we use lazy initialization to initialize the policy population $\Pi$.
 Second, we set $E=1$, and $N$ is regarded as the maximum iteration to optimize the policy.
@@
 Fourth, in each iteration $i$, $h(i)$ represents the specific elements that regret-minimization-based self-play algorithms need to store.
-It’s important to note that the regret-minimization-based series relies heavily on the information contained in $h(i)$.
-For instance, in vanilla Counterfactual Regret Minimization (CFR) [72], it is necessary to store counterfactual regrets for every action in every information set for each player within $h(i)$ and once $h(i)$ is determined, the corresponding policy is also defined through regret matching.
-We will discuss vanilla CFR in detail in Sec.
-III-E2.
-Fifth, the ABR operator is described in Algo.
-4, with the key point being to incorporate new regret minimization-based information into $h(i)$.
-Notably, while the original CFR updates regrets for all players simultaneously, this oracle (Algo.
-4) updates the regrets of each player sequentially.
+Note that this category relies heavily on the information in $h(i)$.
+For instance, in Counterfactual Regret Minimization (CFR) [152], it is necessary to store counterfactual regrets for every action in every information set for each player within $h(i)$.
+Once $h(i)$ is determined, the corresponding policy is also defined.
+We will discuss CFR in detail in Sec.
+3.5.2.
+Fifth, the ABR operator (Algo.
+4) incorporates regret-related information into $h(i)$.
+Unlike the original CFR, Algo.
+4 updates the regrets of each player sequentially.
 This means that the regrets of player 2 are updated after considering the already-updated regrets of player 1.
-Namely, $h(i)$ is changing during the iteration $i$.
-This adjustment has not only been shown to empirically accelerate convergence but also possesses a theoretical error bound [73].
+This adjustment has been shown to accelerate convergence empirically and possesses a theoretical error bound [22].
 Additionally, each $\pi_{i}^{h}$ represents the strategies for all players, and in iteration $i$, player $j$ uses the strategy $\pi_{i}^{h}(j)$.
-Moreover, the MSS of the regret-minimization-based series is the vanilla MSS, as described in Equ. (33). We can derive Collory 4.
+Lastly, the MSS of the regret-minimization-based series is the vanilla MSS, as described in Equ.
+(5).
+Also, we can derive Collory 4.
 Corollary 4.
@@
 Regret minimization-based algorithms always use the latest strategy for training.
-In other words, in iteration $i$, $\pi_{i-1}^{h}$ is always chosen as the opponent policy.
+In other words, at iteration $i$, $\pi_{i-1}^{h}$ is consistently chosen as the opponent policy.
 Consequently, the interaction matrix $\Sigma$ is a unit lower shift matrix.
 ∎
-III-E2 Vanilla CFR
-Regret measures the difference between the actual payoff and the best possible payoff that could have been achieved with a different strategy.
-The regret matching algorithm [74] in game theory optimizes decisions in iterative games by learning from past outcomes.
-More specifically, it involves selecting strategies based on accumulated positive overall regrets.
-Strategies with higher overall regret are generally more likely to be chosen, as players seek to correct past underperformance.
+3.5.2 Vanilla CFR
+Regret measures the difference between the best possible payoff and the actual payoff.
+The regret matching algorithm [49] optimizes decisions by selecting strategies based on accumulated positive overall regrets, with higher overall regret strategies being more likely chosen to correct past underperformance.
 After each round, players update the overall regret values for each strategy.
-When each player aims to reduce their average overall regret, their average strategies will converge towards an approximation of the NE over time.
-However, this traditional regret minimization algorithm applies primarily to normal-form games because computing overall regret in extensive-form games is challenging.
-Although extensive-form games can theoretically be converted into normal-form games, this conversion results in an exponential increase in states, rendering it impractical for complex extensive-form games.
-[72] proposes counterfactual regret (CFR) minimization for extensive-form games with incomplete information.
-Here, it is referred to as vanilla CFR to distinguish it from subsequent advancements in the field.
-Vanilla CFR involves maintaining the strategy and the counterfactual regrets for each information set.
-Theoretically, the immediate counterfactual regret is specific to each information set, and the aggregation of these immediate counterfactual regrets can act as an upper bound for the overall regret.
-Thus, the problem can be simplified from minimizing the overall regret to minimizing the immediate counterfactual regrets in each information set.
-This simplification significantly reduces the computational complexity.
+Strategies will converge to an approximate NE.
+However, this method primarily applies to normal-form games because computing overall regret in extensive-form games is challenging.
+[152] proposes CFR for extensive-form games.
+Here, it is referred to as vanilla CFR to distinguish it from later advancements.
+Vanilla CFR maintains the strategy and the immediate counterfactual regrets for each information set.
+Theoretically, this regret is specific to each information set, and the aggregation of these immediate counterfactual regrets is an upper bound for the overall regret.
+Therefore, the problem can be decomposed into numerous more minor regret minimization problems.
 Next, there is a more detailed description of vanilla CFR.
-In this survey, $\pi_{i}$ is used to denote the strategy at iteration $i$, whereas the original paper uses $\sigma$.
-This change is implemented to maintain consistency throughout the discussion in this survey.
+In this survey, $\pi_{i}$ denotes the strategy at iteration $i$ for consistency, whereas the original paper uses $\sigma$.
 Furthermore, we use $j$ to represent player $j$, and $\pi_{i}(j)$ denotes the policy used by player $j$ at iteration $i$.
@@
 $R_{j}^{T}\leq\sum_{I\in\mathcal{I}_{j}}R_{j,\text{imm}}^{T,+}(I).$ (15)
-Therefore, based on the theory discussed, the overall problem of regret minimization can be decomposed into numerous smaller regret minimization problems.
-This approach makes the problem manageable for extensive-form games of not excessively large sizes.
-In practical applications, the focus is primarily on immediate counterfactual regrets.
-For simplicity, we often drop "immediate" in discussions, thereby referring directly to counterfactual regrets.
-Consequently, the counterfactual regret associated with every action $a$ within every information set $I$ is recorded as follows:
+Thus, the problem can be simplified from minimizing the overall regret to minimizing the immediate counterfactual regrets in each information set separately.
+For simplicity, we often drop "immediate" in discussions and refer directly to counterfactual regrets.
+The counterfactual regret can also be recorded as follows:
 $R_{j}^{T}(I,a)=\frac{1}{T}\sum_{i=1}^{T}\left(v_{j}(\pi_{i}|_{I\to a},I)-v_{j}(\pi_{i},I)\right),$ (16)
@@
 Vanilla CFR has several shortcomings.
-Firstly, it requires traversing the entire game tree in each iteration, which becomes computationally intractable for larger game trees.
+Firstly, it requires traversing the entire game tree in each iteration, computationally intractable for larger game trees.
 Although some efforts have focused on game abstraction to reduce the size of the game tree, greater abstraction can lead to decreased performance.
@@
 These values are stored in $h(i)$ within our proposed framework (Algo.
-1).
-This requirement leads to significant storage challenges.
-III-E3 Time-saving Variants of CFR
-Therefore, numerous studies focus on enhancing the time efficiency of CFR primarily through two main approaches.
+1), leading to significant storage challenges.
+3.5.3 Time-saving Variants of CFR
+Many studies focus on enhancing the time efficiency of CFR.
 The first approach involves modifying the regret calculation to increase its speed.
-For example, CFR+ [75] implements regret-matching+ by storing non-negative regret-like values $R_{j}^{+,i}(I,a)$, rather than $R_{j}^{i}(I,a)$.
-Note that $R_{j}^{+,i}(I,a)$ is computed slightly differently from $R_{j}^{i,+}(I,a)$ in Equ.
-(45).
-Additionally, CFR+ updates the regrets of each player sequentially.
-Furthermore, CFR+ adopts a weighted average strategy instead of a uniform average strategy, assigning a lower weight to the initial strategy for accelerated convergence.
-Beyond employing a weighted average strategy, [76] introduces the concept of weighted regrets and developed Discounted CFR (DCFR).
-Linear CFR (LCFR) is a specific case of DCFR, where LCFR assigns a linear weight of $i$ to regrets at iteration $i$.
-The other approach involves adopting sampling methods.
-Although this approach requires more iterations, the duration of each iteration is reduced, ultimately decreasing the overall time to convergence.
-Monte Carlo CFR (MCCFR) [77] outlines a framework to incorporate sampling into CFR.
-This method divides terminal histories into blocks, with each iteration sampling from these blocks instead of traversing the entire game tree.
-This allows for the calculation of sampled counterfactual values for each player, leading to immediate counterfactual regrets that, in expectation, match those of the Vanilla CFR.
-Vanilla CFR represents a specific case where all histories are divided into just one block.
-MCCFR typically manifests in three forms: outcome-sampling MCCFR, where each block corresponds to a single history; external-sampling MCCFR, which samples opponent and chance nodes; and chance-sampling MCCFR, focusing on chance nodes alone.
-Moreover, [78] expands on chance-sampling into naive chance sampling, opponent/public chance sampling, self/public chance sampling, and public chance sampling.
-These methods differ in how they handle the sampling of public and private chance nodes.
+CFR+ [122] implements regret-matching+ by storing non-negative regret-like values $R_{j}^{+,i}(I,a)$, rather than $R_{j}^{i}(I,a)$.
+Additionally, CFR+ updates the regrets of each player sequentially and adopts a weighted average strategy.
+Moreover, [19] introduces the concept of weighted regrets and develops Linear CFR (LCFR) and Discounted CFR (DCFR).
+The second approach involves adopting sampling methods.
+While it requires more iterations, each iteration is shorter, reducing the overall convergence time.
+Monte Carlo CFR (MCCFR) [72] divides terminal histories into blocks, with each iteration sampling from these blocks instead of traversing the entire game tree.
+This allows for calculating sampled counterfactual values for each player, leading to counterfactual regrets that, in expectation, match those of the Vanilla CFR.
+Vanilla CFR is a specific case where all histories are divided into just one block.
+MCCFR typically manifests in three forms: outcome-sampling MCCFR, where each block corresponds to a single history; external-sampling MCCFR, which samples opponent and chance nodes; and chance-sampling MCCFR, only focusing on chance nodes.
+Moreover, [64] expands on chance-sampling by categorizing based on the handling of public and private chance nodes.
 Some studies also focus on learning how to reduce the variance of MCCFR to speed up convergence [111, 33].
-In addition to these two primary approaches, other studies have identified that warm starting [81] and pruning techniques [82, 83, 84] can accelerate CFR convergence as well.
-III-E4 Space-saving Variants of CFR
-Storing the strategies and regrets of each information set at each iteration requires substantial storage capacity.
-In games with perfect information, decomposition is employed to reduce the scale of problem-solving; that is, we solve subgames instead of the entire game.
-However, this approach faces challenges in imperfect information games.
-The difficulty arises in defining subgames, as they may intersect with the boundaries of information sets, complicating their delineation.
-CFR-D [85] is a pioneering method for decomposing imperfect information games with a theoretical guarantee.
-The game is divided into a main component, called the trunk, and several subgames.
-It posits that in an imperfect information game, a subgame can be conceptualized as a forest of trees that does not divide any information sets.
-In each iteration, CFR is applied within the trunk for both players, accompanied by the use of a solver to determine the counterfactual BR in the subgame.
-The process includes updating the trunk with the counterfactual values from the subgame’s root and updating the average counterfactual values at this root.
-Following these updates, the solution to the subgame is discarded.
-CFR-D minimizes storage needs by only saving values $R^{i}(I_{*},a)$ for information sets located in the trunk and at each subgame’s root, which is denoted by $I_{*}$ at each iteration $i$, trading off storage efficiency against the time required to resolve subgames.
+In addition to these two primary approaches, other studies have identified that warm starting [16] and pruning [15, 13, 17] can also accelerate convergence.
+3.5.4 Space-saving Variants of CFR
+In perfect information games, decomposition reduces problem-solving scale by solving subgames.
+However, in imperfect information games, defining subgames is challenging due to their intersection with information set boundaries.
+CFR-D [21] is a pioneering method for decomposing imperfect information games into the main component called the trunk and subgames defined by forests of trees that do not divide any information sets.
+In each iteration, CFR is applied within the trunk for both players, and a solver is used to determine the counterfactual BR in the subgame.
+The process includes updating the trunk with the counterfactual values from the subgame’s root and updating the average counterfactual values at this root, while the solution to the subgame is discarded.
+CFR-D minimizes storage needs by only saving $R^{i}(I_{*},a)$ for information sets in the trunk and at each subgame’s root, which is denoted by $I_{*}$, trading off storage efficiency against the time required to resolve subgames.
 Similar thoughts are echoed in the Continue-Resolving technique used by DeepStack [95] and the Safe and Nested Subgame Solving technique used by Libratus [18].
 We will discuss these approaches in Sec.
-IV-B1, where we explore their application to the specific context of Texas Hold’em.
-III-E5 Estimation Variants of CFR
-Although the above variants of CFR advance the field, they still fail to solve large imperfect-information extensive-form games directly.
-This limitation largely stems from the reliance on tabular representations for strategies.
-Typically, the process involves abstracting the original large game into a simpler form, applying the above CFR variants to this abstracted game, and then translating the developed strategies back to the original game.
-However, this abstraction process is highly game-specific and is heavily based on domain knowledge.
-Moreover, abstracting the game to a smaller scale often leads to suboptimal results compared to a more comprehensive abstraction, making the choice of abstraction scale critical for performance.
-Given these challenges, there is a shift towards estimation techniques.
-[87] introduces Regression CFR (RCFR), which employs a shared regressor $\varphi(I,a)$ to estimate counterfactual regrets.
-Nevertheless, the use of regression trees as the regressor limits RCFR’s applicability to smaller games, and the necessity for manually crafted features remains a drawback.
+4.2.1, exploring their application to Texas Hold’em.
+3.5.5 Estimation Variants of CFR
+Although these CFR variants advance the field, they can’t directly solve large imperfect-information extensive-form games due to their reliance on tabular representations.
+The typical approach involves abstracting the original game, applying CFR to the abstracted game, and translating strategies back to the original.
+This abstraction is game-specific and relies heavily on domain knowledge.
+Additionally, smaller abstractions often yield suboptimal results.
+Given these challenges, [133] introduces Regression CFR (RCFR), which employs a shared regressor $\varphi(I,a)$ to estimate counterfactual regrets.
+Nevertheless, using regression trees as the regressor limits RCFR’s applicability to miniature games, and the necessity for manually crafted features remains a drawback.
 After advantage-based regret minimization (ARM) [62] merges CFR with deep RL in single-agent scenarios, a growing body of research has focused on applying CFR in conjunction with neural networks to multi-agent scenarios.
@@
 Since these two networks are trained in sequence rather than concurrently, the strategy for each intermediate iteration remains conditioned on the output of the advantage network: $h(i)=V(I,a|\theta_{p})$.
-Despite similarities, Deep CFR distinguishes itself from Double Neural CFR through its data collection and its proven effectiveness in larger-scale poker games.
-Moreover, Single Deep CFR (SD-CFR) [91] demonstrates that training an average strategy network is unnecessary, with only an advantage network required.
+Despite similarities, Deep CFR distinguishes itself from Double Neural CFR through its data collection and proven effectiveness in larger-scale poker games.
+Single Deep CFR (SD-CFR) [118] demonstrates training an average strategy network is unnecessary, with only an advantage network required.
 Building on the foundation of SD-CFR, DREAM [119] utilizes a learned baseline to maintain low variance in a model-free setting when only one action is sampled at each decision point.
 Moreover, advantage regret-matching actor-critic (ARMAC) [46] combines the thought of retrospective policy improvement with CFR.
-III-F Reassessment of the Framework
-After introducing these four categories of self-play algorithms, we will further compare them in this section and illustrate the differences between our framework and other frameworks to demonstrate why our proposed framework is more general.
-We also summarize the key points in Table II.
+Table 1: Overview of Representative Self-play Algorithms within Our Framework.
+Algorithms MSS h(i) Categories $E$ Initialization of $\Pi$ Initialization of $\pi_{i}^{h}$ Suitable Game
+Vanilla Self-play [108] Equ. (5) $\emptyset$ Traditional Self-play $E=1$ Lazy General Stage Game
+Fictitious Self-play [50] Equ. (6)
+$\delta$-Uniform Self-play [5] Equ. (7)
+PFSP [129] Algo. 5
+DO [92] NE-based (Algo. 6) $\emptyset$ PSRO Series $E=1$ Lazy General Stage Game
+PSRO [73] General
+$\alpha$-PSRO [96] $\alpha$-rank-based
+JPSRO [86] (C)CE-based
+R-NaD [105] Equ. (5)
+FTW [60] Equ. (10) $\emptyset$ Ongoing-training-based $E>1$ Immediate $\pi_{i}^{h}\leftarrow\pi_{i}(\cdot|h(i))$ Stage Game
+NeuPL [82] General $\sigma_{i}$
+Simplex-NeuPL [80] General $\sigma_{i}$
+Vanilla CFR [152] Equ. (5) $R^{i}(I,a)$ Regret-minimization-based $E=1$ Lazy $\pi_{i}^{h}\leftarrow\pi_{i-1}(\cdot|h(i-1))$ Repeated Game
+CFR+ [122] $R^{+,i}(I,a)$
+CFR-D [21] $R^{i}(I_{*},a)$
+RCFR[133] $\varphi(I,a)$
+Deep CFR[133] $V(I,a|\theta_{p})$
+3.6 Reassessment of the Framework
+After introducing these four categories of self-play algorithms, we will further compare them in this section, explain the rationale behind categorizing the algorithms, and summarize the representative algorithms in Table 1.
+Moreover, we will illustrate the differences between our framework and other frameworks to demonstrate why our proposed framework is more general.
+3.6.1 Categorization Rationale
 Traditional self-play algorithms and the PSRO series share many similarities.
-Initially, they require only one randomly initiated policy, and as training progresses, the policy population expands.
-Therefore, in our framework, we use placeholder initialization to initialize the policy population and set $E=1$ for these two categories.
+Initially, they require only one randomly initiated policy, and the policy population expands as training progresses.
+Therefore, in our framework, we use lazy initialization to initialize the policy population and set $E=1$ for these two categories.
 The interaction matrix is typically a lower triangular matrix (Corollary 1 and Corollary 2).
-The primary difference between the PSRO series and traditional self-play algorithms is that the PSRO series employs more complex MSSes, designed to handle more intricate tasks.
-For example, $\alpha$-PSRO [46] specifically utilizes an $\alpha$-rank-based MSS to address multi-player general-sum games.
-Unlike the two previously mentioned categories, the ongoing-training-based series employs a different paradigm by training the entire policy population together.
-This approach aims to strengthen all policies simultaneously at each epoch, rather than expanding the policy population and relying on newer policies to be stronger.
-As a result, actual initialization is used for the policy population, and $\pi_{i}(\cdot|h(i))$ is utilized to initialize $\pi_{i}^{h}$ to ensure that policy updates are self-referential.
-Consequently, the interaction matrix is generally not a lower triangular matrix (Corollary 3).
-Lastly, the regret-minimization-based series focuses on the overall performance over time rather than on a single episode.
-For example, it is well-suited for games like Texas Hold’em, which require strategies involving deception, concealment, and bluffing.
-The main goal of the training process in this series is to update the regrets associated with different strategies.
-In our framework, we use $h(i)$ to store this information.
+The primary difference between the PSRO series and traditional self-play algorithms is that the PSRO series employs more complex MSSes to handle tasks with intricate game-theoretical requirements.
+For example, $\alpha$-PSRO [96] specifically utilizes an $\alpha$-rank-based MSS to tackle multi-player general-sum games.
+In other tasks, traditional self-play algorithms are more commonly used to reduce the computational cost associated with complex MSSes in the PSRO series.
+Unlike the two previously mentioned categories, the ongoing-training-based series adopts a different paradigm.
+Instead of gradually expanding the policy population and relying on newer policies to be stronger, this approach strengthens all policies simultaneously at each epoch.
+This method alleviates issues such as early truncation and repeated skill relearning that occur in the above two categories.
+To integrate this category into our framework, immediate initialization is used for the policy population, and $\pi_{i}(\cdot|h(i))$ is utilized to initialize $\pi_{i}^{h}$ to ensure that policy updates are self-referential.
+Also, the interaction matrix is generally not a lower triangular matrix (Corollary 3).
+Lastly, the regret-minimization-based series focuses on overall performance over time rather than on individual episodes, making it particularly suitable for repeated games.
+For example, Texas Hold’em is a classic repeated game where players adjust their strategies based on past interactions and use tactics involving deception and bluffing.
+The main goal of this training process is to update the regrets associated with different strategies, which in turn demands significant storage resources.
+Our framework uses $h(i)$ to store this information.
 Since the policies are determined by $h(i)$, only the most recent policy is relevant.
@@
 We also do not need actually to initialize the whole policy population and only need to use $\pi_{i-1}(\cdot|h(i-1))$ to initialize $\pi_{i}^{h}$ in the training process.
+3.6.2 Comparing with Existing Frameworks
 Our framework is built upon PSRO [73] and NeuPL [82].
-Here, we outline the differences between our framework and these two existing frameworks.
+Here, we outline the differences between our framework and these existing frameworks.
 The primary distinction between our framework and PSRO is the use of an interaction matrix $\Sigma:=\{\sigma_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{1}}$ to represent the opponent sampling strategy, allowing for the integration of more complex competitive dynamics.
-Moreover, in our framework, $\sigma_{i}$ denotes the opponent sampling strategy, which specifies how to sample opponents’ policies against policy $i$, rather than being the meta-strategy of policy $i$.
+Moreover, in our framework, $\sigma_{i}$ denotes the opponent sampling strategy, which specifies how to sample opponents’ policies against policy $i$ rather than being the meta-strategy of policy $i$.
 Additionally, our framework incorporates a policy condition function $h(i)$, making it more general than NeuPL, where policies are conditioned on $\sigma_{i}$.
 This enhancement gives our framework greater expressiveness.
-Furthermore, we describe how to compute the oracle (Line 4 in Alg.
+Furthermore, we describe how to compute the oracle (Line 5 in Alg.
 1) in three different ways (Alg.
@@
 Also, to the best of our knowledge, our framework is the first self-play framework to integrate the regret-minimization-based series, which is a significant self-play paradigm.
-Table II: Overview of Main Self-play Algorithms within Our Framework.
-Algorithms MSS h(i) Categories $E$ Initialization of $\Pi$ Initialization of $\pi_{i}^{h}$
-Vanilla Self-play [6] Equ. (33) $\emptyset$ Traditional Self-play $E=1$ Placeholder General
-Fictitious Self-play [13] Equ. (34)
-$\delta$-Uniform Self-play [7] Equ. (35)
-PFSP [15] Algo. 5
-Independent RL Equ. (37)
-DO [45] NE-based (Algo. 6) $\emptyset$ PSRO Series $E=1$ Placeholder General
-PSRO [42] General
-$\alpha$-PSRO [46] $\alpha$-rank-based
-JPSRO [47] (C)CE-based
-R-NaD [54] Equ. (33)
-FTW [70] Equ. (38) $\emptyset$ Ongoing-training-based $E>1$ Actual $\pi_{i}^{h}\leftarrow\pi_{i}(\cdot|h(i))$
-NeuPL [43] General $\sigma_{i}$
-Simplex-NeuPL [71] General $\sigma_{i}$
-Vanilla CFR [72] Equ. (33) $R^{i}(I,a)$ Regret-minimization-based $E=1$ Placeholder $\pi_{i}^{h}\leftarrow\pi_{i-1}(\cdot|h(i-1))$
-CFR+ [75] $R^{+,i}(I,a)$
-CFR-D [85] $R^{i}(I_{*},a)$
-RCFR[87] $\varphi(I,a)$
-Deep CFR[87] $V(I,a|\theta_{p})$
-IV Empirical Analysis
-In this section, we introduce iconic applications of self-play by categorizing the scenarios into three distinct groups: board games, which typically involve perfect information; card games and Mahjong, which usually involve imperfect information; and video games, which feature real-time actions rather than turn-based play.
-We then illustrate how self-play is applied in each of these complex scenarios and provide a comparative analysis of these applications in Table III.
-IV-A Board Games
-The landscape of board games, the majority of which are perfect information games, was previously revolutionized by the introduction of two key techniques: position evaluation and Monte Carlo tree search (MCTS) [94, 95].
+4 Empirical Analysis
+In this section, we introduce iconic applications of self-play by categorizing the scenarios into three distinct groups: board games, which typically involve perfect information; card games, which usually involve imperfect information; and video games, which feature real-time actions rather than turn-based play.
+We then illustrate how self-play is applied in these complex scenarios and provide a comparative analysis of these applications in Table 2.
+4.1 Board Games
+Board games, the majority of which are perfect information games, were previously revolutionized by the introduction of two essential techniques: position evaluation and Monte Carlo tree search (MCTS) [29, 67].
 These methodologies, with minor modifications, demonstrated superhuman effectiveness in solving board games such as chess [24], checkers [110], othello [23], backgammon [126], and Scrabble [114].
@@
 In light of this, our discussion will specifically focus on the game of Go to illustrate the application of self-play.
-In addition to Go, we will broaden our exploration to include Stratego, a board game characterized by imperfect information, contrasting the majority of board games that are based on perfect information.
-IV-A1 Go
-The paradigm of Go is revolutionized with the launch of DeepMind’s AlphaGo series [8, 9, 10, 11], which leveraged the power of self-play to significantly elevate performance, setting a new benchmark in the field of Go.
+In addition to Go, we will explore Stratego, a board game with imperfect information, in contrast to most board games that only involve perfect information.
+4.1.1 Go
+Go is an ancient board game, which is played on a grid of 19x19 lines, where two players alternately place black and white stones aiming to control the largest territory.
+The paradigm of Go artificial intelligence (AI) is revolutionized with the launch of DeepMind’s AlphaGo series [115, 117, 116, 112], which leveraged the power of self-play to significantly elevate performance.
 In AlphaGo [115], the training regime can be split into three stages.
-In the first stage, supervised learning with expert data trains a fast policy network $p_{\pi}(a|s)$ for rollouts in MCTS expansion and a precise policy network $p_{\sigma}(a|s)$.
-The second stage employs RL to refine the policy network $p_{\rho}(a|s)$ based on $p_{\sigma}(a|s)$ and subsequently trains a value network $v_{\theta}(s)$.
-Self-play is instrumental here.
-More specifically, $p_{\rho}(a|s)$ is refined by competing against a randomly chosen historical version $p_{\rho^{-}}(a|s)$, akin to the MSS shown in Equ.
-(34).
-Afterwards, $v_{\theta}(s)$ is trained using the samples from the games in which the trained policy network $p_{\rho}(a|s)$ competes against itself.
-In the final stage, MCTS integrates the policy and value networks to inform action selection.
-Unlike AlphaGo, AlphaGo Zero [9] does not require any expert data except game rules and instead learns through self-play.
-It utilizes only one network $f_{\theta}(s)$ to concurrently predict the action probabilities $p=\text{Pr}(a|s)$ and the state value $v$.
-Each move is generated by MCTS with guidance from $f_{\theta}(s)$ to aid MCTS expansion, as opposed to the rollouts used in AlphaGo.
+In the first stage, supervised learning with expert data trains a fast policy network $p_{\pi}(a|s)$ for rollouts in MCTS and a precise policy network $p_{\sigma}(a|s)$.
+The second stage employs self-play combined with RL to get a refined policy network $p_{\rho}(a|s)$ based on $p_{\sigma}(a|s)$ and subsequently trains a value network $v_{\theta}(s)$.
+Specifically, $p_{\rho}(a|s)$ is refined by competing against a randomly selected historical version $p_{\rho^{-}}(a|s)$, similar to the MSS shown in Equ.
+(6), while $v_{\theta}(s)$ is trained using the game samples collected by self-play of $p_{\rho}(a|s)$.
+In the third stage, MCTS integrates the policy and value networks to select actions.
+Based on AlphaGo, AlphaGo Zero [117] does not require any expert data except game rules.
+It utilizes only one network $f_{\theta}(s)$ to concurrently predict the action probabilities and the state value.
 Self-play is employed to generate data and refine $f_{\theta}(s)$ with the current best policy competing against itself, a process analogous to the MSS referenced in Equ.
 (5).
-Furthermore, for a new policy to be incorporated into the policy pool, it must surpass a 55 percent win rate against its predecessor, aligning with the stipulations set in Algo.
-2 at Line 1.
-AlphaZero [10] extends AlphaGo Zero to include games beyond Go, such as Chess and Shogi, with some modifications.
-Notably, a draw is introduced as an additional expected outcome, and data augmentation is omitted due to the asymmetry of Chess and Shogi.
-Concerning the self-play procedure, the only difference between AlphaZero and AlphaGo Zero is that AlphaZero utilizes the newly updated network without the validation process present in AlphaGo Zero.
-Building upon AlphaZero, MuZero [11] takes the concept of learning from scratch to the next level, even operating without predefined game rules.
-MuZero incorporates ideas from model-based RL to model the dynamics of games.
-More concretely, in addition to the prediction network $f$ (similar to the networks in AlphaGo Zero and AlphaZero), MuZero introduces a dynamics network $g$ to model the MDP and a representation network $h$ to map observations to hidden states.
-These three networks are trained jointly.
-Similarly to AlphaGo Zero and AlphaZero, MuZero employs MCTS guided by the three aforementioned networks to make decisions.
+A new policy to be incorporated into the policy pool must surpass a 55 percent win rate against its predecessor.
+AlphaZero [116] extends AlphaGo Zero to include games beyond Go, such as Chess and Shogi.
+Concerning the self-play procedure, the only difference between AlphaZero and AlphaGo Zero is that AlphaZero utilizes the newly updated network to generate samples by self-play without validation.
+Building upon AlphaZero, MuZero [112] takes learning from scratch to the next level, even operating without predefined game rules.
 The self-play process in MuZero operates similarly to that in AlphaZero.
-In practice, in addition to excelling in board games like Go, MuZero also achieves state-of-the-art performance in Atari games.
-Furthermore, several studies have extended the capabilities of MuZero.
-For example, Stochastic MuZero [106] learns a stochastic model instead of a deterministic model to enhance its performance in more complex scenarios.
-Sampled MuZero [107] makes learning feasible in games with intricate action spaces.
-IV-A2 Stratego
-Unlike most board games which are perfect information games, Stratego, a two-player imperfect information board game, distinguishes itself by incorporating elements of memory, deduction, and bluffing.
-This complexity is further amplified by the game’s long episode length and a large number of potential game states, estimated $10^{535}$ [108].
-The game is divided into two phases: the deployment phase, where players secretly arrange their units, setting the stage for strategic depth, and the game-play phase, where the objective is to deduce the opponent’s setup and capture their flag.
+In practice, in addition to excelling in board games, MuZero also achieves state-of-the-art performance in Atari games.
+4.1.2 Stratego
+Unlike most board games, which are perfect information games, Stratego, a two-player imperfect information board game, distinguishes itself by incorporating elements of memory, deduction, and bluffing.
+This complexity is further amplified by the game’s long episode length and many potential game states, estimated $10^{535}$ [104].
+The game is divided into two phases: the deployment phase, where players secretly arrange their units, and the game-play phase, where the objective is to deduce the opponent’s setup and capture their flag.
 The depth and computational complexity of the game remained a challenge until breakthroughs such as DeepNash [104] showed promising advances in AI’s ability to tackle it.
-DeepNash scales up R-NaD [54] to neural R-NaD.
+DeepNash scales up evolution-theory-based self-play method R-NaD [105] (discussed in Sec.
+3.3.5) to neural R-NaD.
 It employs a neural network with four heads: one for value prediction, one for the deployment phase, one for piece selection, and one for piece displacement.
-In the dynamics stage, Neural Replicator Dynamics [109] is utilized to obtain the approximate fixed-point policy effectively.
-Additionally, DeepNash incorporates train-time fine-tuning and test-time post-processing methods to eliminate unreliable actions, thereby enhancing the robustness and accuracy of its output.
-DeepNash holds the third-place ranking among all Gravon Stratego players and secures victories in almost every match against existing Stratego bots.
-IV-B Card Games and Mahjong
-IV-B1 Texas Hold’em
+Neural Replicator Dynamics [52] is utilized to obtain the approximate fixed-point policy.
+DeepNash holds the third-place ranking among all professional Gravon Stratego players and wins nearly every match against existing Stratego bots.
+4.2 Card Games
+Unlike board games, card games are typically imperfect information games with a larger state space and greater complexity.
+Here, we introduce three representative card games: Texas Hold’em, DouDiZhu, and Mahjong, to illustrate how self-play is utilized in card games.
+Texas Hold’em is a repeated game rather than a stage game, so the focus is on the overall win-loss outcome rather than on individual rounds.
+Consequently, it is particularly well-suited for regret-minimization-based self-play methods, which aim to minimize overall regret rather than regret on a per-round basis.
+In contrast, DouDiZhu and Mahjong are typically regarded as stage games, so studies often employ self-play methods that do not rely on regret minimization.
+4.2.1 Texas Hold’em
 Texas Hold’em, a popular poker game with 2-10 players, is known for its strategic depth and bluffing elements.
-In its two-player variant, it’s referred to as heads-up Texas Hold’em.
-The game becomes more complex with additional players.
+The two-player variant is called heads-up Texas Hold’em.
 The gameplay begins with each player receiving two private cards (hole cards), followed by a round of betting.
 Subsequently, three community cards (the flop) are revealed, leading to another betting round.
-This is followed by the dealing of a fourth (the turn) and a fifth community card (the river), each accompanied by further betting rounds.
+This is followed by dealing a fourth (the turn) and a fifth community card (the river), each accompanied by further betting rounds.
 The objective is to construct the best five-card poker hand from any combination of hole cards and community cards.
-The betting proceeds until a showdown occurs, where the remaining players disclose their cards.
-The individual with the superior hand claims the pot.
-Moreover, Texas Hold’em is distinguished by its three betting formats: no-limit, fixed-limit, and pot-limit.
-Each format offers a different risk profile and complexity in terms of game state sizes.
-Among these, the no-limit format is particularly noted for its complexity, allowing players to bet any amount up to their entire stack of chips.
-Its popularity is partly due to its frequent appearance in tournaments and media, making it a staple in modern poker culture.
-While simplified versions such as Kuhn Poker and Leduc Poker serve valuable roles in theoretical analysis and educational contexts within game theory, our focus will focus on the development of algorithms designed to compete in real-world Texas Hold’em.
-In poker, especially Texas Hold’em, abstraction is crucial to simplifying the inherent complexity of the game [110].
-This methodology reduces the vast state space of the game to a manageable size, thus allowing algorithms to calculate strategies more efficiently.
-There are two main types of abstraction in Texas Hold’em: action abstraction, which simplifies the range of possible moves, and information abstraction, which groups similar hand strengths or pot sizes, streamlining the decision-making process.
-Although heads-up limit Texas Hold’em (HULHE), the simplest form of Texas Hold’em with approximately $3.16\times 10^{17}$ game states, was not solved until the introduction of Cepheus [111].
-It utilizes fixed-point arithmetic with compression and CFR+ [75] to address the issues of storage and computation, respectively, thus solving HULHE without involving any neural networks.
-This resulted in superhuman performance and surpassed previous agents.
-Moreover, NSFP [56] introduces the use of end-to-end RL through self-play training, achieving competitive performance in HULHE.
-Similarly, Poker-CNN [112], which utilizes a 3D-vector representation of HULHE combined with convolutional networks, learned through self-play and also achieved competitive performance.
-Deep CFR [90], combines deep neural networks with CFR to surpass the performance of NFSP.
-However, it’s noteworthy that despite advancements, network-based methods haven’t surpassed the regret-based method used by Cepheus, despite its substantial computational demands.
-After the solution of HULHE, research focus shifted to heads-up no-limit Texas Hold’em (HUNL), representing a more complex challenge due to its significantly larger space of game states, approximately $10^{164}$ [113].
-Thus, it is impossible to traverse the game tree as Cepheus does in HULHE.
-The primary technique employed by DeepStack [12] is continual re-solving.
-This method involves dynamically computing the strategy in real-time when it is the program’s turn to act, focusing on a subtree of limited depth and breadth.
-To estimate the outcomes of the furthest reaches of this subtree, DeepStack utilizes neural networks.
-These networks take into account both the player’s range and the opponent’s counterfactual values, which are monitored throughout the entire game.
-The neural networks employed include the turn network, the flop network, and the auxiliary network, all of which are pre-trained using 10 million randomly generated poker games.
-The target counterfactual values for these generated poker situations are derived through thousands of iterations using CFR+ [75].
-DeepStack has demonstrated competitive performance against professional poker players by integrating these approaches.
-Libratus [86] employs self-play to develop its blueprint strategy based on an abstraction of the game, leveraging an enhanced version of MCCFR [77].
-Initially, the blueprint strategy is deployed for decision-making in the first two rounds, benefiting from the detailed abstraction.
-In the later rounds, the blueprint strategy aids in nested safe subgame solving, akin to DeepStack’s continual re-solving aiming to get a more precise strategy for subgames in real-time.
-Additionally, Libratus features a self-improvement component designed to enhance the blueprint strategy by addressing any missing branches.
-Through these strategies, Libratus has proven capable of defeating professional HUNL players.
-ReBel [114] expands the concept of a game state in poker into a public belief state (PBS), thereby transforming the imperfect-information game into a perfect-information game with continuous state space.
-ReBel adopts the methodology of AlphaZero [10], which combines RL with a search algorithm to train both the value network and the policy network by self-play.
-To mitigate heavy computation and storage demands, AlphaHoldem [115] employs an end-to-end self-play RL method augmented with several additional techniques, to outperform its reimplementation of DeepStack.
-During training, the system employs the $K$-best Self-play, where it selects the top $K$ agents based on their ELO ratings [35] and the newly initialized agent.
-It then generates experience replays from competitions among these $K+1$ agents.
-When there are more than two players in Texas Hold’em, the game becomes significantly more complex.
-Pluribus [116] based on the framework of Libratus [86], addresses this complexity in six-player no-limit Texas Hold’em.
-It employs both action abstraction and information abstraction to simplify the intricate dynamics of the original game.
-Similarly to Libratus, Pluribus utilizes MCCFR [77] in self-play to develop its blueprint strategy.
-Building on this blueprint strategy, it conducts a depth-limited search before executing actions.
-More specifically, diverging from Libratus, Pluribus maintains a streamlined policy pool consisting of only four strategies.
-This approach assumes that its opponents might adjust their strategies during gameplay among these four strategies, which allows Pluribus to manage complexity more efficiently.
-Despite its strategy not adapting to the observed tendencies of its opponents, and lacking theoretical proofs of achieving an NE in six-player scenarios, Pluribus has demonstrated its capability by outperforming top professional players in six-player no-limit Texas Hold’em.
-IV-B2 DouDiZhu
+Texas Hold’em offers two betting formats: limited betting and no-limit betting.
+The latter one is noted for its complexity, allowing players to bet any amount up to their entire stack of chips.
+While simplified versions such as Kuhn Poker and Leduc Poker serve valuable roles in theoretical analysis, we will focus on the algorithms designed to compete in full Texas Hold’em.
+Heads-up limit Texas Hold’em (HULHE), the simplest form with approximately $3.16\times 10^{17}$ game states, was not solved until the introduction of Cepheus [10].
+It utilizes fixed-point arithmetic with compression and regret-minimization-based self-play method CFR+ [122] to address the issues of storage and computation, respectively, resulting in superhuman performance.
+Deep CFR [14] combines regret-minimization-based self-play method CFR with deep neural networks.
+Furthermore, some studies do not adopt the regret-minimization-based series of self-play; instead, they use the traditional self-play method.
+NSFP [51] introduces a self-play method combined with end-to-end RL training, also achieving competitive performance in HULHE.
+Poker-CNN [140] utilizes self-play to learn convolutional networks to solve a video version of HULHE.
+After solving HULHE, research focus shifts to heads-up no-limit Texas Hold’em (HUNL) with significantly larger game states, approximately $10^{164}$ [63].
+Thus, traversing the game tree as Cepheus does in HULHE is impossible.
+DeepStack [95] is a regret-minimization-based self-play method combined with continual re-solving.
+This method focuses on a subtree of limited depth and breadth and estimates the outcomes of the furthest reaches of this subtree.
+Libratus [18] develops a blueprint strategy, leveraging an enhanced version of regret-minimization-based self-play method MCCFR [72].
+Recursive Belief-based Learning (ReBeL) [12] introduces the public belief state (PBS), transforming the imperfect-information game into a perfect-information game with continuous state space.
+ReBeL utilizes the regret-minimization-based self-play method CFR-D, combined with RL and search, to train both the value and policy networks.
+Unlike the algorithms mentioned above, AlphaHoldem [149] does not use regret-minimization-based self-play.
+Instead, it proposes a $K$-best self-play method that always selects the top $K$ agents, based on their ELO ratings [36], to generate samples.
+Pluribus [20], based on Libratus, addresses six-player no-limit Texas Hold’em.
+Similar to Libratus, Pluribus utilizes a regret-minimization-based self-play method MCCFR [72] to develop its blueprint strategy.
+It conducts a depth-limited search before executing actions.
+Different from Libratus, Pluribus maintains a streamlined policy pool of only four strategies, assuming that its opponents might adjust their strategies during gameplay among these four strategies, which allows Pluribus to manage complexity more efficiently.
+4.2.2 DouDiZhu
 DouDizhu (a.k.a.
-Fight the Landlord) is a strategic Chinese card game for three players.
-In this game, one player takes on the role of the landlord and competes against the other two players, who are the peasants.
-These two roles have distinct objectives: the landlord’s goal is to outplay the peasants, whereas the peasants work together to defeat the landlord.
-The essence of the game lies in its blend of strategy, cooperation among peasants, and competition between the two factions.
+Fight the Landlord) is a three-player strategic card game.
+In this game, one player takes on the role of the landlord and competes against the other two players, the peasants.
 The game is played in two main stages: the bidding stage and the card play stage.
-During the bidding stage, players vie to become the landlord through a process that involves assessing the strength of their cards.
-The outcome of this stage determines the game’s dynamics, with the landlord playing against the peasants.
-The card play stage is where the main action unfolds.
-Players take turns playing cards in various combinations, intending to be the first to empty their hands.
-The game’s complexity arises from its requirement for strategic thinking, both in playing cards effectively and in the peasants’ need to cooperate seamlessly.
-DouDiZhu is characterized by imperfect information, as players can only see their own cards and the cards played, adding a layer of unpredictability and strategic depth to the game.
-Moreover, with an estimated $10^{76}\sim 10^{106}$ possible game states [117] and an action space comprising $27472$ possible moves [118], DouDiZhu presents a vast landscape of strategic possibilities.
-DeltaDou [119], built upon the AlphaZero framework [10], is the first algorithm to achieve expert-level performance in DouDizhu although it only considers the card play stage.
-It enhances the original framework by incorporating asymmetric MCTS to improve exploration in games characterized by imperfect information.
-Additionally, DeltaDou employs policy-value networks specially designed with input and output encodings to accelerate the training process.
-A notable feature of DeltaDou is its ability to infer the cards of other players using Bayesian methods with the help of the policy-value networks.
-DouZero [120] reduces training costs by opting for a sampling method over the tree search approach.
-It utilizes Monte Carlo methods combined with deep neural networks.
-More specifically, it works similarly to independent RL self-play to generate data, which is then stored in buffers.
-Subsequently, data is sampled from the buffers to update the neural networks.
-Remarkably, DouZero surpasses DeltaDou in performance, demonstrating the effectiveness of its methodology in mastering.
-Additionally, DouZero employs supervised learning to train a bidding network used during the bidding stage, utilizing expert data for this purpose.
-Based on DouZero, DouZero+ [121] enhances the original algorithm by incorporating opponent modeling, which assists the agent in making more informed decisions.
-Additionally, DouZero+ introduces coach-guided learning, aiming to ensure a balanced intensity of initial hand cards among the three players to further accelerate the training process.
-PerfectDou [118] leverages perfect information distillation within a Perfect-Training-Imperfect-Execution (PTIE) framework.
-In this approach, the value network is trained using both perfect and imperfect features, while the policy network is trained solely on imperfect features.
-The networks are refined through independent RL self-play, utilizing Proximal Policy Optimization (PPO) [122] with Generalized Advantage Estimation (GAE) [123].
-Notably, PerfectDou does not incorporate the bidding stage into its model, thereby eliminating the need for expert data in this phase.
-PerfectDou achieves better performance than DouZero.
-IV-B3 Mahjong
-Mahjong, renowned for its complex interplay of skill, strategy, and chance, has evolved into various global variants, including the popular Japanese version known as Riichi Mahjong.
-This game is typically played by four participants who must navigate both the visible aspects of the game, such as discarded tiles, and the hidden elements, like their own hand and the unseen public tiles.
-The strategic depth and complexity of Mahjong pose significant challenges for artificial intelligence.
-Despite ongoing research, many AI implementations have yet to reach expert human levels [124, 125, 126].
-AI systems must make decisions based on incomplete information and adapt strategies dynamically in response to the actions of multiple opponents, which complicates the decision-making process.
-Additionally, the game’s intricate winning rules and the enormous number of possible game states, estimated at around $10^{169}$ [117], add layers of complexity.
-These factors make Mahjong a particularly challenging domain for AI development, requiring sophisticated approaches to handle its unique demands effectively.
-Suphx [127] is recognized as one of the first algorithms to master Mahjong, achieving a performance level comparable to expert human players, specifically 10 dan on Tenhou [128], the most popular online Mahjong platform.
+During the bidding stage, players vie to become the landlord.
+During the card play stage, players take turns playing cards in various combinations, intending to be the first to empty their hands.
+DouDiZhu is characterized by imperfect information, as players can only see their own cards and the cards played.
+The essence of the game lies in cooperation among peasants and competition between the two factions.
+It has an estimated $10^{76}\sim 10^{106}$ possible game states [147] and an action space comprising $27472$ possible moves [141].
+DeltaDou [61] is the first algorithm to achieve expert-level performance in DouDizhu.
+Similar to AlphaZero [116], DeltaDou utilizes vanilla self-play to generate samples to train policy and value networks.
+DouZero [148] also uses vanilla self-play to generate data.
+It reduces training costs by opting for a sampling method over the tree search approach.
+Based on DouZero, DouZero+ [150] incorporates opponent modeling, assisting the agent in making more informed decisions.
+PerfectDou [141] utilizes vanilla self-play to train a Perfect-Training-Imperfect-Execution framework.
+4.2.3 Mahjong
+Mahjong has evolved into various global variants, including the famous Japanese version known as Riichi Mahjong.
+This game is typically played by four players who must navigate both the visible aspects of the game, such as discarded tiles, and the hidden elements, like their own hand and the unseen public tiles.
+The strategic depth and complexity of Mahjong pose significant challenges.
+Despite ongoing research, these findings have yet to reach expert human levels [71, 40, 93].
+Difficulties are navigating incomplete information, dynamically adapting strategies to multiple opponents, and contending with complex winning rules and an enormous number of possible game states, estimated at $10^{169}$ [147].
+Suphx [77] is recognized as one of the first algorithms to master Mahjong, achieving a performance level comparable to expert human players, precisely 10 dan on Tenhou [127], the most popular online Mahjong platform.
 Initially, Suphx employs supervised learning, utilizing expert data to train its model.
-It then advances its capabilities through self-play RL, employing several innovative techniques to enhance its effectiveness.
-These techniques include global reward prediction, which provides a continuous reward throughout the game, and oracle guiding, which offers perfect information at the start of self-play to guide strategy development.
-Additionally, parametric Monte Carlo policy adaptation is used to refine the policy further.
+It then advances its capabilities through vanilla self-play combined with RL.
 Similarly, NAGA [128], developed by Dwango Media Village, and LuckyJ [125], designed by Tencent, have also achieved the rank of 10 dan on Tenhou.
-Furthermore, LuckyJ has even defeated human professional players,
-IV-C Video Games
-In contrast to traditional board games and card games, video games often feature real-time actions, long time horizons and a higher level of complexity due to the broader range of possible actions and observations.
+Furthermore, LuckyJ has even defeated human professional players.
+4.3 Video Games
+In contrast to traditional board games and card games, video games often feature real-time actions, long trajectories, and increased complexity stemming from a broader range of actions and observations.
 Here, we illustrate some representative video games to demonstrate how self-play has advanced AI within these games.
-IV-C1 StarCraft II
-StarCraft is a real-time strategy (RTS) game developed and published by Blizzard Entertainment.
-It showcases three distinct species: the Terrans, Zerg, and Protoss, each with unique units and strategic options that enhance the gameplay experience.
-Renowned for its balanced gameplay, strategic depth, and challenging multiplayer features, the game tasks players with gathering resources, constructing bases, and building armies.
+4.3.1 StarCraft II
+StarCraft is a real-time strategy (RTS) game.
+It has three distinct species: the Terrans, Zerg, and Protoss, each with unique units and strategic options that enhance the gameplay complexity.
+Renowned for its balanced gameplay, strategic depth, and intense competitiveness, the game challenges players to gather resources, construct bases, and build armies.
 Victory requires meticulous planning and tactical execution, with defeat occurring when a player loses all their buildings.
-AlphaStar [15], a significant advancement in AI, dominates the 1v1 mode competitions in StarCraft II and has defeated professional players.
-Its framework bears similarities to that of AlphaGo [8], utilizing supervised learning initially to train the policy with data from human experts.
-Subsequently, it uses end-to-end RL and a hierarchical self-play method to further train the networks.
-More specifically, it divides all the agents into three types: main agents, league exploiters, and main exploiters.
-Additionally, it maintains a policy pool of past players that records all these types of agents.
-Main agents engage in both FSP and PFSP, competing against main agents themselves and other agents in the policy pool.
+AlphaStar [129] dominates the 1v1 mode competitions in StarCraft II and has defeated professional players.
+Its framework is similar to AlphaGo [115], initially utilizing supervised learning to train the policy with expert data.
+Subsequently, it uses a hierarchical self-play method combined with end-to-end RL to train the networks.
+More specifically, The proposed self-play method divides all the agents into three types: main agents, league exploiters, and main exploiters.
+It maintains a policy pool of past players that records all these types of agents.
+Main agents engage in traditional self-play algorithms like FSP and PFSP, competing against main agents themselves and other agents in the policy pool.
 They are periodically added to the pool and never reset.
-League exploiters use PFSP to play against all policy pool agents, added to the pool if they show a high win rate and potentially reset to expose global blind spots.
-Main exploiters only compete with main agents to improve their robustness and are added to the pool after achieving a high win rate or certain training steps, and are reset upon each addition.
-Most importantly, among those three agent types, the main agent is the core agent and embodies the final AlphaStar strategy.
-However, the training computation for AlphaStar is extensive.
-Further studies [131, 132, 133] have enhanced the league self-play training procedure by introducing innovative techniques to reduce computations.
-IV-C2 MOBA Games
-Multiplayer Online Battle Arena games (MOBA) are a popular genre of video games that blend RTS with role-playing elements.
-In typical MOBA games, two teams of players control their own unique characters, known as heroes and compete to destroy the opposing team’s main structure, often called the base.
-Each hero has distinct abilities and plays a specific role within the team, such as Warrior, Tank or Support.
-Managing multiple lanes and battling under the fog of war, which obscures parts of the map are key aspects of gameplay.
-Popular examples of MOBA games include League of Legends, Dota 2 and Honor of Kings.
-These games are known for their complex strategic depth, decision-making under conditions of imperfect information and emphasis on teamwork and player skill.
+League exploiters also use a traditional self-play method PFSP to play against all policy pool agents, added to the pool if they show a high win rate and potentially reset to expose global blind spots.
+Main exploiters only compete with main agents to improve their robustness, are added to the pool after achieving a high win rate or specific training steps, and are reset upon each addition.
+Among those three agent types, the main agent is the core agent and embodies the final AlphaStar strategy.
+However, the introduction of three agent types significantly increases the training computation.
+Further studies [48, 131, 57] have enhanced the league self-play training procedure.
+4.3.2 MOBA Games
+Multiplayer Online Battle Arena (MOBA) games are a popular video game genre that blends RTS with role-playing elements.
+In typical MOBA games, two teams of players control their unique characters, known as heroes, and compete to destroy the opposing team’s main structure, often referred to as the base.
+Each hero has distinct abilities and plays a specific role within the team, such as Warrior, Tank, or Support.
+Managing multiple lanes and battling under the fog of war, which obscures parts of the map, are critical aspects of the gameplay.
+Popular examples of MOBA games include League of Legends, Dota 2, and Honor of Kings.
 OpenAI Five [7] defeated the world champion team in a simplified version of Dota 2 that featured a limited pool of heroes and certain banned items.
-It employs distributed RL using PPO [122] along with GAE [123] to scale up training.
-Each player on a team shares the same policy network and receives shared observation inputs, with the only difference being a unique identity input for each player.
-The training process utilizes self-play, with an 80% chance of engaging in vanilla self-play and a 20% chance of using a self-play technique similar to PFSP used in AlphaStar [15].
-Here, the probabilities of selecting each policy from the policy pool vary according to its quality score, which is continually updated based on competition results throughout the training.
+The training process introduces a self-play method that combines two traditional self-play algorithms: with an 80% probability of engaging in vanilla self-play and a 20% probability of employing a technique similar to the PFSP used in AlphaStar [129].
+This technique selects each policy from the policy pool based on its quality score, which is continuously updated from competition results throughout training.
 Higher quality scores increase the likelihood of a policy being selected.
-A new agent is added to the pool every 10 iterations.
-Similarly to AlphaStar, OpenAI Five requires extensive training resources.
-Thus, they developed surgery tools to resume training with minimal performance loss to manage continual changes in the environment and codes during development.
-Additionally, during training, OpenAI Five employs TrueSkill [40] to evaluate agent performance.
-Another notable MOBA game, especially popular in China, is Honor of Kings.
+OpenAI Five also requires extensive training resources.
+Another notable MOBA game, especially viral in China, is Honor of Kings.
 The 1v1 mode is conquered by [144], which boasts a significant winning rate against top professional players.
-This success can be attributed to well-designed systems tailored for large-scale RL training and sophisticated algorithm designs that handle complex control challenges effectively.
-A key element of its RL training methodology is the use of $\delta$-uniform Self-play.
+It utilizes a traditional self-play method $\delta$-uniform self-play to train the RL networks.
 The 5v5 mode was later mastered by [143] as well.
-Unlike OpenAI Five, this work expands the hero pool to 40 heroes, which exponentially increases the possible combinations of lineups and poses a learning collapse problem in RL training.
-It proposes curriculum self-play learning (CSPL) to mitigate this issue.
-Specifically, the training process is divided into three stages, drawing on insights from curriculum learning.
-The first stage involves training fixed lineups through self-play and utilizing human data to keep the two teams balanced to aid policy improvements.
+Unlike OpenAI Five, this work expands the hero pool to 40 heroes, exponentially increasing the possible combinations of lineups.
+It proposes a new self-play method referred to as curriculum self-play learning (CSPL).
+Specifically, the training process is divided into three stages.
+The first stage involves training fixed lineups through self-play and utilizing human data to balance the two teams to aid policy improvements.
 The second stage employs multi-teacher policy distillation to produce a distilled model.
-The final stage uses this distilled model as the initial model for self-play with randomly picked lineups.
+The final stage uses this distilled model as the initial model for another round of self-play with randomly picked lineups.
 This approach defeats professional player teams.
-Additionally, the self-play generated data is used to learn effective lineup drafting by utilizing MCTS and neural networks [136].
-IV-C3 Google Research Football
-Google Research Football (GRF) [137] is an open-source football simulator that emphasizes high-level actions.
+The self-play generated data is also used to learn compelling lineup drafting by utilizing MCTS and neural networks [25].
+4.3.3 Google Research Football
+Google Research Football (GRF) [70] is an open-source football simulator emphasizing high-level actions.
 It initially offers two scenarios: the football benchmark and the football academy with 11 specific tasks.
-Here, we focus exclusively on the football benchmark because it presents a more complex scenario that better demonstrates the effects of self-play, as opposed to testing on specially designed tasks.
+Here, we focus exclusively on the football benchmark because it presents a more complex scenario that better demonstrates the effects of self-play.
 GRF is particularly challenging due to the need for cooperation among teammates and competition against opposing teams.
-Additionally, it features long trajectories with 3000 steps per round, stochastic transitions, and sparse rewards, all of which contribute to its complexity.
-WeKick [138] claimed victory in the GRF competition on Kaggle [139], which modified the game dynamics by allowing competitors to control just one player, either the ball carrier on offense or the nearest defender on defense.
+It features long trajectories with 3000 steps per round, stochastic transitions, and sparse rewards.
+WeKick [153] claimed victory in the GRF competition on Kaggle [45], which simplifies the game dynamics by allowing competitors to control only one player, either the ball carrier on offense or the nearest defender on defense.
 WeKick employed self-play strategies similar to those used in league training [129].
-It initializes its opponent policy pool using strategies developed through reward shaping via RL, and through Generative Adversarial Imitation Learning [140], which facilitated learning from the tactics of other teams.
-Further research delves into the full football game rather than controlling only one player in the team.
-Team-PSRO [141] extends PSRO [42] to team games, demonstrating approximate TMECor [34] and outperforming self-play in the 4v4 version of the full GRF.
-In the context of the 11v11 version of the full GRF, where the goalkeeper is rule-based controlled, TiKick [142] employs imitation learning to glean insights from data gathered by WeKick’s self-play, resulting in a multi-agent AI through distributed offline RL.
-Moreover, Fictitious Cross-Play (FXP) [143] proposes two populations: the main population and the counter population.
+It initializes its opponent policy pool using strategies developed through RL and Generative Adversarial Imitation Learning (GAIL) [55] to facilitate the training process.
+Further research delves into the full football game rather than the simplified version.
+Team-PSRO [91] extends PSRO series of self-play algorithms to team games, outperforming baselines in the 4v4 version of the full GRF.
+In the context of the 11v11 version, where the goalkeeper is rule-based controlled, TiKick [58] utilizes vanilla self-play to collect samples and then employs imitation learning.
+Fictitious Cross-Play (FXP) [137] proposes a new self-play method similar to AlphaStar [129].
+It introduces two populations: the main population and the counter population.
 Policies in the counter population improve solely by cross-playing with policies in the main population as opponents, while policies in the main population engage in playing with policies from both populations.
-FXP achieves a win rate of over 94% against TiKick in the 11v11 version of the full GRF.
+FXP achieves a win rate of over 94% against TiKick.
 TiZero [78], a follow-up to TiKick, combines curriculum learning with FSP and PFSP [129] to avoid expert data reliance, achieving a higher TrueSkill rating [53] than TiKick.
+Table 2: Overview of Representative Studies in Empirical Analysis.
 Category Game Number of Perfect Number of Algorithms Self-play Method Search Expert Data
 Players Information Game States
-Board Game Chess 2 $\checkmark$ $10^{45}$ AlphaZero [10] Independent RL $\checkmark$ $\times$
-MuZero [11] Independent RL $\checkmark$ $\times$
+Board Game Chess 2 $\checkmark$ $10^{45}$ AlphaZero [116] Vanilla SP $\checkmark$ $\times$
+MuZero [112] Vanilla SP $\checkmark$ $\times$
 Go 2 $\checkmark$ $10^{360}$ AlphaGo [115] FSP $\checkmark$ $\checkmark$
-AlphaGo Zero [9] Independent RL $\checkmark$ $\times$
-AlphaZero [10] Independent RL $\checkmark$ $\times$
-MuZero [11] Independent RL $\checkmark$ $\times$
+AlphaGo Zero [117] Vanilla SP $\checkmark$ $\times$
+AlphaZero [116] Vanilla SP $\checkmark$ $\times$
+MuZero [112] Vanilla SP $\checkmark$ $\times$
 Stratego 2 $\times$ $10^{535}$ DeepNash [104] R-NaD $\times$ $\times$
@@
 Six-player No-limit Texas Hold’em 6 $\times$ $>10^{164}$ Pluribus [20] MCCFR $\checkmark$ $\times$
-DouDiZhu 3 $\times$ $10^{76}\sim 10^{106}$ DeltaDou [119] Independent RL $\checkmark$ $\times$
-DouZero [120] Independent RL $\times$ $\checkmark$
-PerfectDou [118] Independent RL $\times$ $\times$
-Mahjong 4 $\times$ $10^{169}$ Suphx [127] Independent RL $\checkmark$ $\checkmark$
+DouDiZhu 3 $\times$ $10^{76}\sim 10^{106}$ DeltaDou [61] Vanilla SP $\checkmark$ $\times$
+DouZero [148] Vanilla SP $\times$ $\checkmark$
+PerfectDou [141] Vanilla SP $\times$ $\times$
+Mahjong 4 $\times$ $10^{169}$ Suphx [77] Vanilla SP $\checkmark$ $\checkmark$
 Video Game StarCraft II 2 $\times$ / AlphaStar [129] FSP & PFSP $\times$ $\checkmark$
-Dota 2 10 $\times$ OpenAI Five [14] Naive SP & PFSP $\times$ $\times$
+Dota 2 10 $\times$ OpenAI Five [7] Vanilla SP & PFSP $\times$ $\times$
 Honor of Kings 10 $\times$ MOBA AI [143] $\delta$-uniform SP $\checkmark$ $\checkmark$
 Google Research Football 22 $\checkmark$ TiZero [78] FSP & PFSP $\times$ $\times$
-Table III: Overview of Representative Studies in Empirical Analysis.
-V Open Problems and Future Work
-Self-play approaches have demonstrated superior performance due to their unique iterative learning processes and ability to adapt to complex environments.
-However, there are still several areas that require further research and development.
-V-A Theoretical Foundation
-Although NE has been shown to exist in games with finite players and finite actions [145], computing NE in larger games remains challenging and consequently, many studies aim to achieve approximate NE [146].
+5 Open Problems and Future Work
+Self-play methods have shown remarkable performance by leveraging iterative learning and adapting to complex environments.
+However, significant challenges and open questions remain, presenting opportunities for further research and development.
+5.1 Theoretical Foundation
+Although NE has been shown to exist in games with finite players and finite actions [100], computing NE with self-play algorithms in larger games remains challenging and consequently, many studies aim to achieve approximate NE [76].
 However, in some cases, even computing an approximate NE is difficult [31].
 Some research has resorted to higher levels of equilibrium, such as CE [86] and $\alpha$-rank [96].
-Although many algorithms are developed with theoretical foundations in game theory, there is often a gap when it comes to applying these algorithms to complex real-world scenarios.
-For example, although AlphaGo [8], AlphaStar [147] and OpenAI Five [14] have achieved empirical success, they lack formal game theory proofs behind their effectiveness.
-Future work should focus on bridging this gap by combining empirical success with theoretical validation.
-This could involve developing new algorithms that are both theoretically sound and practically effective in complex scenarios or proving the theoretical underpinnings of existing successful algorithms in complex environments.
-V-B Non-stationarity of the Environment
-The strategies of the opponent players evolve as training progresses, and the opponents are a vital factor of the environment in the self-play framework.
+Although many algorithms are developed with theoretical foundations, there is often a gap in complex real-world scenarios.
+For example, although self-play algorithms like AlphaGo [115], AlphaStar [109], and OpenAI Five [7] achieve empirical success, they lack formal game theory proofs behind their effectiveness.
+Future work should focus on bridging this gap by developing new theoretically sound and practically effective algorithms or proving the theoretical underpinnings of existing successful self-play algorithms in complex environments.
+5.2 Non-stationarity of the Environment
+In the self-play framework, the opponents are a vital component of the environment, and the strategies of the opponent players evolve as training progresses.
 This evolution can cause the same strategy to lead to different results over time, creating a non-stationary environment.
 This problem is also shared by the MARL area.
-Future research should aim to develop algorithms that are more robust and can adapt to changing conditions.
-For example, incorporating opponent modeling into self-play [121] can help agents anticipate changes in opponent strategies and adjust their own strategies proactively, making them more robust to environmental changes.
-V-C Scalability and Training Efficiency
+Future research should aim to develop self-play algorithms that are more robust and can adapt to changing conditions.
+For example, incorporating opponent modeling [150] into self-play can help agents anticipate changes in opponent strategies and adjust their own strategies proactively, making them more robust to environmental changes.
+5.3 Scalability and Training Efficiency
 The scalability of self-play methods faces significant challenges as the number of teams and players within those teams increases.
@@
 MOBA AI [143] extends this to a 40-hero pool with the help of curriculum learning, but it still cannot cover the entire hero pool available in the actual game.
-One potential solution is to leverage the inherent connections among players to optimize the learning process.
+One potential solution is leveraging players’ inherent connections to optimize the learning process.
 For instance, using graph-based models to represent and exploit the relationships between players can help manage and reduce the complexity of large-scale multi-agent environments.
-These scalability issues are fundamentally rooted in the limited training efficiency of self-play methods, which involve both computational and storage aspects.
-Computational efficiency is a limiting factor due to the iterative nature of self-play, where agents repeatedly play against themselves or past versions, requiring extensive computational resources.
+These scalability issues are fundamentally rooted in the limited training efficiency of self-play methods from two aspects including computation and storage.
+The first issue is computational efficiency, which is induced by the iterative nature of self-play where agents repeatedly play against themselves or past versions.
 Moreover, although forming more complex populations and competitive mechanisms [109] can enhance the intensity and quality of training, it further exacerbates the demand for computational resources.
 Techniques such as parallel computing, distributed learning, and more efficient neural network architectures could be explored to address these challenges.
-Additionally, self-play is storage-intensive, as it requires maintaining a policy pool.
+The second issue is storage because self-play requires maintaining a policy pool.
 Even when using a shared network architecture, storing the parameters of large models can be problematic.
-This issue is particularly pronounced in regret-minimization-based self-play algorithms, which need to store the regrets for each information set and potential action.
+This issue is particularly pronounced in regret-minimization-based self-play algorithms, which must store the regrets for each information set and potential action.
 Managing both the computational load and storage requirements is essential for improving the overall training efficiency and scalability of self-play methods.
-V-D With Large Language Models
+5.4 With Large Language Models
 With their remarkable capability and emergent generalizability, large language models (LLMs) have been regarded as a potential foundation for achieving human-level intelligence [1], and self-play methods have been proposed to fine-tune LLMs, enhance LLMs’ reasoning performance, and build LLM-based agents with strong decision-making abilities.
-Post-training fine-tuning [149, 150] is a key step in aligning LLM with more desired behaviors, but requires a huge amount of human-annotated data.
+Post-training fine-tuning [102, 3] is a key step in aligning LLM with more desired behaviors, but it requires a huge amount of human-annotated data.
 To reduce reliance on human-annotated data, Self-play fIne-tuNing (SPIN) [26] introduces a self-play mechanism to generate training data using the LLM itself and fine-tuning the LLM by distinguishing self-generated responses from human-annotated data.
+Another line of work [98, 121, 135] formulate the alignment problem as a two-player constant-sum game and use self-play methods to solve the game.
 Some other work [56, 146] also utilizes model-generated data to fine-tune LLMs with minimal human annotations.
 The idea of self-improvement has also been applied to improve the reasoning ability of LLMs.
-SPAG [154] observes that self-play in a two-player adversarial language game called Adversarial Taboo can boost the LLM’s performance on a wide range of reasoning benchmarks.
+Self-Play of Adversarial Game (SPAG) [27] observes that self-play in a two-player adversarial language game called Adversarial Taboo can boost the LLM’s performance on various reasoning benchmarks.
 Besides improving the capability of LLMs, self-play methods have also contributed to building LLM-based agents with strong strategic abilities.
@@
 Another work [138] also combines LLM with self-play policy but takes a different approach by first prompting the LLM to propose multiple action candidates and then using the self-play policy to produce the action distribution over these candidates.
-Despite recent progress, applying self-play with LLMs is still underexplored and requires further research efforts.
-V-E Reality Applications
+Despite recent progress, applying self-play with LLMs is still underexplored and requires further research.
+5.5 Real-World Applications
 Self-play is a powerful technique with widespread applications across various domains.
@@
 Moreover, EvoPlay [132] leverages self-play to design protein sequences, utilizing the advanced AlphaFold2 simulator [65] to narrow the Sim2Real gap.
-Similarly, in heterogeneous multi-robot systems, self-play is utilized for adversarial catching tasks, with significant efforts dedicated to Sim2Real transitions for real-world success [165].
-VI Conclusion
+Similarly, in heterogeneous multi-robot systems, self-play is utilized for competitive sports tasks like humanoid football [81, 47], quadruped competition [136], and multi-drone volleyball [139], with significant efforts dedicated to Sim2Real transitions for real-world success [41].
+6 Conclusion
 The burgeoning field of self-play in RL has been systematically explored in this survey.
-The core idea of self-play, in which agents interact with copies or past versions of themselves, has proven to be a powerful approach for developing robust strategies across various domains.
-Before delving into the specifics of self-play, this paper first elucidates the MARL framework and introduces the background of self-play.
+The core idea of self-play that agents interact with copies or past versions of themselves has proven to be a powerful approach for developing robust strategies across various domains.
+Before delving into the specifics of self-play, this paper first elucidates the MARL framework and introduces the game theory background of self-play.
 Moreover, the paper presents a unified framework and categorizes existing self-play algorithms into four main categories: traditional self-play algorithms, the PSRO series, the ongoing-training-based series, and the regret-minimization-based series.
-Details are meticulously provided on how these four groups are seamlessly integrated into our proposed framework, ensuring a comprehensive understanding of their functionalities.
-The transition from theory to application is underscored by a rigorous analysis of self-play’s role within complex scenarios, such as board games, card games, and video games.
+We provide detailed explanations on how these four groups are seamlessly integrated into our proposed framework, ensuring a comprehensive understanding of their functionalities.
+The transition from theory to application is underscored by a rigorous analysis of self-play’s role within various scenarios, such as board games, card games, and video games.
 Despite the successes of self-play in many areas, challenges remain, such as convergence to suboptimal strategies and substantial computational demands.
-Future work may focus on solving these problems, integrating self-play with LLMs, and achieving real-world applications.
-In conclusion, self-play stands as a cornerstone of modern RL research, offering profound insights and tools for developing advanced AI systems.
-This survey serves as an essential guide for researchers and practitioners, paving the way for further advancements in this dynamic and evolving field.
+Also, future work can focus on how to integrate self-play with LLMs and achieve real-world applications.
+In conclusion, self-play is vital to modern RL research, offering key insights and tools for developing advanced AI systems.
+This survey serves as an essential guide for researchers and practitioners, paving the way for further advancements in this field.
 Acknowledgment
-This research was supported by National Natural Science Foundation of China (No.62325405, U19B2019, M-0248), Tsinghua University Initiative Scientific Research Program, Tsinghua-Meituan Joint Institute for Digital Life, Beijing National Research Center for Information Science, Technology (BNRist) and Beijing Innovation Center for Future Chips, Postdoctoral Fellowship Program of CPSF under Grant Number GZC20240830, China Postdoctoral Science Special Foundation 2024T170496.
+This research was supported by National Natural Science Foundation of China (No.62406159, 62325405), Postdoctoral Fellowship Program of CPSF under Grant Number (GZC20240830, 2024M761676), China Postdoctoral Science Special Foundation 2024T170496.
+Appendix A Appendix
+A.1 Abbreviations in Alphabetical Order
+ABR Approximate Best Response
+AI Artificial Intelligence
+ARM Advantage-based Regret Minimization
+ARMAC Advantage Regret-Matching Actor-Critic
+BR Best Response
+CCE Coarse Correlated Equilibrium
+CE Correlated Equilibrium
+CFR Counterfactual Regret Minimization
+COP Combinatorial Optimization Problem
+CSPL Curriculum Self-Play Learning
+CVRP Capacitated Vehicle Routing Problem
+DCFR Discounted Counterfactual Regret Minimization
+DO Double Oracle
+EGTA Empirical Game-Theoretic Analysis
+FP Fictitious Play
+FSP Fictitious Self-Play
+FTW For The Win
+FXP Fictitious Cross-Play
+GAIL Generative Adversarial Imitation Learning
+GRF Google Research Football
+HULHE Heads-Up Limit Texas Hold’em
+HUNL Heads-Up No-Limit Texas Hold’em
+JPSRO Joint Policy-Space Response Oracle
+LCFR Linear Counterfactual Regret Minimization
+LLM Large Language Model
+MARL Multi-Agent Reinforcement Learning
+MCCFR Monte Carlo Counterfactual Regret Minimization
+MCTS Monte Carlo Tree Search
+MDP Markov Decision Process
+MG Markov Game
+ML Machine Learning
+MOBA Multiplayer Online Battle Arena
+MSS Meta-Strategy Solver
+NE Nash Equilibrium
+NeuPL Neural Population Learning
+NFSP Neural Fictitious Self-Play
+ODO Online Double Oracle
+PBR Preference-based Best Response
+PBS Public Belief State
+PFSP Prioritized Fictitious Self-Play
+POMG Partially Observable Markov Game
+PSRO Policy-Space Response Oracle
+RCFR Regression Counterfactual Regret Minimization
+ReBeL Recursive Belief-based Learning
+RL Reinforcement Learning
+R-NaD Regularized Nash Dynamics
+RTS Real-Time Strategy
+SD-CFR Single Deep Counterfactual Regret Minimization
+SPAG Self-Play of Adversarial Game
+SPIN Self-Play fIne-tuNing
+TSP Traveling Salesman Problem
+URR UnRestricted-Restricted
+XDO Extensive-Form Double Oracle

### v3

@@
-
+A Survey on Self-play Methods in Reinforcement Learning
 Abstract
@@
 This paper is an essential guide map for understanding the multifaceted landscape of self-play in RL.
-keywords:
-[…]
+11footnotetext: Corresponding authors.
 1 Introduction
-Reinforcement learning (RL) represents a significant paradigm [120] within machine learning (ML), focused on optimizing decision processes through interaction with an environment.
+Reinforcement learning (RL) represents a significant paradigm (Sutton & Barto, 2018) within machine learning (ML), focused on optimizing decision processes through interaction with an environment.
 In RL, the environment is modeled as a Markov decision process (MDP), where agents observe states, take actions, receive rewards, and cause transitions to new states.
 The primary goal of RL algorithms is to derive the optimal policy that yields the maximum expected accumulated reward over time.
-Deep RL extends traditional RL by employing deep neural networks as function approximators to handle high-dimensional state spaces, contributing to breakthroughs in various complex tasks [94].
-Moreover, transitioning from single-agent to multi-agent reinforcement learning (MARL) introduces complex dynamics [107, 85, 145].
+Deep RL extends traditional RL by employing deep neural networks as function approximators to handle high-dimensional state spaces, contributing to breakthroughs in various complex tasks (Mnih et al., 2013).
+Moreover, transitioning from single-agent to multi-agent reinforcement learning (MARL) introduces complex dynamics (Rashid et al., 2018; Mahajan et al., 2019; Yu et al., 2022).
 In MARL, the interdependence of agents’ actions presents significant challenges, as the environment appears non-stationary to each agent.
@@
 With the help of game theory, a mathematical framework that models the interactions between multiple decision-makers, self-play emerges as an elegant solution to some inherent challenges in MARL.
-Self-play provides an approach where an agent interacts with copies or past versions of itself [108, 5].
+Self-play provides an approach where an agent interacts with copies or past versions of itself (Samuel, 2000; Bansal et al., 2018).
 This method promises a more stable and manageable learning process.
-The capabilities of self-play extend to a wide range of scenarios, including its high-profile applications in Go [115, 117, 116, 112], chess [116, 112], poker [95, 50], and video games [7, 129].
+The capabilities of self-play extend to a wide range of scenarios, including its high-profile applications in Go (Silver et al., 2016; 2017; 2018; Schrittwieser et al., 2020), chess (Silver et al., 2018; Schrittwieser et al., 2020), poker (Moravčík et al., 2017; Heinrich et al., 2015), and video games (Berner et al., 2019; Vinyals et al., 2019).
 In these scenarios, it has developed strategies that surpass human expertise.
-Although the application of self-play is extensive and promising, it is accompanied by limitations, such as the potential convergence to suboptimal strategies and significant computational requirements[115, 116].
-Although some research adopts a broad perspective through empirical game-theoretic analysis (EGTA)[134], relatively few comprehensive surveys focus solely on self-play algorithms.
-Among them, a study develops an algorithmic framework for self-play, but it does not incorporate the Policy-Space Response Oracle (PSRO) series of algorithms [54], while a separate study focuses exclusively on PSRO [8], without considering other self-play algorithms.
-Another study examines the theoretical safety of self-play [34].
+Although the application of self-play is extensive and promising, it is accompanied by limitations, such as the potential convergence to suboptimal strategies and significant computational requirements(Silver et al., 2016; 2018).
+Although some research adopts a broad perspective through empirical game-theoretic analysis (EGTA)(Wellman et al., 2025), relatively few comprehensive surveys focus solely on self-play algorithms.
+Among them, a study develops an algorithmic framework for self-play, but it does not incorporate the Policy-Space Response Oracle (PSRO) series of algorithms (Hernandez et al., 2021), while a separate study focuses exclusively on PSRO (Bighashdel et al., 2024), without considering other self-play algorithms.
+Another study examines the theoretical safety of self-play (DiGiovanni & Zell, 2021).
 While these studies are valuable, they do not provide a comprehensive perspective that fully captures the breadth and depth of self-play.
@@
 An agent interacts with the environment by taking actions, which result in different states and associated rewards.
-MDPs can be extended to multi-agent settings, known as Markov games (MGs) [79] or stochastic games [113].
+MDPs can be extended to multi-agent settings, known as Markov games (MGs) (Littman, 1994) or stochastic games (Shapley, 1953).
 We focus on the most general form: partially observable Markov games (POMGs), where multiple agents interact with the environment, and each agent receives only individual observations instead of the full state, meaning they have limited access to the environment’s state.
@@
 $\rho:\mathcal{S}\rightarrow[0,1]$ describes initial state distribution.
-If it is a fully cooperative setting, agents can share the same reward function [145, 106, 39, 68].
+If it is a fully cooperative setting, agents can share the same reward function (Yu et al., 2022; Rashid et al., 2020; Foerster et al., 2018; Kuba et al., 2021).
 When $n=1,\mathcal{O}_{i}=\mathcal{S}$, the POMG returns to the MDP.
@@
 Each player comprehensively understands the current game state, the complete history of moves made, and the potential future outcomes.
-If these conditions are not met, the game is considered to have imperfect information [84, 99].
+If these conditions are not met, the game is considered to have imperfect information (Lucchetti, 2011; Mycielski, 1992).
 An example of a game with perfect information is Go.
@@
 This transitive property simplifies the strategic landscape, allowing for an ordinal ranking of strategies.
-Conversely, in a non-transitive game, $\exists\leavevmode\nobreak\ \pi_{i},\pi_{j},\pi_{k}\in\Pi$ such that $u(\pi_{i},\pi_{j})>0$ and $u(\pi_{j},\pi_{k})>0$, but $u(\pi_{i},\pi_{k})\leq 0$.
+Conversely, in a non-transitive game, $\exists~{}\pi_{i},\pi_{j},\pi_{k}\in\Pi$ such that $u(\pi_{i},\pi_{j})>0$ and $u(\pi_{j},\pi_{k})>0$, but $u(\pi_{i},\pi_{k})\leq 0$.
 This introduces a cyclic relationship among strategies, thereby complicating the game.
@@
 meaning that no player can increase their payoff by more than $\epsilon$ by unilaterally changing their strategy.
-However, computing NE is generally intractable in complex games, so some researchers utilize $\alpha$-Rank [101] and Correlated Equilibrium (CE) [2] as alternatives.
-Some studies also resort to Replicator Dynamics [124] to analyze strategy evolution.
+However, computing NE is generally intractable in complex games, so some researchers utilize $\alpha$-Rank (Omidshafiei et al., 2019) and Correlated Equilibrium (CE) (Aumann, 1974) as alternatives.
+Some studies also resort to Replicator Dynamics (Taylor & Jonker, 1978) to analyze strategy evolution.
 3 Algorithms
-Based on existing self-play work [54, 73, 82, 42], we propose a self-play framework (Algo.
+Based on existing self-play work (Hernandez et al., 2021; Lanctot et al., 2017; Liu et al., 2022c; Garnelo et al., 2021), we propose a self-play framework (Algo.
 1) that boasts enhanced expressivity and superior generalization capabilities.
@@
 Among these, traditional self-play, the PSRO series, and the regret-minimization-based series utilize lazy initialization and standard training, whereas the ongoing-training-based series employs immediate initialization and ongoing training.
-For a more detailed description and analysis, please refer to Sec.3 and Table1.
-Algorithm 1 the Framework of Self-play.
+For a detailed description and analysis, please refer to Sec.3 and Table1.
+Algorithm 1 A unified framework of self-play.
 1:  Initialize $\Pi,\Sigma$
@@
 In Algo.
-1, we define a unified self-play framework based on [82, 73, 42, 54].
+1, we define a unified self-play framework based on Liu et al.
+(2022c); Lanctot et al.
+(2017); Garnelo et al.
+(2021); Hernandez et al.
+(2021).
 We combine visual illustrations (Fig.
@@
 Next, we provide an in-depth explanation of these processes:
-Figure 4: Interaction matrix $\Sigma$ examples.
-When $C_{1}=N$ and the opponent sampling strategy $\sigma_{ij}$ represents the probability that policy $i$ is optimized against policy $j$, the interaction matrix $\Sigma$ can be depicted in directed interaction graphs.
-Here, we consider three representative self-play algorithms.
-In the Top section, we define the interaction matrix $\Sigma\in\mathbb{R}^{3\times 3}$ as $\{\sigma_{i}\}_{i=1}^{3}$.
-In the Bottom section, we present directed interaction graphs where the outgoing edges from each node are equally weighted, and their weights collectively sum to one.
-The relationship between the Top and Bottom sections is established through directed edges: an edge directed from node $i$ to node $j$ with a weight of $\sigma_{ij}$ signifies that policy $i$ is optimized against policy $j$ with a probability of $\sigma_{ij}$.
-Note that this figure is reproduced from [82] and this concept is initially proposed by [42].
   • $\Pi:=\{\pi_{i}(\cdot|h(i))\}_{i=1}^{N}$: Each policy $\pi_{i}$ in the policy population $\Pi$ is conditioned on a policy condition function $h(i)$, which provides supplementary information for certain algorithms.
@@
 1): The initialization of $\pi_{i}^{h}$ can vary depending on the algorithm being used.
-It can be initialized randomly or by leveraging pre-trained models [115, 129] or through a recently updated policy.
+It can be initialized randomly or by leveraging pre-trained models (Silver et al., 2016; Vinyals et al., 2019) or through a recently updated policy.
   • $\text{ORACLE}(\pi_{i},\sigma_{i},\Pi)$ (Line 5 in Algo.
@@
 Here, we divide ORACLE into three types.
-(1) One type is the BR oracle, which is designed to identify the optimal counter-strategies against an opponent’s strategy, including finding NE [92].
+(1) One type is the BR oracle, which is designed to identify the optimal counter-strategies against an opponent’s strategy, including finding NE (McMahan et al., 2003).
 However, it often requires considerable computational effort.
@@
 4).
-(3) Other specially crafted ORACLEs are either tailored to specific meta-strategy solvers (MSSes) [96, 86] or introduced to enhance diversity [4, 103, 83].
+(3) Other specially crafted ORACLEs are either tailored to specific meta-strategy solvers (MSSes) (Muller et al., 2020; Marris et al., 2021) or introduced to enhance diversity (Balduzzi et al., 2019; Perez-Nieves et al., 2021; Liu et al., 2021).
 Some details of Algo.
@@
 Algo.
-3 is specifically utilized by Regularized Nash Dynamics (R-NaD) [105] and Algo.
+3 is specifically utilized by Regularized Nash Dynamics (R-NaD) (Perolat et al., 2021) and Algo.
 4 is specifically utilized by the regret-minimization-based series.
@@
 1) can be skipped to reduce computations.
-    Moreover, if $\Pi$ uses immediate initialization, $\Sigma$ will be initialized (Line 1 in Algo. 1) by first evaluating the performance and then applying the MSS.
+Moreover, if $\Pi$ uses immediate initialization, $\Sigma$ will be initialized (Line 1 in Algo.
+1) by first evaluating the performance and then applying the MSS.
+Figure 4: Interaction matrix $\Sigma$ examples.
+When $C_{1}=N$ and the opponent sampling strategy $\sigma_{ij}$ represents the probability that policy $i$ is optimized against policy $j$, the interaction matrix $\Sigma$ can be depicted in directed interaction graphs.
+Here, we consider three representative self-play algorithms.
+In the Top section, we define the interaction matrix $\Sigma\in\mathbb{R}^{3\times 3}$ as $\{\sigma_{i}\}_{i=1}^{3}$.
+In the Bottom section, we present directed interaction graphs where the outgoing edges from each node are equally weighted, and their weights collectively sum to one.
+The relationship between the Top and Bottom sections is established through directed edges: an edge directed from node $i$ to node $j$ with a weight of $\sigma_{ij}$ signifies that policy $i$ is optimized against policy $j$ with a probability of $\sigma_{ij}$.
+Note that this figure is reproduced from Liu et al.
+(2022c) and this concept is initially proposed by Garnelo et al.
+(2021).
 Algorithm 2 Compute the oracle in RL.
@@
 3:     $\bm{\pi}=(\pi_{i}^{h}(j),\bm{\pi_{opp}^{h}}(-j))$
-4:     Use regret matching to play the game and obtain new regret minimization information added to $h(i)$.
+4:     Use regret matching to play the game and update $h(i)$ with new regret minimization information.
 5:  end for
@@
 Only the MSS for prioritized fictitious self-play (described in Sec.
-3.2.5) requires the performance matrix; nonetheless, it does not need to solve for complex game outcomes such as NE.
+3.2.5) requires the performance matrix; nonetheless, it does not need to solve for complex game outcomes like NE.
 Next, we outline the traditional self-play schemes. For simplicity, we operate under the following assumption:
 Assumption 1.
-In a two-player symmetric game, $C_{1}=N$, with $\sigma_{ij}$ denoting the probability that policy $i$ is optimized in response to policy $j$ which leads to $\sum_{j=1}^{N}\sigma_{ij}=1,\leavevmode\nobreak\ \forall i$.
+In a two-player symmetric game, $C_{1}=N$, with $\sigma_{ij}$ denoting the probability that policy $i$ is optimized in response to policy $j$ which leads to $\sum_{j=1}^{N}\sigma_{ij}=1,~{}\forall i$.
 Based on Assumption 1, we can further deduce the following important corollary:
@@
 The policy population gradually increases over time.
-Consequently, when policy $i$ is selected for training, only policy $j$ $(j\leq i)$ has already been trained and holds meaningful outcomes.
+When policy $i$ is selected for training, only policy $j$ $(j\leq i)$ has already been trained and holds meaningful outcomes.
 Other policies are inactive policies.
@@
 3.2.2 Vanilla Self-play
-In vanilla self-play [108], agents are trained by competing against their latest versions. Thus, the MSS of vanilla self-play:
-$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}1,\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ j=i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right..$ (5)
+In vanilla self-play (Samuel, 2000), agents are trained by competing against their latest versions. Thus, the MSS of vanilla self-play:
+$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}1,~{}\text{if}~{}j=i-1\vspace{4pt}\\ 0,~{}\text{otherwise}\end{array}\right..$ (5)
 Regardless of $\mathcal{P}$, the MSS produces the same interaction matrix.
@@
 3.2.3 Fictitious Self-play
-Fictitious Play (FP) [11] is a learning algorithm in game theory where each player best responds to the empirical frequency of the strategies used by the opponent.
+Fictitious Play (FP) (Brown, 1951) is a learning algorithm in game theory where each player best responds to the empirical frequency of the strategies used by the opponent.
 If the opponent’s strategy is static, FP can find the NE.
-Based on FP intuition, Fictitious Self-Play (FSP) [50] is introduced to make agents play against past versions of themselves to learn optimal strategies to improve the robustness of vanilla self-play.
-Neural Fictitious Self-Play (NFSP) [51] is a modern variant that combines FSP with deep learning techniques.
+Based on FP intuition, Fictitious Self-Play (FSP) (Heinrich et al., 2015) is introduced to make agents play against past versions of themselves to learn optimal strategies to improve the robustness of vanilla self-play.
+Neural Fictitious Self-Play (NFSP) (Heinrich & Silver, 2016) is a modern variant that combines FSP with deep learning techniques.
 It uses neural networks to approximate the BRs.
 In the original versions of NFSP and FSP, two distinct types of agent memory are used: one to record the agent’s own behavior and the other to capture the opponent’s behavior.
-In more recent approaches [73], random sampling is frequently employed to approximate the opponents’ average strategy, eliminating the need to maintain two separate types of agent memory.
+In more recent approaches (Lanctot et al., 2017), random sampling is frequently employed to approximate the opponents’ average strategy, eliminating the need to maintain two separate types of agent memory.
 Thus, the MSS of FSP:
-$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{i-1},\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ j\leq i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right..$ (6)
+$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{i-1},~{}\text{if}~{}j\leq i-1\vspace{4pt}\\ 0,~{}\text{otherwise}\end{array}\right..$ (6)
 In FSP, the MSS continues to generate a constant interaction matrix.
@@
 […]
-$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{f(i)},\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \lceil\delta(i-1)\rceil\leq j\leq i-1\vspace{4pt}\\ 0,\leavevmode\nobreak\ \text{otherwise}\end{array}\right.,$ (7)
+$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{f(i)},~{}\text{if}~{}\lceil\delta(i-1)\rceil\leq j\leq i-1\vspace{4pt}\\ 0,~{}\text{otherwise}\end{array}\right.,$ (7)
 $f(i)=i-\lceil\delta(i-1)\rceil.$ (8)
@@
 3.2.5 Prioritized Fictitious Self-play
-Prioritized Fictitious Self-Play (PFSP)  [129] leverages a priority function to allocate a higher probability of selection to agents with higher priorities.
-Here, $\mathcal{P}$ represents the winning rates, specifically defined as $\mathcal{P}_{ij}=\text{Prob}(\pi_{i}\leavevmode\nobreak\ \text{beats}\leavevmode\nobreak\ \pi_{j})$.
+Prioritized Fictitious Self-Play (PFSP)  (Vinyals et al., 2019) leverages a priority function to allocate a higher probability of selection to agents with higher priorities.
+Here, $\mathcal{P}$ represents the winning rates, specifically defined as $\mathcal{P}_{ij}=\text{Prob}(\pi_{i}~{}\text{beats}~{}\pi_{j})$.
 The MSS of PFSP is given by Algo.
@@
 Additionally, in broader terms, $\mathcal{P}$ can also be assessed by other metrics.
-A similar MSS in  [7] can be utilized to allocate higher probabilities to strategies that perform better or are more challenging to defeat.
-Algorithm 5 the PFSP Meta-strategy Solver.
+A similar MSS in  Berner et al.
+(2019) can be utilized to allocate higher probabilities to strategies that perform better or are more challenging to defeat.
+Algorithm 5 PFSP meta-strategy solver.
 1:  function $\mathcal{F}(\mathcal{P})$
-2:   $\sigma_{i+1,j}\leftarrow\frac{f(\mathcal{P}_{i,j})}{\sum_{j\leq i}f(\mathcal{P}_{i,j})},\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ j\leq i$
+2:   $\sigma_{i+1,j}\leftarrow\frac{f(\mathcal{P}_{i,j})}{\sum_{j\leq i}f(\mathcal{P}_{i,j})},~{}\text{if}~{}j\leq i$
 3:   Append zeros to $\sigma_{i+1}$ until its length is N.
@@
 Fourth, we set $h(i)=\emptyset$ since the PSRO series of algorithms do not use any policy condition function for their policies.
-Fifth, it’s crucial to highlight that our framework diverges from the traditional PSRO model [73] in how $\sigma_{i}$ is defined.
+Fifth, it’s crucial to highlight that our framework diverges from the traditional PSRO model (Lanctot et al., 2017) in how $\sigma_{i}$ is defined.
 In contrast to being the meta-strategy for policy, in our framework, $\sigma_{i}$ is the opponent sampling strategy.
@@
 Sixth, compared with traditional self-play methods, the MSSes of the PSRO series are often more complex.
-For example, some MSSes incorporate concepts from different types of game equilibria [92, 96, 86].
+For example, some MSSes incorporate concepts from different types of game equilibria (McMahan et al., 2003; Muller et al., 2020; Marris et al., 2021).
 Lastly, we also simply follow Assumption 1.
@@
 3.3.2 Double Oracle
-Double Oracle (DO) [92] can be only applied to two-player normal-form games.
+Double Oracle (DO) (McMahan et al., 2003) can be only applied to two-player normal-form games.
 In this context, we can utilize the payoff matrix as the evaluation matrix.
@@
 In the context of two-player normal-form games, DO theoretically can achieve the NE of the full game.
-Algorithm 6 the NE-based Meta-strategy Solver.
+Algorithm 6 NE-based meta-strategy solver.
 1:  function $\mathcal{F}(\mathcal{P})$
@@
 3.3.3 PSRO
-PSRO [73] extends DO to more complex games beyond just two-player normal-form games.
+PSRO (Lanctot et al., 2017) extends DO to more complex games beyond just two-player normal-form games.
 This approach first introduces the concept of the MSS, which is a broader concept than simply computing NE.
@@
 Some studies focus on making PSRO more computationally tractable in different scenarios or achieving fast convergence.
-By establishing a hierarchical structure of RL workers, Pipeline PSRO [87] achieves the parallelization of PSRO and simultaneously provides guarantees of convergence.
-Efficient PSRO [151] introduces the unrestricted-restricted (URR) game to narrow the selection of opponent policies, thereby preventing the need for meta-game simulations in traditional PSRO.
+By establishing a hierarchical structure of RL workers, Pipeline PSRO (McAleer et al., 2020) achieves the parallelization of PSRO and simultaneously provides guarantees of convergence.
+Efficient PSRO (Zhou et al., 2022) introduces the unrestricted-restricted (URR) game to narrow the selection of opponent policies, thereby preventing the need for meta-game simulations in traditional PSRO.
 Similar to Pipeline PSRO, Efficient PSRO is equipped to solve URR in parallel.
-In addition, unlike traditional PSRO, which confines the integration of population strategies to the game’s initial state, Extensive-Form Double Oracle (XDO) [89] allows this integration across all information states.
+In addition, unlike traditional PSRO, which confines the integration of population strategies to the game’s initial state, Extensive-Form Double Oracle (XDO) (McAleer et al., 2021) allows this integration across all information states.
 It ensures a linear convergence to an approximate NE based on the number of information states, enhancing its tractability in extensive-form games.
-Online Double Oracle (ODO) [35] integrates the no-regret analysis of online learning with DO to improve both the convergence rate to an NE and the average payoff.
-Anytime PSRO [90] and Self-play PSRO [88] are designed to incorporate less exploitable policies into the policy population, facilitating faster convergence.
+Online Double Oracle (ODO) (Dinh et al., 2022) integrates the no-regret analysis of online learning with DO to improve both the convergence rate to an NE and the average payoff.
+Anytime PSRO (McAleer et al., 2022b) and Self-play PSRO (McAleer et al., 2022a) are designed to incorporate less exploitable policies into the policy population, facilitating faster convergence.
 Moreover, as the number of agents grows, determining BRs becomes exponentially challenging.
-Mean-field PSRO [97] has been introduced to address this complexity in mean-field games.
-In addition, due to the computational intractability of solving NE in multi-player general-sum games [32, 31] and the selection problem in NE [44], Müller et al.
-put forward $\alpha$-PSRO [96].
-Instead of NE, $\alpha$-rank [101], which is unique and polynomial-time solvable in multi-player general-sum games, is introduced in the MSS and a preference-based best response (PBR) oracle is incorporated in the approach.
-Similar to $\alpha$-PSRO, [86] proposes Joint Policy-Space Response Oracle (JPSRO) to tackle multi-player general-sum games.
+Mean-field PSRO (Muller et al., 2022) has been introduced to address this complexity in mean-field games.
+In addition, due to the computational intractability of solving NE in multi-player general-sum games (Daskalakis et al., 2009; Daskalakis, 2013) and the selection problem in NE (Goldberg et al., 2013), Müller et al.
+put forward $\alpha$-PSRO (Muller et al., 2020).
+Instead of NE, $\alpha$-rank (Omidshafiei et al., 2019), which is unique and polynomial-time solvable in multi-player general-sum games, is introduced in the MSS and a preference-based best response (PBR) oracle is incorporated in the approach.
+Similar to $\alpha$-PSRO, Marris et al.
+(2021) proposes Joint Policy-Space Response Oracle (JPSRO) to tackle multi-player general-sum games.
 It utilizes CE and coarse correlated equilibrium (CCE) as alternatives to NE to make it computationally tractable in multi-player general-sum games and designs the CE and CCE-based MSS.
 Other research focuses on policy diversification, as deriving a single policy in transitive games is less meaningful.
-[4] introduces an open-ended framework in two-player zero-sum games.
+Balduzzi et al.
+(2019) introduces an open-ended framework in two-player zero-sum games.
 This framework enhances the diversity of the strategy population and introduces a gamescape, which geometrically represents the latent objectives in games for open-ended learning.
@@
 Compared to $\text{PSRO}_{\text{N}}$, $\text{PSRO}_{\text{rN}}$ incorporates an additional step within the ABR oracle to focus on the opponents they beat or tie and ignore the opponents they lose to.
-[103] uses determinantal point process [69] to assess diversity and introduces diverse PSRO by incorporating a diversity term into the PSRO oracle.
+Perez-Nieves et al.
+(2021) uses determinantal point process (Kulesza et al., 2012) to assess diversity and introduces diverse PSRO by incorporating a diversity term into the PSRO oracle.
 This modification can also be easily implemented in FP and $\alpha$-PSRO.
-Similarly,  [83] introduces the concepts of behavioral diversity and response diversity and incorporates them into the PSRO oracle.
-Policy Space Diversity PSRO [142] defines a diversity metric named population exploitability that helps to achieve a full-game NE.
+Similarly,  Liu et al.
+(2021) introduces the concepts of behavioral diversity and response diversity and incorporates them into the PSRO oracle.
+Policy Space Diversity PSRO (Yao et al., 2023) defines a diversity metric named population exploitability that helps to achieve a full-game NE.
 3.3.5 R-NaD
-Although R-NaD [105] is initially described as leveraging evolutionary game theory with a regularization component.
+Although R-NaD (Perolat et al., 2021) is initially described as leveraging evolutionary game theory with a regularization component.
 Here, we categorized it into the PSRO series with a unique oracle computation technique (Algo.
-3), executed in two stages: the first stage involves transforming the reward based on the regularization policy to make it policy-dependent, and the second stage applies replicator dynamics [124] until convergence to a fixed point.
+3), executed in two stages: the first stage involves transforming the reward based on the regularization policy to make it policy-dependent, and the second stage applies replicator dynamics (Taylor & Jonker, 1978) until convergence to a fixed point.
 The oracle added to the policy population $\Pi$ in each iteration is derived from the reward-transformed game, not the original problem’s oracle.
@@
 First, truncating the ABR operators during each iteration is often necessary when operating with a limited budget, introducing sub-optimally trained responses into the population.
-Second, relearning basic skills in every iteration is redundant and becomes untenable when confronted with increasingly formidable opponents [82].
+Second, relearning basic skills in every iteration is redundant and becomes untenable when confronted with increasingly formidable opponents (Liu et al., 2022c).
 To address these challenges, the ongoing-training-based series of algorithms emphasizes the repeated ongoing training of all policies.
@@
 Quake III Arena Capture the Flag is a renowned 3D multi-player first-person video game where two teams vie to capture as many flags as possible.
-The For The Win (FTW) agent [60] is designed to perform at human-level proficiency in this game.
+The For The Win (FTW) agent (Jaderberg et al., 2019) is designed to perform at human-level proficiency in this game.
 A pivotal aspect of FTW is its employment of the ongoing-training-based self-play in RL.
@@
 3.4.3 NeuPL
-Neural Population Learning (NeuPL) [82] introduces another critical innovation: it employs a unified conditional network, where each policy is adjusted against a specific meta-game mixture strategy.
+Neural Population Learning (NeuPL) (Liu et al., 2022c) introduces another critical innovation: it employs a unified conditional network, where each policy is adjusted against a specific meta-game mixture strategy.
 This is instrumental in enabling transfer learning across policies.
 Owing to NeuPL’s reliance on a unified conditional network parameterized by $\theta$, $\pi_{i}(\cdot|h(i))$ can be succinctly represented as $\pi_{\theta}(\cdot|h(i))$.
-Moreover, given that the policies in NeuPL are contingent on the opponent sampling strategy $\sigma_{i}$, it is apt to define $h(i)=\sigma_{i}$.
+Since NeuPL policies depend on the opponent sampling strategy $\sigma_{i}$, we define $h(i)=\sigma_{i}$.
 3.4.4 Simplex-NeuPL
-Simplex-NeuPL [80], which builds on NeuPL, is designed to achieve any-mixture optimality, which signifies that the formulated policy should exhibit flexibility across a diverse spectrum of adversaries, including those that might not present equivalent competitive prowess.
+Simplex-NeuPL (Liu et al., 2022a), which builds on NeuPL, is designed to achieve any-mixture optimality, which signifies that the formulated policy should exhibit flexibility across a diverse spectrum of adversaries, including those that might not present equivalent competitive prowess.
 To model the population learning process from a geometric perspective, Simplex-NeuPL introduces the concept of the population simplex.
@@
 Note that this category relies heavily on the information in $h(i)$.
-For instance, in Counterfactual Regret Minimization (CFR) [152], it is necessary to store counterfactual regrets for every action in every information set for each player within $h(i)$.
+For instance, in Counterfactual Regret Minimization (CFR) (Zinkevich et al., 2007), it is necessary to store counterfactual regrets for every action in every information set for each player within $h(i)$.
 Once $h(i)$ is determined, the corresponding policy is also defined.
@@
 This means that the regrets of player 2 are updated after considering the already-updated regrets of player 1.
-This adjustment has been shown to accelerate convergence empirically and possesses a theoretical error bound [22].
+This adjustment has been shown to accelerate convergence empirically and possesses a theoretical error bound (Burch et al., 2019).
 Additionally, each $\pi_{i}^{h}$ represents the strategies for all players, and in iteration $i$, player $j$ uses the strategy $\pi_{i}^{h}(j)$.
@@
 Regret measures the difference between the best possible payoff and the actual payoff.
-The regret matching algorithm [49] optimizes decisions by selecting strategies based on accumulated positive overall regrets, with higher overall regret strategies being more likely chosen to correct past underperformance.
+The regret matching algorithm (Hart & Mas-Colell, 2000) optimizes decisions by selecting strategies based on accumulated positive overall regrets, with higher overall regret strategies being more likely chosen to correct past underperformance.
 After each round, players update the overall regret values for each strategy.
@@
 However, this method primarily applies to normal-form games because computing overall regret in extensive-form games is challenging.
-[152] proposes CFR for extensive-form games.
+Zinkevich et al.
+(2007) proposes CFR for extensive-form games.
 Here, it is referred to as vanilla CFR to distinguish it from later advancements.
@@
 $R_{j}^{T}=\dfrac{1}{T}\max_{\pi(j)^{*}\in\Pi(j)}\sum_{i=1}^{T}\left(u_{j}(\pi(j)^{*},\pi_{i}(-j))-u_{j}(\pi_{i})\right).$ (14)
-And theoretically, average overall regret $R_{j}^{T}$ is bounded by the sum of positive immediate counterfactual regrets:
+Theoretically, average overall regret $R_{j}^{T}$ is bounded by the sum of positive immediate counterfactual regrets:
 $R_{j}^{T}\leq\sum_{I\in\mathcal{I}_{j}}R_{j,\text{imm}}^{T,+}(I).$ (15)
@@
 $R_{j}^{T,+}(I,a)=\max\{R_{j}^{T}(I,a),0\}.$ (17)
-The regret matching algorithm [49] is used to decide the strategy in each information set:
+The regret matching algorithm (Hart & Mas-Colell, 2000) determines the strategy at each information set:
 $\left.\pi_{j}^{T+1}(I)(a)=\left\{\begin{array}[]{ll}\frac{R_{j}^{T,+}(I,a)}{\sum_{a\in A(I)}R_{j}^{T,+}(I,a)}&\text{ if denominator}>0\\ \frac{1}{|A(I)|}&\text{ otherwise}\end{array}\right.\right..$ (18)
@@
 The first approach involves modifying the regret calculation to increase its speed.
-CFR+ [122] implements regret-matching+ by storing non-negative regret-like values $R_{j}^{+,i}(I,a)$, rather than $R_{j}^{i}(I,a)$.
+CFR+ (Tammelin, 2014) implements regret-matching+ by storing non-negative regret-like values $R_{j}^{+,i}(I,a)$, rather than $R_{j}^{i}(I,a)$.
 Additionally, CFR+ updates the regrets of each player sequentially and adopts a weighted average strategy.
-Moreover, [19] introduces the concept of weighted regrets and develops Linear CFR (LCFR) and Discounted CFR (DCFR).
+Moreover, Brown & Sandholm (2019a) introduces the concept of weighted regrets and develops Linear CFR (LCFR) and Discounted CFR (DCFR).
 The second approach involves adopting sampling methods.
 While it requires more iterations, each iteration is shorter, reducing the overall convergence time.
-Monte Carlo CFR (MCCFR) [72] divides terminal histories into blocks, with each iteration sampling from these blocks instead of traversing the entire game tree.
+Monte Carlo CFR (MCCFR) (Lanctot et al., 2009) divides terminal histories into blocks, with each iteration sampling from these blocks instead of traversing the entire game tree.
 This allows for calculating sampled counterfactual values for each player, leading to counterfactual regrets that, in expectation, match those of the Vanilla CFR.
@@
 MCCFR typically manifests in three forms: outcome-sampling MCCFR, where each block corresponds to a single history; external-sampling MCCFR, which samples opponent and chance nodes; and chance-sampling MCCFR, only focusing on chance nodes.
-Moreover, [64] expands on chance-sampling by categorizing based on the handling of public and private chance nodes.
-Some studies also focus on learning how to reduce the variance of MCCFR to speed up convergence [111, 33].
-In addition to these two primary approaches, other studies have identified that warm starting [16] and pruning [15, 13, 17] can also accelerate convergence.
+Moreover, Johanson et al.
+(2012) expands on chance-sampling by categorizing based on the handling of public and private chance nodes.
+Some studies also focus on learning how to reduce the variance of MCCFR to speed up convergence (Schmid et al., 2019; Davis et al., 2020).
+In addition to these two primary approaches, other studies have identified that warm starting (Brown & Sandholm, 2016) and pruning (Brown & Sandholm, 2015; Brown et al., 2017; Brown & Sandholm, 2017) can also accelerate convergence.
 3.5.4 Space-saving Variants of CFR
@@
 However, in imperfect information games, defining subgames is challenging due to their intersection with information set boundaries.
-CFR-D [21] is a pioneering method for decomposing imperfect information games into the main component called the trunk and subgames defined by forests of trees that do not divide any information sets.
+CFR-D (Burch et al., 2014) is a pioneering method for decomposing imperfect information games into the main component called the trunk and subgames defined by forests of trees that do not divide any information sets.
 In each iteration, CFR is applied within the trunk for both players, and a solver is used to determine the counterfactual BR in the subgame.
@@
 CFR-D minimizes storage needs by only saving $R^{i}(I_{*},a)$ for information sets in the trunk and at each subgame’s root, which is denoted by $I_{*}$, trading off storage efficiency against the time required to resolve subgames.
-Similar thoughts are echoed in the Continue-Resolving technique used by DeepStack [95] and the Safe and Nested Subgame Solving technique used by Libratus [18].
+Similar thoughts are echoed in the Continue-Resolving technique used by DeepStack (Moravčík et al., 2017) and the Safe and Nested Subgame Solving technique used by Libratus (Brown & Sandholm, 2018).
 We will discuss these approaches in Sec.
@@
 Additionally, smaller abstractions often yield suboptimal results.
-Given these challenges, [133] introduces Regression CFR (RCFR), which employs a shared regressor $\varphi(I,a)$ to estimate counterfactual regrets.
+Given these challenges, Waugh et al.
+(2015) introduces Regression CFR (RCFR), which employs a shared regressor $\varphi(I,a)$ to estimate counterfactual regrets.
 Nevertheless, using regression trees as the regressor limits RCFR’s applicability to miniature games, and the necessity for manually crafted features remains a drawback.
-After advantage-based regret minimization (ARM) [62] merges CFR with deep RL in single-agent scenarios, a growing body of research has focused on applying CFR in conjunction with neural networks to multi-agent scenarios.
-Double Neural CFR [75] utilizes two neural networks: one for estimating counterfactual regrets and another for approximating the average strategy.
-In a similar vein, Deep CFR [14] leverages an advantage network $V(I,a|\theta_{p})$ to estimate counterfactual regrets with each player having a distinct hyperparameter $\theta_{p}$ and employs $\pi(I,a|\theta_{\pi})$ for strategy estimation after the training process of the advantage network.
+After advantage-based regret minimization (ARM) (Jin et al., 2018) merges CFR with deep RL in single-agent scenarios, a growing body of research has focused on applying CFR in conjunction with neural networks to multi-agent scenarios.
+Double Neural CFR (Li et al., 2018) utilizes two neural networks: one for estimating counterfactual regrets and another for approximating the average strategy.
+In a similar vein, Deep CFR (Brown et al., 2019) leverages an advantage network $V(I,a|\theta_{p})$ to estimate counterfactual regrets with each player having a distinct hyperparameter $\theta_{p}$ and employs $\pi(I,a|\theta_{\pi})$ for strategy estimation after the training process of the advantage network.
 Since these two networks are trained in sequence rather than concurrently, the strategy for each intermediate iteration remains conditioned on the output of the advantage network: $h(i)=V(I,a|\theta_{p})$.
 Despite similarities, Deep CFR distinguishes itself from Double Neural CFR through its data collection and proven effectiveness in larger-scale poker games.
-Single Deep CFR (SD-CFR) [118] demonstrates training an average strategy network is unnecessary, with only an advantage network required.
-Building on the foundation of SD-CFR, DREAM [119] utilizes a learned baseline to maintain low variance in a model-free setting when only one action is sampled at each decision point.
-Moreover, advantage regret-matching actor-critic (ARMAC) [46] combines the thought of retrospective policy improvement with CFR.
+Single Deep CFR (SD-CFR) (Steinberger, 2019) demonstrates training an average strategy network is unnecessary, with only an advantage network required.
+Building on the foundation of SD-CFR, DREAM (Steinberger et al., 2020) utilizes a learned baseline to maintain low variance in a model-free setting when only one action is sampled at each decision point.
+Moreover, advantage regret-matching actor-critic (ARMAC) (Gruslys et al., 2020) incorporates the thought of retrospective policy improvement.
 Table 1: Overview of representative self-play algorithms within our framework.
 Algorithms MSS h(i) Categories $E$ Initialization of $\Pi$ Initialization of $\pi_{i}^{h}$ Suitable Game
-Vanilla Self-play [108] Equ. (5) $\emptyset$ Traditional Self-play $E=1$ Lazy General Stage Game
-Fictitious Self-play [50] Equ. (6)
-$\delta$-Uniform Self-play [5] Equ. (7)
-PFSP [129] Algo. 5
-DO [92] NE-based (Algo. 6) $\emptyset$ PSRO Series $E=1$ Lazy General Stage Game
-PSRO [73] General
-$\alpha$-PSRO [96] $\alpha$-rank-based
-JPSRO [86] (C)CE-based
-R-NaD [105] Equ. (5)
-FTW [60] Equ. (10) $\emptyset$ Ongoing-training-based $E>1$ Immediate $\pi_{i}^{h}\leftarrow\pi_{i}(\cdot|h(i))$ Stage Game
-NeuPL [82] General $\sigma_{i}$
-Simplex-NeuPL [80] General $\sigma_{i}$
-Vanilla CFR [152] Equ. (5) $R^{i}(I,a)$ Regret-minimization-based $E=1$ Lazy $\pi_{i}^{h}\leftarrow\pi_{i-1}(\cdot|h(i-1))$ Repeated Game
-CFR+ [122] $R^{+,i}(I,a)$
-CFR-D [21] $R^{i}(I_{*},a)$
-RCFR[133] $\varphi(I,a)$
-Deep CFR[133] $V(I,a|\theta_{p})$
+Vanilla Self-play Equ. (5) $\emptyset$ Traditional Self-play $E=1$ Lazy General Stage Game
+FSP Equ. (6)
+$\delta$-Uniform Self-play Equ. (7)
+PFSP Algo. 5
+DO NE-based (Algo. 6) $\emptyset$ PSRO Series $E=1$ Lazy General Stage Game
+PSRO General
+$\alpha$-PSRO $\alpha$-rank-based
+JPSRO (C)CE-based
+R-NaD Equ. (5)
+FTW Equ. (10) $\emptyset$ Ongoing-training-based $E>1$ Immediate $\pi_{i}^{h}\leftarrow\pi_{i}(\cdot|h(i))$ Stage Game
+NeuPL General $\sigma_{i}$
+Simplex-NeuPL General $\sigma_{i}$
+Vanilla CFR Equ. (5) $R^{i}(I,a)$ Regret-minimization-based $E=1$ Lazy $\pi_{i}^{h}\leftarrow\pi_{i-1}(\cdot|h(i-1))$ Repeated Game
+CFR+ $R^{+,i}(I,a)$
+CFR-D $R^{i}(I_{*},a)$
+RCFR $\varphi(I,a)$
+Deep CFR $V(I,a|\theta_{p})$
 3.6 Reassessment of the Framework
@@
 The primary difference between the PSRO series and traditional self-play algorithms is that the PSRO series employs more complex MSSes to handle tasks with intricate game-theoretical requirements.
-For example, $\alpha$-PSRO [96] specifically utilizes an $\alpha$-rank-based MSS to tackle multi-player general-sum games.
+For example, $\alpha$-PSRO (Muller et al., 2020) specifically utilizes an $\alpha$-rank-based MSS to tackle multi-player general-sum games.
 In other tasks, traditional self-play algorithms are more commonly used to reduce the computational cost associated with complex MSSes in the PSRO series.
@@
 3.6.2 Comparing with Existing Frameworks
-Our framework is built upon PSRO [73] and NeuPL [82].
+Our framework is built upon PSRO (Lanctot et al., 2017) and NeuPL (Liu et al., 2022c).
 Here, we outline the differences between our framework and these existing frameworks.
@@
 4.1 Board Games
-Board games, the majority of which are perfect information games, were previously revolutionized by the introduction of two essential techniques: position evaluation and Monte Carlo tree search (MCTS) [29, 67].
-These methodologies, with minor modifications, demonstrated superhuman effectiveness in solving board games such as chess [24], checkers [110], othello [23], backgammon [126], and Scrabble [114].
-In contrast, the application of these techniques to the game of Go with an estimated $2.1\times 10^{170}$ legal board configurations, only enabled performance at the amateur level [9, 30, 6, 37, 43].
+Board games, the majority of which are perfect information games, were previously revolutionized by the introduction of two essential techniques: position evaluation and Monte Carlo tree search (MCTS) (Coulom, 2006; Kocsis & Szepesvári, 2006).
+These methodologies, with minor modifications, demonstrated superhuman effectiveness in solving board games such as chess (Campbell et al., 2002), checkers (Schaeffer et al., 1992), othello (Buro, 1999), backgammon (Tesauro & Galperin, 1996), and Scrabble (Sheppard, 2002).
+In contrast, the application of these techniques to the game of Go with an estimated $2.1\times 10^{170}$ legal board configurations, only enabled performance at the amateur level (Bouzy & Helmstetter, 2004; Coulom, 2007; Baudiš & Gailly, 2011; Enzenberger et al., 2010; Gelly & Silver, 2007).
 In light of this, our discussion will specifically focus on the game of Go to illustrate the application of self-play.
@@
 Go is an ancient board game, which is played on a grid of 19x19 lines, where two players alternately place black and white stones aiming to control the largest territory.
-The paradigm of Go artificial intelligence (AI) is revolutionized with the launch of DeepMind’s AlphaGo series [115, 117, 116, 112], which leveraged the power of self-play to significantly elevate performance.
-In AlphaGo [115], the training regime can be split into three stages.
+The paradigm of Go artificial intelligence (AI) is revolutionized with the launch of DeepMind’s AlphaGo series (Silver et al., 2016; 2017; 2018; Schrittwieser et al., 2020), which leveraged the power of self-play to significantly elevate performance.
+In AlphaGo (Silver et al., 2016), the training regime can be split into three stages.
 In the first stage, supervised learning with expert data trains a fast policy network $p_{\pi}(a|s)$ for rollouts in MCTS and a precise policy network $p_{\sigma}(a|s)$.
@@
 In the third stage, MCTS integrates the policy and value networks to select actions.
-Based on AlphaGo, AlphaGo Zero [117] does not require any expert data except game rules.
+Based on AlphaGo, AlphaGo Zero (Silver et al., 2017) does not require any expert data except game rules.
 It utilizes only one network $f_{\theta}(s)$ to concurrently predict the action probabilities and the state value.
@@
 A new policy to be incorporated into the policy pool must surpass a 55 percent win rate against its predecessor.
-AlphaZero [116] extends AlphaGo Zero to include games beyond Go, such as Chess and Shogi.
+AlphaZero (Silver et al., 2018) extends AlphaGo Zero to include games beyond Go, such as Chess and Shogi.
 Concerning the self-play procedure, the only difference between AlphaZero and AlphaGo Zero is that AlphaZero utilizes the newly updated network to generate samples by self-play without validation.
-Building upon AlphaZero, MuZero [112] takes learning from scratch to the next level, even operating without predefined game rules.
+Building upon AlphaZero, MuZero (Schrittwieser et al., 2020) takes learning from scratch to the next level, even operating without predefined game rules.
 The self-play process in MuZero operates similarly to that in AlphaZero.
@@
 Unlike most board games, which are perfect information games, Stratego, a two-player imperfect information board game, distinguishes itself by incorporating elements of memory, deduction, and bluffing.
-This complexity is further amplified by the game’s long episode length and many potential game states, estimated $10^{535}$ [104].
+This complexity is further amplified by the game’s long episode length and many potential game states, estimated $10^{535}$ (Perolat et al., 2022).
 The game is divided into two phases: the deployment phase, where players secretly arrange their units, and the game-play phase, where the objective is to deduce the opponent’s setup and capture their flag.
-The depth and computational complexity of the game remained a challenge until breakthroughs such as DeepNash [104] showed promising advances in AI’s ability to tackle it.
-DeepNash scales up evolution-theory-based self-play method R-NaD [105] (discussed in Sec.
+The depth and computational complexity of the game remained a challenge until breakthroughs such as DeepNash (Perolat et al., 2022) showed promising advances in AI’s ability to tackle it.
+DeepNash scales up evolution-theory-based self-play method R-NaD (Perolat et al., 2021) (discussed in Sec.
 3.3.5) to neural R-NaD.
 It employs a neural network with four heads: one for value prediction, one for the deployment phase, one for piece selection, and one for piece displacement.
-Neural Replicator Dynamics [52] is utilized to obtain the approximate fixed-point policy.
+Neural Replicator Dynamics (Hennes et al., 2020) is utilized to obtain the approximate fixed-point policy.
 DeepNash holds the third-place ranking among all professional Gravon Stratego players and wins nearly every match against existing Stratego bots.
@@
 While simplified versions such as Kuhn Poker and Leduc Poker serve valuable roles in theoretical analysis, we will focus on the algorithms designed to compete in full Texas Hold’em.
-Heads-up limit Texas Hold’em (HULHE), the simplest form with approximately $3.16\times 10^{17}$ game states, was not solved until the introduction of Cepheus [10].
-It utilizes fixed-point arithmetic with compression and regret-minimization-based self-play method CFR+ [122] to address the issues of storage and computation, respectively, resulting in superhuman performance.
-Deep CFR [14] combines regret-minimization-based self-play method CFR with deep neural networks.
+Heads-up limit Texas Hold’em (HULHE), the simplest form with approximately $3.16\times 10^{17}$ game states, was not solved until the introduction of Cepheus (Bowling et al., 2015).
+It utilizes fixed-point arithmetic with compression and regret-minimization-based self-play method CFR+ (Tammelin, 2014) to address the issues of storage and computation, respectively, resulting in superhuman performance.
+Deep CFR (Brown et al., 2019) combines regret-minimization-based self-play method CFR with deep neural networks.
 Furthermore, some studies do not adopt the regret-minimization-based series of self-play; instead, they use the traditional self-play method.
-NSFP [51] introduces a self-play method combined with end-to-end RL training, also achieving competitive performance in HULHE.
-Poker-CNN [140] utilizes self-play to learn convolutional networks to solve a video version of HULHE.
-After solving HULHE, research focus shifts to heads-up no-limit Texas Hold’em (HUNL) with significantly larger game states, approximately $10^{164}$ [63].
+NSFP (Heinrich & Silver, 2016) introduces a self-play method combined with end-to-end RL training, also achieving competitive performance in HULHE.
+Poker-CNN (Yakovenko et al., 2016) utilizes self-play to learn convolutional networks to solve a video version of HULHE.
+After solving HULHE, research focus shifts to heads-up no-limit Texas Hold’em (HUNL) with significantly larger game states, approximately $10^{164}$ (Johanson, 2013).
 Thus, traversing the game tree as Cepheus does in HULHE is impossible.
-DeepStack [95] is a regret-minimization-based self-play method combined with continual re-solving.
+DeepStack (Moravčík et al., 2017) is a regret-minimization-based self-play method combined with continual re-solving.
 This method focuses on a subtree of limited depth and breadth and estimates the outcomes of the furthest reaches of this subtree.
-Libratus [18] develops a blueprint strategy, leveraging an enhanced version of regret-minimization-based self-play method MCCFR [72].
-Recursive Belief-based Learning (ReBeL) [12] introduces the public belief state (PBS), transforming the imperfect-information game into a perfect-information game with continuous state space.
+Libratus (Brown & Sandholm, 2018) develops a blueprint strategy, leveraging an enhanced version of regret-minimization-based self-play method MCCFR (Lanctot et al., 2009).
+Recursive Belief-based Learning (ReBeL) (Brown et al., 2020) introduces the public belief state (PBS), transforming the imperfect-information game into a perfect-information game with continuous state space.
 ReBeL utilizes the regret-minimization-based self-play method CFR-D, combined with RL and search, to train both the value and policy networks.
-Unlike the algorithms mentioned above, AlphaHoldem [149] does not use regret-minimization-based self-play.
-Instead, it proposes a $K$-best self-play method that always selects the top $K$ agents, based on their ELO ratings [36], to generate samples.
-Pluribus [20], based on Libratus, addresses six-player no-limit Texas Hold’em.
-Similar to Libratus, Pluribus utilizes a regret-minimization-based self-play method MCCFR [72] to develop its blueprint strategy.
+Unlike the algorithms mentioned above, AlphaHoldem (Zhao et al., 2022a) does not use regret-minimization-based self-play.
+Instead, it proposes a $K$-best self-play method that always selects the top $K$ agents, based on their ELO ratings (Elo & Sloan, 1978), to generate samples.
+Pluribus (Brown & Sandholm, 2019b), based on Libratus, addresses six-player no-limit Texas Hold’em.
+Similar to Libratus, Pluribus utilizes a regret-minimization-based self-play method MCCFR (Lanctot et al., 2009) to develop its blueprint strategy.
 It conducts a depth-limited search before executing actions.
@@
 The essence of the game lies in cooperation among peasants and competition between the two factions.
-It has an estimated $10^{76}\sim 10^{106}$ possible game states [147] and an action space comprising $27472$ possible moves [141].
-DeltaDou [61] is the first algorithm to achieve expert-level performance in DouDizhu.
-Similar to AlphaZero [116], DeltaDou utilizes vanilla self-play to generate samples to train policy and value networks.
-DouZero [148] also uses vanilla self-play to generate data.
+It has an estimated $10^{76}\sim 10^{106}$ possible game states (Zha et al., 2019) and an action space comprising $27472$ possible moves (Yang et al., 2022).
+DeltaDou (Jiang et al., 2019) is the first algorithm to achieve expert-level performance in DouDizhu.
+Similar to AlphaZero (Silver et al., 2018), DeltaDou utilizes vanilla self-play to generate samples to train policy and value networks.
+DouZero (Zha et al., 2021) also uses vanilla self-play to generate data.
 It reduces training costs by opting for a sampling method over the tree search approach.
-Based on DouZero, DouZero+ [150] incorporates opponent modeling, assisting the agent in making more informed decisions.
-PerfectDou [141] utilizes vanilla self-play to train a Perfect-Training-Imperfect-Execution framework.
+Based on DouZero, DouZero+ (Zhao et al., 2022b) incorporates opponent modeling, assisting the agent in making more informed decisions.
+PerfectDou (Yang et al., 2022) utilizes vanilla self-play under a Perfect-Training-Imperfect-Execution framework.
 4.2.3 Mahjong
@@
 The strategic depth and complexity of Mahjong pose significant challenges.
-Despite ongoing research, these findings have yet to reach expert human levels [71, 40, 93].
-Difficulties are navigating incomplete information, dynamically adapting strategies to multiple opponents, and contending with complex winning rules and an enormous number of possible game states, estimated at $10^{169}$ [147].
-Suphx [77] is recognized as one of the first algorithms to master Mahjong, achieving a performance level comparable to expert human players, precisely 10 dan on Tenhou [127], the most popular online Mahjong platform.
+Despite ongoing research, these findings have yet to reach expert human levels (Kurita & Hoki, 2020; Gao et al., 2019; Mizukami & Tsuruoka, 2015).
+Difficulties are navigating incomplete information, dynamically adapting strategies to multiple opponents, and contending with complex winning rules and an enormous number of possible game states, estimated at $10^{169}$ (Zha et al., 2019).
+Suphx (Li et al., 2020) is recognized as one of the first algorithms to master Mahjong, achieving a performance level comparable to expert human players, precisely 10 dan on Tenhou (Tsunoda, ), the most popular online Mahjong platform.
 Initially, Suphx employs supervised learning, utilizing expert data to train its model.
 It then advances its capabilities through vanilla self-play combined with RL.
-Similarly, NAGA [128], developed by Dwango Media Village, and LuckyJ [125], designed by Tencent, have also achieved the rank of 10 dan on Tenhou.
+Similarly, NAGA (Village, ), developed by Dwango Media Village, and LuckyJ (Tencent, ), designed by Tencent, have also achieved the rank of 10 dan on Tenhou.
 Furthermore, LuckyJ has even defeated human professional players.
@@
 In contrast to traditional board games and card games, video games often feature real-time actions, long trajectories, and increased complexity stemming from a broader range of actions and observations.
-Here, we illustrate some representative video games to demonstrate how self-play has advanced AI within these games.
+We illustrate representative video games showcasing self-play’s impact.
 4.3.1 StarCraft II
@@
 Victory requires meticulous planning and tactical execution, with defeat occurring when a player loses all their buildings.
-AlphaStar [129] dominates the 1v1 mode competitions in StarCraft II and has defeated professional players.
-Its framework is similar to AlphaGo [115], initially utilizing supervised learning to train the policy with expert data.
+AlphaStar (Vinyals et al., 2019) dominates the 1v1 mode competitions in StarCraft II and has defeated professional players.
+Its framework is similar to AlphaGo (Silver et al., 2016), initially utilizing supervised learning to train the policy with expert data.
 Subsequently, it uses a hierarchical self-play method combined with end-to-end RL to train the networks.
@@
 However, the introduction of three agent types significantly increases the training computation.
-Further studies [48, 131, 57] have enhanced the league self-play training procedure.
+Further studies (Han et al., 2020; Wang et al., 2021; Huang et al., 2024) have enhanced the league self-play training procedure.
 4.3.2 MOBA Games
@@
 Managing multiple lanes and battling under the fog of war, which obscures parts of the map, are critical aspects of the gameplay.
-Popular examples of MOBA games include League of Legends, Dota 2, and Honor of Kings.
-OpenAI Five [7] defeated the world champion team in a simplified version of Dota 2 that featured a limited pool of heroes and certain banned items.
-The training process introduces a self-play method that combines two traditional self-play algorithms: with an 80% probability of engaging in vanilla self-play and a 20% probability of employing a technique similar to the PFSP used in AlphaStar [129].
+OpenAI Five (Berner et al., 2019) defeated the world champion team in a simplified version of Dota 2 that featured a limited pool of heroes and certain banned items.
+The training process introduces a self-play method that combines two traditional self-play algorithms: with an 80% probability of engaging in vanilla self-play and a 20% probability of employing a technique similar to the PFSP used in AlphaStar (Vinyals et al., 2019).
 This technique selects each policy from the policy pool based on its quality score, which is continuously updated from competition results throughout training.
@@
 Another notable MOBA game, especially viral in China, is Honor of Kings.
-The 1v1 mode is conquered by [144], which boasts a significant winning rate against top professional players.
+The 1v1 mode is conquered by Ye et al.
+(2020b), which boasts a significant winning rate against top professional players.
 It utilizes a traditional self-play method $\delta$-uniform self-play to train the RL networks.
-The 5v5 mode was later mastered by [143] as well.
+The 5v5 mode was later mastered by Ye et al.
+(2020a) as well.
 Unlike OpenAI Five, this work expands the hero pool to 40 heroes, exponentially increasing the possible combinations of lineups.
@@
 This approach defeats professional player teams.
-The self-play generated data is also used to learn compelling lineup drafting by utilizing MCTS and neural networks [25].
+The self-play generated data is also used to learn compelling lineup drafting by utilizing MCTS and neural networks (Chen et al., 2021).
 4.3.3 Google Research Football
-Google Research Football (GRF) [70] is an open-source football simulator emphasizing high-level actions.
+Google Research Football (GRF) (Kurach et al., 2020) is an open-source football simulator emphasizing high-level actions.
 It initially offers two scenarios: the football benchmark and the football academy with 11 specific tasks.
@@
 It features long trajectories with 3000 steps per round, stochastic transitions, and sparse rewards.
-WeKick [153] claimed victory in the GRF competition on Kaggle [45], which simplifies the game dynamics by allowing competitors to control only one player, either the ball carrier on offense or the nearest defender on defense.
-WeKick employed self-play strategies similar to those used in league training [129].
-It initializes its opponent policy pool using strategies developed through RL and Generative Adversarial Imitation Learning (GAIL) [55] to facilitate the training process.
+WeKick (Ziyang Li, 2020) claimed victory in the GRF competition on Kaggle (Google, 2020), which simplifies the game dynamics by allowing competitors to control only one player, either the ball carrier on offense or the nearest defender on defense.
+It employed self-play strategies similar to those used in league training (Vinyals et al., 2019).
+It initializes its opponent policy pool using strategies developed through RL and Generative Adversarial Imitation Learning (GAIL) (Ho & Ermon, 2016) to facilitate the training process.
 Further research delves into the full football game rather than the simplified version.
-Team-PSRO [91] extends PSRO series of self-play algorithms to team games, outperforming baselines in the 4v4 version of the full GRF.
-In the context of the 11v11 version, where the goalkeeper is rule-based controlled, TiKick [58] utilizes vanilla self-play to collect samples and then employs imitation learning.
-Fictitious Cross-Play (FXP) [137] proposes a new self-play method similar to AlphaStar [129].
+Team-PSRO (McAleer et al., 2023) extends PSRO series of self-play algorithms to team games, outperforming baselines in the 4v4 version of the full GRF.
+In the context of the 11v11 version, where the goalkeeper is rule-based controlled, TiKick (Huang et al., 2021a) utilizes vanilla self-play to collect samples and then employs imitation learning.
+Fictitious Cross-Play (FXP) (Xu et al., 2023) proposes a new self-play method similar to AlphaStar (Vinyals et al., 2019).
 It introduces two populations: the main population and the counter population.
@@
 FXP achieves a win rate of over 94% against TiKick.
-TiZero [78], a follow-up to TiKick, combines curriculum learning with FSP and PFSP [129] to avoid expert data reliance, achieving a higher TrueSkill rating [53] than TiKick.
+TiZero (Lin et al., 2023), a follow-up to TiKick, combines curriculum learning with FSP and PFSP (Vinyals et al., 2019) to avoid expert data reliance, achieving a higher TrueSkill rating (Herbrich et al., 2006) than TiKick.
 Table 2: Overview of representative studies in empirical analysis.
@@
 Players Information Game States
-Board Game Chess 2 $\checkmark$ $10^{45}$ AlphaZero [116] Vanilla SP $\checkmark$ $\times$
-MuZero [112] Vanilla SP $\checkmark$ $\times$
-Go 2 $\checkmark$ $10^{360}$ AlphaGo [115] FSP $\checkmark$ $\checkmark$
-AlphaGo Zero [117] Vanilla SP $\checkmark$ $\times$
-AlphaZero [116] Vanilla SP $\checkmark$ $\times$
-MuZero [112] Vanilla SP $\checkmark$ $\times$
-Stratego 2 $\times$ $10^{535}$ DeepNash [104] R-NaD $\times$ $\times$
-Card Game HULHE 2 $\times$ $10^{17}$ Cepheus [10] CFR+ $\checkmark$ $\times$
-NFSP [51] NFSP $\times$ $\times$
-HUNL 2 $\times$ $10^{164}$ DeepStack [95] CFR+ $\checkmark$ $\times$
-Libratus [18] MCCFR $\checkmark$ $\times$
-ReBel [12] CFR-D / FSP $\checkmark$ $\times$
-Six-player No-limit Texas Hold’em 6 $\times$ $>10^{164}$ Pluribus [20] MCCFR $\checkmark$ $\times$
-DouDiZhu 3 $\times$ $10^{76}\sim 10^{106}$ DeltaDou [61] Vanilla SP $\checkmark$ $\times$
-DouZero [148] Vanilla SP $\times$ $\checkmark$
-PerfectDou [141] Vanilla SP $\times$ $\times$
-Mahjong 4 $\times$ $10^{169}$ Suphx [77] Vanilla SP $\checkmark$ $\checkmark$
-Video Game StarCraft II 2 $\times$ / AlphaStar [129] FSP & PFSP $\times$ $\checkmark$
-Dota 2 10 $\times$ OpenAI Five [7] Vanilla SP & PFSP $\times$ $\times$
-Honor of Kings 10 $\times$ MOBA AI [143] $\delta$-uniform SP $\checkmark$ $\checkmark$
-Google Research Football 22 $\checkmark$ TiZero [78] FSP & PFSP $\times$ $\times$
+Board Game Chess 2 $\checkmark$ $10^{45}$ AlphaZero Vanilla SP $\checkmark$ $\times$
+MuZero Vanilla SP $\checkmark$ $\times$
+Go 2 $\checkmark$ $10^{360}$ AlphaGo FSP $\checkmark$ $\checkmark$
+AlphaGo Zero Vanilla SP $\checkmark$ $\times$
+AlphaZero Vanilla SP $\checkmark$ $\times$
+MuZero Vanilla SP $\checkmark$ $\times$
+Stratego 2 $\times$ $10^{535}$ DeepNash R-NaD $\times$ $\times$
+Card Game HULHE 2 $\times$ $10^{17}$ Cepheus CFR+ $\checkmark$ $\times$
+NFSP NFSP $\times$ $\times$
+HUNL 2 $\times$ $10^{164}$ DeepStack CFR+ $\checkmark$ $\times$
+Libratus MCCFR $\checkmark$ $\times$
+ReBel CFR-D / FSP $\checkmark$ $\times$
+Six-player No-limit Texas Hold’em 6 $\times$ $>10^{164}$ Pluribus MCCFR $\checkmark$ $\times$
+DouDiZhu 3 $\times$ $10^{76}\sim 10^{106}$ DeltaDou Vanilla SP $\checkmark$ $\times$
+DouZero Vanilla SP $\times$ $\checkmark$
+PerfectDou Vanilla SP $\times$ $\times$
+Mahjong 4 $\times$ $10^{169}$ Suphx Vanilla SP $\checkmark$ $\checkmark$
+Video Game StarCraft II 2 $\times$ / AlphaStar FSP & PFSP $\times$ $\checkmark$
+Dota 2 10 $\times$ OpenAI Five Vanilla SP & PFSP $\times$ $\times$
+Honor of Kings 10 $\times$ MOBA AI $\delta$-uniform SP $\checkmark$ $\checkmark$
+Google Research Football 22 $\checkmark$ TiZero FSP & PFSP $\times$ $\times$
 5 Open Problems and Future Work
@@
 5.1 Theoretical Foundation
-Although NE has been shown to exist in games with finite players and finite actions [100], computing NE with self-play algorithms in larger games remains challenging and consequently, many studies aim to achieve approximate NE [76].
-However, in some cases, even computing an approximate NE is difficult [31].
-Some research has resorted to higher levels of equilibrium, such as CE [86] and $\alpha$-rank [96].
+Although NE has been shown to exist in games with finite players and finite actions (Nash et al., 1950), computing NE with self-play algorithms in larger games remains challenging and consequently, many studies aim to achieve approximate NE (Li et al., 2024).
+However, in some cases, even computing an approximate NE is difficult (Daskalakis, 2013).
+Some research has resorted to higher levels of equilibrium, such as CE (Marris et al., 2021) and $\alpha$-rank (Muller et al., 2020).
 Although many algorithms are developed with theoretical foundations, there is often a gap in complex real-world scenarios.
-For example, although self-play algorithms like AlphaGo [115], AlphaStar [109], and OpenAI Five [7] achieve empirical success, they lack formal game theory proofs behind their effectiveness.
+For example, although self-play algorithms like AlphaGo (Silver et al., 2016), AlphaStar (Samvelyan et al., 2019), and OpenAI Five (Berner et al., 2019) achieve empirical success, they lack formal game theory proofs behind their effectiveness.
 Future work should focus on bridging this gap by developing new theoretically sound and practically effective algorithms or proving the theoretical underpinnings of existing successful self-play algorithms in complex environments.
@@
 Future research should aim to develop self-play algorithms that are more robust and can adapt to changing conditions.
-For example, incorporating opponent modeling [150] into self-play can help agents anticipate changes in opponent strategies and adjust their own strategies proactively, making them more robust to environmental changes.
+For example, incorporating opponent modeling (Zhao et al., 2022b) into self-play can help agents anticipate changes in opponent strategies and adjust their own strategies proactively, making them more robust to environmental changes.
 5.3 Scalability and Training Efficiency
@@
 As the number of participants grows, the complexity of interactions explodes.
-For example, in OpenAI Five [7], the hero pool size is limited to only 17 heroes.
-MOBA AI [143] extends this to a 40-hero pool with the help of curriculum learning, but it still cannot cover the entire hero pool available in the actual game.
+For example, in OpenAI Five (Berner et al., 2019), the hero pool size is limited to only 17 heroes.
+MOBA AI (Ye et al., 2020a) extends this to a 40-hero pool with the help of curriculum learning, but it still cannot cover the entire hero pool available in the actual game.
 One potential solution is leveraging players’ inherent connections to optimize the learning process.
@@
 The first issue is computational efficiency, which is induced by the iterative nature of self-play where agents repeatedly play against themselves or past versions.
-Moreover, although forming more complex populations and competitive mechanisms [109] can enhance the intensity and quality of training, it further exacerbates the demand for computational resources.
+Moreover, although forming more complex populations and competitive mechanisms (Samvelyan et al., 2019) can enhance the intensity and quality of training, it further exacerbates the demand for computational resources.
 Techniques such as parallel computing, distributed learning, and more efficient neural network architectures could be explored to address these challenges.
@@
 5.4 With Large Language Models
-With their remarkable capability and emergent generalizability, large language models (LLMs) have been regarded as a potential foundation for achieving human-level intelligence [1], and self-play methods have been proposed to fine-tune LLMs, enhance LLMs’ reasoning performance, and build LLM-based agents with strong decision-making abilities.
-Post-training fine-tuning [102, 3] is a key step in aligning LLM with more desired behaviors, but it requires a huge amount of human-annotated data.
-To reduce reliance on human-annotated data, Self-play fIne-tuNing (SPIN) [26] introduces a self-play mechanism to generate training data using the LLM itself and fine-tuning the LLM by distinguishing self-generated responses from human-annotated data.
-Another line of work [98, 121, 135] formulate the alignment problem as a two-player constant-sum game and use self-play methods to solve the game.
-Some other work [56, 146] also utilizes model-generated data to fine-tune LLMs with minimal human annotations.
+With their remarkable capability and emergent generalizability, large language models (LLMs) have been regarded as a potential foundation for achieving human-level intelligence (Achiam et al., 2023), and self-play methods have been proposed to fine-tune LLMs, enhance LLMs’ reasoning performance, and build LLM-based agents with strong decision-making abilities.
+Post-training fine-tuning (Ouyang et al., 2022; Bai et al., 2022) is a key step in aligning LLM with more desired behaviors, but it requires a huge amount of human-annotated data.
+To reduce reliance on human-annotated data, Self-play fIne-tuNing (SPIN) (Chen et al., 2024) introduces a self-play mechanism to generate training data using the LLM itself and fine-tuning the LLM by distinguishing self-generated responses from human-annotated data.
+Another line of work (Munos et al., 2024; Swamy et al., 2024; Wu et al., 2024) formulate the alignment problem as a two-player constant-sum game and use self-play methods to solve the game.
+Some other work (Huang et al., 2022; Yuan et al., 2024) also utilizes model-generated data to fine-tune LLMs with minimal human annotations.
 The idea of self-improvement has also been applied to improve the reasoning ability of LLMs.
-Self-Play of Adversarial Game (SPAG) [27] observes that self-play in a two-player adversarial language game called Adversarial Taboo can boost the LLM’s performance on various reasoning benchmarks.
+Self-Play of Adversarial Game (SPAG) (Cheng et al., 2025) observes that self-play in a two-player adversarial language game called Adversarial Taboo can boost the LLM’s performance on various reasoning benchmarks.
 Besides improving the capability of LLMs, self-play methods have also contributed to building LLM-based agents with strong strategic abilities.
-A representative work is Cicero [38] which achieves human-level play in the game of Diplomacy by combining language models with an RL policy trained by self-play.
+A representative work is Cicero († et al.(2022)(FAIR)†, Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al., FAIR) which achieves human-level play in the game of Diplomacy by combining language models with an RL policy trained by self-play.
 Cicero uses the self-play policy to produce an intended action and prompts the language model to generate languages conditioned on the policy’s intent.
-Another work [138] also combines LLM with self-play policy but takes a different approach by first prompting the LLM to propose multiple action candidates and then using the self-play policy to produce the action distribution over these candidates.
+Another work (Xu et al., 2024) also combines LLM with self-play policy but takes a different approach by first prompting the LLM to propose multiple action candidates and then using the self-play policy to produce the action distribution over these candidates.
 Despite recent progress, applying self-play with LLMs is still underexplored and requires further research.
 5.5 Real-World Applications
-Self-play is a powerful technique with widespread applications across various domains.
-It is particularly effective in addressing some problems abstracted from real-world situations through its iterative learning approach.
-In the field of economics, for instance, self-play is used to enhance supervised learning models in multi-issue bargaining tasks [74].
-Furthermore, self-play proves beneficial in solving combinatorial optimization problems (COPs) such as the traveling salesman problem (TSP) and the capacitated vehicle routing problem (CVRP) [130].
-Within the domain of traffic, self-play facilitates the development of human-like autonomous driving behaviors [28] and enables vehicles to learn negotiation strategies to merge on or off roads [123], although currently within 2D simulators.
+Self-play is particularly effective in addressing some problems abstracted from real-world situations through its iterative learning approach.
+In the field of economics, for instance, self-play is used to enhance supervised learning models in multi-issue bargaining tasks (Lewis et al., 2017).
+Furthermore, self-play proves beneficial in solving combinatorial optimization problems (COPs) such as the traveling salesman problem (TSP) and the capacitated vehicle routing problem (CVRP) (Wang et al., 2024).
+Within the domain of traffic, self-play facilitates the development of human-like autonomous driving behaviors (Cornelisse & Vinitsky, 2024) and enables vehicles to learn negotiation strategies to merge on or off roads (Tang, 2019), although currently within 2D simulators.
 However, a notable challenge with self-play is its impracticality for direct application in real-world scenarios.
@@
 The effective deployment of self-play in real-world applications boils down to overcoming the Sim2Real gap.
-For example, for tasks where the Sim2Real gap is less pronounced, self-play can be well employed to enhance video streaming capabilities [59], and to address the image retargeting problem [66].
-Moreover, EvoPlay [132] leverages self-play to design protein sequences, utilizing the advanced AlphaFold2 simulator [65] to narrow the Sim2Real gap.
-Similarly, in heterogeneous multi-robot systems, self-play is utilized for competitive sports tasks like humanoid football [81, 47], quadruped competition [136], and multi-drone volleyball [139], with significant efforts dedicated to Sim2Real transitions for real-world success [41].
+For example, for tasks where the Sim2Real gap is less pronounced, self-play can be well employed to enhance video streaming capabilities (Huang et al., 2021b), and to address the image retargeting problem (Kajiura et al., 2020).
+Moreover, EvoPlay (Wang et al., 2023) leverages self-play to design protein sequences, utilizing the advanced AlphaFold2 simulator (Jumper et al., 2021) to narrow the Sim2Real gap.
+Similarly, in heterogeneous multi-robot systems, self-play is utilized for competitive sports tasks like humanoid football (Liu et al., 2022b; Haarnoja et al., 2024), quadruped competition (Xiong et al., 2024), and multi-drone volleyball (Xu et al., 2025), with significant efforts dedicated to Sim2Real transitions for real-world success (Gao et al., 2023).
 6 Conclusion
@@
 This survey serves as an essential guide for researchers and practitioners, paving the way for further advancements in this field.
-Acknowledgment
-This research was supported by National Natural Science Foundation of China (No.62406159, 62325405), Postdoctoral Fellowship Program of CPSF under Grant Number (GZC20240830, 2024M761676), China Postdoctoral Science Special Foundation 2024T170496.
 Appendix A Appendix

### v4

@@
 A Survey on Self-Play Methods in Reinforcement Learning
-Abstract
-Self-play, characterized by agents’ interactions with copies or past versions of themselves, has recently gained prominence in reinforcement learning (RL).
-This paper first clarifies the preliminaries of self-play, including the multi-agent reinforcement learning framework and basic game theory concepts.
+Abstract.
+Self-play, a learning paradigm where agents iteratively refine their policies by interacting with historical or concurrent versions of themselves or other evolving agents, has shown remarkable success in solving complex non-cooperative multi-agent tasks.
+Despite its growing prominence in multi-agent reinforcement learning (MARL), such as Go, poker, and video games, a comprehensive and structured understanding of self-play remains lacking.
+This survey fills this gap by offering a comprehensive roadmap to the diverse landscape of self-play methods.
+We begin by introducing the necessary preliminaries, including the MARL framework and basic game theory concepts.
 Then, it provides a unified framework and classifies existing self-play algorithms within this framework.
-Moreover, the paper bridges the gap between the algorithms and their practical implications by illustrating the role of self-play in different scenarios.
+Moreover, the paper bridges the gap between the algorithms and their practical implications by illustrating the role of self-play in different non-cooperative scenarios.
 Finally, the survey highlights open challenges and future research directions in self-play.
-This paper is an essential guide map for understanding the multifaceted landscape of self-play in RL.
-11footnotetext: Corresponding authors.
-1 Introduction
-Reinforcement learning (RL) represents a significant paradigm (Sutton & Barto, 2018) within machine learning (ML), focused on optimizing decision processes through interaction with an environment.
+Self-play, reinforcement learning, game theory
+† Corresponding authors.
+This research was supported by National Natural Science Foundation of China (No.62406159, 62325405), Postdoctoral Fellowship Program of CPSF under Grant Number (GZC20240830, 2024M761676), China Postdoctoral Science Special Foundation 2024T170496.
+copyright: noneccs: Computing methodologies Multi-agent systemsccs: General and reference Surveys and overviews
+1. Introduction
+Figure 1. Overview of our survey.
+Reinforcement learning (RL) represents a significant paradigm (Sutton and Barto, 2018) within machine learning (ML), focused on optimizing decision processes through interaction with an environment.
 In RL, the environment is modeled as a Markov decision process (MDP), where agents observe states, take actions, receive rewards, and cause transitions to new states.
@@
 Deep RL extends traditional RL by employing deep neural networks as function approximators to handle high-dimensional state spaces, contributing to breakthroughs in various complex tasks (Mnih et al., 2013).
-Moreover, transitioning from single-agent to multi-agent reinforcement learning (MARL) introduces complex dynamics (Rashid et al., 2018; Mahajan et al., 2019; Yu et al., 2022).
+Transitioning from single-agent to multi-agent reinforcement learning (MARL) introduces complex dynamics (Rashid et al., 2018; Mahajan et al., 2019; Yu et al., 2022).
 In MARL, the interdependence of agents’ actions presents significant challenges, as the environment appears non-stationary to each agent.
-The main issues in MARL are coordination, communication, and equilibrium selection, particularly in competitive scenarios.
-These challenges often lead to difficulties in achieving convergence, maintaining stability, and efficiently exploring the solution space.
+In the context of MARL, environments can generally be categorized into cooperative and non-cooperative settings.
+In cooperative scenarios, all agents share a common objective and are incentivized to work together, often leading to formulations that maximize a joint reward.
+In contrast, non-cooperative settings involve agents with individual or even conflicting goals, where the success of one agent may come at the expense of others.
+In MARL, non-cooperative tasks are generally harder than cooperative ones because the absence of a shared reward prevents the effective use of centralized training, a key technique for mitigating non-stationarity in cooperative settings (Foerster et al., 2016, 2018; Lowe et al., 2017; Yu et al., 2022), thereby further aggravating non-stationarity.
 With the help of game theory, a mathematical framework that models the interactions between multiple decision-makers, self-play emerges as an elegant solution to some inherent challenges in MARL.
-Self-play provides an approach where an agent interacts with copies or past versions of itself (Samuel, 2000; Bansal et al., 2018).
-This method promises a more stable and manageable learning process.
-The capabilities of self-play extend to a wide range of scenarios, including its high-profile applications in Go (Silver et al., 2016; 2017; 2018; Schrittwieser et al., 2020), chess (Silver et al., 2018; Schrittwieser et al., 2020), poker (Moravčík et al., 2017; Heinrich et al., 2015), and video games (Berner et al., 2019; Vinyals et al., 2019).
+Vanilla self-play refers to a setting in which an agent continuously trains by competing against the most recent version of itself, as famously demonstrated for the symmetric zero-sum game of checkers (Samuel, 1959).
+Self-play, however, encompasses far more than this canonical example.
+More broadly, self-play refers to a class of learning methods where at least one agent iteratively improves its policy by interacting with a distribution of evolving opponents, which may include historical or concurrent versions of itself or other agents.
+It is worth noting that although some MARL studies have applied the concept of self-play in cooperative tasks such as Hanabi (Cui et al., 2021) and Overcooked (Strouse et al., 2021), our survey is conducted within non-cooperative settings.
+Self‑play has demonstrated remarkable success in non-cooperative domains such as Go (Silver et al., 2016, 2017, 2018; Schrittwieser et al., 2020), chess (Silver et al., 2018; Schrittwieser et al., 2020), poker (Moravčík et al., 2017; Heinrich et al., 2015), and complex video games (Berner et al., 2019; Vinyals et al., 2019).
 In these scenarios, it has developed strategies that surpass human expertise.
-Although the application of self-play is extensive and promising, it is accompanied by limitations, such as the potential convergence to suboptimal strategies and significant computational requirements(Silver et al., 2016; 2018).
-Although some research adopts a broad perspective through empirical game-theoretic analysis (EGTA)(Wellman et al., 2025), relatively few comprehensive surveys focus solely on self-play algorithms.
-Among them, a study develops an algorithmic framework for self-play, but it does not incorporate the Policy-Space Response Oracle (PSRO) series of algorithms (Hernandez et al., 2021), while a separate study focuses exclusively on PSRO (Bighashdel et al., 2024), without considering other self-play algorithms.
-Another study examines the theoretical safety of self-play (DiGiovanni & Zell, 2021).
+Although the application of self-play is extensive and promising, it is accompanied by limitations, such as the potential convergence to suboptimal strategies and significant computational requirements (Silver et al., 2016, 2018).
+Although some research (Wellman et al., 2025) adopts a different perspective through empirical game-theoretic analysis (EGTA), relatively few comprehensive surveys focus on self-play.
+Among them, a study (Hernandez et al., 2021) develops an algorithmic framework for self-play, but it does not incorporate the Policy-Space Response Oracle (PSRO) (Lanctot et al., 2017) series of algorithms, while a separate study (Bighashdel et al., 2024) focuses exclusively on PSRO, without considering other self-play algorithms.
+Another study (DiGiovanni and Zell, 2021) examines the theoretical safety of self-play.
 While these studies are valuable, they do not provide a comprehensive perspective that fully captures the breadth and depth of self-play.
-Therefore, this survey aims to fill this gap.
+To address this gap, this survey aims to offer a systematic roadmap to the diverse landscape of self-play algorithms in MARL within non-cooperative settings.
 The survey is organized as follows.
 Sec.
-2 introduces the background of self-play, including the RL framework and game theory concepts.
+2 introduces the background of self-play, including the MARL framework and game theory concepts.
 Sec.
@@
 1, which further illustrates the relationships among the subsections.
-Figure 1: Overview of our survey.
-2 Preliminaries
-This section first introduces the framework of RL, where an agent learns to make decisions by interacting with an environment to maximize cumulative rewards.
-We then present basic game theory concepts, which analyze strategic interactions between rational decision-makers and provide tools to study outcomes based on players’ strategies.
-2.1 RL Framework
+2. Preliminaries
+This section first introduces the framework of MARL, where agents interact with an environment to learn policies that maximize cumulative rewards.
+We then present key concepts from non-cooperative game theory, which provides tools to analyze strategic interactions among rational decision-makers with individual objectives.
+2.1. MARL Framework
 In RL, the environment is modeled as an MDP, where the Markovian assumption states that the environment’s evolution is fully determined by its current state, eliminating the need to consider past states.
@@
 MDPs can be extended to multi-agent settings, known as Markov games (MGs) (Littman, 1994) or stochastic games (Shapley, 1953).
-We focus on the most general form: partially observable Markov games (POMGs), where multiple agents interact with the environment, and each agent receives only individual observations instead of the full state, meaning they have limited access to the environment’s state.
+We focus on the most general form: partially observable Markov games (POMGs), where multiple agents interact with the environment, and each agent receives only individual observations instead of the full state, thus having limited access to the environment’s state.
 A POMG $\mathcal{G}$ is defined by $\mathcal{G}=(\mathcal{N},\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{P},\mathcal{R},\gamma,\rho)$.
@@
 $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ denotes the transition probability from one state to another given the actions of each agent.
-=$\mathcal{R}=\{\mathcal{R}_{1},\cdots,\mathcal{R}_{n}\}$, where $\mathcal{R}_{i}:\mathcal{S}\times\mathcal{A}_{i}\rightarrow\mathbb{R}$ denotes the reward function of agent $i$.
+$\mathcal{R}=\{\mathcal{R}_{1},\cdots,\mathcal{R}_{n}\}$, where $\mathcal{R}_{i}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ denotes the reward function of agent $i$.
 $\gamma\in[0,1]$ is the discount factor.
 $\rho:\mathcal{S}\rightarrow[0,1]$ describes initial state distribution.
-If it is a fully cooperative setting, agents can share the same reward function (Yu et al., 2022; Rashid et al., 2020; Foerster et al., 2018; Kuba et al., 2021).
-When $n=1,\mathcal{O}_{i}=\mathcal{S}$, the POMG returns to the MDP.
-More concretely, in RL, agents interact with the environment through the following procedure: At each discrete time step $t$, each agent $i$ receives an observation $o_{i,t}$ from the environment and selects an action based on a stochastic policy $\pi_{\theta_{i}}:\mathcal{O}_{i}\times\mathcal{A}_{i}\rightarrow[0,1]$, where $\theta_{i}$ represents the parameters.
+More concretely, in RL, agents interact with the environment through the following procedure: At each discrete time step $t$, each agent $i$ receives an observation $o_{i,t}$ from the environment and selects an action $a_{i,t}$ based on a stochastic policy $\pi_{i}:\mathcal{O}_{i}\times\mathcal{A}_{i}\rightarrow[0,1]$.
 After receiving the joint actions $\mathbf{a}_{t}=(a_{1,t},\cdots,a_{n,t})$, the environment transitions from the current state $s_{t}$ to a subsequent state $s_{t+1}$ according to the transition function $\mathcal{P}$ and sends a reward $r_{i,t+1}$ to every agent $i$.
-The ultimate goal of agent $i$ is to maximize the expected discounted cumulative rewards: $\mathbb{E}_{\pi_{\theta_{i}}}[\sum_{t=0}^{\infty}\gamma^{t}r_{i,t}]$.
-2.2 Game Theory Concepts
-2.2.1 Perfect Information and Imperfect Information
-In a game with perfect information, only one player moves at a time.
-Each player comprehensively understands the current game state, the complete history of moves made, and the potential future outcomes.
-If these conditions are not met, the game is considered to have imperfect information (Lucchetti, 2011; Mycielski, 1992).
-An example of a game with perfect information is Go.
-In Go, both players have full access to the game state at all times, including the positions of all stones and the identity of the next player.
-In contrast, Texas Hold’em poker is a game with imperfect information.
-Players cannot see their opponents’ private cards, creating uncertainty and requiring decisions based on probabilistic reasoning, bluffing, and interpreting others’ actions.
-2.2.2 Normal-Form and Extensive-Form
+The ultimate goal of agent $i$ is to learn a policy $\pi_{i}$ that maximize the expected discounted cumulative rewards: $\mathbb{E}_{(\pi_{i},\pi_{-i})}[\sum_{t=0}^{\infty}\gamma^{t}r_{i,t+1}]$, where $\pi_{-i}$ denotes the other agents’ policies.
+2.2. Game Theory Concepts
+2.2.1. Normal-Form and Extensive-Form
 The normal form and extensive form are two distinct representations of games in game theory.
@@
 The set $\mathbf{\Pi}=\Pi_{1}\times\cdots\times\Pi_{n}$ represents the pure strategy space for all players.
-A pure strategy specifies a particular and deterministic action for a player in the game.
-A vector $\bm{\pi}=(\pi_{1},\dots,\pi_{n})\in\mathbf{\Pi}$ is called a strategy profile.
-A mixed strategy assigns a probability distribution over the set of pure strategies.
+A pure strategy specifies a deterministic action choice for each point in the game.
+A vector $\boldsymbol{\pi}=(\pi_{1},\dots,\pi_{n})\in\mathbf{\Pi}$ is called a strategy profile.
+A mixed strategy $\boldsymbol{\sigma}$ assigns a probability distribution over the set of pure strategies.
 For player $i$, a mixed strategy is represented by $\sigma_{i}\in\Delta(\Pi_{i})$, where $\Delta$ denotes the probability simplex.
-$\mathbf{u}=(u_{1},\cdots,u_{n})$, where $u_{i}:\mathbf{\Pi}\rightarrow\mathbb{R}$, is a utility function that assigns a real-valued payoff to each player $i$.
-If $\forall\bm{\pi}\in\mathbf{\Pi},\sum_{i}u_{i}(\bm{\pi})=0$, the game is a zero-sum game, otherwise it is a general-sum game.
-If $\Pi_{1}=\cdots=\Pi_{n}$ and the payoffs are invariant under any permutation of the players’ strategies, the game is a symmetric game.
-If a finite set of players (especially two players) are involved and each player has a finite set of strategies, a normal-form game can be directly depicted in a matrix.
-In two-player zero-sum symmetric normal-form games, both players share the same pure strategy space, denoted by $\Pi$, such that $\Pi=\Pi_{1}=\Pi_{2}$.
+$\mathbf{u}=(u_{1},\cdots,u_{n})$, where $u_{i}:\mathbf{\Pi}\rightarrow\mathbb{R}$, is a utility function that assigns the expected payoff to each player $i$.
+If $\exists C\in\mathbb{R},\forall\boldsymbol{\pi}\in\mathbf{\Pi},\sum_{i}u_{i}(\boldsymbol{\pi})=C$, the game is a constant-sum game.
+Specially, if $C=0$, it is a zero-sum games, in which every gain by one player is exactly offset by losses to the others.
+If no such constant $C$ exists, it is a general-sum game.
+Moreover, if $\Pi_{1}=\cdots=\Pi_{n}$ and for any permutation $\tau$ and any strategy profile $(\pi_{1},\dots,\pi_{n})$, $u_{i}(\pi_{1},\cdots,\pi_{n})=u_{\tau(i)}(\pi_{\tau(1)},\cdots,\pi_{\tau(n)})$, the game is a symmetric game.
+If a finite set of players is involved and each player has a finite set of strategies, a normal-form game can be depicted in a tensor $T\in\mathbb{R}^{|\Pi_{1}|\times\cdots\times|\Pi_{n}|\times n}$, where $|\Pi_{i}|$ is the size of player $i$’s strategy space and the final dimension indexes each player’s payoff.
+Specifically, in two-player zero-sum symmetric normal-form games, both players share the same pure strategy space, denoted by $\Pi$, such that $\Pi=\Pi_{1}=\Pi_{2}$.
 Since the utility function satisfies $u_{1}(\pi_{i},\pi_{j})=-u_{2}(\pi_{i},\pi_{j})$, we can simplify the utility to a single function $u$, where for $\pi_{i},\pi_{j}\in\Pi$, if $\pi_{i}$ beats $\pi_{j}$, then $u(\pi_{i},\pi_{j})=-u(\pi_{j},\pi_{i})>0$.
-The evaluation matrix captures the game outcomes by detailing the results of different strategies when they are played against each other: $A_{\Pi}=\{u(\pi_{i},\pi_{j}):\pi_{i},\pi_{j}\in\Pi\times\Pi\}$.
+This game can be directly depicted in an evaluation matrix $M_{\Pi}\in\mathbb{R}^{|\Pi|\times|\Pi|}$.
+It captures the game outcomes by detailing the results of different strategies when they are played against each other: $M_{\Pi}=\{u(\pi_{i},\pi_{j}):\pi_{i},\pi_{j}\in\Pi\times\Pi\}$.
 If a game is represented in the extensive-form, it is expressed sequentially, illustrating the sequence of moves, choices made by the players, and the information available to each player during decision-making.
@@
 This tree demonstrates the sequential and potentially conditional nature of decisions.
-Moreover, if player $i$ has perfect recall, it means that player $i$ remembers which action they have taken in the past.
 A game $\mathcal{G}$ represented in the extensive-form can be expressed by $\mathcal{G}=(\mathcal{N}\cup\{c\},H,Z,P,\mathcal{I},A,\mathbf{u})$.
 $\mathcal{N}=\{1,2,\cdots,n\}$ denotes a set of players.
-$c$ is chance node and can be regarded as a special player.
+$c$ is the chance node and can be regarded as a special player.
 $H$ represents a set of possible histories and $Z\subseteq H$ is a set of terminal histories.
@@
 $\mathcal{I}$ denotes information set partitions and $\mathcal{I}_{i}$ denotes the information set partitions for player $i$.
-This implies that in an imperfect information game, if player $i$ reaches a history $h\in I_{i}$, where $I_{i}\in\mathcal{I}_{i}$ is a specific information set, player $i$ cannot distinguish which particular history $h\in I_{i}$ it is encountering.
+If player $i$ reaches a history $h\in I_{i}$, where $I_{i}\in\mathcal{I}_{i}$ is a specific information set for player $i$, all decision nodes in $I_{i}$ are observationally equivalent to player $i$, and he cannot tell which specific node within $I_{i}$ has been reached.
+A game has perfect information precisely when every information set of every player is a singleton.
+Formally, $\forall i,\ \forall I_{i}\in\mathcal{I}_{i},\ |I_{i}|=1$.
+In such games, whenever a player is to move, he knows exactly which decision node has been reached and, equivalently, the full sequence of past actions and observed states.
+Otherwise the game is considered to have imperfect information.
 Action space is represented by $A(h)$ for a non-terminal history $h\in H$.
@@
 Together, these components define the structure and dynamics of an extensive-form game.
-Moreover, A strategy profile in an extensive-form game can be expressed by $\bm{\pi}=(\pi_{1},\cdots,\pi_{n})$, where $\pi_{i}$ maps each $I_{i}\in\mathcal{I}_{i}$ to a probability distribution over $A(I_{i})$.
-A subgame of an extensive-form game is a portion of the game that starts from a single initial node, includes all successors of any node within the subgame, and contains all nodes in the same information set as any node in the subgame.
-The Prisoner’s Dilemma is a classic example in game theory. The game outcomes are as follows:
-  • If one player confesses (C) and the other lies (L), the confessor will serve 1 year in jail, while the liar will serve 8 years.
-  • If both players choose to confess, they will each serve 7 years behind bars.
-  • If both players choose to lie, they will each serve only 2 years behind bars.
-The classic scenario is known as the simultaneous Prisoner’s Dilemma, where two players must simultaneously decide whether to confess or lie without knowing the other’s choice.
-Games played in this manner are referred to as static games.
-The normal-form representation is suitable for them, as it captures the simultaneous nature of decision-making (as depicted in Fig.
-2a).
-Another variant is the sequential Prisoner’s Dilemma, where the second player decides with knowledge of the first player’s action.
-Games of this nature are called dynamic games.
-The extensive-form representation is well-suited for them, as it can clearly illustrate the sequence of moves and the information available to each player at each decision point (as shown in Fig.
-2b).
-Figure 2: The example of Prisoner’s Dilemma.
-(a) Matrix representation of simultaneous Prisoner’s Dilemma in normal-form.
-(b) Game tree representation of sequential Prisoner’s Dilemma in extensive-form.
+Moreover, a behavior strategy profile can be expressed by $\boldsymbol{\beta}=(\beta_{1},\cdots,\beta_{n})$, where $\beta_{i}$ maps each $I_{i}\in\mathcal{I}_{i}$ to a probability distribution over $A(I_{i})$.
+A game satisfies perfect recall if each player $i$ remembers the entire sequence of his past actions and all of the information he had at each of his previous decision points.
+In any finite extensive-form game with perfect recall, for every mixed strategy $\boldsymbol{\sigma}$, there exists an equivalent behavior strategy $\boldsymbol{\beta}$, and vice versa, such that both strategies induce the same distribution over terminal histories (i.e., the same outcomes in the game tree).
+A subgame of an extensive-form game is any subset of the game tree that (i) has exactly one initial decision node, (ii) whenever it contains a node it also contains all of its descendants, and (iii) whenever it contains a node it contains every other node in that node’s information set.
 Beyond normal-form and extensive-form games, the analysis of complex games often involves a higher-level abstraction: the meta-game.
-The meta-game can be viewed as an advanced form of a normal-form game, but it focuses on the selection of policies (also called strategies) rather than isolated actions.
-Each player chooses from a set of policies, rather than a set of actions.
-This set of policies is referred to as the policy population.
-Meta-strategies are mixed strategies that assign probabilities to the policies within the policy population in the meta-game.
-2.2.3 Transitive Game and Non-transitive Game
-For the sake of simplicity, we restrict our focus to two-player zero-sum symmetric games.
+A meta-game provides an abstraction over the original game by modeling interactions between entire policies rather than individual actions.
+While commonly instantiated as a normal-form game, a meta-game can, in general, take any game-theoretic form (e.g., extensive-form) as long as it captures strategic interactions over a policy population.
+In this setting, players choose policies instead of individual actions, and meta-strategies correspond to mixed strategies defined over these policies.
+2.2.2. Transitive Game and Non-Transitive Game
+To provide a clearer introduction to transitive and non-transitive games, in this section, we confine our analysis to two-player zero-sum symmetric games.
 In a transitive game, the strategies or outcomes follow a transitive relationship.
@@
 A classical example of a non-transitive game is Rock-Paper-Scissors, in which no single strategy uniformly dominates all others.
-2.2.4 Stage Game and Repeated Game
+2.2.3. Stage Game and Repeated Game
 A stage game (or one-shot game) is a game that is played only once, namely a one-shot interaction between players.
@@
 An example of a repeated game is Texas Hold’em, where players engage in multiple rounds, and each player’s strategy may evolve based on previous rounds, betting patterns, and the history of interactions among players.
-2.2.5 Nash Equilibrium
+2.2.4. Nash Equilibrium
 For simplicity, $\pi_{i}$ denotes the strategy of player $i$, and $\pi_{-i}$ denotes the strategies of all players other than player $i$.
 Given $\pi_{-i}$, player $i$’s best response (BR) is the strategy that maximizes player $i$’s payoff:
-$BR_{i}(\pi_{-i})=\arg\max_{\pi_{i}}u_{i}(\pi_{i},\pi_{-i}).$ (1)
-A strategy $\pi_{i}^{*}$ is an $\bm{\epsilon}$-BR to strategies $\pi_{-i}$ if:
-$u_{i}(\pi_{i}^{*},\pi_{-i})\geq u_{i}(BR_{i}(\pi_{-i}),\pi_{-i})-\epsilon,$ (2)
+(1) $BR_{i}(\pi_{-i})=\arg\max_{\pi_{i}}u_{i}(\pi_{i},\pi_{-i}).$
+A strategy $\pi_{i}^{*}$ is an $\boldsymbol{\epsilon}$-BR to strategies $\pi_{-i}$ if:
+(2) $u_{i}(\pi_{i}^{*},\pi_{-i})\geq u_{i}(BR_{i}(\pi_{-i}),\pi_{-i})-\epsilon,$
 where $\epsilon$ is a pre-specified threshold.
 A strategy profile $(\pi_{1}^{*},\pi_{2}^{*},...,\pi_{n}^{*})$ is a Nash equilibrium (NE) if, for each player $i$:
-$u_{i}(\pi_{i}^{*},\pi_{-i}^{*})\geq u_{i}(\pi_{i},\pi_{-i}^{*}),\forall\pi_{i},$ (3)
+(3) $u_{i}(\pi_{i}^{*},\pi_{-i}^{*})\geq u_{i}(\pi_{i},\pi_{-i}^{*}),\forall\pi_{i},$
 meaning that no player can benefit by changing their strategy unilaterally, given the strategies of others.
 In other words, an NE is a situation where each player’s strategy is a BR to the others’ strategies.
-A strategy profile $(\pi_{1}^{*},\pi_{2}^{*},...,\pi_{n}^{*})$ is an $\bm{\epsilon}$-NE if, for each player $i$:
-$u_{i}(\pi_{i}^{*},\pi_{-i}^{*})\geq u_{i}(\pi_{i},\pi_{-i}^{*})-\epsilon,\forall\pi_{i},$ (4)
+A strategy profile $(\pi_{1}^{*},\pi_{2}^{*},...,\pi_{n}^{*})$ is an $\boldsymbol{\epsilon}$-NE if, for each player $i$:
+(4) $u_{i}(\pi_{i}^{*},\pi_{-i}^{*})\geq u_{i}(\pi_{i},\pi_{-i}^{*})-\epsilon,\forall\pi_{i},$
 meaning that no player can increase their payoff by more than $\epsilon$ by unilaterally changing their strategy.
 However, computing NE is generally intractable in complex games, so some researchers utilize $\alpha$-Rank (Omidshafiei et al., 2019) and Correlated Equilibrium (CE) (Aumann, 1974) as alternatives.
-Some studies also resort to Replicator Dynamics (Taylor & Jonker, 1978) to analyze strategy evolution.
-3 Algorithms
+Some studies also resort to Replicator Dynamics (Taylor and Jonker, 1978) to analyze strategy evolution.
+2.2.5. Static Game and Dynamic Game
+The market entry game is a classic example in game theory.
+There are two players: Firm 1 (the entrant), which is considering whether to enter a market, and Firm 2 (the incumbent), which currently dominates that market.
+The game outcomes are as follows:
+  • If the entrant stays out of the market, the entrant earns a payoff of 4, and the incumbent earns 10.
+  • If the entrant enters and the incumbent chooses to attack, both firms receive a payoff of 3.
+  • If the entrant enters and the incumbent chooses not to attack, both firms receive a payoff of 6.
+Figure 2.
+The example of the market entry game.
+(a) Matrix representation of simultaneous market entry game in normal-form.
+(b) Game tree representation of sequential market entry game in extensive-form.
+In the simultaneous version of the market entry game, both players choose their strategies at the same time.
+The entrant decides whether to enter, and the incumbent simultaneously decides whether to attack, without knowing the entrant’s decision.
+Games of this kind are referred to as static games.
+Although a static game can be represented in either normal- or extensive-form, it is conventionally shown in the normal form because it succinctly captures the strategic structure of simultaneous moves (as shown in Fig.
+2a).
+In the sequential version, the entrant moves first, and the incumbent responds after observing the entrant’s decision.
+Games of this nature are called dynamic games.
+Although any dynamic game admits a normal-form representation, that representation scales exponentially with the number of information sets.
+Consequently, dynamic games are typically illustrated in extensive form, whose game tree compactly encodes the order of moves and the information available to each player (as depicted in Fig.
+2b).
+The market entry game demonstrates how the structure of a game influences equilibrium outcomes.
+In the static version, non-credible threats—such as the incumbent’s threat to attack even when it is not in its best interest—may support a Nash equilibrium where the entrant stays out.
+In contrast, the dynamic version allows for refinement through subgame perfect Nash equilibrium (SPNE), which requires that players’ strategies constitute a Nash equilibrium in every subgame.
+Under this refinement, the threat to attack is ruled out as non-credible, and the predicted outcome is that the entrant will enter and the incumbent will accommodate, since attacking would harm the incumbent itself.
+3. Algorithms
 Based on existing self-play work (Hernandez et al., 2021; Lanctot et al., 2017; Liu et al., 2022c; Garnelo et al., 2021), we propose a self-play framework (Algo.
@@
 In each iteration, a newly initialized policy is trained, while opponent policies are sampled from the population.
-After training, the new policy is added to the population.
+After training, the new policy is added to the population, possibly replacing an existing policy.
 The updated policy population is then evaluated, and then the opponent sampling strategy is recalculated for the next iteration.
@@
 To make it more straightforward, we highlight the representative classic algorithms from these and present them in Table 1 for comparison within our framework.
-In Sec.
-3.6, we compare the four categories, explain our classification rationale, and highlight the distinctions between our framework and existing ones to demonstrate its greater generality.
-3.1 Framework Definition
-Figure 3: Illustration of our framework.
+Sec.
+3.6 compares the four categories, explains our classification rationale, contrasts the framework with prior ones to demonstrate its greater generality, and highlights the design insights it offers for future self-play research.
+3.1. Framework Definition
+Figure 3.
+Illustration of our framework.
 The top row shows two different initialization methods for the policy population $\Pi$ and the interaction matrix $\Sigma$: lazy initialization and immediate initialization (corresponding to Line 1 in Algo.
@@
 2:for $e\in[[E]]$ do
-3:     for $\sigma_{i}\in\Sigma$ do
-4:        Initialize $\pi_{i}^{h}$
-5:        $\pi_{i}^{h}\leftarrow\textit{ORACLE}(\pi_{i}^{h},\sigma_{i},\Pi)$$\triangleright$ Compute the oracle.
-6:        $\mathcal{P}\leftarrow\textit{EVAL}(\Pi)$triangleright Get policies’ performance.
-7:        $\Sigma\leftarrow\textit{MSS}(\mathcal{P})$$\triangleright$ Update the interaction matrix.
+3:  for $\sigma_{[k]}\in[[K]]$ do
+4:   Initialize $\pi_{{[k]}}^{h}$
+5:   $\pi_{[k]}^{h}\leftarrow\textit{ORACLE}(\pi_{[k]}^{h},\sigma_{[k]},\Pi)\triangleright$ Compute the oracle.
+6:   $\mathcal{P}\leftarrow\textit{EVAL}(\Pi)$ $\triangleright$ Get policies’ performance (optional).
+7:   $\Sigma\leftarrow\textit{MSS}(\mathcal{P})\triangleright$ Update the interaction matrix (optional).
 8:  end for
@@
 Next, we provide an in-depth explanation of these processes:
-  • $\Pi:=\{\pi_{i}(\cdot|h(i))\}_{i=1}^{N}$: Each policy $\pi_{i}$ in the policy population $\Pi$ is conditioned on a policy condition function $h(i)$, which provides supplementary information for certain algorithms.
-We denote the policy $\pi_{i}(\cdot|h(i))$ as $\pi_{i}^{h}$.
-Note that $i$ refers to the $i$th policy in the population, not the $i$th player, and $N$ denotes the policy population size, not the number of players.
+  • $\Pi:=\{\pi_{[k]}(\cdot|h(k)\}_{k=1}^{K}$: Each policy $\pi_{[k]}$ in the policy population $\Pi$ is conditioned on a policy condition function $h(k)$, which provides supplementary information for certain algorithms.
+We denote the policy $\pi_{[k]}(\cdot|h(k))$ as $\pi_{[k]}^{h}$.
+Note that $\pi_{[k]}$ refers to the $k$th policy in the population, and $K$ denotes the policy population size.
 $\Pi$ can be initialized (Line 1 in Algo.
 1) in two ways: lazy initialization and immediate initialization.
-In lazy initialization, $\Pi$ starts with $N$ placeholder policies, with the actual policies being initialized during the training iterations (Line 4 in Algo.
+In lazy initialization, $\Pi$ starts with $K$ placeholder policies, with the actual policies being initialized during the training iterations (Line 4 in Algo.
 1).
-In immediate initialization, $\Pi$ is initialized upfront with $N$ real policies, which may be randomly generated or pre-trained models.
-A policy $\pi_{i}^{h}$ is considered inactive if it is a placeholder; otherwise, it is considered active.
-  • $\Sigma:=\{\sigma_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{1}}$: The interaction matrix $\Sigma$ consists of rows $\sigma_{i}$, where each $\sigma_{i}\in\mathbb{R}^{C_{1}}$ represents the opponent sampling strategy of policy $i$, indicating how to sample opponents’ policies.
-The dimension $C_{1}$ varies according to the sampling method.
-For instance, $\sigma_{i}$ can represent the probability distribution over opponent policies in the policy population $\Pi$, where $C_{1}=N^{n-1}$ and $n$ denotes the number of players.
-Alternatively, $\sigma_{i}$ can be viewed as parameters of a sampling network.
-Specially, in a two-player game, if $C_{1}=N$ and $\sigma_{ij}$ represents the probability that policy $i$ is optimized against policy $j$, $\Sigma$ can be depicted in directed interaction graphs (Fig.
+In immediate initialization, $\Pi$ is initialized upfront with $K$ real policies, which may be randomly generated or pre-trained models.
+A policy $\pi_{[k]}^{h}$ is considered inactive if it is a placeholder; otherwise, it is considered active.
+  • $\Sigma:=\{\sigma_{[k]}\}_{k=1}^{K+1}\in\mathbb{R}^{K\times C_{1}}$: The interaction matrix $\Sigma$ consists of rows $\sigma_{[k]}$.
+Each $\sigma_{[k]}\in\mathbb{R}^{C_{1}}$ represents the probability distribution over opponent policies in the policy population $\Pi$, where $C_{1}=K^{n-1}$ and $n$ denotes the number of players.
+Specially, in a two-player game, if $C_{1}=K$ and $\Sigma_{mn}$ represents the probability that policy $m$ is optimized against policy $n$, $\Sigma$ can be depicted in directed interaction graphs (Fig.
 4).
-$\Sigma$ will be initialized (Line 1 in Algo.
-1) based on the initialization of $\Pi$.
-If $\Pi$ uses lazy initialization, $\Sigma$ will be initialized as an empty matrix.
-If $\Pi$ uses immediate initialization, $\Sigma$ will be calculated based on $\Pi$.
-We will discuss this in more detail later.
+$\Sigma$ will be initialized with all zeros (Line 1 in Algo.
+1), providing a starting point before the meta-strategy solver updates it using the first payoff information.
   • For epoch $e\in[[E]]$ (Line 2 in Algo.
@@
 For example, if the algorithm only introduces new policies into the population without updating existing ones, then $E=1$.
-This implies that only the inactive policy is likely to be chosen for training in each iteration and turned into an active one, while activa policies remain unchanged.
+This implies that only the inactive policy is likely to be chosen for training in each iteration and turned into an active one, while active policies remain unchanged.
 Conversely, if active policies are updated multiple times throughout the algorithm, then $E>1$.
-Indeed, $E$ accurately reflects the number of updates performed.
-  • Initialize $\pi_{i}^{h}$ (Line 4 in Algo.
-1): The initialization of $\pi_{i}^{h}$ can vary depending on the algorithm being used.
+Indeed, $E$ reflects the number of updates performed.
+  • Initialize $\pi_{[k]}^{h}$ (Line 4 in Algo.
+1): The initialization of $\pi_{[k]}^{h}$ can vary depending on the algorithm being used.
 It can be initialized randomly or by leveraging pre-trained models (Silver et al., 2016; Vinyals et al., 2019) or through a recently updated policy.
-  • $\text{ORACLE}(\pi_{i},\sigma_{i},\Pi)$ (Line 5 in Algo.
+  • $\text{ORACLE}(\pi_{[k]},\sigma_{[k]},\Pi)$ (Line 5 in Algo.
 1): The ORACLE is an abstract computational entity that returns a new policy adhering to specific criteria.
@@
 4).
-(3) Other specially crafted ORACLEs are either tailored to specific meta-strategy solvers (MSSes) (Muller et al., 2020; Marris et al., 2021) or introduced to enhance diversity (Balduzzi et al., 2019; Perez-Nieves et al., 2021; Liu et al., 2021).
+(3) Other specially crafted ORACLEs are either tailored to specific meta-strategy solvers (MSSs) (Muller et al., 2020; Marris et al., 2021) or introduced to enhance diversity (Balduzzi et al., 2019; Perez-Nieves et al., 2021; Liu et al., 2021).
 Some details of Algo.
@@
 2 is a RL method that can be widely used across different categories.
-Moreover, off-policy RL algorithms typically require a replay buffer to gather samples; however, they are not explicitly included for simplicity.
+Moreover, on-policy RL algorithms do not require a sample buffer to collect trajectories.
+In contrast, only off-policy methods rely on a sample buffer to store and reuse past experiences (Line 11 and 15 in Algo.
+2).
 Algo.
@@
 4 is specifically utilized by the regret-minimization-based series.
-They will be introduced in detail in Sec.
+Details are provided in Sec.
 3.3.5 and Sec.
-3.5 respectively.
+3.5, respectively.
   • $\text{EVAL}(\Pi)$ (Line 6 in Algo.
@@
 There are multiple evaluation metrics available to assess the performance of each policy.
-The performance matrix is represented as $\mathcal{P}:=\{p_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{2}}$.
-$p_{i}$ is the performance of policy i, and the dimension $C_{2}$ depends on the evaluation metric used.
-For instance, $p_{i}$ can be depicted as the relative skill like Elo ratings ($C_{2}=1$) or can be depicted as the payoff tensor ($C_{2}=N^{n-1}$, where n is the number of players).
+The performance matrix is represented as $\mathcal{P}:=\{p_{k}\}_{k=1}^{K}\in\mathbb{R}^{K\times C_{2}}$.
+$p_{k}$ is the performance of policy k, and the dimension $C_{2}$ depends on the evaluation metric used.
+For instance, $p_{k}$ can be depicted as the relative skill like Elo ratings ($C_{2}=1$) or can be depicted as the payoff tensor ($C_{2}=K^{n-1}$, where n is the number of players).
 Specially, in two-player symmetric zero-sum games, the expected payoffs can serve as the evaluation metric.
-In such cases, $\mathcal{P}$ is a square matrix ($C_{2}=N$).
-  • $\textit{MSS}:\mathbb{R}^{N\times C_{2}}\rightarrow\mathbb{R}^{N\times C_{1}}$ (Line 7 in Algo.
-1): A meta-strategy solver (MSS) MSS takes the performance matrix $\mathcal{P}$ as its input and produces a new interaction matrix $\Sigma$ as its output.
+In such cases, $\mathcal{P}$ is a square matrix ($C_{2}=K$).
+  • $\textit{MSS}:\mathbb{R}^{K\times C_{2}}\rightarrow\mathbb{R}^{K\times C_{1}}$ (Line 7 in Algo.
+1): A meta-strategy solver (MSS) takes the performance matrix $\mathcal{P}$ as its input and produces a new interaction matrix $\Sigma$ as its output.
 If the MSS produces a constant matrix irrespective of the input, the evaluation step (Line 6 in Algo.
 1) can be skipped to reduce computations.
+In this case, the MSS step also can be skipped.
+Additionally, if it is the final iteration, the MSS step should also be skipped, as the resulting meta-strategy will not be used.
 Moreover, if $\Pi$ uses immediate initialization, $\Sigma$ will be initialized (Line 1 in Algo.
 1) by first evaluating the performance and then applying the MSS.
-Figure 4: Interaction matrix $\Sigma$ examples.
-When $C_{1}=N$ and the opponent sampling strategy $\sigma_{ij}$ represents the probability that policy $i$ is optimized against policy $j$, the interaction matrix $\Sigma$ can be depicted in directed interaction graphs.
+Figure 4.
+Interaction matrix $\Sigma$ examples.
+When $C_{1}=K$ and the opponent sampling strategy $\Sigma_{mn}$ represents the probability that policy $m$ is optimized against policy $n$, the interaction matrix $\Sigma$ can be depicted in directed interaction graphs.
 Here, we consider three representative self-play algorithms.
-In the Top section, we define the interaction matrix $\Sigma\in\mathbb{R}^{3\times 3}$ as $\{\sigma_{i}\}_{i=1}^{3}$.
+In the Top section, we define the interaction matrix $\Sigma\in\mathbb{R}^{3\times 3}$ as $\{\sigma_{[k]}\}_{k=1}^{3}$.
 In the Bottom section, we present directed interaction graphs where the outgoing edges from each node are equally weighted, and their weights collectively sum to one.
-The relationship between the Top and Bottom sections is established through directed edges: an edge directed from node $i$ to node $j$ with a weight of $\sigma_{ij}$ signifies that policy $i$ is optimized against policy $j$ with a probability of $\sigma_{ij}$.
+The relationship between the Top and Bottom sections is established through directed edges: an edge directed from node $m$ to node $n$ with a weight of $\Sigma_{mn}$ signifies that policy $m$ is optimized against policy $n$ with a probability of $\Sigma_{mn}$.
 Note that this figure is reproduced from Liu et al.
@@
 Algorithm 2 Compute the oracle in RL.
-0:  $\pi_{i}^{h}$$\triangleright$ Policy $i$ is being trained.
-0:  $\sigma_{i}$$\triangleright$ Opponent sampling strategy of policy $i$.
-0:  $\Pi$$\triangleright$ Policy population.
-1:  while $\pi_{i}^{h}$ is not valid do
+[…]
+[…]
+[…]
+0:$B$ $\triangleright$ Sample buffer (optional).
+1:while $\pi_{[k]}^{h}$ is not valid do
 2:  for trajectory $\tau\in[[\tau_{max}]]$ do
-3:        sample $\bm{\pi_{opp}^{h}}\sim P(\sigma_{i})$$\triangleright$ Policies of opponents.
-4:        $\bm{\pi}=(\pi_{i}^{h},\bm{\pi_{opp}^{h}})$
-5:        $s_{0},\bm{o_{0}}\sim\rho$triangleright Initial state.
+3:   sample $\boldsymbol{\pi_{opp}^{h}}\sim P(\sigma_{[k]})\triangleright$ Policies of opponents.
+4:   $\boldsymbol{\pi}=(\pi_{[k]}^{h},\boldsymbol{\pi_{opp}^{h}})$
+5:   $s_{0},\boldsymbol{o_{0}}\sim\rho$ $\triangleright$ Initial state.
 6:   for $t\in[[t_{max}]]$ do
-7:           $\bm{a_{t}}\sim\bm{\pi}(\bm{o_{t}})$
-8:           $s_{t+1},\bm{o_{t+1}}\sim P(s_{t},\bm{a_{t}})$
-9:           $\bm{r_{t}}\leftarrow\bm{R}(s_{t},\bm{a_{t}})$
+7:    $\boldsymbol{a_{t}}\sim\boldsymbol{\pi}(\boldsymbol{o_{t}})$
+8:    $s_{t+1},\boldsymbol{o_{t+1}}\sim P(s_{t},\boldsymbol{a_{t}})$
+9:    $\boldsymbol{r_{t}}\leftarrow\boldsymbol{R}(s_{t},\boldsymbol{a_{t}})$
 10:   end for
-11:        $\pi_{i}^{h}\leftarrow\text{update}(\pi_{i}^{h})$$\triangleright$ Using RL algorithms
-12:     end for
-13:  end while
-14:  return  $\pi_{i}^{h}$
+11:   $B\leftarrow B\cup\{\tau\}$ $\triangleright$ Store the trajectory in the buffer (optional).
+12:   $\pi_{[k]}^{h}\leftarrow\text{update}(\pi_{[k]}^{h})\triangleright$ Use RL algorithms to train the policy $k$.
+13:  end for
+14:end while
+15:return $\pi_{[k]}^{h},B$ $\triangleright$ Return the buffer (optional).
 Algorithm 3 Compute the oracle in evolution theory.
-0:  $\pi_{i}^{h}$$\triangleright$ Policy $i$ is being trained.
-0:  $\sigma_{i}$$\triangleright$ Opponent sampling strategy of policy $i$.
-0:  $\Pi$$\triangleright$ Policy population.
-1:  sample $\bm{\pi_{opp}^{h}}\sim P(\sigma_{i})$$\triangleright$ Policies of opponents.
-2:  $\bm{\pi_{\text{reg}}}=(\pi_{i}^{h},\bm{\pi_{opp}^{h}})$$\triangleright$ Regularization policies.
-3:  Transformed the reward according to $\bm{\pi_{\text{reg}}}$.
-4:  $\pi_{i}^{h}\leftarrow$ Use replicator dynamics to play the reward-transformed game until convergence.
-5:  return  $\pi_{i}^{h}$
+[…]
+[…]
+[…]
+1: sample $\boldsymbol{\pi_{opp}^{h}}\sim P(\sigma_{[k]})\triangleright$ Policies of opponents.
+[…]
+3: Transform the reward by adding a policy-dependent reward according to $\boldsymbol{\pi_{\text{reg}}}$ (Perolat et al., 2021).
+[…]
+5:return $\pi_{[k]}^{h}$
 Algorithm 4 Compute the oracle in regret matching.
-0:  $\pi_{i}^{h}$$\triangleright$ Policy $i$ is being trained.
-0:  $\sigma_{i}$$\triangleright$ Opponent sampling strategy of policy $i$.
-0:  $\Pi$$\triangleright$ Policy population.
-1:  for each player j do
-2:     sample $\bm{\pi_{opp}^{h}}\sim P(\sigma_{i})$$\triangleright$ Policies of opponents.
-3:     $\bm{\pi}=(\pi_{i}^{h}(j),\bm{\pi_{opp}^{h}}(-j))$
-4:     Use regret matching to play the game and update $h(i)$ with new regret minimization information.
+[…]
+[…]
+[…]
+1:for each player i do
+2:  sample $\boldsymbol{\pi_{opp}^{h}}\sim P(\sigma_{[k]})\triangleright$ Policies of opponents.
+3:  $\boldsymbol{\pi}=(\pi_{[k]}^{h}(i),\boldsymbol{\pi_{opp}^{h}}(-i))$
+4:  Use regret matching to play the game and update $h(k)$ with new regret minimization information.
 5:end for
-6:  return  $\pi_{i}^{h}$
-3.2 Traditional Self-play Algorithms
+6:return $\pi_{[k]}^{h}$
+3.2. Traditional Self-Play Algorithms
 Traditional self-play algorithms involve agents improving their strategies by repeatedly playing against themselves, allowing them to explore various strategies and enhance their decision-making abilities without external input.
@@
 This section will explain how traditional self-play algorithms fit into our framework and introduce representative traditional self-play methods, ranging from simpler forms to more complex ones.
-3.2.1 Integration into Our Framework
+3.2.1. Integration into Our Framework
 Traditional self-play algorithms can be incorporated into our proposed framework (Algo.
@@
 Second, we set $E=1$ because only the inactive policy can be trained in each iteration, then turning into an active policy.
-Here, the policy population size $N$ serves as the upper limit for the number of active policies in the population.
-In other words, we use $N$ iterations to optimize the policy.
-Third, the strategy being trained $\pi_{i}^{h}$ can be initialized in a general manner.
+Here, the policy population size $K$ serves as the upper limit for the number of active policies in the population.
+In other words, we use $K$ iterations to optimize the policy.
+Third, the strategy being trained $\pi_{[k]}^{h}$ can be initialized in a general manner.
 For instance, the strategy can be initialized randomly, learning from scratch.
-More often, $\pi_{i}^{h}$ is initialized by $\pi_{i-1}(\cdot|h(i-1))$, allowing incremental learning and adaptation based on the most current trained policy to accelerate convergence.
-Fourth, the policies in traditional self-play algorithms don’t need supplementary information, we set the policy condition function $h(i)=\emptyset$.
-Fifth, the MSSes of traditional self-play algorithms are straightforward, often yielding a constant interaction matrix $\Sigma$ that eliminates performance evaluation.
+More often, $\pi_{[k]}^{h}$ is initialized by $\pi_{[k-1]}(\cdot|h(k-1))$, allowing incremental learning and adaptation based on the most current trained policy to accelerate convergence.
+Fourth, the policies in traditional self-play algorithms don’t need supplementary information, we set the policy condition function $h(k)=\emptyset$.
+Fifth, the MSSs of traditional self-play algorithms are straightforward, often yielding a constant interaction matrix $\Sigma$ that eliminates performance evaluation.
 Only the MSS for prioritized fictitious self-play (described in Sec.
@@
 Assumption 1.
-In a two-player symmetric game, $C_{1}=N$, with $\sigma_{ij}$ denoting the probability that policy $i$ is optimized in response to policy $j$ which leads to $\sum_{j=1}^{N}\sigma_{ij}=1,~{}\forall i$.
+In a two-player symmetric game, $C_{1}=K$, with $\Sigma_{mn}$ denoting the probability that policy $m$ is optimized in response to policy $n$ which leads to $\sum_{n=1}^{K}\Sigma_{mn}=1,~\forall m$.
 Based on Assumption 1, we can further deduce the following important corollary:
@@
 The policy population gradually increases over time.
-When policy $i$ is selected for training, only policy $j$ $(j\leq i)$ has already been trained and holds meaningful outcomes.
+When policy $m$ is selected for training, only policy $n$ $(n\leq m)$ has already been trained and holds meaningful outcomes.
 Other policies are inactive policies.
-As a result, we exclusively select policy $j$, where $j\leq i$, to serve as the opponent for policy $i$.
+As a result, we exclusively select policy $n$, where $n\leq m$, to serve as the opponent for policy $m$.
 ∎
-3.2.2 Vanilla Self-play
-In vanilla self-play (Samuel, 2000), agents are trained by competing against their latest versions. Thus, the MSS of vanilla self-play:
-$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}1,~{}\text{if}~{}j=i-1\vspace{4pt}\\ 0,~{}\text{otherwise}\end{array}\right..$ (5)
+3.2.2. Vanilla Self-Play
+In vanilla self-play (Samuel, 1959), agents are trained by competing against their latest versions. Thus, the MSS of vanilla self-play:
+(5) $\textit{MSS}(\mathcal{P})_{mn}=\left\{\begin{array}[]{l}1,~\text{if}~n=m-1\\ 0,~\text{otherwise}\end{array}\right.$.
 Regardless of $\mathcal{P}$, the MSS produces the same interaction matrix.
@@
 Vanilla self-play is effective in transitive games, but it can lead the agent to cyclic learning patterns in non-transitive games.
-It may also lead to an unstationary problem, causing the system to get stuck in a local optimum.
+It may also lead to a nonstationary problem, causing the system to get stuck in a local optimum.
 A more modern version of vanilla self-play is to combine it with off-policy RL.
-In each iteration, the policy to be trained, $\pi_{i}^{h}$, is initialized using the previous policy $\pi_{i-1}(\cdot|h(i-1))$.
+In each iteration, the policy to be trained, $\pi_{[k]}^{h}$, is initialized using the previous policy $\pi_{[k-1]}(\cdot|h(k-1))$.
 Typically, each player’s policy is fixed, and samples are collected into a buffer by vanilla self-play.
@@
 Once the policy is updated, it will be added to the policy population.
-3.2.3 Fictitious Self-play
+Moreover, vanilla self-play is frequently used in conjunction with Monte Carlo Tree Search (MCTS), and we provide a detailed discussion in Sec.
+4.1.1, using the game of Go as a representative application.
+3.2.3. Fictitious Play
 Fictitious Play (FP) (Brown, 1951) is a learning algorithm in game theory where each player best responds to the empirical frequency of the strategies used by the opponent.
@@
 Based on FP intuition, Fictitious Self-Play (FSP) (Heinrich et al., 2015) is introduced to make agents play against past versions of themselves to learn optimal strategies to improve the robustness of vanilla self-play.
-Neural Fictitious Self-Play (NFSP) (Heinrich & Silver, 2016) is a modern variant that combines FSP with deep learning techniques.
+Neural Fictitious Self-Play (NFSP) (Heinrich and Silver, 2016) is a modern variant that combines FSP with deep learning techniques.
 It uses neural networks to approximate the BRs.
-In the original versions of NFSP and FSP, two distinct types of agent memory are used: one to record the agent’s own behavior and the other to capture the opponent’s behavior.
-In more recent approaches (Lanctot et al., 2017), random sampling is frequently employed to approximate the opponents’ average strategy, eliminating the need to maintain two separate types of agent memory.
-Thus, the MSS of FSP:
-$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{i-1},~{}\text{if}~{}j\leq i-1\vspace{4pt}\\ 0,~{}\text{otherwise}\end{array}\right..$ (6)
-In FSP, the MSS continues to generate a constant interaction matrix.
+Both NFSP and FSP utilize two distinct types of memory: one to store the agent’s own behavioral data, and the other to record observed opponent behavior.
+More recent approaches (Lanctot et al., 2017) adopt FP in meta-games, where the opponent’s average strategy is approximated directly, eliminating the need to maintain two separate types of memory.
+In this setting, the MSS of FP is defined as:
+(6) $\textit{MSS}(\mathcal{P})_{mn}=\left\{\begin{array}[]{l}\frac{1}{m-1},~\text{if}~n\leq m-1\\ 0,~\text{otherwise}\end{array}\right.$.
+In FP, the MSS continues to generate a constant interaction matrix.
 Compared to vanilla self-play, this approach enhances the robustness of the policy by sampling older versions of its own policies to be used as opponent strategies.
-3.2.4 $\delta$-uniform Self-play
-[…]
-$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{f(i)},~{}\text{if}~{}\lceil\delta(i-1)\rceil\leq j\leq i-1\vspace{4pt}\\ 0,~{}\text{otherwise}\end{array}\right.,$ (7)
-$f(i)=i-\lceil\delta(i-1)\rceil.$ (8)
-In $\delta$-uniform self-play, The MSS generates a constant interaction matrix. Specially, if $\delta=0$, it corresponds to FSP, and if $\delta=1$, it corresponds to vanilla self-play.
-3.2.5 Prioritized Fictitious Self-play
+3.2.4. $\delta$-Uniform Self-Play
+$\delta$-uniform self-play, introduced by Bansal et al.
+(2018), uses the hyper-parameter $\delta$, ranging from 0 to 1, to select the most recent $1-\delta$ percentage of policies for uniform sampling to generate opponent policies.
+When policy $m$ is in the training phase, according to Corollary 3.1, only the previous m1m-1 policies are relevant.
+To select opponents for policy $m$, we sample from the range of policies between $\lceil\delta(m-1)\rceil,m-1]$, where $\lceil\cdot\rceil$ denotes the ceiling function, which rounds up the input to the nearest integer greater than or equal to it.
+When $\delta=0$, the system retains the complete historical memory, whereas $\delta=1$ implies that only the most recent policy is utilized.
+Thus, the MSS of $\delta$-uniform self-play:
+(7) $\textit{MSS}(\mathcal{P})_{mn}=\left\{\begin{array}[]{l}\frac{1}{f(m)},~\text{if}~\lceil\delta(m-1)\rceil\leq n\leq m-1\\ 0,~\text{otherwise}\end{array}\right$.,
+(8) $f(m)=m-\lceil\delta(m-1)\rceil.$
+In $\delta$-uniform self-play, The MSS generates a constant interaction matrix. Specially, if $\delta=0$, it corresponds to FP, and if $\delta=1$, it corresponds to vanilla self-play.
+3.2.5. Prioritized Fictitious Self-Play
 Prioritized Fictitious Self-Play (PFSP)  (Vinyals et al., 2019) leverages a priority function to allocate a higher probability of selection to agents with higher priorities.
-Here, $\mathcal{P}$ represents the winning rates, specifically defined as $\mathcal{P}_{ij}=\text{Prob}(\pi_{i}~{}\text{beats}~{}\pi_{j})$.
+Here, $\mathcal{P}$ represents the winning rates, specifically defined as $\mathcal{P}_{mn}=\text{Prob}(\pi_{[m]}~\text{beats}~\pi_{[n]})$.
 The MSS of PFSP is given by Algo.
@@
 For example, $f(x)=(1-x)^{p}$ with $p>0$ indicates that stronger policies against the currently being trained policy have a higher chance of being chosen as opponents.
-Alternatively, $f=x(1-x)$ implies that players of similar levels are more likely to be chosen as opponents.
+Alternatively, f=x(1x)f=x(1-x) implies that players of similar levels are more likely to be chosen as opponents.
 Additionally, in broader terms, $\mathcal{P}$ can also be assessed by other metrics.
@@
 1:function $\mathcal{F}(\mathcal{P})$
-2:   $\sigma_{i+1,j}\leftarrow\frac{f(\mathcal{P}_{i,j})}{\sum_{j\leq i}f(\mathcal{P}_{i,j})},~{}\text{if}~{}j\leq i$
-3:   Append zeros to $\sigma_{i+1}$ until its length is N.
+2:  $\sigma_{m+1,n}\leftarrow\frac{f(\mathcal{P}_{m,n})}{\sum_{n\leq m}f(\mathcal{P}_{m,n})},~\text{if}~n\leq m$
+3:  Append zeros to $\sigma_{m+1}$ until its length is K.
 4:return $\Sigma$
-3.3 PSRO Series
+3.3. PSRO Series
 Similar to traditional self-play algorithms, the PSRO series of algorithms starts with a single policy and gradually expands the policy space by incorporating new oracles.
 These oracles are policies that approximate optimal responses to the current meta-strategies of other agents.
-3.3.1 Integration into Our Framework
+3.3.1. Integration into Our Framework
 The PSRO series of algorithms can also be integrated into our proposed framework (Algo.
@@
 First, similar to traditional self-play algorithms, we also utilize lazy initialization to initialize $\Pi$.
-Second, we also set $E=1$ and $N$ can be considered as the upper limit for the policy population size in the original PSRO algorithms.
-Third, in the context of the PSRO series of algorithms, the strategy of our player $\pi_{i}^{h}$ can also be initialized in a general manner.
-Fourth, we set $h(i)=\emptyset$ since the PSRO series of algorithms do not use any policy condition function for their policies.
-Fifth, it’s crucial to highlight that our framework diverges from the traditional PSRO model (Lanctot et al., 2017) in how $\sigma_{i}$ is defined.
-In contrast to being the meta-strategy for policy, in our framework, $\sigma_{i}$ is the opponent sampling strategy.
-It means that $\sigma_{i}$ here represents the opponent’s meta-strategy against policy $i$ for the PSRO series of algorithms.
-Sixth, compared with traditional self-play methods, the MSSes of the PSRO series are often more complex.
-For example, some MSSes incorporate concepts from different types of game equilibria (McMahan et al., 2003; Muller et al., 2020; Marris et al., 2021).
+Second, we also set $E=1$ and $K$ can be considered as the upper limit for the policy population size in the original PSRO algorithms.
+Third, in the context of the PSRO series of algorithms, the strategy of our player $\pi_{[k]}^{h}$ can also be initialized in a general manner.
+Fourth, we set $h(k)=\emptyset$ since the PSRO series of algorithms do not use any policy condition function for their policies.
+Fifth, it’s crucial to highlight that our framework diverges from the traditional PSRO model (Lanctot et al., 2017) in how $\sigma_{[k]}$ is defined.
+In contrast to being the meta-strategy for policy, in our framework, $\sigma_{[k]}$ is the opponent sampling strategy.
+It means that $\sigma_{[k]}$ here represents the opponent’s meta-strategy against policy $k$ for the PSRO series of algorithms.
+Sixth, compared with traditional self-play methods, the MSSs of the PSRO series are often more complex.
+For example, some MSSs incorporate concepts from different types of game equilibria (McMahan et al., 2003; Muller et al., 2020; Marris et al., 2021).
 Lastly, we also simply follow Assumption 1.
-We can derive the Corollary 2 using a similar proof as Corollary 1.
+We can derive the Corollary 3.2 using a similar proof as Corollary 3.1.
 Corollary 0.
 In the PSRO series of algorithms, the interaction matrix $\Sigma$ is a lower triangular matrix.
-3.3.2 Double Oracle
+3.3.2. Double Oracle
 Double Oracle (DO) (McMahan et al., 2003) can be only applied to two-player normal-form games.
@@
 6.
-The opponent sampling strategy $\sigma_{i}$ corresponds to the opponent’s NE strategy of the restricted game.
+The opponent sampling strategy $\sigma_{[k]}$ corresponds to the opponent’s NE strategy of the restricted game.
 Therefore, the oracle in DO is a BR rather than an ABR, computing the BR against the current NE opponent strategy of the restricted game.
-In the context of two-player normal-form games, DO theoretically can achieve the NE of the full game.
+In two-player normal-form settings, DO theoretically can achieve the NE of the full game.
 Algorithm 6 NE-based meta-strategy solver.
 1:function $\mathcal{F}(\mathcal{P})$
-2:   $\sigma_{i+1}\leftarrow$ SOLVE-NASH$(\mathcal{P}_{1:i,1:i})$triangleright Opponent’s NE meta-strategy.
-3:   Append zeros to $\sigma_{i+1}$ until its length is N.
+[…]
+3:  Append zeros to $\sigma_{[k+1]}$ until its length is K.
 4:return $\Sigma$
 We take rock-paper-scissors as an example, a symmetric two-player game.
-The policy condition function $h(i)=\emptyset$.
-The policy population size $N=4$.
-The policy population $\Pi$ uses lazy initialization and the interaction matrix $\Sigma$ is initialized with all zeros.
-Assume that in the first iteration, the policy $\pi_{1}$ is set to "rock".
-Obviously, the NE-based MSS returns $\sigma_{2}=(1,0,0,0)$.
-In the second iteration, the BR oracle adds "paper" to $\Pi$ as $\pi_{2}$.
-Consequently, the NE-based MSS yields $\sigma_{3}=(0,1,0,0)$, since "paper" always beats "rock." In the third iteration, noting that $\sigma_{3}$ selects only "paper" and never "rock," the BR oracle adds "scissors" as $\pi_{3}$.
-The NE-based MSS then solves the game and returns $\sigma_{4}=(1/3,1/3,1/3,0)$.
-In the final iteration, the BR oracle adds a mixed strategy that assigns a uniform probability over "rock," "paper," and "scissors" as $\pi_{4}$.
+The policy condition function $h(k)=\emptyset$.
+The policy population size $K=3$.
+The policy population $\Pi$ uses lazy initialization, and the interaction matrix $\Sigma$ is initialized with all zeros.
+Assume that in the first iteration, the policy $\pi_{1}$ is set to rock.
+Obviously, the NE-based MSS returns $\sigma_{2}=(1,0,0)$.
+In the second iteration, the BR oracle adds paper to $\Pi$ as $\pi_{2}$.
+Consequently, the NE-based MSS yields $\sigma_{3}=(0,1,0)$, since paper always beats rock.
+In the third iteration, noting that $\sigma_{3}$ selects only paper and never rock, the BR oracle adds scissors as $\pi_{3}$.
 The final interaction matrix is:
-$\Sigma=\begin{bmatrix}0&0&0&0\\ 1&0&0&0\\ 0&1&0&0\\ \frac{1}{3}&\frac{1}{3}&\frac{1}{3}&0\end{bmatrix}$ (9)
+(9) $\Sigma=\begin{bmatrix}0&0&0\\ 1&0&0\\ 0&1&0\\ \end{bmatrix}$
+Finally, the NE over this policy population becomes the uniform distribution over rock, paper, and scissors, indicating convergence.
 This process demonstrates how the DO algorithm iteratively expands the strategy set within our framework.
-Also, using DO as an example, we show that, compared with traditional self-play algorithms, the PSRO series employs more sophisticated MSSes to better capture the essence of the game, albeit at a higher computational cost.
-3.3.3 PSRO
+Also, using DO as an example, we show that, compared with traditional self-play algorithms, the PSRO series employs more sophisticated MSSs to better capture the essence of the game, albeit at a higher computational cost.
+3.3.3. PSRO
 PSRO (Lanctot et al., 2017) extends DO to more complex games beyond just two-player normal-form games.
@@
 The MSS framework allows for a more flexible representation of strategic solutions in various game settings.
-Many variants of PSRO focus on designing new MSSes to better capture different aspects of strategic play in these more complex games.
+Many variants of PSRO focus on designing new MSSs to better capture different aspects of strategic play in these more complex games.
 In the original PSRO framework, the oracle is computed using RL techniques, similar to those described in Algorithm 2.
 This allows the algorithm to effectively handle large and intricate strategy spaces, making it applicable to various game scenarios.
-3.3.4 PSRO Variants
-The traditional PSRO algorithm has been the subject of numerous extensions in recent research.
-Some studies focus on making PSRO more computationally tractable in different scenarios or achieving fast convergence.
+3.3.4. PSRO Variants
+The traditional PSRO algorithm has inspired numerous extensions in recent research.
+Although the original PSRO aligns well with our proposed framework, its variants may fall outside its scope; however, they remain noteworthy and merit discussion.
+The first line of PSRO variants focuses on accelerating convergence.
 By establishing a hierarchical structure of RL workers, Pipeline PSRO (McAleer et al., 2020) achieves the parallelization of PSRO and simultaneously provides guarantees of convergence.
@@
 Anytime PSRO (McAleer et al., 2022b) and Self-play PSRO (McAleer et al., 2022a) are designed to incorporate less exploitable policies into the policy population, facilitating faster convergence.
-Moreover, as the number of agents grows, determining BRs becomes exponentially challenging.
+In parallel, other works aim to make PSRO more computationally tractable in different game settings.
+As the number of agents grows, determining BRs becomes exponentially challenging.
 Mean-field PSRO (Muller et al., 2022) has been introduced to address this complexity in mean-field games.
-In addition, due to the computational intractability of solving NE in multi-player general-sum games (Daskalakis et al., 2009; Daskalakis, 2013) and the selection problem in NE (Goldberg et al., 2013), Müller et al.
-put forward $\alpha$-PSRO (Muller et al., 2020).
-Instead of NE, $\alpha$-rank (Omidshafiei et al., 2019), which is unique and polynomial-time solvable in multi-player general-sum games, is introduced in the MSS and a preference-based best response (PBR) oracle is incorporated in the approach.
+In addition, due to the computational intractability of solving NE in multi-player general-sum games (Daskalakis et al., 2009; Daskalakis, 2013) and the selection problem in NE (Goldberg et al., 2013)Muller et al.
+(2020) put forward $\alpha$-PSRO.
+Instead of NE, $\alpha$-rank (Omidshafiei et al., 2019), which is unique and polynomial-time solvable in multi-player general-sum games, is introduced in the MSS, and a preference-based best response (PBR) oracle is incorporated in the approach.
 Similar to $\alpha$-PSRO, Marris et al.
-(2021) proposes Joint Policy-Space Response Oracle (JPSRO) to tackle multi-player general-sum games.
+(2021) proposes the Joint Policy-Space Response Oracle (JPSRO) to tackle multi-player general-sum games.
 It utilizes CE and coarse correlated equilibrium (CCE) as alternatives to NE to make it computationally tractable in multi-player general-sum games and designs the CE and CCE-based MSS.
-Other research focuses on policy diversification, as deriving a single policy in transitive games is less meaningful.
+Other lines of research focus on promoting policy diversity, especially in non-transitive games where a single policy is often insufficient.
 Balduzzi et al.
@@
 Policy Space Diversity PSRO (Yao et al., 2023) defines a diversity metric named population exploitability that helps to achieve a full-game NE.
-3.3.5 R-NaD
+3.3.5. R-NaD
 Although R-NaD (Perolat et al., 2021) is initially described as leveraging evolutionary game theory with a regularization component.
 Here, we categorized it into the PSRO series with a unique oracle computation technique (Algo.
-3), executed in two stages: the first stage involves transforming the reward based on the regularization policy to make it policy-dependent, and the second stage applies replicator dynamics (Taylor & Jonker, 1978) until convergence to a fixed point.
+3), executed in two stages: the first stage involves transforming the reward based on the regularization policy to make it policy-dependent, and the second stage applies replicator dynamics (Taylor and Jonker, 1978) until convergence to a fixed point.
 The oracle added to the policy population $\Pi$ in each iteration is derived from the reward-transformed game, not the original problem’s oracle.
@@
 This equation illustrates that the fixed point reached in each iteration, the oracle, is utilized as the regularization policy for the next iteration.
-3.4 Ongoing-training-based Series
+3.4. Ongoing-Training-Based Series
 In the PSRO series of algorithms, two key challenges arise.
@@
 In other words, instead of only inactive policies being selected for training, all active policies are also likely to be chosen for training.
-3.4.1 Integration into Our Framework
+3.4.1. Integration into Our Framework
 We can incorporate these ongoing-training-based series of algorithms into our proposed framework (Algo.
@@
 In other words, each unique policy undergoes iterative training for $E_{max}$ times.
-Third, since each policy undergoes training for $E_{max}$ times, we utilize $\pi_{i}(\cdot|h(i))$ to initialize $\pi_{i}^{h}$.
+Third, since each policy undergoes training for $E_{max}$ times, we utilize $\pi_{[k]}(\cdot|h(k))$ to initialize $\pi_{[k]}^{h}$.
 This means that policy updates are self-referential.
-Lastly, under Assumption 1, different from Collory 1 and 2, due to the continuous training process of all policies, we derive Collory 3.
+Lastly, under Assumption 1, different from Collory 3.1 and 3.2, due to the continuous training process of all policies, we derive Collory 3.3.
 Corollary 0.
@@
 Proof.
-When policy $i$ is selected for training, policy $k$ $(k\geq i)$ has already been actually initialized and is therefore considered an active policy.
-Furthermore, in epoch $e$ $(e>1)$, policy $k$ has already been updated and holds significant meaning.
-As a result, policy $k$, where $k\geq i$, is likely to be chosen as the opponent for policy $i$.
+When policy $m$ is selected for training, policy $n$ $(n\geq m)$ has already been actually initialized and is therefore considered an active policy.
+Furthermore, in epoch $e$ $(e>1)$, policy $n$ has already been updated and holds significant meaning.
+As a result, policy $n$, where $n\geq m$, is likely to be chosen as the opponent for policy $m$.
 Consequently, the interaction matrix $\Sigma$ is generally not a lower triangular matrix.
 ∎
-3.4.2 FTW
+3.4.2. FTW
 Quake III Arena Capture the Flag is a renowned 3D multi-player first-person video game where two teams vie to capture as many flags as possible.
@@
 A pivotal aspect of FTW is its employment of the ongoing-training-based self-play in RL.
-Specifically, it trains a population of $N$ different policies in parallel, which compete and collaborate with each other.
-When policy $i$ is undergoing training, FTW samples both its teammates and adversaries from the population.
+Specifically, it trains a population of $K$ different policies in parallel, which compete and collaborate with each other.
+When policy $k$ is undergoing training, FTW samples both its teammates and adversaries from the population.
 Specially, in scenarios where each team comprises a single member, it can be seamlessly integrated into our framework using the subsequent MSS:
-$\textit{MSS}(\mathcal{P})_{ij}=\frac{1}{N}.$ (10)
+(10) $\textit{MSS}(\mathcal{P})_{mn}=\frac{1}{K}.$
 This essentially means that the interaction graph is densely connected.
 Moreover, all policies draw upon a unified policy network parameterized by $\phi$.
-Hence, $\pi_{i}(\cdot|h(i))$ can be aptly depicted as $\pi_{\phi}(\cdot|h(i))$.
-Furthermore, since these policies are not conditioned on external parameters, it is straightforward to represent the conditioning function $h(i)=\emptyset$.
-3.4.3 NeuPL
+Hence, $\pi_{[k]}(\cdot|h(k))$ can be aptly depicted as $\pi_{\phi}(\cdot|h(k))$.
+Furthermore, since these policies are not conditioned on external parameters, it is straightforward to represent the conditioning function $h(k)=\emptyset$.
+3.4.3. NeuPL
 Neural Population Learning (NeuPL) (Liu et al., 2022c) introduces another critical innovation: it employs a unified conditional network, where each policy is adjusted against a specific meta-game mixture strategy.
 This is instrumental in enabling transfer learning across policies.
-Owing to NeuPL’s reliance on a unified conditional network parameterized by $\theta$, $\pi_{i}(\cdot|h(i))$ can be succinctly represented as $\pi_{\theta}(\cdot|h(i))$.
-Since NeuPL policies depend on the opponent sampling strategy $\sigma_{i}$, we define $h(i)=\sigma_{i}$.
-3.4.4 Simplex-NeuPL
+Owing to NeuPL’s reliance on a unified conditional network parameterized by $\theta$, $\pi_{[k]}(\cdot|h(k))$ can be succinctly represented as $\pi_{\theta}(\cdot|h(k))$.
+Since NeuPL policies depend on the opponent sampling strategy $\sigma_{[k]}$, we define $h(k)=\sigma_{[k]}$.
+3.4.4. Simplex-NeuPL
 Simplex-NeuPL (Liu et al., 2022a), which builds on NeuPL, is designed to achieve any-mixture optimality, which signifies that the formulated policy should exhibit flexibility across a diverse spectrum of adversaries, including those that might not present equivalent competitive prowess.
 To model the population learning process from a geometric perspective, Simplex-NeuPL introduces the concept of the population simplex.
-Analogously to its predecessor, Simplex-NeuPL integrates a conditional network to characterize the policy, represented as $\pi_{\theta}(\cdot|h(i))$ conditioned on the opponent sampling strategy $h(i)=\sigma_{i}$.
-Intriguingly, there is a slight possibility that $\sigma_{i}$ does not originate from the MSS.
-Instead, it is drawn as a sample from the population simplex.
-This sampling mechanism results in greater robustness.
-3.5 Regret-minimization-based Series
-Another line of self-play algorithms is based on regret minimization.
-Unlike other categories, this series prioritizes accumulated payoff over time rather than focusing solely on a single episode, which is common in repeated games.
-For example, in games like Texas Hold’em or Werewolf, players must use deception, concealment, and bluffing to aim for overall victories rather than just winning a single game.
+Analogously to its predecessor, Simplex-NeuPL integrates a conditional network to characterize the policy, represented as $\pi_{\theta}(\cdot|h(k))$ conditioned on the opponent sampling strategy $h(k)=\sigma_{[k]}$.
+Notably, $\sigma_{[k]}$ may be sampled from the population simplex rather than the MSS, enhancing robustness.
+3.5. Regret-Minimization-Based Series
+Another line of self-play algorithms is based on regret minimization (RM).
+Unlike other categories, this series aims to minimize cumulative regret over time, a concept rooted in online learning theory.
+This distinction is crucial in two-player zero-sum games, where these methods enjoy strong theoretical guarantees.
 Traditional regret-minimization-based self-play typically doesn’t use RL, but later studies have combined them for improved performance.
-This section will also discuss traditional regret-minimization methods to lay the groundwork for understanding their integration with RL.
-3.5.1 Integration into Our Framework
+This section will also discuss traditional regret-minimization methods to lay the ground for understanding their integration with RL.
+3.5.1. Integration into Our Framework
 We can also integrate the regret-minimization-based series of algorithms into our proposed framework (Algo.
 1) with the following settings: First, similar to traditional self-play algorithms and the PSRO series, we use lazy initialization to initialize the policy population $\Pi$.
-Second, we set $E=1$, and $N$ is regarded as the maximum iteration to optimize the policy.
-Third, we initialize $\pi_{i}^{h}$ using $\pi_{i-1}(\cdot|h(i-1))$ to utilize the most current trained policy.
-More specifically, $h(i)=h(i-1)$ and $\pi_{i}^{h}=\pi_{i-1}^{h}$.
-Fourth, in each iteration $i$, $h(i)$ represents the specific elements that regret-minimization-based self-play algorithms need to store.
-Note that this category relies heavily on the information in $h(i)$.
-For instance, in Counterfactual Regret Minimization (CFR) (Zinkevich et al., 2007), it is necessary to store counterfactual regrets for every action in every information set for each player within $h(i)$.
-Once $h(i)$ is determined, the corresponding policy is also defined.
+Second, we set $E=1$, and $K$ is regarded as the maximum iteration to optimize the policy.
+Third, in each iteration $k$, $h(k)$ represents the specific elements that regret-minimization-based self-play algorithms need to store.
+Note that this category relies heavily on the information in $h(k)$.
+For instance, in Counterfactual Regret Minimization (CFR) (Zinkevich et al., 2007), it is necessary to store counterfactual regrets for every action in every information set for each player within $h(k)$.
+Once $h(k)$ is determined, the corresponding policy is also defined.
 We will discuss CFR in detail in Sec.
 3.5.2.
+Fourth, we initialize $\pi_{[k]}^{h}$ using $\pi_{[k-1]}(\cdot|h(k-1))$ to utilize the most current trained policy.
+More specifically, $h(k)$ is initialized by h(k1)h(k-1) and $\pi_{[k]}^{h}$ is initialized by $\pi_{[k-1]}^{h}$.
+In each new iteration, regret-minimization information is continuously added to $h(k)$ (Line 4 in Algo.
+4).
+As a result, $h$ accumulates information across all iterations until the end of training.
 Fifth, the ABR operator (Algo.
-4) incorporates regret-related information into $h(i)$.
+4) incorporates regret-related information into $h(k)$.
 Unlike the original CFR, Algo.
@@
 This adjustment has been shown to accelerate convergence empirically and possesses a theoretical error bound (Burch et al., 2019).
-Additionally, each $\pi_{i}^{h}$ represents the strategies for all players, and in iteration $i$, player $j$ uses the strategy $\pi_{i}^{h}(j)$.
+Each $\pi_{[k]}^{h}$ represents the strategies for all players, and in iteration $k$, player $i$ uses the strategy $\pi_{[k]}^{h}(i)$.
 Lastly, the MSS of the regret-minimization-based series is the vanilla MSS, as described in Equ.
 (5).
-Also, we can derive Collory 4.
+Also, we can derive Collory 3.4.
 Corollary 0.
@@
 Proof.
-Regret minimization-based algorithms always use the latest strategy for training.
-In other words, at iteration $i$, $\pi_{i-1}^{h}$ is consistently chosen as the opponent policy.
-Consequently, the interaction matrix $\Sigma$ is a unit lower shift matrix.
+Regret-minimization-based algorithms always train against the latest strategy; that is, at iteration $k$, the opponent policy is $\pi_{[k-1]}^{h}$.
+As a result, the interaction matrix $\Sigma$ becomes a unit lower shift matrix.
 ∎
-3.5.2 Vanilla CFR
+3.5.2. Vanilla CFR
 Regret measures the difference between the best possible payoff and the actual payoff.
-The regret matching algorithm (Hart & Mas-Colell, 2000) optimizes decisions by selecting strategies based on accumulated positive overall regrets, with higher overall regret strategies being more likely chosen to correct past underperformance.
+The regret matching algorithm (Hart and Mas-Colell, 2000) optimizes decisions by selecting strategies based on accumulated positive overall regrets, with higher overall regret strategies being more likely chosen to correct past underperformance.
 After each round, players update the overall regret values for each strategy.
-Strategies will converge to an approximate NE.
+The algorithm is guaranteed to converge to the set of NE, and any finite stopping point yields an approximate NE.
 However, this method primarily applies to normal-form games because computing overall regret in extensive-form games is challenging.
@@
 (2007) proposes CFR for extensive-form games.
-Here, it is referred to as vanilla CFR to distinguish it from later advancements.
-Vanilla CFR maintains the strategy and the immediate counterfactual regrets for each information set.
-Theoretically, this regret is specific to each information set, and the aggregation of these immediate counterfactual regrets is an upper bound for the overall regret.
-Therefore, the problem can be decomposed into numerous more minor regret minimization problems.
-Next, there is a more detailed description of vanilla CFR.
-In this survey, $\pi_{i}$ denotes the strategy at iteration $i$ for consistency, whereas the original paper uses $\sigma$.
-Furthermore, we use $j$ to represent player $j$, and $\pi_{i}(j)$ denotes the policy used by player $j$ at iteration $i$.
-The counterfactual value $v_{j}(\pi,I)$, which represents the expected value upon reaching the information set $I$ when all players, except for player $j$, adhere to the strategy $\pi$:
-$v_{j}(\pi,I)=\sum_{h\in I,h^{\prime}\in Z}\text{Prob}_{-j}^{\pi}(h)\text{Prob}^{\pi}(h,h^{\prime})u_{j}(h^{\prime}).$ (11)
-It is an unnormalized form of the counterfactual utility in the original paper. Based on this counterfactual value, the immediate counterfactual regret is defined as:
-$R_{j,\text{imm}}^{T}(I)=\frac{1}{T}\max_{a\in A(I)}\sum_{i=1}^{T}\left(v_{j}(\pi_{i}|_{I\to a},I)-v_{j}(\pi_{i},I)\right),$ (12)
-where $\pi_{i}|_{I\to a}$ denotes player $j$ choose action $a$ with probability $1$ at iteration $i$. Moreover, the positive immediate counterfactual regret is:
-$R_{j,\mathrm{imm}}^{T,+}(I)=\mathrm{max}(R_{j,\mathrm{imm}}^{T}(I),0).$ (13)
-The player $j$’s average overall regret $R_{j}^{T}$ is defined as:
-$R_{j}^{T}=\dfrac{1}{T}\max_{\pi(j)^{*}\in\Pi(j)}\sum_{i=1}^{T}\left(u_{j}(\pi(j)^{*},\pi_{i}(-j))-u_{j}(\pi_{i})\right).$ (14)
-Theoretically, average overall regret $R_{j}^{T}$ is bounded by the sum of positive immediate counterfactual regrets:
-$R_{j}^{T}\leq\sum_{I\in\mathcal{I}_{j}}R_{j,\text{imm}}^{T,+}(I).$ (15)
-Thus, the problem can be simplified from minimizing the overall regret to minimizing the immediate counterfactual regrets in each information set separately.
-For simplicity, we often drop "immediate" in discussions and refer directly to counterfactual regrets.
-The counterfactual regret can also be recorded as follows:
-$R_{j}^{T}(I,a)=\frac{1}{T}\sum_{i=1}^{T}\left(v_{j}(\pi_{i}|_{I\to a},I)-v_{j}(\pi_{i},I)\right),$ (16)
-$R_{j}^{T,+}(I,a)=\max\{R_{j}^{T}(I,a),0\}.$ (17)
-The regret matching algorithm (Hart & Mas-Colell, 2000) determines the strategy at each information set:
-$\left.\pi_{j}^{T+1}(I)(a)=\left\{\begin{array}[]{ll}\frac{R_{j}^{T,+}(I,a)}{\sum_{a\in A(I)}R_{j}^{T,+}(I,a)}&\text{ if denominator}>0\\ \frac{1}{|A(I)|}&\text{ otherwise}\end{array}\right.\right..$ (18)
-It’s worth noting that in some studies, the normalization factor $\frac{1}{T}$ is omitted in Equ. (12), (14) and (16).
+In this survey, we refer to it as vanilla CFR to distinguish it from later advancements.
+Vanilla CFR maintains both the strategy and the counterfactual regret values for each information set.
+Theoretically, the sum of these local regrets provides an upper bound on the overall regret.
+Therefore, the problem can be decomposed into many smaller regret minimization subproblems.
+CFR focuses on minimizing regret locally at each information set, making the approach computationally tractable.
+At each iteration, the strategy is updated using regret matching, and the final strategy is obtained through strategy averaging, where the strategies played across iterations are averaged over time.
+This averaged strategy is proven to converge to an NE in two-player zero-sum games.
 Vanilla CFR has several shortcomings.
@@
 Although some efforts have focused on game abstraction to reduce the size of the game tree, greater abstraction can lead to decreased performance.
-Second, it requires storing counterfactual regrets $R^{i}(I,a)$ for every action $a$ in every information $I$ at each iteration $i$.
-These values are stored in $h(i)$ within our proposed framework (Algo.
+Second, it requires storing counterfactual regrets $R^{k}(I,a)$ for every action $a$ in every information $I$ at each iteration $k$.
+These values are stored in $h(k)$ within our proposed framework (Algo.
 1), leading to significant storage challenges.
-3.5.3 Time-saving Variants of CFR
+3.5.3. Time-Saving Variants of CFR
 Many studies focus on enhancing the time efficiency of CFR.
 The first approach involves modifying the regret calculation to increase its speed.
-CFR+ (Tammelin, 2014) implements regret-matching+ by storing non-negative regret-like values $R_{j}^{+,i}(I,a)$, rather than $R_{j}^{i}(I,a)$.
+CFR+ (Tammelin, 2014) implements regret-matching+ by storing non-negative regret-like values $R^{+,k}(I,a)$, rather than $R^{k}(I,a)$ in $k^{th}$ iteration.
 Additionally, CFR+ updates the regrets of each player sequentially and adopts a weighted average strategy.
-Moreover, Brown & Sandholm (2019a) introduces the concept of weighted regrets and develops Linear CFR (LCFR) and Discounted CFR (DCFR).
+Moreover, Brown and Sandholm (2019a) introduces the concept of weighted regrets and develops Linear CFR (LCFR) and Discounted CFR (DCFR).
 The second approach involves adopting sampling methods.
@@
 Moreover, Johanson et al.
-(2012) expands on chance-sampling by categorizing based on the handling of public and private chance nodes.
+(2012b) expands on chance-sampling by categorizing based on the handling of public and private chance nodes.
 Some studies also focus on learning how to reduce the variance of MCCFR to speed up convergence (Schmid et al., 2019; Davis et al., 2020).
-In addition to these two primary approaches, other studies have identified that warm starting (Brown & Sandholm, 2016) and pruning (Brown & Sandholm, 2015; Brown et al., 2017; Brown & Sandholm, 2017) can also accelerate convergence.
-3.5.4 Space-saving Variants of CFR
+In addition to these two primary approaches, other studies have shown that warm starting (Brown and Sandholm, 2016) and pruning (Brown and Sandholm, 2015; Brown et al., 2017; Brown and Sandholm, 2017) can also accelerate convergence.
+3.5.4. Space-Saving Variants of CFR
 In perfect information games, decomposition reduces problem-solving scale by solving subgames.
@@
 The process includes updating the trunk with the counterfactual values from the subgame’s root and updating the average counterfactual values at this root, while the solution to the subgame is discarded.
-CFR-D minimizes storage needs by only saving $R^{i}(I_{*},a)$ for information sets in the trunk and at each subgame’s root, which is denoted by $I_{*}$, trading off storage efficiency against the time required to resolve subgames.
-Similar thoughts are echoed in the Continue-Resolving technique used by DeepStack (Moravčík et al., 2017) and the Safe and Nested Subgame Solving technique used by Libratus (Brown & Sandholm, 2018).
+CFR-D minimizes storage needs by only saving $R^{k}(I_{*},a)$ for information sets in the trunk and at each subgame’s root, which is denoted by $I_{*}$, trading off storage efficiency against the time required to resolve subgames.
+Similar thoughts are echoed in the Continue-Resolving technique used by DeepStack (Moravčík et al., 2017) and the Safe and Nested Subgame Solving technique used by Libratus (Brown and Sandholm, 2018).
 We will discuss these approaches in Sec.
 4.2.1, exploring their application to Texas Hold’em.
-3.5.5 Estimation Variants of CFR
+3.5.5. Estimation Variants of CFR
 Although these CFR variants advance the field, they can’t directly solve large imperfect-information extensive-form games due to their reliance on tabular representations.
@@
 After advantage-based regret minimization (ARM) (Jin et al., 2018) merges CFR with deep RL in single-agent scenarios, a growing body of research has focused on applying CFR in conjunction with neural networks to multi-agent scenarios.
-Double Neural CFR (Li et al., 2018) utilizes two neural networks: one for estimating counterfactual regrets and another for approximating the average strategy.
+Double Neural CFR (Li et al., 2020a) utilizes two neural networks: one for estimating counterfactual regrets and another for approximating the average strategy.
 In a similar vein, Deep CFR (Brown et al., 2019) leverages an advantage network $V(I,a|\theta_{p})$ to estimate counterfactual regrets with each player having a distinct hyperparameter $\theta_{p}$ and employs $\pi(I,a|\theta_{\pi})$ for strategy estimation after the training process of the advantage network.
-Since these two networks are trained in sequence rather than concurrently, the strategy for each intermediate iteration remains conditioned on the output of the advantage network: $h(i)=V(I,a|\theta_{p})$.
+Since these two networks are trained in sequence rather than concurrently, the strategy for each intermediate iteration remains conditioned on the output of the advantage network: $h(k)=V(I,a|\theta_{p})$.
 Despite similarities, Deep CFR distinguishes itself from Double Neural CFR through its data collection and proven effectiveness in larger-scale poker games.
@@
 Moreover, advantage regret-matching actor-critic (ARMAC) (Gruslys et al., 2020) incorporates the thought of retrospective policy improvement.
-Table 1: Overview of representative self-play algorithms within our framework.
-Algorithms MSS h(i) Categories $E$ Initialization of $\Pi$ Initialization of $\pi_{i}^{h}$ Suitable Game
-Vanilla Self-play Equ. (5) $\emptyset$ Traditional Self-play $E=1$ Lazy General Stage Game
-FSP Equ. (6)
+Table 1. Overview of representative self-play algorithms within our framework.
+Algorithms MSS h(k) Categories $E$ Initialization of $\Pi$ Initialization of $\pi_{[k]}^{h}$
+Vanilla Self-Play Equ. (5) $\emptyset$ Traditional Self-Play $E=1$ Lazy General
+FP Equ. (6)
 $\delta$-Uniform Self-Play Equ. (7)
 PFSP Algo. 5
-DO NE-based (Algo. 6) $\emptyset$ PSRO Series $E=1$ Lazy General Stage Game
+DO NE-Based (Algo. 6) $\emptyset$ PSRO Series $E=1$ Lazy General
 PSRO General
@@
 R-NaD Equ. (5)
-FTW Equ. (10) $\emptyset$ Ongoing-training-based $E>1$ Immediate $\pi_{i}^{h}\leftarrow\pi_{i}(\cdot|h(i))$ Stage Game
-NeuPL General $\sigma_{i}$
-Simplex-NeuPL General $\sigma_{i}$
-Vanilla CFR Equ. (5) $R^{i}(I,a)$ Regret-minimization-based $E=1$ Lazy $\pi_{i}^{h}\leftarrow\pi_{i-1}(\cdot|h(i-1))$ Repeated Game
-CFR+ $R^{+,i}(I,a)$
-CFR-D $R^{i}(I_{*},a)$
+FTW Equ. (10) $\emptyset$ Ongoing-Training-Based $E>1$ Immediate $\pi_{[k]}^{h}\leftarrow\pi_{[k]}(\cdot|h(k))$
+NeuPL General $\sigma_{[k]}$
+Simplex-NeuPL General $\sigma_{[k]}$
+Vanilla CFR Equ. (5) $R^{k}(I,a)$ Regret-Minimization-Based $E=1$ Lazy $\pi_{[k]}^{h}\leftarrow\pi_{[k-1]}(\cdot|h(k-1))$
+CFR+ $R^{+,k}(I,a)$
+CFR-D $R^{k}(I_{*},a)$
 RCFR $\varphi(I,a)$
 Deep CFR $V(I,a|\theta_{p})$
-3.6 Reassessment of the Framework
+3.6. Reassessment of the Framework
 After introducing these four categories of self-play algorithms, we will further compare them in this section, explain the rationale behind categorizing the algorithms, and summarize the representative algorithms in Table 1.
 Moreover, we will illustrate the differences between our framework and other frameworks to demonstrate why our proposed framework is more general.
-3.6.1 Categorization Rationale
+Finally, we discuss how our framework can guide the development of future self-play algorithms by offering a unified perspective on their design and evaluation.
+3.6.1. Categorization Rationale
 Traditional self-play algorithms and the PSRO series share many similarities.
@@
 Therefore, in our framework, we use lazy initialization to initialize the policy population and set $E=1$ for these two categories.
-The interaction matrix is typically a lower triangular matrix (Corollary 1 and Corollary 2).
-The primary difference between the PSRO series and traditional self-play algorithms is that the PSRO series employs more complex MSSes to handle tasks with intricate game-theoretical requirements.
+The interaction matrix is typically a lower triangular matrix (Corollary 3.1 and Corollary 3.2).
+The primary difference between the PSRO series and traditional self-play algorithms is that the PSRO series employs more complex MSSs to handle tasks with intricate game-theoretical requirements.
 For example, $\alpha$-PSRO (Muller et al., 2020) specifically utilizes an $\alpha$-rank-based MSS to tackle multi-player general-sum games.
-In other tasks, traditional self-play algorithms are more commonly used to reduce the computational cost associated with complex MSSes in the PSRO series.
+In other tasks, traditional self-play algorithms are more commonly used to reduce the computational cost associated with complex MSSs in the PSRO series.
 Unlike the two previously mentioned categories, the ongoing-training-based series adopts a different paradigm.
@@
 This method alleviates issues such as early truncation and repeated skill relearning that occur in the above two categories.
-To integrate this category into our framework, immediate initialization is used for the policy population, and $\pi_{i}(\cdot|h(i))$ is utilized to initialize $\pi_{i}^{h}$ to ensure that policy updates are self-referential.
-Also, the interaction matrix is generally not a lower triangular matrix (Corollary 3).
-Lastly, the regret-minimization-based series focuses on overall performance over time rather than on individual episodes, making it particularly suitable for repeated games.
-For example, Texas Hold’em is a classic repeated game where players adjust their strategies based on past interactions and use tactics involving deception and bluffing.
-The main goal of this training process is to update the regrets associated with different strategies, which in turn demands significant storage resources.
-Our framework uses $h(i)$ to store this information.
-Since the policies are determined by $h(i)$, only the most recent policy is relevant.
-Therefore, the interaction matrix is a unit lower shift matrix (Corollary 4).
-We also do not need actually to initialize the whole policy population and only need to use $\pi_{i-1}(\cdot|h(i-1))$ to initialize $\pi_{i}^{h}$ in the training process.
-3.6.2 Comparing with Existing Frameworks
+To integrate this category into our framework, immediate initialization is used for the policy population, and $\pi_{[k]}(\cdot|h(k))$ is utilized to initialize $\pi_{[k]}^{h}$ to ensure that policy updates are self-referential.
+Also, the interaction matrix is generally not a lower triangular matrix (Corollary 3.3).
+Lastly, the regret-minimization-based series follows an online learning paradigm that focuses on optimizing overall performance across time rather than on individual episodes.
+Representative methods such as CFR and its variants are specifically designed for extensive-form imperfect information games, making them particularly effective in such settings.
+The core mechanism of regret-minimization-based training is to iteratively update the regrets associated with different strategies.
+However, tracking and accumulating regret for every information set incurs substantial memory overhead, especially in large game trees.
+Furthermore, available theoretical guarantees of convergence to a Nash equilibrium have so far been established only for two-player zero-sum games; achieving comparable assurances in multiplayer or general-sum settings remains an open research problem.
+Our framework uses $h(k)$ to store this information.
+Since the policies are determined by $h(k)$, only the most recent policy is relevant.
+Therefore, the interaction matrix is a unit lower shift matrix (Corollary 3.4).
+We also do not need to actually initialize the whole policy population and only need to use $\pi_{[k-1]}(\cdot|h(k-1))$ to initialize $\pi_{[k]}^{h}$ in the training process.
+3.6.2. Comparing with Existing Frameworks
 Our framework is built upon PSRO (Lanctot et al., 2017) and NeuPL (Liu et al., 2022c).
 Here, we outline the differences between our framework and these existing frameworks.
-The primary distinction between our framework and PSRO is the use of an interaction matrix $\Sigma:=\{\sigma_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{1}}$ to represent the opponent sampling strategy, allowing for the integration of more complex competitive dynamics.
-Moreover, in our framework, $\sigma_{i}$ denotes the opponent sampling strategy, which specifies how to sample opponents’ policies against policy $i$ rather than being the meta-strategy of policy $i$.
-Additionally, our framework incorporates a policy condition function $h(i)$, making it more general than NeuPL, where policies are conditioned on $\sigma_{i}$.
+The primary distinction between our framework and PSRO is the use of an interaction matrix $\Sigma:=\{\sigma_{[k]}\}_{k=1}^{K}\in\mathbb{R}^{K\times C_{1}}$ to represent the opponent sampling strategy, allowing for the integration of more complex competitive dynamics.
+Moreover, in our framework, $\sigma_{[k]}$ denotes the opponent sampling strategy, which specifies how to sample opponents’ policies against policy $k$ rather than being the meta-strategy of policy $k$.
+Additionally, our framework incorporates a policy condition function $h(k)$, making it more general than NeuPL, where policies are conditioned on $\sigma_{[k]}$.
 This enhancement gives our framework greater expressiveness.
@@
 Also, to the best of our knowledge, our framework is the first self-play framework to integrate the regret-minimization-based series, which is a significant self-play paradigm.
-4 Empirical Analysis
+3.6.3. Implications for Future Self-Play Algorithm Design
+Beyond merely organizing existing algorithms, our framework provides insights into the design of future self-play methods.
+First, the four algorithmic categories we identify are not isolated; mechanisms developed in one family can be transplanted into another.
+Second, although most previous work has concentrated on the ORACLE and MSS, our unified framework uncovers additional, under-explored dimensions that warrant further investigation.
+The remainder of this section first illustrates how techniques can migrate across categories and then presents two specific research directions that arise when the full framework design space is considered.
+Cross-category inspiration.
+Recent advances in self-play have explicitly blended techniques across different algorithmic categories, yielding hybrid methods that transcend traditional boundaries.
+RM-BR (Johanson et al., 2012a) pairs RM for one player with a BR oracle for the other, leveraging RM’s equilibrium guarantees and exploitative strength.
+Anytime PSRO (McAleer et al., 2022b) proposes RM-BR DO by incorporating a regret-minimization component into DO to ensure that exploitability monotonically decreases as new strategies are added.
+ODO (Dinh et al., 2022) fuses another online no-regret learning method multiplicative weights update (MWU) (Freund and Schapire, 1999) with DO.
+Our framework naturally subsumes ODO.
+Policies are still added iteratively, and the auxiliary information $h(k)$ records the average regret-based losses.
+MWU acts as the ORACLE.
+Crucially, unlike the original DO, this MWU oracle need not compute a full Nash equilibrium, reducing computational cost; yet ODO still provably converges to a Nash equilibrium in two-player zero-sum normal-form games while preserving rationality, allowing agents to exploit sub-optimal opponents.
+Taken together, these studies show that innovation thrives when regret-minimization-based series intersect with PSRO series, and they suggest that integrating other categories could be just as fruitful.
+Prioritised re-training when $E\gg 1$.
+When each policy is reused across many encounters ($E\gg 1$), repeatedly updating all members becomes computationally expensive.
+A natural open question is: which policies are worth further improvement?
+For instance, we can formalize this as a multi-armed bandit: give every policy a score that estimates how much new information an additional update would add (e.g., its expected information gain), and let that score decide where to spend oracle calls.
+This bandit scheduling automatically balances exploitation (further refining the strongest policies) with exploration (re-examining under-trained or atypical policies that might expose blind spots).
+Auxiliary information $h(k)$ and knowledge transfer.
+The history kernel $h(k)$ need not serve merely as a passive log of past interactions; instead, it can be designed to encode rich auxiliary domain knowledge, such as exploitable behavioral patterns, partial payoff structures, or latent embeddings of opponents.
+By treating $h(k)$ as a learnable memory object that can be inherited or shared across subsets of agents, we enable a form of continual-learning self-play.
+In this setting, knowledge acquired by one agent can propagate throughout the population organically, without requiring intervention from an external oracle.
+This facilitates more efficient collective learning and strategic evolution within multi-agent systems.
+4. Empirical Analysis
 In this section, we introduce iconic applications of self-play by categorizing the scenarios into three distinct groups: board games, which typically involve perfect information; card games, which usually involve imperfect information; and video games, which feature real-time actions rather than turn-based play.
 We then illustrate how self-play is applied in these complex scenarios and provide a comparative analysis of these applications in Table 2.
-4.1 Board Games
-Board games, the majority of which are perfect information games, were previously revolutionized by the introduction of two essential techniques: position evaluation and Monte Carlo tree search (MCTS) (Coulom, 2006; Kocsis & Szepesvári, 2006).
-These methodologies, with minor modifications, demonstrated superhuman effectiveness in solving board games such as chess (Campbell et al., 2002), checkers (Schaeffer et al., 1992), othello (Buro, 1999), backgammon (Tesauro & Galperin, 1996), and Scrabble (Sheppard, 2002).
-In contrast, the application of these techniques to the game of Go with an estimated $2.1\times 10^{170}$ legal board configurations, only enabled performance at the amateur level (Bouzy & Helmstetter, 2004; Coulom, 2007; Baudiš & Gailly, 2011; Enzenberger et al., 2010; Gelly & Silver, 2007).
-In light of this, our discussion will specifically focus on the game of Go to illustrate the application of self-play.
+4.1. Board Games
+Board games, the majority of which are perfect information games, were previously revolutionized by the introduction of two essential techniques: position evaluation and MCTS (Coulom, 2006; Kocsis and Szepesvári, 2006).
+These methodologies, with minor modifications, demonstrated superhuman effectiveness in solving board games such as chess (Campbell et al., 2002), checkers (Schaeffer et al., 1992), othello (Buro, 1999), backgammon (Tesauro and Galperin, 1996), and Scrabble (Sheppard, 2002).
+In contrast, the application of these techniques to the game of Go with an estimated $2.1\times 10^{170}$ legal board configurations, only enabled performance at the amateur level (Bouzy and Helmstetter, 2004; Coulom, 2007; Baudiš and Gailly, 2011; Enzenberger et al., 2010; Gelly and Silver, 2007).
+In light of this, our discussion will specifically focus on the game of Go to illustrate the application of vanilla self-play.
 In addition to Go, we will explore Stratego, a board game with imperfect information, in contrast to most board games that only involve perfect information.
-4.1.1 Go
+4.1.1. Go
 Go is an ancient board game, which is played on a grid of 19x19 lines, where two players alternately place black and white stones aiming to control the largest territory.
-The paradigm of Go artificial intelligence (AI) is revolutionized with the launch of DeepMind’s AlphaGo series (Silver et al., 2016; 2017; 2018; Schrittwieser et al., 2020), which leveraged the power of self-play to significantly elevate performance.
+The paradigm of Go artificial intelligence (AI) is revolutionized with the launch of DeepMind’s AlphaGo series (Silver et al., 2016, 2017, 2018; Schrittwieser et al., 2020), which leveraged the power of vanilla self-play to significantly elevate performance.
 In AlphaGo (Silver et al., 2016), the training regime can be split into three stages.
 In the first stage, supervised learning with expert data trains a fast policy network $p_{\pi}(a|s)$ for rollouts in MCTS and a precise policy network $p_{\sigma}(a|s)$.
-The second stage employs self-play combined with RL to get a refined policy network $p_{\rho}(a|s)$ based on $p_{\sigma}(a|s)$ and subsequently trains a value network $v_{\theta}(s)$.
+The second stage employs vanilla self-play combined with RL to get a refined policy network $p_{\rho}(a|s)$ based on $p_{\sigma}(a|s)$ and subsequently trains a value network $v_{\theta}(s)$.
 Specifically, $p_{\rho}(a|s)$ is refined by competing against a randomly selected historical version $p_{\rho^{-}}(a|s)$, similar to the MSS shown in Equ.
-(6), while $v_{\theta}(s)$ is trained using the game samples collected by self-play of $p_{\rho}(a|s)$.
+(6), while $v_{\theta}(s)$ is trained using the game samples collected by vanilla self-play of $p_{\rho}(a|s)$.
 In the third stage, MCTS integrates the policy and value networks to select actions.
 Based on AlphaGo, AlphaGo Zero (Silver et al., 2017) does not require any expert data except game rules.
-It utilizes only one network $f_{\theta}(s)$ to concurrently predict the action probabilities and the state value.
-Self-play is employed to generate data and refine $f_{\theta}(s)$ with the current best policy competing against itself, a process analogous to the MSS referenced in Equ.
+It utilizes only one network $f_{\theta}(s)$ to concurrently predict the action probabilities $\text{Pr}(a|s)$ and the state value $v$.
+Each move is generated by MCTS with guidance from $f_{\theta}(s)$ to aid MCTS expansion, unlike the rollouts used in AlphaGo.
+Vanilla self-play is employed to generate data and refine $f_{\theta}(s)$ with the current best policy competing against itself, a process analogous to the MSS referenced in Equ.
 (5).
-A new policy to be incorporated into the policy pool must surpass a 55 percent win rate against its predecessor.
+A new policy to be incorporated into the policy pool must surpass a 55 percent win rate against its predecessor, aligning with the stipulation set in Algo.
+2 at Line 1.
 AlphaZero (Silver et al., 2018) extends AlphaGo Zero to include games beyond Go, such as Chess and Shogi.
-Concerning the self-play procedure, the only difference between AlphaZero and AlphaGo Zero is that AlphaZero utilizes the newly updated network to generate samples by self-play without validation.
+Notably, a draw is introduced as an additional expected outcome, and data augmentation is omitted due to the asymmetry of Chess and Shogi.
+Concerning the vanilla self-play procedure, the only difference between AlphaZero and AlphaGo Zero is that AlphaZero utilizes the newly updated network without the validation process present in AlphaGo Zero.
 Building upon AlphaZero, MuZero (Schrittwieser et al., 2020) takes learning from scratch to the next level, even operating without predefined game rules.
-The self-play process in MuZero operates similarly to that in AlphaZero.
-In practice, in addition to excelling in board games, MuZero also achieves state-of-the-art performance in Atari games.
-4.1.2 Stratego
+MuZero incorporates ideas from model-based RL to model the dynamics of games.
+In addition to the prediction network $f$ (similar to the networks in AlphaGo Zero and AlphaZero), MuZero introduces a dynamics network $g$ to model the MDP and a representation network $h$ to map observations to hidden states.
+These three networks are trained jointly.
+MuZero employs MCTS guided by the three networks above to make decisions.
+The vanilla self-play in MuZero operates similarly to that in AlphaZero.
+In practice, in addition to excelling in board games like Go, MuZero also achieves state-of-the-art performance in Atari games.
+4.1.2. Stratego
 Unlike most board games, which are perfect information games, Stratego, a two-player imperfect information board game, distinguishes itself by incorporating elements of memory, deduction, and bluffing.
@@
 DeepNash holds the third-place ranking among all professional Gravon Stratego players and wins nearly every match against existing Stratego bots.
-4.2 Card Games
+4.2. Card Games
 Unlike board games, card games are typically imperfect information games with a larger state space and greater complexity.
-Here, we introduce three representative card games: Texas Hold’em, DouDiZhu, and Mahjong, to illustrate how self-play is utilized in card games.
-Texas Hold’em is a repeated game rather than a stage game, so the focus is on the overall win-loss outcome rather than on individual rounds.
-Consequently, it is particularly well-suited for regret-minimization-based self-play methods, which aim to minimize overall regret rather than regret on a per-round basis.
-In contrast, DouDiZhu and Mahjong are typically regarded as stage games, so studies often employ self-play methods that do not rely on regret minimization.
-4.2.1 Texas Hold’em
+Some regret-minimization-based self-play methods, particularly CFR and its variants, are well-suited for handling imperfect information and have achieved strong results in card games.
+However, recent studies have shown that other self-play methods without regret minimization can also perform well in such settings.
+We introduce three representative card games: Texas Hold’em, DouDiZhu, and Mahjong, to illustrate how self-play is utilized in card games.
+4.2.1. Texas Hold’em
 Texas Hold’em, a popular poker game with 2-10 players, is known for its strategic depth and bluffing elements.
@@
 Furthermore, some studies do not adopt the regret-minimization-based series of self-play; instead, they use the traditional self-play method.
-NSFP (Heinrich & Silver, 2016) introduces a self-play method combined with end-to-end RL training, also achieving competitive performance in HULHE.
+NSFP (Heinrich and Silver, 2016) introduces a self-play method combined with end-to-end RL training, also achieving competitive performance in HULHE.
 Poker-CNN (Yakovenko et al., 2016) utilizes self-play to learn convolutional networks to solve a video version of HULHE.
@@
 This method focuses on a subtree of limited depth and breadth and estimates the outcomes of the furthest reaches of this subtree.
-Libratus (Brown & Sandholm, 2018) develops a blueprint strategy, leveraging an enhanced version of regret-minimization-based self-play method MCCFR (Lanctot et al., 2009).
+Libratus (Brown and Sandholm, 2018) develops a blueprint strategy, leveraging an enhanced version of regret-minimization-based self-play method MCCFR (Lanctot et al., 2009).
 Recursive Belief-based Learning (ReBeL) (Brown et al., 2020) introduces the public belief state (PBS), transforming the imperfect-information game into a perfect-information game with continuous state space.
@@
 Unlike the algorithms mentioned above, AlphaHoldem (Zhao et al., 2022a) does not use regret-minimization-based self-play.
-Instead, it proposes a $K$-best self-play method that always selects the top $K$ agents, based on their ELO ratings (Elo & Sloan, 1978), to generate samples.
-Pluribus (Brown & Sandholm, 2019b), based on Libratus, addresses six-player no-limit Texas Hold’em.
+Instead, it proposes $K$-best self-play that selects the top $K$ policies, based on their ELO ratings (Elo and Sloan, 1978), to generate samples.
+Pluribus (Brown and Sandholm, 2019b), based on Libratus, addresses six-player no-limit Texas Hold’em.
 Similar to Libratus, Pluribus utilizes a regret-minimization-based self-play method MCCFR (Lanctot et al., 2009) to develop its blueprint strategy.
@@
 Different from Libratus, Pluribus maintains a streamlined policy pool of only four strategies, assuming that its opponents might adjust their strategies during gameplay among these four strategies, which allows Pluribus to manage complexity more efficiently.
-4.2.2 DouDiZhu
+4.2.2. DouDiZhu
 DouDizhu (a.k.a.
@@
 PerfectDou (Yang et al., 2022) utilizes vanilla self-play under a Perfect-Training-Imperfect-Execution framework.
-4.2.3 Mahjong
+4.2.3. Mahjong
 Mahjong has evolved into various global variants, including the famous Japanese version known as Riichi Mahjong.
@@
 The strategic depth and complexity of Mahjong pose significant challenges.
-Despite ongoing research, these findings have yet to reach expert human levels (Kurita & Hoki, 2020; Gao et al., 2019; Mizukami & Tsuruoka, 2015).
+Despite ongoing research, these findings have yet to reach expert human levels (Kurita and Hoki, 2020; Gao et al., 2019; Mizukami and Tsuruoka, 2015).
 Difficulties are navigating incomplete information, dynamically adapting strategies to multiple opponents, and contending with complex winning rules and an enormous number of possible game states, estimated at $10^{169}$ (Zha et al., 2019).
-Suphx (Li et al., 2020) is recognized as one of the first algorithms to master Mahjong, achieving a performance level comparable to expert human players, precisely 10 dan on Tenhou (Tsunoda, ), the most popular online Mahjong platform.
+Suphx (Li et al., 2020b) is recognized as one of the first algorithms to master Mahjong, achieving a performance level comparable to expert human players, precisely 10 dan on Tenhou (Tsunoda, [n.
+d.]), the most popular online Mahjong platform.
 Initially, Suphx employs supervised learning, utilizing expert data to train its model.
 It then advances its capabilities through vanilla self-play combined with RL.
-Similarly, NAGA (Village, ), developed by Dwango Media Village, and LuckyJ (Tencent, ), designed by Tencent, have also achieved the rank of 10 dan on Tenhou.
+Similarly, NAGA (Village, [n.
+d.]), developed by Dwango Media Village, and LuckyJ (Tencent, [n.
+d.]), designed by Tencent, have also achieved the rank of 10 dan on Tenhou.
 Furthermore, LuckyJ has even defeated human professional players.
-4.3 Video Games
+4.3. Video Games
 In contrast to traditional board games and card games, video games often feature real-time actions, long trajectories, and increased complexity stemming from a broader range of actions and observations.
 We illustrate representative video games showcasing self-play’s impact.
-4.3.1 StarCraft II
+4.3.1. StarCraft II
 StarCraft is a real-time strategy (RTS) game.
@@
 It maintains a policy pool of past players that records all these types of agents.
-Main agents engage in traditional self-play algorithms like FSP and PFSP, competing against main agents themselves and other agents in the policy pool.
+Main agents engage in traditional self-play algorithms like FP and PFSP, competing against main agents themselves and other agents in the policy pool.
 They are periodically added to the pool and never reset.
@@
 Further studies (Han et al., 2020; Wang et al., 2021; Huang et al., 2024) have enhanced the league self-play training procedure.
-4.3.2 MOBA Games
+4.3.2. MOBA Games
 Multiplayer Online Battle Arena (MOBA) games are a popular video game genre that blends RTS with role-playing elements.
@@
 Managing multiple lanes and battling under the fog of war, which obscures parts of the map, are critical aspects of the gameplay.
-OpenAI Five (Berner et al., 2019) defeated the world champion team in a simplified version of Dota 2 that featured a limited pool of heroes and certain banned items.
+OpenAI Five (Berner et al., 2019) defeated the world champion team in a simplified version of Dota 2, which supported only 17 heroes and excluded complex features like summons and illusions.
 The training process introduces a self-play method that combines two traditional self-play algorithms: with an 80% probability of engaging in vanilla self-play and a 20% probability of employing a technique similar to the PFSP used in AlphaStar (Vinyals et al., 2019).
@@
 (2020a) as well.
-Unlike OpenAI Five, this work expands the hero pool to 40 heroes, exponentially increasing the possible combinations of lineups.
+Unlike OpenAI Five, which supported only 17 heroes, this work expands the hero pool to 40 heroes, substantially increasing the number of possible lineup combinations.
 It proposes a new self-play method referred to as curriculum self-play learning (CSPL).
@@
 The self-play generated data is also used to learn compelling lineup drafting by utilizing MCTS and neural networks (Chen et al., 2021).
-4.3.3 Google Research Football
+4.3.3. Google Research Football
 Google Research Football (GRF) (Kurach et al., 2020) is an open-source football simulator emphasizing high-level actions.
@@
 It features long trajectories with 3000 steps per round, stochastic transitions, and sparse rewards.
-WeKick (Ziyang Li, 2020) claimed victory in the GRF competition on Kaggle (Google, 2020), which simplifies the game dynamics by allowing competitors to control only one player, either the ball carrier on offense or the nearest defender on defense.
+WeKick (Ziyang Li, 2020) wins the GRF competition on Kaggle (Google, 2020), which simplifies the game dynamics by allowing competitors to control only one player, either the ball carrier on offense or the nearest defender on defense.
 It employed self-play strategies similar to those used in league training (Vinyals et al., 2019).
-It initializes its opponent policy pool using strategies developed through RL and Generative Adversarial Imitation Learning (GAIL) (Ho & Ermon, 2016) to facilitate the training process.
+It initializes its opponent policy pool using strategies developed through RL and Generative Adversarial Imitation Learning (GAIL) (Ho and Ermon, 2016) to facilitate the training process.
 Further research delves into the full football game rather than the simplified version.
@@
 FXP achieves a win rate of over 94% against TiKick.
-TiZero (Lin et al., 2023), a follow-up to TiKick, combines curriculum learning with FSP and PFSP (Vinyals et al., 2019) to avoid expert data reliance, achieving a higher TrueSkill rating (Herbrich et al., 2006) than TiKick.
-Table 2: Overview of representative studies in empirical analysis.
-Category Game Number of Perfect Number of Algorithms Self-play Method Search Expert Data
-Players Information Game States
-Board Game Chess 2 $\checkmark$ $10^{45}$ AlphaZero Vanilla SP $\checkmark$ $\times$
-MuZero Vanilla SP $\checkmark$ $\times$
-Go 2 $\checkmark$ $10^{360}$ AlphaGo FSP $\checkmark$ $\checkmark$
-AlphaGo Zero Vanilla SP $\checkmark$ $\times$
-AlphaZero Vanilla SP $\checkmark$ $\times$
-MuZero Vanilla SP $\checkmark$ $\times$
-Stratego 2 $\times$ $10^{535}$ DeepNash R-NaD $\times$ $\times$
-Card Game HULHE 2 $\times$ $10^{17}$ Cepheus CFR+ $\checkmark$ $\times$
-NFSP NFSP $\times$ $\times$
-HUNL 2 $\times$ $10^{164}$ DeepStack CFR+ $\checkmark$ $\times$
-Libratus MCCFR $\checkmark$ $\times$
-ReBel CFR-D / FSP $\checkmark$ $\times$
-Six-player No-limit Texas Hold’em 6 $\times$ $>10^{164}$ Pluribus MCCFR $\checkmark$ $\times$
-DouDiZhu 3 $\times$ $10^{76}\sim 10^{106}$ DeltaDou Vanilla SP $\checkmark$ $\times$
-DouZero Vanilla SP $\times$ $\checkmark$
-PerfectDou Vanilla SP $\times$ $\times$
-Mahjong 4 $\times$ $10^{169}$ Suphx Vanilla SP $\checkmark$ $\checkmark$
-Video Game StarCraft II 2 $\times$ / AlphaStar FSP & PFSP $\times$ $\checkmark$
-Dota 2 10 $\times$ OpenAI Five Vanilla SP & PFSP $\times$ $\times$
-Honor of Kings 10 $\times$ MOBA AI $\delta$-uniform SP $\checkmark$ $\checkmark$
-Google Research Football 22 $\checkmark$ TiZero FSP & PFSP $\times$ $\times$
-5 Open Problems and Future Work
+TiZero (Lin et al., 2023), a follow-up to TiKick, combines curriculum learning with FP and PFSP (Vinyals et al., 2019) to avoid expert data reliance, achieving a higher TrueSkill rating (Herbrich et al., 2006) than TiKick.
+Table 2.
+Overview of representative studies in empirical analysis.
+Category Game Number of Perfect Number of Algorithms Self-play Method Search Expert Data Players Information Game States Board Game Chess 2 $\checkmark$ $10^{45}$ AlphaZero Vanilla SP $\checkmark$ $\times$ MuZero Vanilla SP $\checkmark$ $\times$ Go 2 $\checkmark$ $10^{360}$ AlphaGo FP $\checkmark$ $\checkmark$ AlphaGo Zero Vanilla SP $\checkmark$ $\times$ AlphaZero Vanilla SP $\checkmark$ $\times$ MuZero Vanilla SP $\checkmark$ $\times$ Stratego 2 $\times$ $10^{535}$ DeepNash R-NaD $\times$ $\times$ Card Game HULHE 2 $\times$ $10^{17}$ Cepheus CFR+ $\checkmark$ $\times$ NFSP NFSP $\times$ $\times$ HUNL 2 $\times$ $10^{164}$ DeepStack CFR+ $\checkmark$ $\times$ Libratus MCCFR $\checkmark$ $\times$ ReBel CFR-D / FP $\checkmark$ $\times$ Six-player No-limit Texas Hold’em 6 $\times$ $10^{164}$ Pluribus MCCFR $\checkmark$ $\times$ DouDiZhu 3 $\times$ $10^{76}\sim 10^{106}$ DeltaDou Vanilla SP $\checkmark$ $\times$ DouZero Vanilla SP $\times$ $\checkmark$ PerfectDou Vanilla SP $\times$ $\times$ Mahjong 4 $\times$ $10^{169}$ Suphx Vanilla SP $\checkmark$ $\checkmark$ Video Game StarCraft II 2 $\times$ / AlphaStar FP & PFSP $\times$ $\checkmark$ Dota 2 10 $\times$ OpenAI Five Vanilla SP & PFSP $\times$ $\times$ Honor of Kings 10 $\times$ MOBA AI $\delta$-uniform SP $\checkmark$ $\checkmark$ Google Research Football 22 $\checkmark$ TiZero FP & PFSP $\times$ $\times$
+5. Open Problems and Future Work
 Self-play methods have shown remarkable performance by leveraging iterative learning and adapting to complex environments.
-However, significant challenges and open questions remain, presenting opportunities for further research and development.
-5.1 Theoretical Foundation
+However, significant challenges and open questions remain, presenting opportunities for further research.
+5.1. Theoretical Foundation
 Although NE has been shown to exist in games with finite players and finite actions (Nash et al., 1950), computing NE with self-play algorithms in larger games remains challenging and consequently, many studies aim to achieve approximate NE (Li et al., 2024).
 However, in some cases, even computing an approximate NE is difficult (Daskalakis, 2013).
-Some research has resorted to higher levels of equilibrium, such as CE (Marris et al., 2021) and $\alpha$-rank (Muller et al., 2020).
-Although many algorithms are developed with theoretical foundations, there is often a gap in complex real-world scenarios.
-For example, although self-play algorithms like AlphaGo (Silver et al., 2016), AlphaStar (Samvelyan et al., 2019), and OpenAI Five (Berner et al., 2019) achieve empirical success, they lack formal game theory proofs behind their effectiveness.
-Future work should focus on bridging this gap by developing new theoretically sound and practically effective algorithms or proving the theoretical underpinnings of existing successful self-play algorithms in complex environments.
-5.2 Non-stationarity of the Environment
+Some research has resorted to alternative solution concepts, such as CE (Marris et al., 2021) and $\alpha$-rank (Muller et al., 2020).
+While many successful self-play algorithms are grounded in theoretical intuition, formal game-theoretic guarantees for their effectiveness in large, imperfect-information games remain limited.
+For instance, approaches such as AlphaGo (Silver et al., 2016), AlphaStar (Samvelyan et al., 2019), and OpenAI Five (Berner et al., 2019) achieve empirical success.
+However, under realistic constraints such as finite computational resources and limited exploration, their convergence or optimality remains difficult to formally characterize within standard game-theoretic frameworks.
+Future work may benefit from further narrowing the gap between empirical success and theoretical understanding, either by extending existing theoretical tools to more complex settings or by designing new self-play methods with provable guarantees under realistic assumptions.
+5.2. Non-stationarity of the Environment
 In the self-play framework, the opponents are a vital component of the environment, and the strategies of the opponent players evolve as training progresses.
@@
 For example, incorporating opponent modeling (Zhao et al., 2022b) into self-play can help agents anticipate changes in opponent strategies and adjust their own strategies proactively, making them more robust to environmental changes.
-5.3 Scalability and Training Efficiency
+5.3. Scalability and Training Efficiency
 The scalability of self-play methods faces significant challenges as the number of teams and players within those teams increases.
@@
 This issue is particularly pronounced in regret-minimization-based self-play algorithms, which must store the regrets for each information set and potential action.
-Managing both the computational load and storage requirements is essential for improving the overall training efficiency and scalability of self-play methods.
-5.4 With Large Language Models
+Managing both the computational load and storage requirements is essential for improving the training efficiency and scalability.
+5.4. With Large Language Models
 With their remarkable capability and emergent generalizability, large language models (LLMs) have been regarded as a potential foundation for achieving human-level intelligence (Achiam et al., 2023), and self-play methods have been proposed to fine-tune LLMs, enhance LLMs’ reasoning performance, and build LLM-based agents with strong decision-making abilities.
@@
 Another line of work (Munos et al., 2024; Swamy et al., 2024; Wu et al., 2024) formulate the alignment problem as a two-player constant-sum game and use self-play methods to solve the game.
-Some other work (Huang et al., 2022; Yuan et al., 2024) also utilizes model-generated data to fine-tune LLMs with minimal human annotations.
+Some other work (Huang et al., 2023; Yuan et al., 2024) also utilizes model-generated data to fine-tune LLMs with minimal human annotations.
 The idea of self-improvement has also been applied to improve the reasoning ability of LLMs.
-Self-Play of Adversarial Game (SPAG) (Cheng et al., 2025) observes that self-play in a two-player adversarial language game called Adversarial Taboo can boost the LLM’s performance on various reasoning benchmarks.
+Self-Play of Adversarial Game (SPAG) (Cheng et al., 2024) observes that self-play in a two-player adversarial language game called Adversarial Taboo can boost the LLM’s performance on various reasoning benchmarks.
 Besides improving the capability of LLMs, self-play methods have also contributed to building LLM-based agents with strong strategic abilities.
-A representative work is Cicero († et al.(2022)(FAIR)†, Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al., FAIR) which achieves human-level play in the game of Diplomacy by combining language models with an RL policy trained by self-play.
+A representative work is Cicero (Meta et al., 2022), which achieves human-level play in the game of Diplomacy by combining language models with an RL policy trained by self-play.
 Cicero uses the self-play policy to produce an intended action and prompts the language model to generate languages conditioned on the policy’s intent.
-Another work (Xu et al., 2024) also combines LLM with self-play policy but takes a different approach by first prompting the LLM to propose multiple action candidates and then using the self-play policy to produce the action distribution over these candidates.
+Another line of work (Xu et al., 2024, 2025a) also combines LLM with self-play to train strategic language agents in the Werewolf game using reinforcement learning or counterfactual regret minimization.
 Despite recent progress, applying self-play with LLMs is still underexplored and requires further research.
-5.5 Real-World Applications
-Self-play is particularly effective in addressing some problems abstracted from real-world situations through its iterative learning approach.
+5.5. Real-World Applications
+Self-play is effective in addressing some problems abstracted from real-world situations through its iterative learning approach.
 In the field of economics, for instance, self-play is used to enhance supervised learning models in multi-issue bargaining tasks (Lewis et al., 2017).
 Furthermore, self-play proves beneficial in solving combinatorial optimization problems (COPs) such as the traveling salesman problem (TSP) and the capacitated vehicle routing problem (CVRP) (Wang et al., 2024).
-Within the domain of traffic, self-play facilitates the development of human-like autonomous driving behaviors (Cornelisse & Vinitsky, 2024) and enables vehicles to learn negotiation strategies to merge on or off roads (Tang, 2019), although currently within 2D simulators.
-However, a notable challenge with self-play is its impracticality for direct application in real-world scenarios.
-Its iterative approach requires extensive trial and error, which is computationally demanding and often involves actions that are impractical or unsafe outside of a controlled simulation.
-Therefore, self-play is often carried out in simulators.
-The effective deployment of self-play in real-world applications boils down to overcoming the Sim2Real gap.
-For example, for tasks where the Sim2Real gap is less pronounced, self-play can be well employed to enhance video streaming capabilities (Huang et al., 2021b), and to address the image retargeting problem (Kajiura et al., 2020).
+Within the domain of traffic, self-play facilitates the development of human-like autonomous driving behaviors (Cornelisse and Vinitsky, 2024) and enables vehicles to learn negotiation strategies to merge on or off roads (Tang, 2019), although currently within 2D simulators.
+Deploying self-play directly in the physical world remains challenging.
+Its iterative, trial‑and‑error nature is computationally expensive and can produce behaviors that are impractical or even unsafe outside a controlled environment.
+Consequently, most recent work performs self‑play in simulators, and practical adoption hinges on bridging the simulation‑to‑reality (Sim2Real) gap.
+When this gap is small, self‑play has already proved useful.
+For instance, self-play can be well employed to enhance video streaming capabilities (Huang et al., 2021b), and to address the image retargeting problem (Kajiura et al., 2020).
 Moreover, EvoPlay (Wang et al., 2023) leverages self-play to design protein sequences, utilizing the advanced AlphaFold2 simulator (Jumper et al., 2021) to narrow the Sim2Real gap.
-Similarly, in heterogeneous multi-robot systems, self-play is utilized for competitive sports tasks like humanoid football (Liu et al., 2022b; Haarnoja et al., 2024), quadruped competition (Xiong et al., 2024), and multi-drone volleyball (Xu et al., 2025), with significant efforts dedicated to Sim2Real transitions for real-world success (Gao et al., 2023).
-6 Conclusion
+In multi-robot systems, self-play has been recently utilized for competitive robot sports tasks like quadruped and humanoid football (Liu et al., 2022b; Haarnoja et al., 2024; Xiong et al., 2024; Su et al., 2025a), robot-arm and humanoid table tennis (DAmbrosio et al., 2025; Su et al., 2025b), and multi-drone volleyball (Xu et al., 2025b; Zhang et al., 2025), with substantial efforts dedicated to Sim2Real transitions for real-world success.
+A complementary research direction is to boost the sample efficiency of self-play so that policies can be trained online, reducing or even eliminating the need for elaborate Sim2Real pipelines in real-world deployments.
+6. Conclusion
 The burgeoning field of self-play in RL has been systematically explored in this survey.
-The core idea of self-play that agents interact with copies or past versions of themselves has proven to be a powerful approach for developing robust strategies across various domains.
+The core idea of self-play, that agents iteratively refine their policies by interacting with historical or concurrent versions of themselves or other evolving agents, has proven to be a powerful approach for developing robust strategies across various domains.
 Before delving into the specifics of self-play, this paper first elucidates the MARL framework and introduces the game theory background of self-play.
@@
 We provide detailed explanations on how these four groups are seamlessly integrated into our proposed framework, ensuring a comprehensive understanding of their functionalities.
+We then compare the four categories within this common lens and draw design insights that can guide future self-play research.
 The transition from theory to application is underscored by a rigorous analysis of self-play’s role within various scenarios, such as board games, card games, and video games.
@@
 This survey serves as an essential guide for researchers and practitioners, paving the way for further advancements in this field.
-Appendix A Appendix
-A.1 Abbreviations in Alphabetical Order
-ABR Approximate Best Response
-AI Artificial Intelligence
-ARM Advantage-based Regret Minimization
-ARMAC Advantage Regret-Matching Actor-Critic
-BR Best Response
-CCE Coarse Correlated Equilibrium
-CE Correlated Equilibrium
-CFR Counterfactual Regret Minimization
-COP Combinatorial Optimization Problem
-CSPL Curriculum Self-Play Learning
-CVRP Capacitated Vehicle Routing Problem
-DCFR Discounted Counterfactual Regret Minimization
-DO Double Oracle
-EGTA Empirical Game-Theoretic Analysis
-FP Fictitious Play
-FSP Fictitious Self-Play
-FTW For The Win
-FXP Fictitious Cross-Play
-GAIL Generative Adversarial Imitation Learning
-GRF Google Research Football
-HULHE Heads-Up Limit Texas Hold’em
-HUNL Heads-Up No-Limit Texas Hold’em
-JPSRO Joint Policy-Space Response Oracle
-LCFR Linear Counterfactual Regret Minimization
-LLM Large Language Model
-MARL Multi-Agent Reinforcement Learning
-MCCFR Monte Carlo Counterfactual Regret Minimization
-MCTS Monte Carlo Tree Search
-MDP Markov Decision Process
-MG Markov Game
-ML Machine Learning
-MOBA Multiplayer Online Battle Arena
-MSS Meta-Strategy Solver
-NE Nash Equilibrium
-NeuPL Neural Population Learning
-NFSP Neural Fictitious Self-Play
-ODO Online Double Oracle
-PBR Preference-based Best Response
-PBS Public Belief State
-PFSP Prioritized Fictitious Self-Play
-POMG Partially Observable Markov Game
-PSRO Policy-Space Response Oracle
-RCFR Regression Counterfactual Regret Minimization
-ReBeL Recursive Belief-based Learning
-RL Reinforcement Learning
-R-NaD Regularized Nash Dynamics
-RTS Real-Time Strategy
-SD-CFR Single Deep Counterfactual Regret Minimization
-SPAG Self-Play of Adversarial Game
-SPIN Self-Play fIne-tuNing
-TSP Traveling Salesman Problem
-URR UnRestricted-Restricted
-XDO Extensive-Form Double Oracle